A Complete Bibliography of the Journal of Statistical Software

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

30 November 2022 Version 2.14

Title word cross-reference

(X̄, s_{c}) [ECK04]. (X̄, v_{c}) [ECK04]. (p, q) [IMR17]. 1, 2, 3 [SMDS11]. 2 [EN11, GKSU15, Grö14]. 2^{k} [Law08]. 2^{k−p} [Law08]. 2 × 2 [ILS11]. 3 [AMMRP14, GGK10, LPLPD14]. 4 [HWY18]. $69.99 [Pos15]. α [LPLPD14, dVSWAL17]. F [Ram00]. G [NE20]. K [KSBZ16, PG15, HFKB12, Lei10, MAK06]. N [MMR20, HMR^{+}13]. p [MF15, ZD17]. Q [JWHS16]. T [Pat00, AHvD09, LM13, MP99, Som98, Som01].

-Components [MP99]. -function [PG15]. -Gram [HMR^{+}13]. -Level [Grö14]. -Means [HFKB12, KSBZ16]. -Mixture [MMR20]. -Modeling [NE20]. -Modes [Lei10]. -Parameter [HWY18]. -Shape [LPLPD14]. -Stable [dVSWAL17]. -State [AMMRP14]. -Statistics [JWHS16].

//dsarkar.fhcrc.org/lattice/ [Gro08].

//global.oup.com/academic/product/9780199660346 [How15a]. //ukcatalogue.oup.com/product/9780198729068.do [Hil15b]. //ukcatalogue.oup.com/product/9780199671137.do [Han15]. //ukcatalogue.oup.com/product/9780199827633.do [Hil15a]. //www.amherst.edu/ [Grö15a]. //www.ceremade.dauphine.fr/ [Pos15]. //www.crcpress.com/9781498709576 [Grö16]. //www.crcpress.com/9781498711548 [Gle16]. //www.crcpress.com/9781498712361 [Kha16b]. //www.crcpress.com/9781498715232 [Zei15]. //www.imperial.ac.uk/bio/research/crawley/statistics/ [Hel15]. //www.r [Iac15]. //www.r-datacollection.com/ [Iac15]. //www.sciencedirect.com/science/book/9781785480058 [How15b]. //www.springer.com/9783319140926 [Zei16]. //www.springer.com/9783319234458 [How16b]. //www.springer.com/9783319238821 [Mat16a]. //www.springer.com/book/9781461487746 [Kha15b]. //www.springer.com/book/9781493909827 [Nun15]. //www.springer.com/book/9781493917013 [Mat15]. //www.springer.com/book/9783319144351 [Rus15]. //www.springer.com/book/9788132223399 [Hof15]. //www.wiley.com/WileyCDA/WileyTitle/productCd [But15, Kha15a]. //www.wiley.com/WileyCDA/WileyTitle/productCd-111833258X.html [But15]. //www.wiley.com/WileyCDA/WileyTitle/productCd-1118762614.html [Kha15a].

1.1 [Ber01]. 1.4.2 [Kar06]. 1.7.3 [Ros08]. 111833258X.html [But15]. 1118762614.html [Kha15a]. 13ARIMA [SE18]. 16 [HHJCRB21].

2.0 [GK16, KCM08, SBL^{+}10, Sch06]. 2.5 [Sto12]. 2011 [LCK11]. 2nd [CH11, Dem17, DN17, Dow19, Edd09a, GR17a, GR18a, Grö15a, Hel15, Hel16b, Hel17, Hel18, Kha16c, Kha17a, Kha18c, Kha20, KN17, Lor18, Mai09c, SCD07].

3.0 [WK18, WWH04, Ziv11, Mac12]. 3rd [Pav20, San10a].

4th [Ros17].

6.4 [MLVMY05].

90 [KKH10, KH13, Mil04, MHdL12, Som98, SB01]. 95 [Gen06]. 978-0-19-872906-8 [Hil15b]. 978-0-19-966034-6 [How15a]. 978-0-19-967113-7 [Han15]. 978-0-19-982763-3 [Hil15a].

978-0-387-75968-5 [Gro08]. 978-1-118-33258-0 [But15]. 978-1-118-83481-7 [Iac15]. 978-1-118-94109-6 [Hel15]. 978-1-119-95021-9 [Kha15a]. 978-1-4614-8686-2 [Pos15]. 978-1-4614-8774-6 [Kha15b]. 978-1-4822-3736-8 [Grö15a]. 978-1-4939-0983-4 [Nun15]. 978-1-4939-1701-3 [Mat15]. 978-1-4987-0957-6 [Grö16]. 978-1-4987-1154-8 [Gle16]. 978-1-4987-1236-1 [Kha16b]. 978-1-4987-1523-2 [Zei15]. 978-1-78548-004-1 [How15b]. 978-3-319-14092-6 [Zei16]. 978-3-319-14435-1 [Rus15]. 978-3-319-23446-5 [How16b]. 978-3-319-23882-1 [Mat16a]. 978-81-322-2339-9 [Hof15]. 9780198729068.do [Hil15b]. 9780199660346 [How15a]. 9780199671137.do [Han15]. 9780199827633.do [Hil15a]. 9781461487746 [Kha15b]. 9781493909827 [Nun15]. 9781493917013 [Mat15]. 9781498709576 [Grö16]. 9781498711548 [Gle16]. 9781498712361 [Kha16b]. 9781498715232 [Zei15]. 9781785480058 [How15b]. 9783319140926 [Zei16]. 9783319144351 [Rus15]. 9783319234458 [How16b]. 9783319238821 [Mat16a]. 9788132223399 [Hof15].

A3 [FR15]. Abandoning [dL05d]. Aberration [SSH16]. Absence [FM08]. Abstraction [BLM^{+}15]. Abundance [Fie12, FC11, MMR20]. academic [How15a]. Accelerated [CKY14, HXY12]. Acceleration [DV20]. Accept [Hüs06]. Accept-and-Reject [Hüs06]. Acceptance [CC11, Kie08]. AcceptanceSampling [Kie08]. Accessing [DC09]. Accounting [CKE20]. Accuracies [Ngu07]. Accuracy [ALO^{+}10, NAA17, Whe14, AGM07]. Accurate [AGM07, Hug07, Pat00]. acebayes [OWA20]. Acquiring [Grö18a]. across [GW18]. acs [GR17b]. Action [Edd12b]. Active [dLHM09]. actuar [DGP08]. Actuarial [DGP08, Spe13]. ada [CJM06]. adabag [AGG13]. Adaptive [AHvD09, Cha16, DHM11, FS20, Ham10, Jam99, KPH^{+}21, MR12, MB17, MMB15, MJGM10, PT07, PT09, PPT20, RSW15, WLK08]. Add [But09, KR09, SH17]. Add-In [But09, KR09]. Adding [Kla18]. Additional [CFSR15, Den16]. Additive [BKL05, GR18a, GRS12, Han06a, HC05, KB16, KV13, LM19, LC10, Mai06, Pet97b, Sch11, SSM21, SR07b, UAK^{+}15, Woo16]. ade4 [BD18, DD07, TD07]. ade4TkGUI [TD07]. adehabitat [Cal07]. adimpro [PT07]. Adjacent [dLHM09]. Adjusted [SFC20, WZ16, MAK06]. Adjustment [KK15, KMTS14, SE18]. AdMit [AHvD09]. adoptr [KPH^{+}21]. ads [PG15]. Advanced [Wic10]. Advances [Str10]. Advection [SKS15]. Advection-Diffusion [SKS15]. After [Mar06a]. aftgee [CKY14]. Age [Jon07, MAK06, SH07]. Age-adjusted [MAK06]. Age-Period-Cohort [SH07]. Age-structured [Jon07]. Agent [RGK20, YZZ^{+}20]. Agglomerative [Mül13, Rec10a]. Aggregation [Lan14]. Agricultural [Ban16]. Aid [NV11]. Aided [SH09]. Ajay [Mat15]. Alchemy [Mat16c]. Algebra [BE13, KOS20, Kha16b, dL08]. Algebraic [Ost20]. Algorithm [BM12, Cha03, CLL17, FS13, HC05, HI20, Hüs06, JMD08, KSBZ16, Kon13, LM13, MNA^{+}97, MP99, PK08, RK14, Wol12, Wol15, dLHM09]. Algorithms [BMB16, But15, CBA20, DV20, GFS14, Jur15, MGG^{+}04, MYK07, MCM12, Mon06, NV11, SBL^{+}10, Scr13, Scu17, Tri03, YEL18, ZFZ10]. Aligning [CS12]. Alignment [KSS^{+}07]. Alignments [Gio09]. alphahull [PLRC10]. Alternative [Kam08]. amei [MJGM10]. Amelia [HKB11]. American [GR17b]. amsrpm [KSS^{+}07]. anacor [dLM09c]. Analogue [MHM20, Sim07, Sim07]. Analyses [CCD09, Dha10, HM18, Hil06, Pen08b, Vie10, Wil09]. Analysing [KWE^{+}17]. Analysis [Ach09, ÁBZ20, ASBMB17, Arm19, BM07, BL14, Bai20, Bak20, Ban99, BP12, Bea17, Bec18, BFC02, BBG12, BK17, BER^{+}16, BP19b, BGRR15, BM96, BM97b, BPGC14, BDdM11, BPB09, Bos11, BD18, Bow10, Boy05, Bur12, But08b, Can04a, CFHBK11, CKY14, CFSR15, CSY15, CRW05, CG10, CO16, CL13, DCS20, DN99, DM19, Duo07, DHF15, EE07, EMU20, EL09, Eva11, FBdlF12, Fel12, Fer11, FDGD16, iM16, FF14, FC12, FS20, FM08, FKP17, FGEMI12, GFPB19, GP12, GK16, GKD14, GKSU15, Góm15, GCA12, GDMB08, GU07, GT10, GVM16, Grö10a, Grö15a, GR18b, GV12, HD18, HHB^{+}08b, HSL11, Han05, Har08, HN13, HP09, Hel17, Hil09a, Hil09b, Hil10a, Hla16b, HB17, How16b, HPK18, HSG12, HJS16, JM15, JD15, JCPL19, Jon07, JH16, Kav15, KK15, Kha17b]. Analysis [Kha18a, Kha18b, KE14, KS10, Kle16, KR09, Kro16, KT10, KKH10, Lal17, LDHM21, LCK11, LRGTA12, LCL^{+}08, Law08, LJH08, LMO^{+}19, Lei10, Len09a, LFF17, LZHC17, LL11, LC10, LSL20, LRN18b, dUJ13, LSvdVK09, LSvdV19, Lor18, LV16, Lum98, Lum04, LCSC14, LWOW18, MF14, MRV21, MS19, MdL10, MHM20, MDvdV19, MGG^{+}04, Mat16a, MF15, MCAP19, MYK07, MKC21, Meu13, MD18, MHH17, MIG12, MBK18, LMB^{+}07, NGVCFLS19, NG07, NO06, NGBK12, Nun15, Obe14, Ost20, OK14, PB15, PPGD15, PP18, PV20, RBN19, RJH14, RS05, Riz06, RFKM12, RMG12, RGD12, Röv20, San06, San10a, SBL^{+}10, SMM^{+}15, Sch06, She11, SP10, Spa06, SLMV17, SGS^{+}14, TV11, TPE19, TMN16, TM05, TD07, TYH^{+}14, TF09, TBE^{+}21, ÜS10, VM09a, Ver18, Vin17, WLH^{+}18, WdVW^{+}11]. Analysis [WO97, WS11, WM14, WG10, Wic11, Wil14a, WEH11, YR19, Yee10, ZG11, Zei15, Zha14, ZQS16, Zha07, ZK12, dS11, dL05b, dLM09c, dWFP11, vdA07, vdA12, vdHvB16, VMY02, GR17a, Hel16b, Edd09a, Eks13, Ell09, Fri12, How16a, Mai09c, VM10, dL05a]. Analytics [Kil16, Lal16, Otn17, Rus15]. Analyze [EBO^{+}13, GW18, MdUÁC10]. AnalyzeFMRI [BDdM11]. Analyzing [BT05, BCHY09, BKL05, CB17, DHM11, Eva14, GRMS11, GR16a, GRD13, Gos11, Grö14, KAK05, LX12, MMRP11, Mug10, OO11, RR11, SWAF15, SZ11, Sch16, SM07, dI17, Dow19]. Anchoring [WKL11]. anchors [WKL11]. and/or [KM16]. Anderson [MM04]. Andy [Hil15b]. Angoff [MF14]. Animal [MMR20, MLAN02]. Animated [NL12]. Animating [BdMM08]. animation [Xie13]. Animations [Xie13]. ANOVA [DW17, Gu14, IM05, PKC^{+}20]. ANOVA robust [PKC^{+}20]. Antagonism [BN10]. Anthropometric [Vin17]. Anthropometry [Vin17]. Antony [Zei15]. Any [BM12]. app2web [VFV13]. applet [WO97]. Application [BFRP13, HNCI18, Kel07, MH07, SWAF15, WM18, Wol15]. Applications [Cha16, GR16b, HR14, Hel16b, Jos05, Mül16, dREP12, SBL04, VFV13, dL05c, San09, Wic10]. Applied [Dry09, Sar09, Edd09b, JS13, Kha16a, Lor18, SL09, Zei16]. Apply [Wic11]. Approach [BV17, DKMT11, EN11, FGEMI12, Gou10, Hil07, Lal17, Mat16b, MSCCV05, Mil17, RJH14, Sar09, SCS13, VS14, San10a, Fay18, Kuk09, Otn17]. Approaches [Len09a, PU13]. Approximate [Gra16, OWA20]. Approximation [HH14, KNB^{+}16, LPLPD14]. Arbitrary [BH08a]. ArcGIS [PV14]. Archaeological [PV20]. ArchaeoPhases [PV20]. Archetypal [EL09]. Archimedean [GHH20, HM11, OR14]. archivist [BK17]. Area [HMM21, LW03]. Areal [GF15, LRN18a]. arf3DS4 [WdVW^{+}11]. Arguments [Hof11, Hug07]. ARIMA [Bel11b]. Arithmetic [BH07, BH08b, KJ14]. Arm [JPM19, KWE^{+}17, LZHC17]. Arm-Based [LZHC17]. Array [BS18, LBW18]. Arrow [TKM07]. Art [SME^{+}09, Lan14]. Arthur [Han15]. arules [HGH05]. ASMap [TB17]. aspect [MdL10]. Aspects [MHH08]. Assemblage [Yua07]. Assembly [GC18]. Assess [Bar96]. Assessing [BN10, CBAA19, Kie08, LQC^{+}12, ZFH^{+}20]. Assessment [AACRC21, CB17, JvdBP07, SlRC16, Smi07, USHH18, WEH11]. Assessments [ZUL14]. Asset [TBS14]. Assistant [PE19]. Association [HGH05, HSG12, MF15, SL09, WZL20, ZFZ10, ZPZW21]. Associations [BGM06]. AssocTests [WZL20]. Asymmetry [LS02a]. Asymptotic [BLR97, FS10a, ZFZ10]. Asypow [BLR97]. Attachment [PSS20]. Attributable [LVK^{+}11]. Attributes [BGM06]. Augmentation [TDRD15]. Authoring [Mor18]. Automated [AM11, CdM10, PRS18, Sek11, Iac15]. Automatic [GZ09, HK08, KNB^{+}16, Nas05]. Automatically [HL09]. Autoregression [MZ08]. Autoregressive [ABC19, KLNN20, Lee13, LRN18a, Tri03, Wan13b]. Auxiliary [TKM07]. Average [AW16, HPWdL15]. Averaging [BK15, CN18, WB08, ZF15].

B [Hil15a, WG00]. B/D [WG00]. B2Z [MBR11]. BACCO [Han05]. Back [Kro16, WDT^{+}12, vdHvB16]. Back-End [WDT^{+}12]. Bagging [AGG13]. Balance [Sek11]. Balanced [IM05]. Balancing [HX13]. Ball [ZPZW21]. bamdit [Ver18]. BAMP [SH07]. Band [WZ16]. Bandit [KvEPK20]. Bands [JLZJ05]. Bandwidth [dREP12]. BANOVA [DW17]. Bar [HR14, TKM07]. BARD [AM11]. BARS [WLK08]. BART [SSM21]. bartMachine [KB16]. Base [Hel16a, Wei12]. Baseball [Dow19, Dow18]. Based [AG19, AR14, AWBM18, BBG12, BIG17, BP19b, BYPJ19, Bra15, Bro15, BR17, CCF19, CBA20, CQZ^{+}10, CB21, CYK^{+}09, CFT11, Cul11, DP12, ESKHB20, FHH17, FS13, Fuj17, GRMMR19, GFS14, HPD19, HD12, HMM21, HMR^{+}13, IoG10, JLH^{+}20, KBMB19, Kle16, KK14, KSPR16, Lal17, LIL^{+}15, LSPvdL17, LR20, LZHC17, LBC^{+}16, LG11, MSCCV05, Mel16, MVS13, MNT17, NO06, Nor17, NAA17, PMD19, QY19, RBHB15, RGD12, SL09, SWAF15, San10a, SR18, Scu17, SLS^{+}12, TM05, TFR16, TBE^{+}21, WMS17, Wee10, XTL13, YFK^{+}15, ZK12, BM07, Ban99, FR07, GU07, MLAN02, MAK06, Gum07]. Baseline [KMG^{+}13]. Bases [KT10]. Basic [Aiz12, Fox05b, PLR^{+}16, Kha03]. Basic-Statistics [Fox05b]. Basics [Fox05a, Kha16b]. Basis [Pac07]. BASS [FS20]. Batch [BLM^{+}15, Hof11]. BatchExperiments [BLM^{+}15]. Batching [Hof11]. BatchJobs [BLM^{+}15]. Bayes [JS05, KG17, MHdL12, NE20, Woo05, WG00]. bayesclust [GFC12]. Bayesian [Eks13, Hil07, AGZ19, BS15, BGH^{+}17, BV17, BK15, Bos11, BD03, BKL05, Bür17, Bur12, CF14, Can17, CRW05, CG10, DR20, DW17, DN99, EE07, iM16, FS20, GF15, GG16, GTDT20, Gra07, GR18b, Han05, JHQ^{+}11, JCPL19, KB16, Lee13, LR15a, LBC^{+}16, MS19, MHE20, MSCCV05, MR18, MW19, MBR11, Mur15, ND12, NAA17, OK14, OWA20, Pac07, PPGRMB21, PHF10, Röv20, SP14, Sav16, Sch11, SH07, Scu10, Scu17, She10, SCS13, SSM21, SBMD14, SMS13, TDRD15, TR17, Ver18, VR14, WLK08, WLH^{+}18, WM14, WEH11, YZZ^{+}20, YR19, YMB20, ZF15, Zha14, ZLC18, ZHZ20, dS11, Dem17, Ham10, Hil07, Hil10a, Pos15]. BayesLCA [WM14]. bayesmeta [Röv20]. BayesNetBP [YMB20]. bayesPop [SR16]. bayesQR [BV17]. bayesTFR [SAR11]. BayesX [BKL05, UAK^{+}15]. BB [VG09]. BClass [MSCCV05]. bclust [ND12]. bcp [EE07]. bcrm [SMS13]. BDgraph [MW19]. BDM [MH05]. Beanplot [Kam08]. beanz [WLH^{+}18]. Behavioral [Lan14]. benchden [MW12]. Benchmark [MW12]. Bergm [CF14]. Bernoulli [BS18, MHT07]. Bernstein [Can17]. Bespoke [NRD16]. BeSS [WZQW20]. Best [Day01, Nas14, WZQW20]. Beta [BRC98, CNZ10, GKZ12, MP14, NAA17]. Beta-Binomial [NAA17]. Better [AM11]. Between [SBMG06, WDT^{+}12, WZ16]. Beyond [MDvdV19, PT11, SDJ20, Var14, ZUL14]. BFDA [YR19]. bgmm [BSVT12]. Bi [TV11]. Bi-Parental [TV11]. Bias [CCF19, CHML17, FS12, NGVCFLS19]. Bias-Corrected [CCF19]. Bias-Reduced [FS12]. bibliographical [Coe00]. Bidimensional [Car13]. BiDimRegression [Car13]. BIEMS [MHdL12]. Big [How15a, Iac17]. bild [GCA12]. Binary [GCA12, LP17, MD18, Ngu07, SF19]. Binomial [AH12, DM18, KV13, NAA17, TKM17]. bio [Hel15]. bio.infer [Yua07]. Bioassay [RS05]. Bioconductor [dSdSCLC18, Unw09]. Biodiversity [Han07b]. Bioinformatics [Sch08, dL09b]. Biological [KHLF^{+}10, MSCCV05]. Biologists [Hil15b]. Biology [BFRP13, LCL^{+}08]. BioMark [WF12]. Biometrics [Gow16]. Biostatistics [HGG08]. Bipartite [AH08]. Biplot [LS02b, Udi05]. BiplotGUI [lGlRGL09]. Biplots [lGlRGL09]. Birds [EBO^{+}13]. Bivariate [EC16, GR18b, KN05, Kha16a, MH05, Mey13, MMM12, Nag18, NAA17, PZK^{+}12, SSV14]. Black [SGHY11]. blavaan [MR18]. Blind [GL14, MNT17, NMM^{+}21]. blm [KV13]. blockcluster [BIG17]. Blocked [MvB14]. Blocking [LSC09]. BMS [ZF15]. bnlearn [Scu10, Scu17]. boa [Smi07]. Boca [Gle16, Grö15a, Grö16, Kha16b, Zei15]. BOIN [YZZ^{+}20]. Book [Ach09, Arm19, Ban16, Bar18, Bow09, Bow10, Boy05, Bul06, But15, Chr09, Chr11, CH11, Cox07, Dem11, Dem17, DN17, Dem18a, Dem18b, Der10, Dow16, Dow17, Dow18, Dow19, Edd09b, Edd09a, Edd09c, Edd11, Edd12b, Edd12a, Eks10, Eks13, Ell09, Eva11, Eva14, Fay18, Fox05a, Fri12, Gen06, Gle16, GR16b, GR17a, GR17b, GR18a, Gou05, Gou10, Grö10a, Grö11, Grö13, Grö15b, Grö15a, Grö16, Grö17, Grö18a, Grö19, Grö21, Gro18c, Gro08, Gum07, Ham10, Han15, Har08, Hel15, Hel16a, Hel16b, Hel17, Hel18, Hel20, Hil06, Hil07, Hil09a, Hil09b, Hil10b, Hil10a, Hil15a, Hil15b, Hof15, How11, How15a, How15b, How16a, How16b, Iac15, Iac17, JS13, Kha15b, Kha15a, Kha16a, Kha16b, Kha16c, Kha17a, Kha17b, Kha18a, Kha18b, Kha18c]. Book [Kha20, Kil16, KN17, Kuk09, Lal16, Lal17, Lan17, Len09a, Leo10, Lor17, Lor18, Mai06, Mai07, Mai09a, Mai09b, Mai08, Mai09c, Mal09, Mat13, Mat15, Mat16b, Mat16a, McN14, MT14, Mor18, Mül16, Mun14, Mur05, Mur06, Nun15, Oom10, Otn17, Pav16, Pav20, Pie05, Pos15, Ros09, Ros17, Rui16, Rus15, Sab07, SL09, San06, San07, San09, San10b, San10a, San11, SCD07, Sar09, Sch08, She11, Soe10, Spa06, Str10, Unw09, VM09b, VM09a, VM10, Web16, Wic08b, Wic08a, Wic10, WG10, Wil14a, Wil09, Wil14b, Xu06, Yal10, Zei15, Zei16, Zho10, dL05b, dL05c, dL05a, dL06, dL07, dL08, dL09a, dL09c, dL09b, dL12]. bookdown [Mor18]. Books [Mor18]. books.html [Pos15]. Boosting [AGG13, BRG20, CJM06]. Bootstrap [HH14, LCD18, MAK06, VW13, VdL09]. Bootstrap-based [MAK06]. Boruta [KR10b]. Boundary [DLN17, HS18]. Bounds [Hla16b]. Box [EB03, SGHY11]. Box-Percentile [EB03]. Boxplot [Kam08]. BPEC [MHE20]. Bradley [Fir05, TF12]. BradleyTerry2 [TF12]. Brain [RR11]. Breaks [PRS18]. bridgesampling [GSW20]. Bring [Tiv19]. brms [Bür17]. Brownian [Coe00]. BRugs [KCCG11, MW10]. bsamGP [JCPL19]. bspmma [Bur12]. BSSasymp [MNT17]. btergm [LCD18]. Bubbles [Cas17]. BUGS [Eks13, Cur10, GTDT20, Ntz02]. BugsXLA [Woo05]. Build [Per03]. Building [CG15, HMS16, Kuh08, Lum96, SSV14]. Bundle [Han05]. Burg [Tri03]. Business [Otn17]. BYM [GF15].

C [Kha15b, BS13, BFM18, EF11, Joh09, Nag21, PQM11, dVSWAL17, SCS13, WLK08, WZ17, XGY06, Zho15]. C- [BS13]. C/C [dVSWAL17]. c060 [SHBZ14]. ca [NG07]. Caching [Pen08b]. CADF [Lup09]. Calculate [EC06, HZBB20, PV14, TFR16, VS14, ZFZ10]. Calculating [BL18, MHdL12]. Calculation [LVK^{+}11, New05, Pat00, Som07]. Calculations [BLR97, KSPR16, LQC^{+}12, PLZ^{+}15, Spe13]. Calculator [SWAF15]. Calculus [SH20]. Calibrate [Coo97]. Call [PLLC11]. Calling [But05, EC16]. cancerclass [BKvT^{+}14]. Candidate [AHvD09]. Canonical [GDMB08, LM18, dLM09c]. Capabilities [KK14]. Capability [SFS12]. Capture [BR07, YSH15]. Capture-Recapture [BR07, YSH15]. CAR [LSvdVK09]. CARBayes [Lee13]. CARBayesST [LRN18a]. Carcinogenicity [MLAN02]. caret [Kuh08]. Carlo [BFFN19, BM97a, CGRvD15, CGC11, DK18, GB20, Joh09, KCCG11, LT16, MQP11, She08a, She08b, SP10, War02, Zho15]. Cartograms [Pan18]. Cartographic [TKM07]. Case [ADH11, HSL11, JWHS16, NMB15, Str04, Unw09, ZFZ10]. Case-Control [HSL11, JWHS16, ZFZ10]. Case-Mother [NMB15]. Catastrophe [GvdMW09]. Categorical [BPP17, GT10, Mel16, PU13, Yee10]. Categorization [HMR^{+}13]. catR [MR12, MB17]. Causal [AW16, HPWdL15, HIKS11, KMC^{+}12, RJH14, SH17, SvdLN17, TK17, TYH^{+}14]. causaleffect [TK17]. Cause [MS19]. Cause-Specific [MS19]. CCA [GDMB08]. CDF [DKT20, SS19]. CDF-Quantile [SS19]. cdfquantreg [SS19]. CDM [GRK^{+}16]. CDVine [BS13]. Cell [BFRP13]. cells [MJ00]. cem [IKP09]. Censored [AB17, FS10a, Koe08, MNA^{+}97, OL17, TGJ17]. Censoring [MMM12]. Census [Alm10]. Central [Phi10]. CEoptim [BDKL17]. CFC [MS19]. cg [PO13]. cgam [LM19]. cgeostat [dI17]. CGIwithR [Fir03]. Chain [BFFN19, Ber01, KCCG11, MQP11, She08a, She08b, VV16, War02]. Chained [RW11, vBGO11]. Chains [BMB16, Sch16]. Chan [Kha15b]. Change [EE07, JM15, MKC21, Ros15, ZLHK02]. Change-Point [MKC21]. Changepoint [KE14, KE14]. Changing [DN99]. Chapman [Gle16, Grö16, Kha16b, Rus15]. Characteristic [LW03]. Characteristics [GG16]. Characterization [BLY18]. Characterize [MTPL15]. Chart [EC06, ZP13]. Charts [But09, ECK04, EC06, Grö21, HR14]. Chemical [Guh07, HPK18, SSS^{+}20]. Chemometrics [FR07, MvS07a]. Chichester [But15, Hel15, Iac15, Kha15a]. Choice [Aiz12, Eva14, dREP12, Sar16, Sch11, TSV20]. Chris [Rus15]. Christian [Iac15, Pos15]. Chromatograms [CS12]. Cichosz [But15]. CircNNTSR [FDGD16]. CircStat [Ber09]. Circulant [DB13]. Circular [BFRP13, Ber09, FDGD16, OCRC14, Han15]. cirt [FEvdL07]. Clarify [TWK03]. Class [Bea17, CDP20, HHvdWZ08, Lei04, LHS08, LL11, MD18, PC11, Sho13, WM14]. Classes [KR10a, PLPL17, RK14]. Classical [Len09a, Sch06]. Classification [AGG13, AWBM18, Arc10, CC08, FT11, FR07, GSD12, GZP14, HNCI18, LIL^{+}15, Lee18, Mai09a, MSCCV05, PMD19, RBHB15, TBE^{+}21, Whe14, ZWH17, ZD17]. Classifiers [MLMK12, VR14]. classify [Whe14]. Cleaning [Grö18a]. Clear [FR15]. Click [Dow16]. ClickClust [Mel16]. Clickstream [Sch16, Sch16]. Client [AB12]. Climatology [Hil15a]. Clinical [CKE20, Ham10, JPM19, KPH^{+}21, LQC^{+}12, RSW15, USHH18, WWG09, WCLS20, She11]. CLME [JP16]. clogitL1 [RT14]. Closing [WDT^{+}12]. Cloud [Mat15, MY13]. CLUE [Hor05]. clues [CQZ^{+}10]. clusrank [JLH^{+}20]. Cluster [BPDD08, FGEMI12, GASA15, Gre17, Hof11, IoG10, KH08, MPI18, TSN21, Hor05]. Cluster-Randomized [Gre17]. Cluster-Weighted [MPI18]. Cluster/Multicore [Hof11]. Clustered [HC05, JLH^{+}20, ZKG20]. Clustering [AWBM18, AM14, BBG12, BIG17, BER^{+}16, Bra15, CQZ^{+}10, CKSLS12, EC16, FM18, HD10, HBF17, HPD19, HFKB12, JWHS16, KK14, KSBZ16, LH12, MHE20, MDvdV19, MSCCV05, MCM12, Mel16, MV14, Mül13, ND12, Rec10a, SR18, SHR96, TBE^{+}21, YEL18]. Clusters [CGBN14, GRMMR19, GFC12]. ClustOfVar [CKSLS12]. clustvarsel [SR18]. clValid [BPDD08]. CM [Edl04]. CM-Estimates [Edl04]. CMregr [Edl04]. Co [BIG17, GVM16]. Co-Clustering [BIG17]. Co-Occurrence [GVM16]. Coarse [MBGK18]. Coarse-Grain [MBGK18]. Coarsened [IKP09]. CoCo [Bad01]. Code [Aiz12, AD15, AW16, BDMP15, BK11, Bak20, BS18, Bel11a, CKE20, CG15, Car13, Car15, Cas17, CS12, CBAA19, CO16, Cur10, DMB18, DP12, Dha10, Dry09, EC16, Esm14, FS12, FKP17, GF15, Gos14, GC18, Gre17, GVM16, HF18, Han05, Han06a, HW07, Han07a, HZBB20, Har15, HXY12, Hla16a, Hof11, Hoh18, Hug07, Jal19, Jur15, KM16, Kam08, Kla18, Kon13, LSC09, LW16, Lec14, LR20, LQC^{+}12, LZ17, LG11, LRN18b, LV16, MF14, MB17, MHJS16, MTPL15, MBGK18, MF15, MP14, MHT07, NK06, Nag21, NAA17, OC15, Pan09, Pan18, Pap16, Pav15, PSY20, Pen08a, PZK^{+}12, PKC^{+}20, PV20, Phi10, RBN19, Rec10a, Rec10b, RPC15, RBB18, RGK20, Ruf09, Sac17, SWAF15, SBMG06, SGS^{+}14, TKM07, TR14, VSV09, VS14, VYD^{+}12, VDT14, WWG09, WZ16, WM18]. Code [Wei12, WEH11, Wol15, XMW10, YS12, dSJdSF14, dSdSCLC18, Gro18c]. Coding [dL05b]. Coefficient [BR17]. Coefficient-Wise [BR17]. Coefficients [TR14]. COGARCH [IMR17]. Cognitive [GRK^{+}16, MdlT20]. Coherent [DM19]. Cohort [LVK^{+}11, SH07, Wan13a]. Cohorts [BBGL17]. coin [HHvdWZ08]. Coincidence [EMU20]. Cointegrated [Edd09a]. Collaboration [BBGL17]. Collection [Iac15, Mil04, RG07]. Collocation [HRX16]. CollocInfer [HRX16]. Colors [CBAA19, ZFH^{+}20]. colorspace [ZFH^{+}20]. Combination [CGRvD15, YZZ^{+}20, YEL18]. Combinations [JMD08]. Combinatorics [Han07a]. Combine [Wic11]. Combining [BMGT15, Dry09]. Command [HL09, Hof11, RGK20]. Command-Line [HL09]. Commander [Dow16, Fox05b, FC12]. Comment [LZL^{+}05]. Commerce [Chr11]. Common [DH05, LSM^{+}19, Woo05]. Commonalities [Gow16]. Communicating [PPC15]. Communication [Grö19]. Community [GR17b]. Companion [dL05a, Mat13]. Comparative [ADH11, ABS01, LNY^{+}20, Coe00]. compareGroups [SSV14]. Comparing [BP15]. Comparison [JLZJ05, Kam08, Kla18, OL17, PO13, TM15]. Comparisons [Fot03, HD12, KPSH15, MVS13]. Compensating [CBDM07]. Competing [MS19, Put11, SZ11, dWFP11]. Competing-Risk [MS19]. Complete [OK14]. Completely [JJJ14]. Complex [AT13, FW18, GP12, How15a, Lum04, Mat16c, MN14, Obe14, PSS20, SvdLN17, TMKD17, dI17, Kha18a]. Complex-Valued [dI17]. Complexity [Bra15]. Complexity-Based [Bra15]. Component [Bel11b, BDdM11, Bos11, CBA20, GKD14, MW07]. Component-Based [CBA20]. Components [BdMM15, MP99, PSY20]. Composition [Yua07]. Compound [GC18]. CompPD [LZ15]. CompRandFld [PB15]. Comprehensive [RFKM12]. Computation [Bel11a, Ber01, Dem17, Der10, HD16, How15a, Hug07, KMG^{+}13, KT10, NRG^{+}17, PQM11, dVSWAL17, SB01, Som01, SSM21, Zei06, Kha15b, Eks10].

Computational [BFH^{+}14, Din06, HGH05, Hil07, MAK06, MHH08, WDT^{+}12, Edd09c, Kha18b, Mur05, dL07]. Computationally [BG18, EWMR20]. Computations [DR11, KJ14, Mat16c, WM18]. Compute [Jur15, LSvdVK09, LSvdV19, McB06, NGVCFLS19, Phi10, dSJdSF14]. Computer [DHF15, EC06, Han05, KOS20, MvB14, PPGD15, Pet97a, Pet97b, RGD12, San09, VS14, Hed99]. Computerized [MR12, MB17]. Computing [AGM07, Bai04, Cha10, CSY15, Den16, DKT20, Esm14, FBdlF12, FT08, FMM^{+}20, Gio09, HSH07, JMS^{+}09, KEW13, Kha03, LP13, LCK11, LZ15, Pan09, SFS12, SME^{+}09, SDDD12, SL11, VMY02, WSZ12, Zei04, Udi00, Mat15]. Concave [CGS09, DR11]. Concomitant [GL08]. Concurrent [AD15]. Conditional [Lee13, LRN18a, Mon06, MHT07, OHD17, RT14, YSH15]. Conditioning [Mar06a]. Conditions [Yua07]. Conducting [KWE^{+}17, SvdLN17, Vie10, Zha14]. condvis [OHD17]. Cone [LM14]. coneproj [LM14]. Conference [LCK11]. Confidence [JLZJ05, Kel07, KPSH15, LCD18, New05, SB01, WZ16, MJ00]. Confirmatory [Hua20, KKH10]. Conformal [HZS21]. Conjugate [Sho13]. Connectivity [WdVW^{+}11]. Consistency [LQC^{+}12, Whe14]. Consistent [FR15]. Constant [GL08]. Constants [GSW20]. Constrained [BFRP13, Gre17, LM14, LM19, LC10, Mug10, MHdL12, Rec10a]. Constraint [Hüs06, Scu17]. Constraint-Based [Scu17]. Constraints [Grö10b, JP16]. Constructing [CSNF18, MB15, MJ00]. Construction [Jon07, Kav15, TB17, Udi05, WWG09]. Consumed [PZK^{+}12]. Contact [GW18]. Contaminated [PMM18]. ContaminatedMixt [PMM18]. Content [Mai08]. Context [NMM^{+}21, RG07]. Contiguity [Rec10a]. Contiguity-Constrained [Rec10a]. conting [OK14]. Contingency [LS02a, MZH06, OK14, WH08]. Continual [SMS13]. Continuous [AVD21, CC08, DKT20, DOV17, Hoh18, KY10, KK14, MHM20, Mul14, PJSPC17, SD17, VR14, VGDM^{+}18, Wan13b]. Contrast [WS11]. Contrast-Enhanced [WS11]. Control [ADH11, But09, ECK04, EC06, HSL11, JWHS16, MG09, NMB15, RD21, TKM15, ZFZ10, ZP13, dSdSCLC18]. Control-Mother [NMB15]. Controlled [HZSW10]. Controlling [NGVCFLS19]. Controls [BCAB07]. Convenient [SDJ20]. Convergence [Smi07]. Conversion [Lei13]. Converting [Wan13a]. Convex [FNB20, KM14, PLRC10, Som98]. Convolution [RK14]. convoSPAT [RC17]. cooccur [GVM16]. Cookbook [How11]. Coordinate [FHT10, GRS12, OWA20, SFHT11]. Copula [BER^{+}16, MV17, Nag18, NAA17, WMR16, KY10, Yan07]. Copula-Based [NAA17]. CopulaDTA [NAA17]. Copulae [OR14]. copulaedas [GFS14]. Copulas [BS13, GFS14, GHH20, HD18, HM11, Yan07]. Copy [SIR^{+}11]. Core [Hil07]. corr2D [GFPB19]. Corrected [BK11, CCF19, HS18]. Correlated [AS04, AD15, CNLR19, KM16, KS10, KMM^{+}17, RMG12]. Correlation [BKT14, GFPB19, GDMB08, Gos14, HM18]. Correlations [LH12]. Correspondence [BM97b, LSvdVK09, LSvdV19, NG07, dLM09c, dL05b]. Cost [Bai20]. Cost-Effectiveness [Bai20]. costat [CN13]. Costationarity [CN13]. Count [GF15, HD18, JD15, KM16, KBMB19, LFF17, SSH16, SF16, ZKJ08]. Counterfactuals [SKZ05]. Countr [KBMB19]. Counts [Cam12, DS15, MJ00]. Coupon [FH10]. Course [Hel17, Wic08a]. Courses [Dry09]. COVAR [Pet97a]. Covariable [KK15]. Covariance [CDP20, Mon06, PP17, Pet97a, RC17, Zei04, ZK12]. Covariances [ZKG20]. Covariate [Gre17, HPWdL15, HXY12, KMG^{+}13]. Covariate-Constrained [Gre17]. Covariates [VKVC15]. covatest [CDP20]. Coverage [TKM17]. CovSel [HPWdL15]. Cox [CG15, DPSH18, Gan15, SHBZ14, SFHT11, TDRD13, TDRD15, TR14, WZQW20]. coxphw [DPSH18]. COZIGAM [LC10]. cpm [Ros15]. Cramér [XGY06]. Crawley [Hel15]. CRC [Gle16, Grö15a, Grö16, Kha16b, Zei15]. Creating [CBAA19, Grö14, HL09, Ver12, Xie13]. Creation [NRD16]. credsubs [SFC20]. Criteria [Day01, Esm14, Gos14]. Criterion [KH13]. Critical [McB06, SB01, Som07, VS14]. crlmm [SIR^{+}11]. crmPack [BYPJ19]. Cross [BDKL17, ESO16, KO06, WWG09]. Cross-Entropy [BDKL17]. Cross-Language [ESO16]. Cross-Over [WWG09]. Cross-Validation [KO06]. Csárdi [Nun15]. cts [Wan13b]. ctsem [DOV17]. CUDA [SM12]. cudaBayesreg [dS11]. Cumulative [BH07, BH08a, BH08b, Mey13]. CURAND [SM12]. Cursi [How15b]. Curve [Das06, FH10, LW03, Wil14a]. Curves [KHLF^{+}10, MIG12, OL17, Sac17, TFR16, WZ16]. Cusp [GvdMW09, GvdMW09]. Customizable [MH09]. cutpointr [TH21]. Cutpoints [LRRÁCSGS14, TH21]. CVTresh [KO06]. CVXR [FNB20]. Cycling [FK17].

D [Han15, Nun15, BS13, EN11, FF14, GGK10, GKSU15, LPLPD14, WG00]. D-STEM [FF14]. D-Vine [BS13]. DAD [RG07]. DAKS [ÜS10]. Dale [MH05]. Daniel [Zei16]. Darling [MM04]. DATA [MT14, Alm10, ÁBZ17, ÁBZ20, AD15, AB17, BL14, Bai04, Bak20, BFRP13, Bar96, BPP17, BdMM08, Ber01, BBG12, BK17, BGRR15, BDdM11, Bow10, Boy05, BW01, BWMBC21, BKvT^{+}14, BEHB17, But08a, CB17, CBDM07, CGBN14, CCH15, CYK^{+}09, CBHK21, CNLR19, CG10, CM08, Dow17, Dow19, DM19, Duo07, ESO16, Ell09, ECW^{+}12, FS10a, FBdlF12, Fel12, Fer11, FDGD16, FBG15, FM18, FK17, Fri12, Fri06, Fuj17, GHR18, GASA15, GP12, GF15, GKZ16, GGK10, GR17a, GCA12, GU07, Gos11, GD21, Grö15a, Grö18a, Gro08, HD10, HBF17, HD18, HHB^{+}08b, HK20, Har08, HW20, HM18, Hel17, HH19, HRBP15, HC05, HHJCRB21, HKB11, HCW12, HJS16, IP08, Iac15, Iac17, Jac11, Jal19, JM15, JLH^{+}20, JD15, JH16, KM16, KEW13, Kha18a, KBMB19, KS10, KZ03]. Data [KSS^{+}07, KK14, KT10, Lal16, Lan14, LDHM21, LRGTA12, LMO^{+}19, Lei10, LSPvdL17, LM03, LGF^{+}18, Lor17, LP17, MBGK18, MNA^{+}97, MH05, MSCCV05, MMRP11, MCM12, MD18, MP12, MdUÁC10, MN14, LMB^{+}07, Mun14, Mur03, NGBK12, NL12, NOT08, NRD16, Nun15, OL17, PPGRMB21, PP17, Pav16, Peb12, Pen08a, PE19, PV14, PPC15, PU13, QY19, Riz10, Riz16, RD21, RMG12, RVCG19, San06, SGK03, SH14, Sar09, SP14, Sav16, SZ11, Sch16, She11, SCS13, SYC08, SF16, SF19, SvdLN17, Sól09, Spa06, Su07, TDRD15, TKM15, TMKD17, The02, TD07, ÜS10, VM09a, VM10, Ver18, VS14, Vin17, VYD^{+}12, VGDM^{+}18, WGSL12, WKL11, Wan13a, WdVW^{+}11, WDM^{+}11, WST11, Wic07, Wic10, WG10, Wic11, WCHB11, Wic14, WBEE19, Wol12, WZ17, Xu06, XHW20, YR19, Yee10, YSH15]. Data [ZKJ08, Zei15, ZCI^{+}16, ZQS16, dS11, dL05b, dSdSCLC18, vdLdJ21, vdL21, How15a, Mai09c, Ach09, Arm19, But15, Dem18a, Grö17, Grö21, Hel20, Hil09a, How15a, Kha20, Mai08, San10a, Wic08b]. Data-Driven [LDHM21, WBEE19, How15a]. Database [MBK18]. Database-Inspired [MBK18]. Databases [NRG^{+}17]. datacollection.com/ [Iac15]. DataGraph [Mac12]. dataMaid [PE19]. Datasets [Mur06, RPC15, ZMC21]. Dates [GW11]. DATforDCEMRI [Fer11]. dawai [CFSR15]. Dayal [Hof15]. dbEmpLikeGOF [MVS13]. DBKGrad [MP14]. dbmss [MTPL15]. dbscan [HPD19]. DCE [Fer11]. DCE-MRI [Fer11]. DClusterm [GRMMR19]. ddalpha [PMD19]. deal [BD03]. Dealing [Kas16, Pap16]. Death [BdUÁ18, MMRP11, TGJ17, WCLS20]. DeCo [CGRvD15]. Decomposition [CBA20, DMB18, Lei10, LBW18, LS02b]. Decompositions [EVBK19]. decon [WW11]. Deconvolution [Fer11, NE20, RS07, WW11]. deconvolveR [NE20]. Deducer [Fel12]. Deepayan [Gro08]. Degradation [CP12]. Dehli [Hof15]. Delay [JvdBP07]. Delta [MF14]. deltaPlotR [MF14]. demogR [Jon07]. Demographic [Alm10, Jon07, SM07]. Demonstrating [Xie13]. Demonstration [CCD09, NRG^{+}17]. Dendrograms [SHFB17]. Densities [MW12, Nag18]. Density [AM14, CGRvD15, CGS09, DHM11, DR11, Duo07, DP03, HPD19, HS18, MW12, Pav15, VDT14, Wol12]. Density-Based [HPD19]. DEoptim [MAG^{+}11]. Dependence [BDMP15, BS13, CFHBK11]. Dependent [KO06]. Deployment [VFV13]. depmixS4 [VS10]. Depth [LZ15, PMD19]. Depth-Based [PMD19]. Derivative [WM18]. Derivatives [BRC98, MS11]. Deriving [Arc10]. Descent [FHT10, GRS12, SFHT11]. Description [CA17]. Descriptive [Dha10, Ruf09]. Design [BG21, BPB09, CKE20, DHF15, Esm14, GG16, Gra07, Grö11, HSL11, HR14, HMM21, HZSW10, LVK^{+}11, LW03, MvB14, OWA20, WWG09, Wol15, ZP13, Grö10a, Len09a]. Design-Based [HMM21]. Designed [FBF14]. Designing [JPM19, RSW15, YZZ^{+}20]. Designs [BYPJ19, CBHK21, DR20, FS13, FKP17, Fri10, Gre17, Grö14, HWY18, IM05, Kao09, KPH^{+}21, Law08, LSM^{+}19, NMB15, SlRC16, SMS13, TSV20, Was15, YZZ^{+}20]. deSolve [SPS10]. Detect [ESKHB20]. Detecting [CGC11, ZPZW21]. Detection [GRMMR19, McB06, RG07, Ros15, SSH16]. Determining [CGBN14]. Deterministic [MRC15]. Developed [RG07]. Developer [BG18]. Developer-Friendly [BG18]. Developing [Gen06]. Development [BKvT^{+}14, RFKM12]. Developments [JLZJ05, vdA12]. deviances [Can04a]. deviation [SM12]. Devices [FK17]. DFIT [Cer17]. dglars [AMW14]. Diagnose [HHB^{+}08c]. Diagnosis [GRK^{+}16, MdlT20, TB17]. Diagnostic [Bar96, BKvT^{+}14, GR18b, LRRÁCSGS14, LX12, Ngu07, NAA17, PJSPC17, RG07, Ver18]. Diagnostics [LH14, SGHY11]. Diagonal [KN05]. Diagonalization [MNT17]. Diagram [DD07]. Diagrams [BDMP15, HF18]. DiagTest3Grp [LX12]. Diallelic [Gra15]. DiceDesign [DHF15]. DiceEval [DHF15]. DiceKriging [RGD12]. DiceOptim [RGD12]. Dichotomous [WB08]. DICOM [WST11]. Difference [KM08, ZPZW21]. Differences [Gow16, PKC^{+}20, WZ16]. Differential [BFH^{+}14, Cer17, CGC11, GB20, HRX16, MF14, MAG^{+}11, SPS10]. Differentiation [KNB^{+}16]. Difficult [MT02]. Difficult-to-pass [MT02]. diffIRT [MTvdM15]. Diffusion [MTvdM15, PT09, PT11, SKS15, WS11]. Diffusion-Weighted [PT11, WS11]. Diffusive [BFM18]. Digital [PT07]. Dimension [AR14, BM12, CFT11, MDvdV19, NMM^{+}21, Wei02]. Dimensional [AEL10, BBG12, BKvT^{+}14, CNLR19, Coo97, GFPB19, HW20, HPK18, HCW12, ND12, Pat16, SF18, VG09, WZ17, BW01, NG07, Pen03]. Direct [WZ16]. Directional [KGP^{+}19]. Dirichlet [Han10, LHA^{+}15]. Dirk [Kha15b, Kha15a]. Disaggregated [KPRP^{+}19, RD21]. Disciplined [FNB20]. Disclosure [TKM15]. Discovering [LAF^{+}17, Xu06]. Discovery [DK18, Lal16]. Discrete [DKT20, KK14, MTW04, MP14, Sar16, SD17, TSV20, USA19]. Discrete-Event [USA19]. Discretized [Su07]. Discriminant [BBG12, CFSR15, Duo07, TBE^{+}21]. Disease [AVD21, BVE^{+}15, GRMMR19, JGM18, JWHS16, MGFA^{+}20, WBEE19]. Disequilibria [SBMG06]. Dispersed [SF19]. Display [DN99, Hel17, SBMG06]. Displaying [FT08]. Displays [Fox03, FH09, Fri06, NL12]. Dissimilarity [GU07]. Dissimilarity-based [GU07]. Distance [MRT^{+}19, QY19, SH09]. Distance-Aided [SH09]. Distance-Based [QY19]. Distances [vE17]. Distributed [Gas11, LP13, Mug10, NRG^{+}17, TFH12]. Distributing [Pen08b]. Distribution [AH12, AWBM18, AHvD09, BH07, BH08a, BH08b, Car15, DKT20, DP03, GFS14, Han10, MTW03, MM04, Mar04, McB06, Mey13, MR05, MMM12, Phi10, dREP12, Ram00, SBL^{+}10, SL11, TBE^{+}21, ZQS16, ZPZW21]. Distributions [AHvD09, BS18, BMvS^{+}07, DMD15, Dem18b, Den16, Gil15, HG14, Kam08, KY10, LM13, LM18, LHS08, MNA^{+}97, NK06, NR16, Pan09, PCL13, PMM18, dVSWAL17, Su07, ZFZ10]. distrMod [KR10a]. Divergence [LR20]. Diverging [HR14]. Diverse [FR15]. Diversity [MH15]. Dixon [VS14]. DixonTest.CriticalValues [VS14]. dlmap [HSG12]. dlnm [Gas11]. DNest4 [BFM18]. Document [Dry09]. Documentation [HWPG13]. Documenting [PE19]. Documents [McN14, Mor18]. DoE.base [Grö18b]. Dominic [Iac15]. Donors [SH09]. Dose [BPB09, BYPJ19, CO16, HZSW10, MAK06, SMS13, WB08, WEH11, YZZ^{+}20]. Dose-Finding [BPB09, SMS13, YZZ^{+}20]. Dose-related [MAK06].

Dose-Response [CO16, WEH11]. dosresmeta [CO16]. Douglas [Mat16a]. Down [Som12]. downscale [MBGK18]. Downscaling [MBGK18]. DPB [LP17]. DPpackage [JHQ^{+}11]. Drift [Han07b]. Driven [DS15, LDHM21, WBEE19, How15a]. Drug [BN10, YZZ^{+}20]. Drug-Combination [YZZ^{+}20]. dsarkar.fhcrc.org [Gro08]. DTDA [MdUÁC10]. dti [PT09, PT11]. DTR [TM15]. DTRreg [WMS17]. dtw [Gio09]. dtwSat [MCAP19]. Dual [LM14]. Duality [DD07]. Dummies [SGS^{+}14, Edd12a]. Dumont [Gle16]. Dynamic [BM97b, CN18, CS12, DN99, Gio09, JvdBP07, LP17, MGFA^{+}20, MCAP19, McN14, Oom13, PSY20, Pet10, TM15, VM09a, WMS17, WS11]. dynamicnetwork [BdMM08].

e-Commerce [Chr11]. E-Learning [ZUL14]. Early [Tak16]. Earthquake [CA17, Jal19]. Ease [LT16]. Easy [Bro15, CdM10, GW11, HM16, SMDS11]. ebalance [HX13]. EbayesThresh [JS05]. eBook [Gle16, Zei16]. eco [ILS11]. ecodist [GU07]. Ecological [AB12, GU07, Han07b, ILS11, Kin04, KP07, MHE20, PR07, Sól09]. Ecologists [DD07]. Ecology [Chr09, KP07, Soe10, Hil10a]. Econometric [Zei04]. Econometrics [BP15, CM08, Edd09b, HR08, Yal10, ZK08]. Economic [Bai20, ZP13]. Economics [CN18, Hof15]. ecp [JM15]. ed [Hel18]. edcc [ZP13]. Edge [HL07]. Edges [HD16]. Edition [CH11, Dem17, DN17, Dow19, Edd09a, GR17a, GR18a, Grö15a, Hel15, Hel16b, Hel17, Kha17a, Kha18c, Kha20, KN17, Lor18, Mai09c, Pav20, Ros17, San10a, SCD07, VM09b, Wil09, dL05c, Fox05a, Kha16c]. eDMA [CN18]. Eduardo [How15b]. Educational [CCD09]. Effect [CB21, CLL17, Fox03, FH09, FW18, Kel07, SvdLN17, SLMV17, Cha03]. Effectiveness [Bai20, LNY^{+}20]. Effects [AW16, BMBW15, Bea17, Chr09, Gan15, HPWdL15, HN13, JP16, KCCG11, Kol16, KSPR16, KBC17, Lec14, LQC^{+}12, LG11, Mug10, Röv20, SP14, TKM17, TK17, VDT14, WLH^{+}18, WM18, ZLC18, Hed99, FH09]. Efficient [AHvD09, BG18, BGH^{+}17, CGRvD15, EWMR20, ESO16, New05, RBN19, SSM21, TB17, Lan17]. Efficiently [HZBB20]. Elastic [SHBZ14]. Elastic-Net [SHBZ14]. Elea [Rus15]. Electoral [OO11]. Elegant [BE13, GR17a, VM10]. Elimination [JMD08]. Ellipses [Fri06]. Ellipsoidal [Som01]. Elliptic [Han06b, Han06b]. elrm [ZMG07]. Elsner [Hil15a]. EM-Like [DV20]. Embarrassingly [Mat16c]. Embarrassingly-Parallel [Mat16c]. Embedded [BDdM11]. Embedding [DB13]. emdi [KPRP^{+}19]. emil [BG18]. EMMIX [MP99]. EMMIXcskew [LM18]. EMMIXuskew [LM13]. Emphasis [FS10b]. Empirical [ASB11, ADN15, Cha10, JS05, KG17, NE20]. Employing [ZF15]. Emulator [Han12]. Enabled [FK17]. Enables [Wei12]. End [HM16, WDT^{+}12]. End-to-End [HM16]. End-Users [WDT^{+}12]. Energy [PZK^{+}12]. Enforcing [Oom13]. Enhanced [WS11]. Enhancement [Dry09]. Enhancing [BBGL17]. Enjoy [Yan07]. Ensembles [Fok20, IoG10, Hor05]. entropart [MH15]. Entropy [BDKL17, HX13, MVS13, VdL09]. Enumerative [Han07a]. Envelope [CSY15]. Envelopment [ÁBZ20]. Environment [BM97b, Cha12, HGH05, Mug10, RFKM12, SDDD12, SHR96, Wei12, Udi00]. Environmental [FF14, Yua07, dL09c]. Environments [BLM^{+}15]. envlp [CSY15]. Epidemic [MHH17]. Epidemics [GW18]. Epidemiologic [Hil09b]. Epidemiological [MJGM10, PC11]. Epidemiology [dL09c]. EpiILMCT [AVD21]. EpiModel [JGM18]. epinet [GW18]. Episodically [PZK^{+}12]. Equality [Grö10b, MHdL12]. equate [Alb16]. equateIRT [Bat15]. Equating [Alb16, ABW13, Bat15, Gon14]. Equation [DOV17, HRX16, Hua20, Lum96, MR18, ML12, Obe14, Ros12, UNBK17]. Equations [BFH^{+}14, GB20, HHY06, HH07b, RS08, RW11, SPS10, Tou15, VG09, vBGO11]. Equivalent [LAF^{+}17]. ER-fMRI [Kao09]. ERA [LV16]. ERG [YFK^{+}15]. ergm [HHB^{+}08c]. ergm.graphlets [YFK^{+}15]. ergm.userterms [HGH13]. Eric [Nun15]. eRm [MH07]. Error [HXY12, Mil17, MIG12, PZK^{+}12, WW11]. Errors [BK11, Bel11b, WM18]. Escalation [BYPJ19]. Essentials [Pos15]. Estimate [AMW14, PZK^{+}12]. Estimates [Edl04, Gan15]. Estimating [BP19a, BKT14, Bra17, Das06, DBBD07, Fie12, GSW20, HPWdL15, HHY06, HW20, HH07b, KS14, KPRP^{+}19, LW03, Lum96, MMR20, RS08, SM07, Tou15, You10]. Estimation [AG19, AMYR16, AT13, AW16, ALV07, AMMRP14, AM14, BdUÁ18, BP15, CCF19, CA17, CLL17, Cro20, CGS09, Cul11, DHM11, DBZ^{+}11, DR11, DSH16, Duo07, FDB12, FH10, Fot03, GFS14, GvdL12, HWM16, HS18, Joh07, KL14, Kol16, KSP15, LCD18, LSPvdL17, dUJ13, MMB15, MW12, MGH18, MLAN02, MMM12, Nag18, NE20, Nor17, OC15, PJSPC17, PSS20, PLPL17, dREP12, RC17, RMG12, SBL^{+}10, SH20, Sav16, SIR^{+}11, SVMMRP17, She08a, She10, Sho13, SvdLN17, TSN21, TM15, TH21, TR14, VV16, WMS17, WW11, Wan11, WB08, Wol12, XM14]. Estimations [HMM21]. Estimator [Pav15, SP05]. Estimators [ABS01, CSY15, Jur15, Mil17, Zei04, Zei06]. ETAS [Jal19]. etasFLP [CA17]. etm [ASB11]. EUR [Hof15]. Evaluate [BH07, BH08a, BH08b, DM19]. Evaluating [GG16, JKL19, MTW03, MM04, Mar04, MIG12, Ngu07, SKZ05]. Evaluation [Bai20, Car15, HSL11, Mey13, Som98, WEH11]. Evaluations [TKM17]. Event [CBHK21, GHR18, KMM^{+}17, Riz10, Riz16, RSW15, USA19]. Events [KZ03, KMM^{+}17, MMM12, RBN19]. EViews [Cas17, Van11]. evmix [HS18]. Evolution [MAG^{+}11, NdSL16]. Evolutionary [CBA20, GZP14]. evtree [GZP14]. Exact [BM12, Caf06, FS10a, HGK^{+}20, IKP09, Kha03, MHT07, WH08, ZMG07]. Exact-Like [ZMG07]. exactLoglinTest [Caf06]. Example [Eva11, San10a]. Example-Based [San10a]. Examples [Fri06, Hel17, Grö17, Sar09]. Exams [GZ09, ZUL14]. Excel [Web16, But05, But09, Hla16a, Kla18, KR09, Web16]. Excel\begin{matrix}c \\ ◯ \end{matrix} [LS02b]. Excess [CB21]. Exchange [OWA20]. Exclusion [AT13]. Excursion [BL18]. excursions [BL18]. Execution [SR07a]. Exercises [Kha16c]. Expanded [HO20]. Expansion [MR18]. Expansions [LP13].

Experimental [Grö11, IM05, Kao09, SlRC16, Vig16]. Experimenting [KvEPK20]. Experiments [Aiz12, DHF15, FO15, FBF14, Grö10a, Grö18b, LW16, Len09a, MvB14, MHT07, NGBK12, OWA20, RGD12, SBL04, TP11, TSV20]. Explained [But15]. explorase [LCL^{+}08]. Exploration [SD13, Vig16]. Exploratory [LRGTA12, LCL^{+}08, SGS^{+}14, Eva11]. Explore [HCW12]. Exploring [BW01, Cal07, CCH15, GLC^{+}15, Gra15, NOT08, PPC15, WCHB11]. Exponential [CF14, HHB^{+}08c, LCD18, MR05, MHH08, SL18]. Exponential-Family [HHB^{+}08c, MHH08, SL18]. Exposure [JWHS16]. Expression [Boy05]. Extend [GDMB08]. Extended [GKZ12, LV16, MH07, PLPL17, SHBZ14, ZC10]. Extending [CNLR19, FC12, HGH13, San07]. Extensible [HD10, HBF17, HWPG13, TSH20]. Extension [YEL18]. Extensions [BGRR15, Chr09, FH09, GKSU15]. Extensive [KS14]. extracat [PU13]. Extreme [DGJ^{+}18, GK16, Hla16b, HS18, McB06]. ExtremeBounds [Hla16b]. extremefit [DGJ^{+}18]. extRemes [GK16]. Eyes [Coo97].

F [Hug07]. F-to-z [Hug07]. Facilities [Fir03]. FactoMineR [LJH08]. Factor [BP12, CFHBK11, DM19, DSH16, LS02a, NGVCFLS19, PKC^{+}20]. Factorial [FS13, FKP17, Grö14, Grö18b, NGBK12]. Factorials [LSC09]. Factors [EC06, MHdL12]. Failure [CKY14, HXY12]. FamEvent [CBHK21]. Families [NR16]. Family [ALV07, CBHK21, DP12, HHB^{+}08c, MHH08, PG15, SL18]. FAMT [CFHBK11]. Far [RBB18]. Fast [BE13, Bel11a, DCS20, HPD19, HWM16, LH12, MTW04, Mül13, Pat00, PG15, RS07, dVSWAL17, SLF14, VW13, WZ17]. fastcluster [Mül13]. Faster [GTDT20]. fastnet [DCS20]. fda.usc [FBdlF12]. FDboost [BRG20]. FDR [Som07]. Feature [CNLR19, DK18, KR10b, LAF^{+}17, Meu13, MD18]. fechner [ÜKD09]. Fechnerian [ÜKD09]. Feit [Rus15]. Ferns [Kur14]. Fertility [SAR11]. Festschrift [MM16]. FFT [RK14]. fgui [HL09]. FHtest [OL17]. FIAR [RR11]. Field [Pat16]. Fields [BILL07, BILL10, DB13, FS10b, PB15, SMM^{+}15, dI17]. Fieller [Kha16b]. Figure [GC18]. Files [Dry09]. Filter [Wan13b]. Filtering [KGP^{+}19, Tus11, VYD^{+}12]. Finance [CN18, Mül16]. Financial [SCD07, Spe13]. Finding [BPB09, HWY18, SMS13, Was15, YZZ^{+}20, Edl04]. Fine [MBGK18]. Fine-Grain [MBGK18]. Finite [GL08, Lei04, PCL13]. FinMetrics [Ziv11]. First [Wic08a]. Fisher [HG14, Kha03]. Fit [Bar96, FW18, GHR18, HHB^{+}08c, LR20, Mar05, MVS13, Pav15, PV14, TMN16, ZA17, LT16]. fitdistrplus [DMD15]. Fitness [PSS20]. Fitting [BS18, BMBW15, CAA15, CB21, CKY14, DMD15, Dem18b, FC11, Fok20, FS20, Gil15, GvdMW09, GH11, HHV20, HG14, IV05, Jal19, KHLF^{+}10, KV13, KH08, LM13, LM18, MRC15, McI16, MNA^{+}97, MP99, MH05, MTvdM15, PZK^{+}12, PCL13, PMM18, Riz16, Su07, TGJ17, UNBK17, Van11, ZHZ20]. Fixed [AH08, CC11, DHM11, Wol15, ZF15]. Fixed-Size [Wol15]. flexCWM [MPI18]. Flexible [AHvD09, CB21, FSLZ12, KBMB19, MPI18, SlRC16, SLMV17, VDT14, WK18, ZUL14, ZF15, Bar18, Dem18a, Kha20]. FlexMix [GL08, Lei04]. flexsurv [Jac16]. FLIM [LMB^{+}07]. Fluorescence [LMB^{+}07]. FluxSimulator [BMvS^{+}07]. FME [SP10]. fmixed [LG11]. fMRI [BDdM11, Kao09, MvB14, WdVW^{+}11, WDM^{+}11, dS11, TP11]. Focus [DLC06]. Follow [DR20, PC11]. Follow-Up [DR20, PC11]. Following [LFF17]. Food [WEH11]. Food-Safety [WEH11]. Foods [PZK^{+}12]. forecast [HK08]. Forecasting [HK08, Yaf07]. Forecasts [JKL19]. Forest [CNLR19, HMM21]. forestinventory [HMM21]. Forests [CNLR19, MIG12, WZ17, ZWH17]. Form [KS14]. Format [Dry09, Wee10]. forms [Fir03]. Formulas [ZC10]. Formulating [DLC06]. FORTRAN [Gen06, dL12, Mil04, SB01, Bel11b, KKH10, KH13, MHdL12, Som98]. Forward [RPC15]. Foundations [Lal16]. Fourier [AG19, Pac07]. Fourier-Based [AG19]. Fourth [Fox05a]. FracSim [DC05]. Fraction [LVK^{+}11]. Fractional [BILL07, BILL10, FS13, Grö14, Coe00]. Frailty [BP19a, MGH18, MRL12, RMG12]. frailtyEM [BP19a]. frailtypack [RMG12]. frailtySurv [MGH18]. Frames [HHJCRB21]. Framework [ATF10, BFH^{+}14, CBA20, Cer17, DR20, GTDT20, HBF17, HMS16, Kle16, Lei04, MdL10, MPI18, MZH06, Nor17, NAA17, PR07, TF09, WdVW^{+}11]. FRB [VW13]. frbs [RBHB15]. Free [BFC02, IP08]. French [HJS16]. Frequencies [Meu13]. Frequency [GKZ16, TKM17]. Frequent [HGH05]. FrF2 [Grö14]. Friendly [BG18, Nag21, SGS^{+}14]. FRK [ZMC21]. Frontier [HB17]. Function [BRC98, DBZ^{+}11, MMM12, Pat00, PZK^{+}12, dREP12, SBMG06, VSV09, VG09, PG15]. Functional [Ach09, BRG20, CG10, FBdlF12, HK20, LG11, RR11, Sav16, SCS13, TP11, YR19, YEL18, Bow10]. Functionalities [BMGT15]. Functionality [Guh07]. Functioning [Cer17, CGC11, MF14]. Functions [Aiz12, AW16, BCAB07, BH07, BH08a, BH08b, CDP20, Han06b, LH12, LQC^{+}12, MS19, McB06, Pac07, Phi10, RC17, Ruf09, TKM07, WEH11]. funcy [YEL18]. Fundamental [LM18]. Fusion [SYC08]. Future [Tie05, ZK08]. Fuzzy [CNLR19, RBHB15, VV16]. FuzzyStatProb [VV16].

GA [Scr13]. Gábor [Nun15]. gamboostLSS [HMS16]. Games [KS14]. GAMLSS [Bar18, HMS16, SR07b]. Gamma [MT98, XTL13]. Gap [WDT^{+}12, Zha07]. GARCH [ABB^{+}19]. GAS [ABC19]. Gauss [Das06]. Gaussian [TDRD13, BKT14, BER^{+}16, BP19b, BG21, BILL07, BILL10, DB13, EC16, FS10b, Gra07, GT10, Gra16, HD18, HL07, JCPL19, LHS08, MRC15, MV17, Pac07, PLZ^{+}15, PT11, SR18, SBMD14, TKM17, TDRD15]. Gaussian-Process-Based [BP19b]. GBP [Han15, Iac15, Nun15]. gcKrig [HD18]. GDINA [MdlT20]. gdistance [vE17]. gdpc [PSY20]. geepack [HHY06]. GEEQBOX [RS08]. geex [SH20]. GemIdent [HKL09]. gems [BVE^{+}15]. Gene [Boy05, ESKHB20, GD21, Sab07]. Gene-Based [ESKHB20]. Gene-Gene [ESKHB20, GD21]. Genealogical [RVCG19]. GeneGeneInteR [ESKHB20]. General [CL13, FKP17, Gre17, Kur14, Lei04, LHS08, MdL10, MGH18, PRS18, RK14, She10]. General-Purpose [Kur14]. General-to-Specific [PRS18]. Generalised [Fox03, Lum96]. Generalization [Han10, YEL18]. Generalizations [AH12]. Generalized [AMYR16, ABC19, Arm19, AMW14, BdMM15, CdM10, Can04a, Cha10, Dun99, FHT10, GHR18, Had10, HHY06, HC05, ICL16, KLNN20, KH13, LM19, LFF17, LC10, McI16, MH09, ML11, PSY20, Ram00, RS08, Sch11, SHBZ14, SR07b, Su07, TBE^{+}21, Tou15, VDT14, dSJdSF14, GR18a, Mai06]. Generalizing [PLRC10]. Generate [Ruf09]. Generating [Cha16, CBHK21, Fuj17, GD21, KM16, LS02a, MT98, TSV20, WDM^{+}11, MT00]. Generation [AD15, Cha03, Dry09, FS13, GZ09, HWPG13, MR12, MTW04, VYD^{+}12, YS12, ZUL14]. Generator [PSW07, SM12]. Generators [Bre04, Vig16, XTL13, PL05, SLF14]. Genetic [Gra15, JFDB15, Jur15, MF15, MS11, Scr13, TB17, WZL20, Wol15, ZFZ10, Zha07]. Genetics [SL09]. Genotype [EC16]. Genotyping [SIR^{+}11]. geoCount [JD15]. Geographical [Jos05, Kav15, vE17]. Geographically [GLC^{+}15]. Geospatial [PPC15]. Geostatistical [HD18, JD15, SYC08]. Geostatistics [Bro15, Gum07]. GeoXp [LRGTA12]. GETS [PRS18]. Getting [HHB08a]. GFD [FKP17]. ggenealogy [RVCG19]. ggm [Mar06a]. ggmcmc [iM16]. GGobi [VM09a]. ggplot2 [HF18, GR17a, VM10]. ggtern [HF18]. Gibbs [BTMB13, Ntz02, Woo16]. Gifi [dLM09a, vdHvB16]. Gillespie [PK08]. GillespieSSA [PK08]. glarma [DS15]. GLDEX [Su07]. GLIMMPSE [KMG^{+}13]. GLMM [DP12, XM14]. glmulti [CdM10]. Global [DSH16, HMM21, MAG^{+}11, Mul14]. global.oup.com [How15a]. Globally [GZP14]. Glotaran [SLS^{+}12]. GLUMIP [KCM08]. GMCM [BER^{+}16]. gmnl [SD17]. gMWT [FO15]. GNAR [KLNN20]. GNU [Bak20]. GoFKernel [Pav15]. Gold [Ngu07]. Gompertz [Das06, MNA^{+}97]. Goodness [LT16, LR20, Mar05, MVS13, Pav15]. Goodness-of-Fit [LR20, Mar05, MVS13, Pav15, LT16]. GPareto [BP19b]. GPfit [MRC15]. GPGPU [SM12]. GPS [FK17]. GPS-Enabled [FK17]. Gradient [BFFN19, BMR14]. Graduation [MP14]. Graeme [Han15]. Grain [MBGK18, Høj12]. Gram [HMR^{+}13]. gramEvol [NdSL16]. Grammar [Cox07]. Grammatical [NdSL16]. Graph [Cul11, LCD18, MHH08, SL18]. Graph-Based [Cul11]. gRapHD [AEL10]. Graphical [AEL10, AB12, CCH15, CCD09, DH05, Dun99, Fox05b, HW20, HL09, Høj04, HL07, Høj12, KMC^{+}12, LL10, LSvdV19, LSM^{+}19, Mar06a, MW19, RFKM12, SlRC16, SBMG06, SLS^{+}12, TMW18, VML12, WLH^{+}18, Zei15]. Graphics [GR17a, HSH07, LBC^{+}16, LR15b, Mur09, MG09, NG07, NL12, Pav20, San10a, VM09a, VM10, CH11, Grö15a, Bul06, Cox07, Grö21, Mur06, WG10]. graphicsQC [MG09]. Graphlet [HD16, YFK^{+}15]. Graphs [AH08, CF14, HD16]. Grassmann [ACW12]. GrassmannOptim [ACW12]. gRbase [DH05]. gRc [HL07]. Greedy [Kon13, Wol12]. gret1 [Ros08]. gretl [BK15, Luc11, LP17, YS12]. Grids [vE17]. grImport [Mur09]. groc [BdMM15]. grofit [KHLF^{+}10]. Group [GG16, Was15]. Group-Sequential [Was15]. Grouped [SH14, SF19]. Groups [LX12, PO13]. Growth [KHLF^{+}10, Wil14a]. gsbDesign [GG16]. gss [Gu14]. GUI [Fel12, HCW12]. Guide [Bad01, Ber01, GR17b, Grö16, Grö18a, Hil15a, Iac15, Iac17, Kha18c, Mat16a, Web16, Wic10, Zho10]. gWidgets [HCW12]. gWidgetsWWW [Ver12]. GWmodel [GLC^{+}15].

H [Grö15a, Hil15a, How15b, Iac15, Kha15b, Kha15a, Kha16b]. Habitat [Cal07]. HAC [OR14, Zei04]. HACopula [GHH20]. Hadoop [FM18]. Half [MHD17]. Half-Normal [MHD17]. Hall [Gle16, Grö16, Kha16b]. Hall/CRC [Gle16, Grö16, Kha16b]. Handbook [Hil06, MT14, Wil09, dL08, Dem18b, San06]. Handling [JH16]. Hands [Yal10]. Hands-On [Yal10]. hapassoc [BGM06]. Haplotypes [BGM06]. Hardcover [Zei16]. Hardware [Mur15]. HardyWeinberg [Gra15]. Hazard [CB21, Gan15]. Hazard-Based [CB21]. Hazards [GRS12, SFHT11, WZQW20]. HC [Zei04]. HDclassif [BBG12]. Health [ASBMB17, Bai20, SSH16, VMU05]. Health-Related [ASBMB17]. Heavy [Gil15]. Hector [Hil15b]. hergm [SL18]. Hero [EL09]. Hessian [Bra17]. Hessians [Bra14]. Heterogeneity [GLC^{+}15, SD17]. Heterogeneous [BL14, MSCCV05, WLH^{+}18]. Heteroskedastic [BG21, Pir10]. Heteroskedasticity [SYC08]. hetGP [BG21]. HGLMMM [ML11]. Hidden [HH19, OH11, VS10, Kha17a, San11]. Hierarchical [Arm19, CGC14, DW17, FBC07, FC11, GF15, GHH20, Hil09a, LH12, LH14, ML11, Mül13, OR14, Rec10a, Rec10b, SL18, SDJ20]. Hierarchically [SH14]. High [AEL10, BBG12, BW01, BKvT^{+}14, CNLR19, Coo97, HW20, HPK18, HCW12, LCK11, Mur15, ND12, PSS^{+}17, dVSWAL17, VG09, WZ17, XM14, dSdSCLC18]. High-Dimensional [AEL10, BBG12, BKvT^{+}14, CNLR19, Coo97, HW20, HPK18, HCW12, ND12, VG09, BW01]. High-Performance [Mur15, XM14]. High-Precision [dVSWAL17]. High-Throughput [PSS^{+}17, dSdSCLC18]. Hilbe [Grö16]. HLMdiag [LH14]. hmi [SDJ20]. hmmm [CGC14]. hnp [MHD17]. homals [dLM09a]. Homogeneity [BM96]. Homogeneous [Ber01]. Homoscedasticity [JJJ14]. Honoring [MM16]. Horton [Grö15a]. HPGLIMMIX [XM14]. HPLC [RG07]. HPLC-DAD [RG07]. hsdar [LMO^{+}19]. HTML [Lei13]. httk [PSS^{+}17]. http [But15, Grö15a, Gro08, Han15, Hel15, Hil15a, Hil15b, Hof15, How15b, How16b, Iac15, Kha15b, Kha15a, Mat15, Mat16a, Nun15, Rus15, Zei15, Zei16]. https [Gle16, Grö16, How15a, Kha16b, Pos15]. Hull [PLRC10]. Humanities [Pav16]. Hurricane [Hil15a]. Hybrid [CGC11]. hybridModels [MGFA^{+}20]. Hybrids [BTMB13]. HYDRA [War02]. Hyperbolic [TBE^{+}21]. Hyperspectral [LMO^{+}19]. Hypotheses [Pet97a]. Hypothesis [Caf06, Kha15a].

ibr [CHML17]. ic.infer [Grö10b]. icenReg [AB17]. ICS [NOT08]. idefix [TSV20]. idem [WCLS20]. iDEMO [CP12]. idendr0 [SHFB17]. idendro [SHFB17]. Identification [HY21, LDHM21, SFC20, Coe00]. Identifying [TK17]. idm [DMB18]. II [Wic10, HKB11, KWE^{+}17]. III [How16b]. Illness [BdUÁ18, MMRP11, TGJ17]. Illness-Death [BdUÁ18, MMRP11, TGJ17]. Illumina [EC16]. Illustrating [Gan15]. Illustration [GF15]. Image [BLY18, HKL09, KCCG11, MCAP19, SLGB14]. Images [GC18, PT07]. Imaging [CMS^{+}11, LMB^{+}07, PT09, PT11, TW11, WS11]. IML [AS02, JMS^{+}09, KCM08, NO06, Pan09, SB01, ZK12]. Implement [BMGT15, VDT14, XWHL15]. Implementation [Aiz12, BTMB13, CCD09, CLL17, GFPB19, GF15, GTDT20, JWHS16, KM16, Kel07, Kle16, Kur14, LPLPD14, LZL^{+}05, PG15, Pfa08, RW11, SBL^{+}10, USHH18, WLK08, WZ17, ZKG20, dS11]. Implementations [BP15, HI20, Scu17]. implemented [PSW07]. Implementing [BK11, DD07, EN11, HHvdWZ08, JMD08, LSPvdL17, PK08, SR18, Tri03, Udi00, Wel18, Wil14b, ZMG07]. Import [VYD^{+}12]. Importance [AH08, GT10, Grö06, RJH14]. Importing [Mur09]. Improved [Car15, MZ08, TH21, dSJdSF14]. ImpuR [RG07]. Impurity [RG07]. Imputation [CGK11, CF08, Dem18a, Kha20, KT16, RW11, SH09, SDJ20, SGHY11, Yua11, Yuc11, vBGO11]. IMPUTE [CGK11]. Including [Gan15, JLZJ05]. Incomplete [BRC98, OK14]. Incremental [DMB18]. IndElec [OO11]. Independence [Høj12, SF18, LS02a]. Independencies [Mar06a]. Independent [BDdM11]. Indexed [BILL10, Har10, Sav16]. Indicator [PRS18]. Indicators [AT13, KPRP^{+}19]. Indices [SFS12]. Individual [AVD21, SD17]. Individual-Level [AVD21]. Induced [Mar06a]. Inductive [Sch06]. Industry [Mat16b]. Inequality [Grö10b, JP16, MHdL12]. Infectious [AVD21, JGM18]. Infer.NET [LWOW18]. Inference [BFRP13, BGH^{+}17, BFH^{+}14, BEHB17, BGM06, CCF19, Can17, iM16, GTDT20, Grö10b, GV12, HIKS11, HL07, HRX16, IMR17, ILS11, KMC^{+}12, Kin04, KNI16, MHT07, RJH14, SH14, SH17, SVMMRP17, SHBZ14, Smi07, SBMD14, TDRD13, TDRD15, TR17, VW13, VGDM^{+}18, ZMG07]. Inferences [WCLS20]. inferference [SH17]. Inflated [KN05, LC10]. Informatics [Guh07, Lal16]. Information [BG18, CFSR15, Day01, Jos05, KH13, PV14, SSS^{+}20, Wan13a]. informR [MB15]. Infrastructure [FHM08, FK17, HH07a, TSH20, ZG05, vdLdJ21]. INLA [BGRR15, LR15a, MMR20, PPGRMB21]. inlinedocs [HWPG13]. Innovations [Pir10]. Inputs [GT10]. Inspired [MBK18]. Instrumental [AW16]. Intake [PZK^{+}12]. Intake epis food [PZK^{+}12]. Integer [Han06a]. Integrals [Som98]. Integrated [CP12, Edd09a, RFKM12, WdVW^{+}11]. Integration [Dah20, EF11, HHJCRB21, Kuo03, RR11]. Intensity [RG07]. Interaction [ESKHB20, NMB15]. Interactions [Tak16]. Interactive [BM97b, BCAB07, BGSC10, BW01, DKMT11, Gan15, Grö21, HKL09, Jos05, LLS15, NL12, SHFB17, SD13, The02, TD07, Udi05, Ver12, VYD^{+}12, lGlRGL09, Udi00, VM09a]. interactively [VMY02]. Interdependency [DK18]. Interface [AB12, CCH15, Den16, Dow16, Dun99, Fox05b, FC12, KCCG11, LL10, LSvdV19, MRHA20, RFKM12, SLS^{+}12, SGS^{+}14, UAK^{+}15, WLH^{+}18].

Interfaces [Cha16, HL09, Unw12, VML12]. Interfacing [Woo16]. Interference [SH17]. Intermediate [Har08, Hel17, Yal10]. Internal [KCM08]. International [CB17, LCK11]. Interpretation [San06]. Interpreter [DC09]. Interpreting [TWK03]. Interval [AB17, FS10a, TGJ17, YZZ^{+}20, FS10a]. Interval-Censored [TGJ17]. Intervals [Kel07, Kha18c, KPSH15, LCD18, New05, SB01, You10, MJ00]. Interventions [MJGM10]. Introducing [AH12, BKT14, Han05, Han06b, Han07b, Han12, LC10, LGF^{+}18, WMS17]. Introduction [AFJZ11, Edd09c, Eks13, GR18a, HHB08a, HBF17, Hel15, Hil15b, Hof15, Kha17a, KN17, KP07, Mai06, MvS07a, OHD17, San09, San11, Thi14, Tho11, WWH04, Xu06, dLM07, DN17, Hil10b, Mal09, Pie05, Sch08]. Introductory [Gou05, Kil16, San10b, VM09b]. intsvy [CB17]. Invariant [IP08, RS07]. Inventories [HMM21]. Inverse [LHS08, SP10, VSV09, vdWG11]. ipdpower [KSPR16]. ipw [vdWG11]. iqLearn [LLS15]. Irregular [PG15, ZG05]. Irregular-Shaped [PG15]. Irregularly [MGG^{+}04]. IRT [Bat15, Fox07, MH07, She08a, She08b, TMN16, Wee10]. IRT-Based [Wee10]. IRTrees [DP12]. ISAP [HPK18]. ISBN [But15, Gle16, Grö15a, Grö16, Gro08, Han15, Hel15, Hil15a, Hil15b, Hof15, How15a, How15b, How16b, Iac15, Kha15b, Kha15a, Kha16b, Mat15, Mat16a, Nun15, Pos15, Rus15, Zei15, Zei16]. isocir [BFRP13]. Isosurfaces [FT08]. Isotone [dLHM09]. Isotonic [BFRP13]. Isotopomer [BMvS^{+}07]. Isotropy [Wel18]. Issue [PLR^{+}16, Put11]. ISTE [How15b]. Itô [GB20]. ITA [Sch06]. Item [Cha12, Cha16, CGC11, Cur10, DBZ^{+}11, DP12, HGH05, Joh07, MF14, MPM14, MTvdM15, Riz06, Sch06, Whe14]. Items [ALV07, Cer17]. Iterative [CGC11, CHML17]. Iteratively [Pet97b].

J [Grö15a, Hel15, How15a]. Jackknife [New05, SP05]. Jacquez [JWHS16]. JADE [MNT17]. Jagger [Hil15a]. JAGS [Den16, Woo16]. James [Hil15a]. Jan [HJS16, Kro16, MM16, Tak16, Ylv16]. JASP [LSM^{+}19]. Java [dL05b, BW01, HKL09, Nar05, PSW07, SLS^{+}12, War02, WO97]. Java-Based [SLS^{+}12]. Java-implemented [PSW07]. Jean [Pos15]. JM [Riz10, GHR18]. JMbayes [Riz16]. jmcm [PP17]. JMFit [ZCI^{+}16]. John [But15, Hel15, Iac15, Kha15a]. Joint [DSH16, KMM^{+}17, MDvdV19, MNT17, PP17, Riz10, Riz16, XHW20, ZCI^{+}16]. Jointly [GHR18]. Jointness [BK15]. JOP [KR13]. Joseph [Grö16]. Joshua [Kha15b]. Journal [FL16]. Journalism [Grö19]. Joy [Yan07]. JSM [XHW20]. Judgment [Eva14]. Julia [Hel20]. Just [Woo16].

K-S [BH07, BH08a, BH08b]. Kalman [Tus11, Wan13b]. kamila [FM18]. kdecopula [Nag18]. Ken [Grö15a]. Kenward [HH14]. kequate [ABW13]. kerdiest [dREP12]. Kernel [ABW13, CCF19, DHM11, Duo07, FHH17, HS18, KSHZ04, MP14, MPM14, Nag18, Pav15, dREP12, Scr01, Wol12]. Kernel-Based [CCF19, FHH17]. kernlab [KSHZ04]. KernSmoothIRT [MPM14]. Kleinman [Grö15a]. kml [GASA15]. kml3d [GASA15]. knitr [McN14]. kNN [CF08]. Knowledge [BSVT12, Lal16, ÜS10, Xu06]. kohonen [WB07, WK18]. Kolaczyk [Nun15]. Kolmogorov [Car15, DKT20, MTW03, SL11]. Kriging [RGD12]. Kriging-Based [RGD12]. KRLS [FHH17]. Kroese [Kha15b]. ks [Duo07]. Kuhnt [Kha15a]. Kullback [LR20]. Kutz [How15a].

L1 [NO11]. Label [Pap16, TKM07]. label.switching [Pap16]. Lack [FW18, ZA17]. Lack-of-Fit [ZA17]. laeken [AT13]. Lag [Gas11, Mug10]. laGP [Gra16]. Lambda [Su07]. landsat [Gos11]. Language [CGH^{+}17, ESO16, SDDD12, WG00]. Languages [Wei05]. Laplace [Ger14, KNB^{+}16]. LARF [AW16]. Large [AGZ19, CB17, CFHBK11, DCS20, FBG15, Gra16, Hug07, Mur06, RPC15, VG09, ZMC21, MJ00]. Large-Scale [AGZ19, CB17, CFHBK11, DCS20, Gra16]. Lasso [HHJCRB21, SHBZ14]. Latent [ALV07, BPP17, Bea17, JFDB15, KH08, Lei04, LL11, Meu13, MD18, PLPL17, Riz06, She10, WM14]. latentnet [KH08]. LaTeX [Ruf09]. Lattice [PPGRMB21, SBL04, Gro08, Gro08]. Lattice-Ordered [SBL04]. lavaan [Ros12]. lavaan.survey [Obe14]. Law [HGG08]. lawstat [HGG08]. Lazily [Har15]. LazySorted [Har15]. LC [KSS^{+}07]. LC/MS [KSS^{+}07]. lcmm [PLPL17]. LDheatmap [SBMG06]. LDR [CFT11, AR14]. Leader [TKM07]. Leakage [BG18]. Learning [AGZ19, BD03, GZP14, KB16, Kur14, LLS15, Mai09b, MW19, Sch08, Scu10, Scu17, SSM21, VR14, ZUL14, Hel16a, Mai07, Mai09a]. Least [Das06, FHH17, KS10, Len16, MW07, ML12, Pet97b, RS08]. Least-Squares [Len16]. Leeuw [HJS16, Kro16, MM16, Ylv16]. Leibler [LR20]. Leroux [GF15]. Level [AVD21, Esm14, FS13, Grö14, JMD08, KO06, LSC09, Rec10b, RD21]. Level-Dependent [KO06]. Lévy [DC05]. Lexis [CP11, PC11]. lgcp [TDRD13]. LibBi [Mur15]. libDirectional [KGP^{+}19]. Library [BLR97, LIL^{+}15, MRHA20, PQM11, Scr01, War02, PQM11]. libstable [dVSWAL17]. Life [ASBMB17, SVMMRP17]. lifecontingencies [Spe13]. Lifetime [LMB^{+}07]. Light [TFR16]. Like [DV20, ZMG07]. Likelihood [AR14, BGM06, Cha10, CYK^{+}09, CFT11, CGS09, GvdL12, GV12, Hua20, Joh07, Jur15, KM08, MVS13, NMB15, OC15, RC17, RMG12, Su07, TSN21, YSH15, Yua07]. Likelihood-Based [AR14, CYK^{+}09, CFT11]. Likely [Hot20]. Likert [HR14]. Line [HL09, Hof11]. Linear [Arm19, AMW14, BE13, BMBW15, But05, CdM10, Can04a, Dun99, Fox03, FBF14, FHT10, Fri06, Gas11, Ger14, Grö06, Grö10b, Had10, HH14, Hel16b, HGK^{+}20, JLZJ05, JP16, JMS^{+}09, KL14, Kol16, KMG^{+}13, KBC17, LFF17, LGF^{+}18, LH14, McI16, MYK07, ML11, Pet97a, Pet10, RBN19, Rec10b, San07, SHBZ14, SBMD14, TV11, TM05, VSV09, VDT14, WM18, WZQW20, WG00, ZLHK02, ZLC18, dSJdSF14, Caf06, Pet97b, VMY02]. Link [IoG10]. Link-Based [IoG10]. Linkage [SBMG06, TB17]. LinkCluE [IoG10]. Linked [PMW^{+}15]. Linking [Alb16, Wee10]. Linux [Oom13]. Lion [MM16].

Lisp [Jos05, Nar05, Tie05, VMU05]. LISREL [Kha17b]. List [Har15]. Literate [LH07, SDDD12]. lmdme [FBF14]. lme4 [BMBW15, DBZ^{+}11, DBBD07]. lmer [DBZ^{+}11]. lmerTest [KBC17]. LMest [BPP17]. lmSubsets [HGK^{+}20]. loa [LR15b]. Local [AW16, BKT14, BMB16, CQZ^{+}10, Gra16, RC17]. LocalControl [LNY^{+}20]. localgauss [BKT14]. Localization [WdVW^{+}11]. Locally [CN13, EN11, TPE19]. Location [HN13, Lec14, SR07b]. Log [Caf06, CGS09, DM18, DR11, Pet97a, TDRD13, TDRD15, VMY02]. Log-Binomial [DM18]. Log-Concave [CGS09, DR11]. Log-Gaussian [TDRD13, TDRD15]. Log-Linear [Pet97a, Caf06, VMY02]. Logarithmic [XTL13]. logbin [DM18]. LogConcDEAD [CGS09]. logcondens [DR11]. Loggamma [AMYR16]. Logistic [CGC11, Dha10, FS12, Grö16, Hel16b, HWY18, KZ03, RT14, SED14, WZQW20, ZMG07, Hed99]. Logit [FH09, HWM16, HZSW10, SD17]. Loglinear [BR07]. Logrank [FS10a]. London [How15b]. Long [PC11]. Long-Term [PC11]. Longitudinal [ASBMB17, BPP17, BdMM08, CBDM07, DLC06, GHR18, GASA15, GCA12, KK14, KMM^{+}17, LSPvdL17, NGBK12, OC15, PP17, Riz10, Riz16, SP14, SvdLN17, XHW20, ZCI^{+}16]. Looking [AJ11, vdHvB16]. lordif [CGC11]. Loss [LSPvdL17]. Loss-Based [LSPvdL17]. Lost [Tak16, Wei05]. Lovelace [Gle16]. Low [BM07, Coo97, HCW12]. Low-Dimensional [Coo97, HCW12]. Low-Resolution [BM07]. lp solve [But05]. lqmm [Ger14]. lrasch mml [OC15]. LS2W [EN11]. lslx [Hua20]. lsmeans [Len16]. ltm [Riz06]. ltmle [LSPvdL17]. lubridate [GW11]. Luke [Mat16a]. lumberjack [vdL21]. LyX [Kar06].

M [Grö16, SH20]. M-Estimation [SH20]. m2r [KOS20]. Macaulay2 [KOS20]. Machine [KB16, Kur14, Mai07, Sch08, SSM21]. Machines [KMH06]. Macro [Dha10, FS12, GHR18, Gre17, Hla16a, Jur15, KS10, LG11, LV16, MH05, OC15, Pan09, PKC^{+}20, SH09, SR07a, VDT14, VGDM^{+}18, Wei12, XM14, ZCI^{+}16, ZK12]. Macros [LVK^{+}11, Law08, LS02b, NO06]. Made [GW11, HM16]. Magnetic [CMS^{+}11, KCCG11, TW11, WS11]. Main [Cha03]. Main-effect [Cha03]. Majorization [GvdV16, dLM09b]. Make [Bak20]. Making [Wic10]. Mallows [ICL16]. MAMS [JPM19]. Man [EL09, Woo05]. Manage [Bak20]. Management [BBGL17, Grö15a, MJGM10]. Managing [BdMM08, BK17, But08a, Grö18a]. Manifold [ACW12, BILL10, Mai09a, MRHA20]. ManifoldOptim [MRHA20]. Manipulating [ZFH^{+}20]. Manipulation [Wic08b]. Mann [FO15]. Mantel [BV02]. Many [CBAA19]. Map [LR15b, TB17, Sho13]. Mapping [FF14, KPRP^{+}19, TKM06, Zho10, Sab07]. MapReduce [VR14]. Maps [Grö21, TP11, Ten18, WB07, WK18]. MARCH [Ber01]. Marginal [CGC14, Joh07, KAK05, OC15]. Marginalization [Mar06a]. Marginalized [VGDM^{+}18]. Marginals [AH08]. Margins [KY10]. Marin [Pos15]. Markdown [Mor18]. Marked [Har10]. Markers [Gra15]. Marketing [Rus15]. Markov [Kha17a, San11, ABB^{+}19, BFFN19, BPP17, Ber01, FS10b, HD10, HH19, KCCG11, KNI16, KSP15, Mar06a, MQP11, OH11, Sch16, She08a, She08b, VV16, VS10, War02]. Markov-Switching [ABB^{+}19]. Markus [Han15]. Marries [Thi14]. Marsaglia [Bre04, Vig16]. Mass [CS12, Grö19]. Massive [KEW13]. MATCH [Wie04]. Matching [IP08, IKP09, KSS^{+}07, RBB18, Sek11, Wie04]. MatchIt [HIKS11]. Mateda [SBL^{+}10]. Mateda-2.0 [SBL^{+}10]. Materials [Dry09]. Mathematica [BH07, BH08a, BH08b, HSH07, LP13, Per03]. Mathematical [JGM18, Mil04]. Mathematics [Spe13]. mathStatica [Sto12]. MATLAB [Pie05, ÁBZ17, ÁBZ20, Ber09, CGRvD15, CSY15, Dun99, Esm14, Góm15, GHH20, HPK18, IoG10, Kao09, KT10, LZ17, LZ15, LSvdVK09, LSvdV19, McI16, PA11, RS08, dVSWAL17, RSW15, SBL^{+}10, She08a, She08b, She10, SCL^{+}18, VFV13, YR19, Edl04, How15b, Bow10]. Matrices [Bra17, RG07]. Matrix [ASB11, EVBK19, FS10b, Kha16b, KN03, Mon06, Pet97a, Zei04, dL08, dL08]. MATS [KT10]. Maximum [CGS09, GvdL12, Joh07, Jur15, KM08, NMB15, OC15, Su07, TSN21, VdL09, Yua07]. mbonsai [HNCI18]. MCAR [JJJ14]. McDonnell [Rus15]. mcglm [Bon18]. mclust [FR07]. MCMC [Den16, iM16, FS10b, Had10, Pap16, Riz16, Smi07]. MCMCglmm [Had10]. MCMCpack [MQP11]. MCPMod [BPB09]. MCSim [BM97a]. mdscore [dSJdSF14]. Mean [BM12, Mon06, PP17, PKC^{+}20, SF16, SF19]. Mean-Covariance [PP17]. Means [Bel11a, HFKB12, KSBZ16, Len16, RK14, SBL04]. Measure [MH15]. Measurement [HXY12, PZK^{+}12, Tiv19, WW11]. Measurements [MvS07b]. Measures [BK15, Fri10, KAK05, KT10, LW03, Rec10b]. meboot [VdL09]. Mechanisms [BLM^{+}15]. medflex [SLMV17]. Median [FMM^{+}20]. Mediation [SLMV17, TYH^{+}14, TYH^{+}14]. Medical [Hil09b, JS13, NRG^{+}17, Spa06]. Meet [HM11]. mefa [Sól09]. Meißner [Iac15]. Membership [SP14]. Memetic [BMB16]. Menu [BP12]. merDeriv [WM18]. Meta [BER^{+}16, Bur12, CO16, GR18b, GV12, KR09, LZHC17, LCSC14, MF15, Röv20, Ver18, Vie10, WF12, How16a]. Meta-Analyses [Vie10]. Meta-Analysis [BER^{+}16, Bur12, CO16, GR18b, GV12, KR09, LZHC17, LCSC14, MF15, Röv20, Ver18, How16a]. Meta-Statistics [WF12]. meta4diag [GR18b]. Metabolic [BMvS^{+}07]. MetaEasy [KR09]. metafor [Vie10]. metaLik [GV12]. Metamodeling [RGD12]. Method [ABW13, CC08, Cha10, Das06, Dem11, FR15, HSH07, HL09, Kur14, LZL^{+}05, MT98, NMB15, Pen08a, SMS13, Yua07, MT00]. Methodologists [WDT^{+}12]. Methodology [GR16b, HNCI18, AFJZ11]. Methods [ADH11, BMR14, BP15, Boy05, Chr11, CKO11, DMB18, Den16, Der10, DKMT11, Doa11, DG11, ESKHB20, FR07, FS10b, Gon14, GRS12, GB20, Had10, HH14, Ham10, How15a, Iac17, JvdBP07, KSHZ04, Kha18b, KAK05, KG17, KR10a, Len09b, Luc11, Nas14, Nor17, NO11, OCRC14, Pel11, RD21, Ros17, Scr01, Sim07, Str04, Su07, ÜS10, Wee10, Wic10, Xie13, dL06, dL09c, dLM09a, dLHM09, Har08]. Metric [IP08, ZPZW21]. mexhaz [CB21]. MF [SWAF15]. MFSAS [CC11]. mgcv [Woo16]. mgm [HW20]. mhsmm [OH11]. mi [SGHY11]. MICE [RW11, vBGO11]. Michael [Hel15]. Michel [Pos15]. Micro [RD21, TKM15]. Micro-Data [TKM15]. Micro-Level [RD21]. Microarray [CYK^{+}09, DM19, FO15]. Microarrays [EC16]. Microdata [Kav15]. micromap [PMW^{+}15]. Micromaps [PMW^{+}15]. micromapST [PPC15]. Microscopy [LMB^{+}07]. Microsimulation [Gle16]. Microsoft [KR09]. microsynth [RD21]. MIDAS [SH09]. midasr [GKZ16]. Million [Mur06]. mimR [Høj04]. Minimal [Cha03]. Minimum [FS13, LSPvdL17]. Minimum-Size [FS13]. Mining [But15, Fay18, FHM08, Fuj17, HGH05, Lal16, Mai08, Ros09, TFH12, Wic10, Iac15, Xu06]. Minitab [GL14]. MinPROMEP [Cha03]. mipfp [BS18]. MIRT [She10, Cha12]. Mises [HG14, XGY06]. Missing [CBDM07, CCH15, Dem18a, HKB11, JJJ14, JH16, Kha20]. Missingness [WCLS20]. missMDA [JH16]. MissMech [JJJ14]. MitISEM [BGH^{+}17]. Mix [SH17]. mixAK [KK14]. Mixed [BMBW15, CGK11, CA17, CLL17, DKT20, FS13, FM18, GHR18, Ger14, GKZ16, GKZ12, Had10, HH14, HW20, HN13, HSG12, JP16, Kol16, KSPR16, KBC17, LSC09, Lec14, LG11, MF15, NW04, PLPL17, Sch11, TV11, VDT14, WM18, Wee10, ZLC18, Hed99, Chr09]. Mixed-Effects [BMBW15, HN13, Kol16, Lec14, Hed99]. Mixed-Format [Wee10]. Mixed-Level [FS13, LSC09]. Mixed-Type [FM18]. mixexp [LW16]. MixGHD [TBE^{+}21]. MixMAP [MF15]. Mixmod [LIL^{+}15]. MIXNO [Hed99]. MIXREGLS [HN13, Lec14]. MixSim [MCM12]. mixsmsn [PCL13]. mixtools [BCHY09]. Mixture [AHvD09, BCHY09, BSVT12, BER^{+}16, BMGT15, FSLZ12, HH19, HS18, KG17, LW16, LM18, Lei04, LHA^{+}15, MSCCV05, MMR20, PCL13, Sav16, Su07]. Mixtures [Can17, GL08, HG14, LM13, MNA^{+}97, PMM18]. MLDS [KM08]. MLGA [Jur15]. mlirt [Fox07]. mlogit [Cro20]. mlt [Hot20]. MLwiN [LC13, ZPC^{+}16]. MM [AH12, DV20]. mmeta [LCSC14]. mnlogit [HWM16]. MNM [NO11]. MNP [IV05]. Mode [AACRC21, Meu13, MD18]. Model [AEL10, AMYR16, AWBM18, BdUÁ18, Bar96, BBG12, BIG17, BK15, BYPJ19, Bro15, CdM10, CN18, Den16, DM18, DBBD07, FR07, GRMMR19, GRS12, HD10, HMS16, HZSW10, HSG12, HWY18, Jal19, LIL^{+}15, Lei13, LM19, LC10, MRC15, Mar06a, MH05, MMRP11, Mel16, MGH18, NW04, PZK^{+}12, PT11, San07, Sch11, SR18, She08b, TMW18, TBE^{+}21, WB08, ZC10, ZF15, dS11, IV05, SFHT11, Gum07]. Model-Based [AWBM18, BBG12, BIG17, BYPJ19, Bro15, GRMMR19, LIL^{+}15, Mel16, SR18, TBE^{+}21, FR07, Gum07]. Modeler [Woo16]. Modeling [AB12, BG18, Bai20, BS15, BSVT12, BG21, BS13, Car13, CYK^{+}09, CBHK21, Der10, DOV17, FR15, Fox07, FEvdL07, GF15, GTDT20, Gra16, Grö15b, HD12, Hel16b, How15a, How15b, HS18, Hua20, Jac16, JHQ^{+}11, JGM18, Kha15b, KCCG11, KP07, KY10, KSPR16, KMM^{+}17, Kuk09, LC13, Lee13, LRN18a, MdlT20, MH07, MF15, ML12, MvS07b, NE20, PP17, PR07, PRS18, Riz06, Ros12, SH07, Sel11, SKS15, SCL^{+}18, SYC08, SF16, SF19, Tri03, VPCS14, VKM18, VP18, XHW20, YFK^{+}15, Ziv11, SCD07]. Modelling [Har10, HC05, Høj04, LR15a, Mur15, PHF10, Riz10, SP10].

Models [ASB11, AVD21, AB17, ALV07, AMMRP14, ABC19, ABB^{+}19, Arm19, AMW14, BTMB13, BR07, BP19a, BPP17, BMBW15, Bel11b, BCHY09, BER^{+}16, BVE^{+}15, Bon18, Bos11, BKL05, BRG20, BMGT15, Bür17, Bur12, Caf06, CdM10, Can04a, CG15, CP11, CB21, CP12, CKY14, Chr09, CGC14, CLL17, Cro20, DBZ^{+}11, DP12, Den16, DH05, DLC06, Dun99, Eva14, FDB12, Fie12, FBC07, FBG15, Fir05, FC11, Fot03, Fox03, FH09, FW18, FBF14, FSLZ12, FHT10, Fri06, Gan15, GHR18, Gas11, GRK^{+}16, Ger14, GF15, GKZ16, GLC^{+}15, GR18a, Góm15, Gon14, Gra07, GT10, GH11, Gu14, Had10, HH14, HWM16, HW20, HXY12, Hel16b, HH19, HHV20, Hoh18, HL07, HRX16, HHB^{+}08c, ICL16, IM05, JvdBP07, Jac11, JLZJ05, JFDB15, JP16, JCPL19, Joh07, JMS^{+}09, Jon07, KMC^{+}12]. Models [KN05, KL14, Kha17a, KBMB19, KAK05, KR10a, Kol16, KV13, KMG^{+}13, KH08, KSP15, Kuh08, KBC17, LS02a, LCD18, Lei04, LFF17, LG11, LHA^{+}15, LH14, LP17, Lum96, Mai06, MH07, MPI18, McI16, MSCCV05, MMR20, Men11, MR18, MP12, Mil17, MW19, ML11, MTvdM15, MBR11, MHD17, MHH08, MHdL12, MRL12, NAA17, OH11, OHD17, Obe14, OC15, PPGRMB21, PPGD15, Pen03, PA11, Pet97b, PV14, Pet10, PP11, Pfa08, PKC^{+}20, Pir10, PLPL17, Put11, QY19, Rec10b, Riz16, RST19, RMG12, SF18, San07, San11, Sar16, SD17, SP14, Sch11, SL18, She08a, She10, Sho13, SHBZ14, SR07b, SLMV17, SBMD14, SM07, TKM17, TSN21, TV11, TR17, TM05, TR14, TH08, TGJ17, TF12, UNBK17, UAK^{+}15, Van11, VS10, VDT14, WW11, Wan13b, WM18, WZQW20]. Models [Whe14, WMR16, XWHL15, ZMG07, ZLHK02, ZKJ08, ZCI^{+}16, ZLC18, ZHZ20, dSJdSF14, dWFP11, Hil09a]. Modern [DKMT11, Gou10, Hil15a, Dow17, Mai09a, Rui16, dL05c, dL12]. Modes [Lei10]. Modifed [Tri03]. Modifying [MB15]. Modular [Han06b]. Module [SBMD14]. MOEADr [CBA20]. Mokken [vdA07, vdA12]. Molecular [BKvT^{+}14]. Moment [GB20]. Moments [Cha10, Phi10]. Momocs [BPGC14]. Mondrian [The02]. Monitoring [SSH16, vdL21]. Monkeying [Mar05]. Monogan [How16b]. Monotone [DV20]. Monte [BFFN19, KCCG11, BM97a, CGRvD15, CGC11, DK18, GB20, Joh09, LT16, MQP11, She08a, She08b, SP10, War02, Zho15]. Monty [MT98]. Moodle [ZUL14]. Morgane [Gle16]. Morphometrics [Bow09]. Morse [GP12]. Mortality [CB21, MP14, Mug10, VKM18]. MortalitySmooth [Cam12]. Most [Hot20]. mosum [MKC21]. Mother [NMB15]. motion [Coe00]. Motions [DC05]. Moult [EBO^{+}13, EBO^{+}13]. Moving [LRN18b, MKC21]. movMF [HG14]. MPCI [SFS12]. mplot [TMW18]. MR [EC06]. MRI [Ach09, FT11, Fer11, TP11]. mritc [FT11]. MS [KSS^{+}07]. msBP [Can17]. MSGARCH [ABB^{+}19]. msm [Jac11]. msr [GP12]. msSurv [FDB12]. MST [CSNF18]. mstate [dWFP11]. MTPmle [VGDM^{+}18]. multgee [Tou15]. Multi [ASB11, BP19b, CP11, Had10, Jac11, JPM19, Kao09, KSP15, LQC^{+}12, MvB14, MZH06, MvS07b, MLMK12, Nag21, Pen03, Put11, SDDD12, She08b, dWFP11]. Multi-Arm [JPM19]. Multi-dimensional [Pen03]. Multi-Language [SDDD12]. Multi-Objective [BP19b, Kao09, MLMK12]. Multi-Regional [LQC^{+}12]. Multi-Response [Had10]. Multi-Stage [JPM19]. Multi-State [ASB11, CP11, Jac11, KSP15, Put11, dWFP11]. Multi-Subject [MvB14]. Multi-Threading [Nag21]. Multi-Unidimensional [She08b]. Multi-way [MZH06, MvS07b]. Multiblock [BD18]. MultiBUGS [GTDT20]. Multiclass [CC08]. Multicore [Hof11]. Multidimensional [BS18, Cha12, Cha16, GvdV16, LBW18, dL05c, dLM09b]. Multidisciplinary [Kuk09]. Multifactor [Pet97a]. Multifractional [DC05]. Multilevel [Bür17, CGK11, CC11, DBBD07, Fot03, Fox07, Grö15b, Hil09a, LC13, dS11]. Multilevel/Hierarchical [Hil09a]. multimode [AACRC21]. Multinomial [CGC14, FH09, HWM16, SD17, Tou15, MJ00, IV05]. Multiobjective [CBA20]. multipanelfigure [GC18]. Multiphase [HMM21]. multiPIM [RJH14]. Multiple [ALV07, Bon18, CGK11, CFHBK11, GC18, HWY18, JM15, JLZJ05, Kla18, KPSH15, KT10, KR13, LSvdV19, OH11, Ost20, SP14, SH09, SB01, SDJ20, SGHY11, WGSL12, Yua11, Yuc11, ZC10, RW11]. Multiple-Objective [HWY18]. Multiple-Table [WGSL12]. MultipleCar [LSvdV19]. Multiplex [MRV21, Ost20]. Multiplication [LBW18]. Multiplicative [AH12, Pet97b]. Multiplicity [SFC20]. Multiplicity-Adjusted [SFC20]. Multiresolution [SCS13]. Multiscale [Can17]. Multistate [FDB12]. multitable [WGSL12]. Multivariable [AS02, CCH15]. Multivariate [AG19, AS04, AH12, AWBM18, BS18, BEHB17, CSNF18, CSY15, CHML17, CO16, CGS09, Duo07, Eks10, Eva11, FDGD16, FBC07, FBG15, FBF14, Fri06, GKSU15, Gro08, Han12, HHV20, JM15, JJJ14, JMS^{+}09, JH16, Kha16a, KY10, KK14, LDHM21, LCL^{+}08, LJH08, LM13, LM03, LBC^{+}16, LCSC14, Mai09a, Mai09c, MdL10, NO06, NOT08, NO11, PPGRMB21, Pen08a, Phi10, PMM18, San06, SFS12, SMM^{+}15, Sek11, Som98, Som01, TDRD15, TPE19, TD07, TF09, Tri03, VW13, WCHB11, Zei16, dL05a, vdHvB16, vBGO11, Kha16c, Kha17b, Mül16, Ros17]. multivator [Han12]. multiwave [AG19]. Multiway [Kro16, Lei10]. Munzert [Iac15]. mvLSW [TPE19]. mvord [HHV20]. MXM [LAF^{+}17]. My [Kro16, Tak16].

nacopula [HM11]. Nash [CI17]. Nathan [How15a]. Native [Wei12]. Natter [SLGB14]. Natural [SLGB14, SLMV17]. Navigation [HO20]. NbClust [CGBN14]. Near [RBB18, Was15]. Near-Far [RBB18]. Near-Optimal [Was15]. nearfar [RBB18]. NeD [SGS^{+}14]. Need [Wei05]. Needed [PV14]. Nested [BFM18, HM11]. Nestedness [SGS^{+}14]. Net [SHBZ14]. netCoin [EMU20]. NetLogo [Thi14]. Network [BdMM08, BB12, But08b, ECW^{+}12, EMU20, HHB^{+}08b, KLNN20, LZHC17, Mat16a, Nun15, PV14, RBN19, Scu17, SD13, VR14, But08a]. Networking [LCK11]. Networks [AGZ19, Bec18, BB12, BMvS^{+}07, BD03, DCS20, GD21, GW18, Høj12, HPK18, HHB^{+}08c, JGM18, KH08, MRV21, MGFA^{+}20, Ost20, PSS20, Scu10, VPCS14, YMB20]. networksis [AH08]. Neuhäuser [Han15]. Neural [Bec18, BB12]. NeuralNetTools [Bec18]. neuRosim [WDM^{+}11].

Neutral [Han07b]. Newdistns [NR16]. Newton [Das06, MHJS16]. nhorton [Grö15a]. nhorton/r2/ [Grö15a]. NHPoisson [CAA15]. Nicholas [Grö15a]. Nick [Kha16b]. NIfTI [WST11]. Nineteen [AJ11]. NISS [SGK03]. nlsem [UNBK17]. nlstools [BRC^{+}15]. Node [PSS20]. Nodes [HD16]. nominal [Hed99]. Nomogram [CG15]. Non [Ber01, Cha16, CA17, Gas11, LS02a, MS19, MW10, Ngu07, Per03, Pet97b, PSS20, SP14]. Non-Adaptive [Cha16]. Non-Bayesian [MS19]. Non-Binary [Ngu07]. Non-homogeneous [Ber01]. Non-independence [LS02a]. Non-Linear [Gas11, Pet97b]. Non-Parametric [CA17, Per03, PSS20, SP14]. Non-Standard [MW10]. nonbinROC [Ngu07]. Nonhomogeneous [CAA15]. Nonlinear [BRC^{+}15, BFC02, CLL17, Fot03, Gan15, Gra07, Jam99, PZK^{+}12, UNBK17, VG09, VDT14, vdHvB16, dL09a]. Nonnegative [FDGD16]. Nonparametric [ABS01, AM14, BEHB17, CCF19, Can17, Can04b, CQZ^{+}10, Con03, DLN17, DSD13, FDB12, HR08, HIKS11, JM15, JHQ^{+}11, KK15, KPSH15, MS19, MVS13, MW12, Mül16, NGBK12, OCRC14, PP18, dREP12, Ros15, Sav16, Scr01, SVMMRP17, SSM21, SBL04, Wel18, ZA17, ZK12]. NonpModelCheck [ZA17]. Nonstandard [Gon14]. Nonstationary [Gra07]. nopp [CI17]. Normal [AD15, Eks10, Mar04, Mar06b, MP99, Mey13, MHD17, Phi10, PCL13, PMM18, She08a, Som98, Som01]. Normality [JJJ14]. Normalizing [GSW20]. normalp [MR05]. North [TKM07]. Note [Bre04]. Notes [Tie05]. Novel [Cha03]. November [LCK11]. np [HR08]. nparcomp [KPSH15]. NParCov3 [ZK12]. nparLD [NGBK12]. npbr [DLN17]. NPCirc [OCRC14]. npmv [BEHB17]. npregfast [SVMMRP17]. nprobust [CCF19]. NScluster [TSN21]. Nucleotide [SBMG06]. Null [ZFZ10]. Number [Bre04, CGBN14, Dry09, SIR^{+}11, XTL13, YS12, MJ00, PL05, SM12, SLF14]. Numbers [AS04, SMDS11]. Numerical [ALO^{+}10, BE13, HSH07, Kha18b, Kuo03, Mey13, Pie05, Som01, Var14, Mat16b]. Nutshell [Edd11, Leo10]. NY [Gro08]. Nyhuis [Iac15].

OasisR [Tiv19]. Object [ATF10, BM97b, HK20, Kle16, PR07, SHR96, TF09, Zei06, ZKG20, Udi00]. Object-Oriented [ATF10, BM97b, HK20, Kle16, PR07, SHR96, TF09, ZKG20, Zei06, Udi00]. Objective [BP19b, DR20, HWY18, Kao09, MLMK12, VG09]. Objects [CP11]. Observation [DS15, OH11]. Observation-Driven [DS15]. Observations [TGJ17, VV16]. Observed [Alb16, KNI16]. Observed-Score [Alb16]. OBsMD [DR20]. Occupancy [MBGK18]. Occurrence [FC11, GVM16]. Octave [GHH20]. Odds [FH09, Wan13a]. Off [DV20]. Off-the-Shelf [DV20]. ofw [CC08]. Ogive [She08a]. Ohri [Mat15]. Oja [FMM^{+}20]. OjaNP [FMM^{+}20]. OLAT [ZUL14]. Oncology [ASBMB17, WS11]. One [BH07, BH08a, BH08b, Pat16, PKC^{+}20]. One-Dimensional [Pat16]. One-Factor [PKC^{+}20]. One-sample [BH07]. One-Sided [BH08a, BH07]. OneArmPhaseTwoStudy [KWE^{+}17]. Ones [Mat16c]. Online [Din06, KMG^{+}13, Zha14]. Open [DKMT11, PQM11]. Opening [SGHY11]. OpenMP [TSN21]. Operating [GG16, LW03]. Opportunities [Wei05]. OptGS [Was15]. optim [Var14]. Optimal [CI17, GZP14, HZSW10, HWY18, Kao09, KPH^{+}21, LRRÁCSGS14, MdL10, MvB14, OWA20, TH21, TSV20, TBS14, Was15, YZZ^{+}20, dLM09a, Grö11]. OptimalCutpoints [LRRÁCSGS14]. Optimization [ACW12, Bar14, BDKL17, BP19b, Bra14, FNB20, GT10, KM14, KR13, MRHA20, MS11, MAG^{+}11, Mul14, NV11, Nas14, RGD12, Rui16, Sek11, TSH20, Var14, dLHM09]. Optimizations [MBK18]. Optimized [Scu17]. Optimizing [VG09]. Optimum [PJSPC17]. optimx [NV11]. Orange [HZS21, SD13]. Orbits [HD16]. Order [HHB08a, KH13, RBN19, Tiv19]. Order-Restricted [KH13]. Ordered [MMM12, SBL04]. Ordinal [AD15, Arc10, BP12, CGC11, GSD12, Hel16b, HHV20, LX12, MHM20, MH05]. ordinalCont [MHM20]. OrdNor [AD15]. Organizing [WK18, WB07]. Orientation [Mur03]. Oriented [ATF10, BM97b, HK20, Kle16, PR07, SHR96, TF09, ZKG20, Udi00, Zei06]. Orientlib [Mur03]. Origins [Gow16]. Orthogonal [ADN15, BdMM15, Cha03, FS13]. Oscars [Unw12]. osDesign [HSL11]. Other [BGM06, Dem11, DV20, HR14, HM16, RG07]. Outcomes [KMM^{+}17, RSW15, TM15]. Outlier [Law08, McB06]. Outliers [PRS18]. Outline [BPGC14]. Output [Han05, Lei13, Smi07]. Outputs [MRC15, Pap16]. Over-Dispersed [SF19]. Overall [CB21]. Overdispersed [KM16, MHD17, SF16]. Owen [Pat00]. Ox [Bos11, HB17, Pel11]. Ox/SsfPack [Pel11]. Oxford [Han15, Hil15a, Hil15b, How15a].

P [Gle16, Grö16, Gro08, Han15, Hel15, Hil15a, Hil15b, How15a, Kha15b, Mat15, Mat16a, Nun15, Pos15, Rus15, Cam12]. P-Splines [Cam12]. p3state.msm [MMRP11]. PaCAL [KJ14]. Package [ADH11, AEL10, AH08, ACW12, AR14, AMYR16, Alb16, AGG13, ATF10, AT13, ASB11, AVD21, AH12, AD15, AACRC21, ABW13, AWBM18, ASBMB17, AGZ19, Arc10, AHvD09, ABC19, ABB^{+}19, AMW14, AM14, BG18, BT05, BDMP15, BK11, BFFN19, BP19a, BdUÁ18, Bar14, BFRP13, BS18, BPP17, BGH^{+}17, BE13, Bat15, BRC^{+}15, Bea17, BBGL17, BCHY09, BDKL17, BKT14, BBG12, BMB16, BIG17, BSVT12, BK17, BdMM15, BP19b, BMvS^{+}07, BVE^{+}15, BN10, Bon18, BDdM11, BPB09, BD03, BD18, BCAB07, Bra15, Bra14, Bra17, BS13, BWMBC21, BPDD08, BKvT^{+}14, BEHB17, Bür17, Bur12, But08a, CdM10, CSNF18, Cam12, CKE20, CBA20, Can17, CC08, CDP20, CB17, CN18, CFHBK11, CAA15, Cer17, Cha12, CQZ^{+}10, CGBN14, CB21, CKSLS12, CC11, CA17, CKY14, CGC11, CBHK21, CGC14, CFSR15]. Package [CFT11, CHML17, CO16, CM08, Cro20, CF08, CL13, CGS09, CJM06, CI17, DMB18, DLN17, DBZ^{+}11, DC05, DR20, DMD15, Den16, DH05, DW17, DCS20, DM18, DBBD07, DK18, DD07, DOV17, DV20, DM19, DPSH18, DS15, DGJ^{+}18, DGP08, EWMR20, ESKHB20, EE07, EBO^{+}13, EMU20, Esm14, FS10a, FBdlF12, FT11, FDB12, Fer11, FDGD16, FH10, Fie12, FBC07, FO15, FMM^{+}20, FC11, Fok20, Fox07, FEvdL07, FH09, FC12, FS20, FM08, FSLZ12, FKP17, FGEMI12, FNB20, FS10b, GR16a, GRD13, GSD12, Gan15, Gas11, GGC^{+}15, GRK^{+}16, Ger14, GP12, GG16, GKZ16, GK16, Gil15, GKD14, Gio09, GGK10, GLC^{+}15, GKSU15, GCA12, GFS14, GDMB08, GFC12, GU07, Gos11, Gra15, Gra07, GT10, GvdMW09, GD21, GW18, Grö06, Grö10b, Grö14, Grö18b, GSW20]. Package [GvdL12, GH11, Gu14, GB20, GV12, Had10, HPWdL15, HHB08a, HX13, HHY06, HH14, HNCI18, HD18, HSL11, Han06b, Han07b, Har10, HWM16, HD12, HR08, HXY12, HM18, HP09, HH19, HH07b, HMM21, HHV20, HM11, HO20, HMS16, Hoh18, Høj04, HL07, Høj12, HMR^{+}13, HG14, HHvdWZ08, Hot20, HPK18, HSG12, HGG08, HY21, HHB^{+}08c, HGH13, HK08, HWY18, IP08, IMR17, IoG10, IV05, ILS11, ICL16, JvdBP07, Jac11, JPM19, Jal19, JM15, JJJ14, JP16, JGM18, JLH^{+}20, JD15, JCPL19, Jon07, JH16, KOS20, KMC^{+}12, KSHZ04, KL14, Kas16, Kav15, KK15, KBMB19, Kie08, KWE^{+}17, KE14, KO06, KNI16, Kle16, KLNN20, KN03, KG17, KR10a, KY10, Kol16, KK14, KSBZ16, KPSH15, KJ14, KV13, KT16, KPRP^{+}19, KSP15, Kuh08, KR13, KPH^{+}21, KR10b, KBC17, LAF^{+}17]. Package [LDHM21, LRGTA12, LNY^{+}20, LJH08, LIL^{+}15, Lee13, LM13, LM18, Lee18, LRN18a, LMO^{+}19, Lei10, LHS08, LSPvdL17, Len16, LR20, LBW18, LM14, LM19, LFF17, LZHC17, LL11, LBC^{+}16, LC10, LZ15, LSL20, LHA^{+}15, LRRÁCSGS14, dUJ13, LSvdVK09, LX12, LCSC14, MdlT20, MRC15, MR12, MF14, MB17, MHJS16, MS19, MH07, MdL10, MHE20, MHM20, Mar06a, MH15, MGFA^{+}20, MBGK18, MF15, MP14, MPM14, MYK07, MZ08, MS11, MMR20, MKC21, MCM12, Mel16, MJGM10, Meu13, MW07, MHH17, MVS13, MW12, MR05, MW19, ML11, MTvdM15, MBR11, MV14, MHD17, MdUÁC10, MN14, MvS07b, MAG^{+}11, Mur03, Mur09, MG09, NR16, Nag18, NG07, NMB15, ND12, NGBK12, NOT08, NO11, NAA17, OH11, OHD17, Obe14, OR14, OCRC14, OL17, Oom13, OK14, OWA20, Pac07, PPGD15, PP17, Pan18]. Package [Pap16, PLRC10, PMW^{+}15, PSS^{+}17, PG15, PSY20, PP18, Pet10, Pfa08, PSS20, PV20, PO13, PU13, PMD19, PT07, PT09, PT11, PLPL17, PMM18, QY19, dREP12, RT14, RBB18, RC17, RJH14, Riz06, Riz10, Riz16, RST19, RR11, RMG12, Ros15, Ros12, Röv20, RVCG19, RSW15, SF18, SBL^{+}10, SFS12, Sar16, SD17, SIR^{+}11, SZ11, SED14, SMM^{+}15, Sch16, SlRC16, Scr13, SR18, Scu10, Scu17, SVMMRP17, She08b, Sho13, SS19, SKS15, Sim07, Smi07, SYC08, SLS^{+}12, SP10, SPS10, SvdLN17, Sól09, Som12, SSM21, Spe13, SDJ20, SLMV17, SM07, SLG05, SSV14, SSS^{+}20, TP11, TKM17, TSN21, TM15, TMW18, TV11, TDRD13, TB17, TR17, TPE19, TKM15, TMKD17, TMN16, Thi14, TFR16, TK17, TYH^{+}14, Tiv19, TH08, Tou15, TGJ17, TSV20, TF12]. Package [UNBK17, ÜKD09, ÜS10, USHH18, VW13, VG09, VPCS14, Ver18, VKVC15, Vie10, VV16, VKM18, VdL09, Vin17, VYD^{+}12, VS10, WGSL12, WMS17, WW11, Wan11, Wan13b, WLH^{+}18, WZL20, WCLS20, Was15, Wee10, WB07, WF12, Wel18, WDM^{+}11, WZQW20, Whe14, WM14, Wic07, WCHB11, WBEE19, Wie04, WMR16, XMW10, XWHL15, Xie13, XHW20, Yan07, YZZ^{+}20, YEL18, YFK^{+}15, Yee10, YSH15, You10, YMB20, Yua07, ZA17, ZMC21, ZG11, ZF15, ZPC^{+}16, ZQS16, Zha07, ZWH17, ZP13, ZPZW21, ZD17, dSJdSF14, dLM09a, dLM09c, dSdSCLC18, dWFP11, vE17, vdL21, vdWG11, Edl04, HS18, LM03, Sek11, SAR11, TBE^{+}21, ZLHK02, GR17b]. Packages [Alm10, BdMM08, DHF15, GASA15, HL09, KMTS14, MNT17, RGD12, SHFB17, TD07]. PAFit [PSS20]. Pages [Ver12]. Paired [HD12]. Pairwise [SBMG06]. Palaeoecology [Sim07]. Palettes [CBAA19, ZFH^{+}20]. Palm [TSN21]. Panel [ÁBZ17, BL14, BK11, CM08, Jac11, LP17, MP12, Mil17]. Panel-Corrected [BK11]. Papers [Tak16]. Parallel [CGRvD15, Den16, GTDT20, GB20, HM16, Hof11, Mat16c, dVSWAL17, SMDS11, SME^{+}09, Scu17, Zho15, dS11]. Parallelizing [PLZ^{+}15]. Parameter [CLL17, HWY18, MR18, MLMK12]. Parameterization [Mug10]. Parameters [GL08, RC17, Sar16]. Parameterwise [DSH16]. Parametric [CA17, CL13, Cul11, HH14, HIKS11, Jac16, KL14, KSP15, MS19, MRL12, NMB15, Per03, PKC^{+}20, PSS20, Ros15, SP14, TP11, XHW20]. Parametrical [RMG12]. Parental [TV11]. parfm [MRL12]. parseuas [RGK20]. Parsimonious [PMM18]. Partial [BV02, FW18, MW07, ML12, Wan13a]. Partially [Cha03, Har15, KNI16]. Particle [Nor17]. Particle-Based [Nor17]. Partition [MH15, Sho13]. Partitioned [GKZ12]. Partitioning [GGC^{+}15]. Partitions [Han06a, HW07]. Parts [ZC10]. Party [CI17, OO11]. pass [MT02]. Passing [Hof11]. Past [Tie05, ZK08]. Patch [PPT20]. Patch-Wise [PPT20]. Paths [FHT10, RT14, SFHT11]. Patient [CKE20]. Pattern [BMGT15, Mai07]. Pattern-Mixture [BMGT15]. Patterns [BT05, DM19, GRD13, GR16b, MR12, Mai08, MTPL15, PG15, PPC15]. PAVA [dLHM09]. Paveł [But15]. pbkrtest [HH14]. pcalg [KMC^{+}12]. pcIRT [Hoh18]. pcnetmeta [LZHC17]. PCovR [VKVC15]. pcse [BK11]. pdc [Bra15]. pdfCluster [AM14]. Peak [RG07]. Pearson [Pan09]. Penalized [CRW05, GRS12, Hua20, RMG12, McI16]. Penalties [LGF^{+}18, McI16]. Penalty [CS12]. Percentage [Pan09]. Percentile [EB03]. PerFit [TMN16]. Perform [Can17, GGC^{+}15]. Performance [Esm14, LCK11, MCM12, Mur15, Som12, XM14]. Performing [ABW13, EE07, FS20, GB20, LZHC17, Spe13]. Performs [RS07]. Period [SH07]. Periodograms [TFR16]. Perl [Bai04, Ros09]. PerMallows [ICL16]. Permutation [HHvdWZ08]. Permutations [Hüs06]. Person [TMN16]. Person-Fit [TMN16]. personalized [HY21]. Perspective [Kil16, Mai09b]. Perspectives [BMR14]. Peter [Iac15]. Pewsey [Han15]. Phase [HSL11, KWE^{+}17, SED14, SMS13]. Phases [PV20]. Phenomena [MHH17]. Phonetic [HI20]. Photovoltaic [Lam12]. PHP [MP06, dL97]. phtt [BL14]. Phylogeographic [MHE20]. Pilots [KCM08]. pinktoe [Nas05]. PIRLS [Pet97b]. Pitfalls [GF15]. Planning [KCM08, KWE^{+}17]. Plans [Cha03, Kie08, MMB15]. Platform [AB12, DH05, Jac16]. plink [Wee10]. plm [CM08]. Plot [EB03, MF14]. plotKML [HRBP15]. plotROC [Sac17]. Plots [BFC02, DP03, FW18, Fri06, Fri10, GC18, HO20, MHD17]. Plotting [GRD13, Sac17]. pls [MW07]. Plug [TFH12]. Plug-In [TFH12]. PLUS [dL05a, BLR97, But05, Can04b, Con03, HC05, Kuo03, SP05, ZG11, Wie04, SCD07, Fox05a]. POBE [MvB14]. Point [BT05, BTMB13, Dow16, EE07, FBC07, FBG15, GRD13, GR16b, Har10, JM15, KSS^{+}07, MTPL15, MKC21, PG15, Pen03, TSN21]. Point-and-Click [Dow16]. Point-Referenced [FBG15, FBC07]. Points [LPLPD14, Pan09]. Poisson [Cam12, CAA15, KN05, Pet97b, TKM17]. poLCA [LL11]. Policies [KvEPK20, Oom13]. Policy [HGG08]. ’Political [AFJZ11, How16b]. Poly [MAK06]. Poly- [MAK06]. Polychrome [CBAA19]. Polygenic [ZLC18]. Polygons [LRN18b]. Polyhedral [LPLPD14]. Polymorphisms [SBMG06]. Polynomials [Can17, Dem11]. Polytomous [ALV07, Hoh18, LL11, OC15]. pomp [KNI16]. Pool [dLHM09]. Pool-Adjacent-Violators [dLHM09]. popbio [SM07]. POPS [JFDB15]. Population [Fie12, Hil10a, JFDB15, LVK^{+}11, SL09, SR16]. Population-Based [SL09]. Populations [TV11]. Portable [Dry09]. Positions [CI17]. Possible [Law08]. Posterior [DN99, Smi07]. Potentially [RG07]. Power [BLR97, Dem11, HSH07, JMS^{+}09, KSPR16, KMG^{+}13, LT16, LR15b, MR05, MLAN02, Pet97b, LT16]. poweRlaw [Gil15]. POWERLIB [JMS^{+}09]. pp [But15, Gle16, Grö15a, Grö16, Gro08, Han15, Hel15, Hil15a, Hil15b, Hof15, How15a, How15b, How16b, Iac15, Kha15b, Kha15a, Kha16b, Mat15, Mat16a, Nun15, Pos15, Rus15, Zei15, Zei16]. PPtreeViz [Lee18]. Practical [Eks13, Hil07, Iac15, Iac17, Mat16b, Otn17, Web16, Wic10, Grö16, Grö19, Ros09]. Practice [Fox07, GR18b, Nas14]. Practitioners [CN18, Kha18c, GL14]. pre [Fok20]. Precision [BH08a, dVSWAL17]. Predict [MBGK18]. Predicting [Arc10, Yua07]. Prediction [Fok20, HZS21, JFDB15, KMM^{+}17, MIG12, SMM^{+}15, SH07, ZMC21]. Predictive [BG18, Kuh08]. Predictor [FW18]. Preference [CKE20, CKE20]. Preferences [HD12]. Preferential [PSS20]. prefmod [HD12]. Preliminaries [Pie05]. PReMiuM [LHA^{+}15]. Preparation [Kra07]. Preprocessing [HIKS11]. Presence [FM08]. PresenceAbsence [FM08]. Present [ZK08]. Presenting [TWK03]. PResiduals [LSL20]. Presmoothed [dUJ13]. Press [Grö15a, Han15, Hil15a, Hil15b, How15a, How15b, Zei15]. Pricing [TBS14]. Primal [LM14]. Primer [GvdMW09, Hel18, Soe10, Ros17, Grö13]. Principal [Lei10, MW07, PSY20, VKVC15]. Principles [Kuk09]. Priors [JCPL19, Lee13, LRN18a, ZF15]. Probabilistic [BL18, CGH^{+}17, GR16a, GVM16, JKL19, Meu13, MD18, SAR11, SR16, YMB20]. Probabilities [AMMRP14, BdUÁ18, Pat16, Som01, VV16, Eks10]. Probability [Kha03, KR10a, LQC^{+}12, LSL20, Mal09, vdWG11, San09]. Probability-Scale [LSL20]. Probit [IV05]. Problem [HH07a, Pap16]. Problems [EE07, VSV09, Web16, Wol15]. PROC R [Wei12]. Procedure [LP13]. Procedures [BGH^{+}17, Fot03, Kla18, Som07, Som12, TMW18].

Proceedings [LCK11]. Process [BTMB13, BP19b, BG21, But09, EN11, Gra07, GT10, JCPL19, MRC15, PLZ^{+}15, Pen03, SFS12, SKS15, TSN21]. Processes [CAA15, Gra16, Har10, KBMB19, KNI16, KLNN20, LHA^{+}15, Pac07, PLPL17, TDRD13, TDRD15]. Processing [Bai04, KCCG11, LBC^{+}16, Sól09, Fir03]. Product [Sho13, Han15, Hil15a, Hil15b, How15a]. productCd-111833258X.html [But15]. productCd-1118762614.html [Kha15a]. profdpm [Sho13]. Profile [LHA^{+}15, Wie04]. Program [Bel11b, BM97a, BMGT15, But05, Das06, EC06, HN13, HKB11, HZSW10, Jam99, Kin04, KKH10, KH13, Lan14, LC13, Lec14, MvB14, MHdL12, NE20, Pet97a, Pet97b, Sch06, Som98, XGY06, Hed99, Lum98]. Programming [CGH^{+}17, DN17, Hil10b, LH07, SDDD12, Wei12, Wic08a, WG00, dL09b, KN17, MT14, Lan17]. Programs [Góm15, JS05, McB06, NK06, SB01, SBL04, VGDM^{+}18]. Progression [BVE^{+}15]. Project [NRG^{+}17, BFH^{+}14]. Projected [BMR14]. Projection [Lee18, LZ15]. Projections [Coo97, HCW12, LM14, SAR11, SR16, WCHB11]. Projects [Bak20]. Propensity [HP09, Sek11]. Proportional [FH09, Gan15, TKM06, WZQW20, ZQS16, SFHT11]. Proportional-Odds [FH09]. proportions [MJ00]. Prototype [BdMM08]. Prototypes [Jos05]. Providing [Den16]. Proximities [IP08]. PSAgraphics [HP09]. Pseudo [HZBB20]. Pseudo-Ranks [HZBB20]. Pseudorandom [AS04, SM12, SLF14]. PST [GR16a]. Psychoco [WSZ12]. Psychological [NGVCFLS19]. Psychometric [ECW^{+}12, WSZ12]. Psychometrics [Gow16, dLM07]. psychomix [FSLZ12]. PTAk [Lei10]. PtProcess [Har10]. Public [HGG08, SSH16]. Publication [Hla16a]. Publication-Quality [Hla16a]. Pure [DC09]. Purely [DKT20]. Purpose [Kur14, RK14]. Purposes [MP06]. Pursuit [Lee18]. pvclass [ZD17]. pyJacqQ [JWHS16]. pylearn [LGF^{+}18]. pylearn-simulate [LGF^{+}18]. PyMC [PHF10]. pyParticleEst [Nor17]. PypeR [XMW10]. PySSM [SBMD14]. Python [Mat16b, BFM18, Har15, HHJCRB21, JWHS16, KJ14, MT98, Mül13, Nor17, Pat16, PHF10, SLGB14, SBMD14, XMW10].

Q [LLS15]. Q-Learning [LLS15]. qgraph [ECW^{+}12]. QLS [KS10]. QoLR [ASBMB17]. QTL [Zho10, HSG12, TV11, Zho10]. Qualitative [CBAA19, JvdBP07]. Quality [ASBMB17, Hla16a, MG09, PE19, dSdSCLC18]. qualV [JvdBP07]. Quantification [How15b]. Quantile [BV17, Ger14, Kle16, Koe08, SS19]. Quantile-Based [Kle16]. Quantiles [DGJ^{+}18, Kla18]. Quantitative [Hof15, WS11]. Quantities [BL18]. quantspec [Kle16]. Quasi [Can04a, Fir00, KS10, RS08]. Quasi-deviances [Can04a]. Quasi-Least [KS10, RS08]. Quasi-variances [Fir00]. Questionnaire [Fri12]. Questionnaires [Lal17]. queuecomputer [EWMR20]. Queues [EWMR20]. Quizzes [ZUL14]. QXLA [Kla18].

R [BGRR15, Bow10, DN17, Dow16, Dow19, Fay18, GL14, GR17b, GR18a, Gou05, Grö13, Grö15a, Grö17, Grö19, Gro18c, Hel17, Hel18, Hil15b, Hof15, Iac15, Kha17a, KP07, Lal17, Lan17, Lor18, Mat13, McN14, Mor18, MvS07a, Mül16, Otn17, SL09, San07, San09, San10a, Sar09, VM09a, dL05a, dLM07, ADH11, AEL10, ACW12, AR14, AMYR16, Aiz12, Alb16, AGG13, ATF10, AT13, Alm10, AD15, AACRC21, AB17, AWBM18, ASBMB17, ADN15, Arc10, AHvD09, ABC19, ABB^{+}19, AMW14, AB12, AM14, BM07, BG18, BL14, BT05, BDMP15, BK11, BR07, BS15, BFFN19, BP19a, Bar14, BFRP13, BS18, BPP17, BP12, BGH^{+}17, Bat15, BRC^{+}15, BLY18, Bea17, BBGL17, BCHY09, BDKL17, BKT14, BBG12, BB12, BMB16, BIG17, BSVT12, BK17, BdMM15, BP19b, BG21]. R [BMvS^{+}07, BLM^{+}15, BVE^{+}15, BN10, Bon18, BPGC14, BDdM11, BPB09, BD18, BYPJ19, BCAB07, Bra15, Bra14, Bra17, BS13, BWMBC21, BPDD08, BKvT^{+}14, BMGT15, BEHB17, Bür17, Bur12, But05, But08a, CF14, CdM10, CSNF18, Cam12, CKE20, Can17, CC08, CG15, CDP20, Car13, CB17, CP11, CFHBK11, CAA15, Cer17, Cha12, CQZ^{+}10, CGBN14, CB21, Cha10, CKSLS12, CP12, CC11, CA17, CKY14, CGC11, CBHK21, CMS^{+}11, CS12, CGC14, CLL17, CFSR15, CHML17, CNZ10, CO16, CM08, Cro20, CF08, CGS09, CJM06, Cul11, CI17, DMB18, DC09, Dah20, DLN17, DHM11, DB13, DBZ^{+}11, DC05, DR20, DMD15, Den16, DH05, DLC06, DW17, DCS20, DM18, DK18, DOV17, DV20, DM19, DPSH18, Duo07, DHF15, DGP08, EWMR20, EN11, EF11, ESO16, ESKHB20, EE07, EVBK19, EBO^{+}13]. R [EMU20, EL09, FS10a, FBdlF12, FHM08, Fel12, FT08, FDB12, Fer11, FDGD16, FHH17, Fie12, FBC07, Fir03, Fir05, FO15, FMM^{+}20, FC11, Fok20, FM18, Fox03, Fox05b, FH09, FC12, FL16, FS20, FM08, FBF14, FK17, FKP17, FGEMI12, FNB20, Fuj17, FS10b, GRMS11, GR16a, GRD13, GSD12, Gan15, Gas11, GGC^{+}15, GFPB19, GASA15, GRK^{+}16, GP12, GG16, GKZ16, GK16, GKD14, Gio09, GGK10, GLC^{+}15, GCA12, GFS14, GDMB08, GFC12, Gos11, Gra07, GT10, Gra16, GvdMW09, GVM16, GD21, GW18, Grö06, Grö10b, Grö14, Grö18b, GSW20, GvdL12, GZP14, GZ09, GH11, GKZ12, Gu14, Guh07, GB20, GV12, Had10, HPWdL15, HHB08a, HD10, HBF17, HPD19, HHY06, HH14, HD18, HSL11, Han05, Han06a, Han06b, HW07, Han07a, Han07b, HZBB20, Har10, HWM16, HD12]. R [HXY12, HM18, HP09, HH19, HH07b, HMM21, HHV20, Hla16b, HM11, HM16, HO20, HL09, Hof11, HGK^{+}20, HMS16, Hoh18, Høj04, HL07, Høj12, HMR^{+}13, HG14, HI20, HS18, HSG12, HGG08, HK08, HWY18, IP08, IV05, ILS11, ICL16, JvdBP07, Jac11, Jac16, JPM19, Jal19, JM15, JJJ14, JHQ^{+}11, JP16, JGM18, JLH^{+}20, JD15, JCPL19, Joh07, JS05, Jon07, KOS20, KHLF^{+}10, KMC^{+}12, KSHZ04, KMH06, KN05, KL14, Kas16, Kav15, KK15, KS14, Kie08, KWE^{+}17, KE14, KO06, KNI16, KSS^{+}07, Kle16, KM08, KN03, KM14, KG17, KR10a, KY10, Kol16, KK14, KSBZ16, KPSH15, KV13, KMTS14, KT16, Kra07, KPRP^{+}19, KSP15, Kuh08, KR13, KPH^{+}21, Kuo03, LPLPD14, LT16, LAF^{+}17, Lam12, LDHM21, LH12, LRGTA12, LNY^{+}20, LL10, LW16, LJH08, LIL^{+}15, Lee13]. R [LM13, LM18, Lee18, LRN18a, LMO^{+}19, Lei10, Lei13, Lei04, LHS08, LSPvdL17, Len09b, Len16, LR20, LQC^{+}12, LBW18, LM14, LM19, LFF17, LM03, LZHC17, LR15a, LLS15, LL11, LBC^{+}16, LC10, LSL20, LHA^{+}15, LR15b, LRN18b, LRRÁCSGS14, dUJ13, LH14, LX12, LCSC14, Lup09, MdlT20, MRC15, MR12, MF14, MB17, MRV21, MHJS16, MS19, MH07, MdL10, MMB15, MHE20, MHM20, Mar06a, MTPL15, MH15, MB15, MDvdV19, MGFA^{+}20, MBGK18, MQP11, MRHA20, MV17, MF15, MCAP19, MP14, MPM14, MYK07, MZ08, MS11, MMR20, MCM12, Mel16, MJGM10, Meu13, MW07, MH09, MHH17, MVS13, MNT17, MRT^{+}19, MP12, MP06, MW19, ML11, MTvdM15, MBR11, MV14, MHD17, MdUÁC10, MN14, Mug10, MvS07b, MAG^{+}11, Mul14, Mül13, MRL12, Mur03, Mur09, MG09, NK06, NR16, Nag18, Nag21, Nar05]. R [NV11, Nas14, NG07, ND12, NGBK12, NL12, NdSL16, NMM^{+}21, NRD16, NAA17, OH11, OHD17, Obe14, OCRC14, OL17, Oom13, OK14, OWA20, PLZ^{+}15, PPGD15, PP17, Pan18, Pap16, PLRC10, PSS^{+}17, Peb12, PG15, PSY20, Pen03, Pen08b, PJSPC17, PZK^{+}12, PP18, PE19, Pet10, PP11, PR07, Pfa08, PSS20, PV20, Phi10, PU13, PK08, Pir10, PC11, PMD19, PT07, PT09, PPT20, PLLC11, PLR^{+}16, PLPL17, PMM18, QY19, dREP12, RS07, Rec10b, RBB18, RC17, RG07, RJH14, RS05, RBHB15, Riz06, Riz10, Riz16, RD21, RST19, RFKM12, RR11, RMG12, Ros15, Ros12, RGD12, Röv20, dVSWAL17, Ruf09, RVCG19, SF18, SSH16, SFS12, Sar16, SD17, SH17, SH20, SP14, Sav16, SE18, SIR^{+}11, SZ11, Sch11, SME^{+}09, Sch16, SlRC16, Scr13, SR18]. R [Scu10, Scu17, Sek11, SVMMRP17, SAR11, SBMG06, Sho13, SS19, SHFB17, SKS15, SCS13, Smi07, SYC08, SLS^{+}12, SP10, SPS10, SvdLN17, Sól09, SSM21, Spe13, SDJ20, SR07b, SLMV17, SKZ05, SM07, SLG05, Su07, SGHY11, SSV14, SSS^{+}20, TP11, TW11, TKM17, TSN21, TM15, TKM06, TKM07, TMW18, TV11, TDRD13, TDRD15, TB17, TR17, TPE19, TKM15, TMKD17, TMN16, Ten18, TM05, TFH12, TSH20, Thi14, TH21, TFR16, TD07, TR14, TK17, TYH^{+}14, Tiv19, TH08, TBE^{+}21, Tou15, TGJ17, TSV20, TF12, Tus11, USA19, UNBK17, UAK^{+}15, ÜKD09, ÜS10, USHH18, VML12, VW13, VSV09, VG09, Var14, VFV13, VPCS14, Ver18, VKVC15, Ver12, Vie10, VV16, VKM18, VdL09, Vin17, VYD^{+}12, VS10, WGSL12, WDT^{+}12, WMS17, WLK08, WW11, Wan11, Wan13b, WLH^{+}18, WZL20]. R [WCLS20, Was15, Wee10, WB07, WF12, Wei12, Wei02, Wel18, WDM^{+}11, WZQW20, WS11, WST11, WM14, WSZ12, WCHB11, WBEE19, WMR16, WZ17, XMW10, XWHL15, Xie13, XHW20, YZZ^{+}20, YEL18, You10, YMB20, Yua07, ZA17, ZMC21, ZFZ10, ZLHK02, ZK08, ZKJ08, ZC10, ZUL14, ZKG20, ZF15, ZPC^{+}16, ZQS16, ZWH17, ZHZ20, ZP13, ZPZW21, ZD17, dSJdSF14, dLM09a, dLHM09, dLM09b, dLM09c, dWFP11, lGlRGL09, vBGO11, vE17, vdA07, vdA12, vdLdJ21, vdL21, vdWG11, But15, Dem17, Edd09a, Eva14, Gle16, Grö15b, Gro08, Han15, Hel15, Hil15a, How16b, Kha15a, Kha16b, Mat16a, Nun15, Pos15, VM09b, Wil14a, Zei15, Zei16, Arm19, Bar18, Bow09, Bul06, Chr09, CH11, Dem18b, Dow16, Dow17, Edd09b, Edd09c, Edd11, Edd12b, Edd12a, Eva11, Fri12, GR16b, Gou10]. R [Grö11, Grö18a, Hel16a, Hil06, Hil10b, How11, How16a, Kha18b, Leo10, Lor17, Mai06, Mal09, Mat15, Mun14, Oom10, Pav16, Pav20, Rui16, Rus15, San10b, San11, She11, Soe10, Str10, Wic08b, Wic08a, WG10, Yal10, Zho10, dL05b, dL06, dL09a, dL09c, dL09b]. R-based [BM07]. R-Friendly [Nag21]. R-INLA [BGRR15, LR15a, MMR20]. R-Menu [BP12]. R-php [MP06]. R-Squared [Rec10b]. R/MATLAB [VFV13]. R/qtl [Zho10]. r2 [Grö15a]. R2GUESS [LBC^{+}16]. R2MLwiN [ZPC^{+}16]. R2WinBUGS [SLG05]. Radiation [Lam12]. RAGE [PSW07]. Raju [Cer17]. ramps [SYC08]. Random [Bea17, Bre04, BILL07, CF14, CB21, CNLR19, Cro20, DB13, Dry09, FS10b, JJJ14, KCCG11, KJ14, Kur14, LCD18, MR12, MTW04, MIG12, MHH08, PB15, Pat16, PSW07, Röv20, Sar16, SMM^{+}15, SL18, TKM17, VDT14, WZ17, XTL13, YS12, ZWH17, dI17, MT00, PL05, SMDS11]. Random-Effects [Röv20]. RandomFields [SMM^{+}15]. Randomization [Gre17, USHH18, WWG09, ZK12]. Randomization-Based [ZK12]. Randomized [EVBK19, Gre17, HM18, MHT07, RSW15]. randomizeR [USHH18]. randomLCA [Bea17]. Randomly [MdUÁC10]. Randomness [MT02]. Range [Kla18]. ranger [WZ17]. Rank [Fri06, JLH^{+}20, New05, TM05]. Rank-Based [JLH^{+}20, TM05]. rankdist [QY19]. Ranking [QY19]. Rankings [HD12]. Ranks [HZBB20]. RAppArmor [Oom13]. Rare [KZ03]. RARtool [RSW15]. Rasch [ALV07, DBBD07, FSLZ12, Hoh18, MH07, OC15, VHM07]. Rate [SAR11]. Rates [MP14]. Ratings [HD12]. Ratio [MVS13, Wan13a]. Rational [BH07, BH08b]. Ratios [Mar06b, McB06]. Raton [Gle16, Grö15a, Grö16, Kha16b, Zei15]. RATS [Doa11]. Rcapture [BR07]. Rchoice [Sar16]. RcmdrPlugin.survival [FC12]. Rcpp [EF11]. RcppEigen [BE13]. Ready [MY13, VYD^{+}12]. Ready-to-Use [VYD^{+}12]. REALCOM [CGK11]. REALCOM-IMPUTE [CGK11]. Reasoning [YMB20]. Reassessment [SMS13]. REBayes [KG17]. Recapture [BR07, YSH15]. Receiver [LW03]. Recipe [SH17]. recmap [Pan18]. Recognition [Mai07]. Recognize [Coo97]. Recording [BK17]. Recovering [PLR^{+}16]. Rectangular [Pan18]. Recurrence [BFC02]. Recurrent [KMM^{+}17]. Recursive [Mey13]. Redistricting [AM11]. Reduced [FS12, Fri06]. Reduced-Rank [Fri06]. Reduction [AR14, CFT11, CHML17, MDvdV19, NMM^{+}21, Wei02]. Redundancy [LV16]. Redux [Koe08]. Reference [Kle16]. Referenced [FBG15, FBC07]. REGCMPNT [Bel11b]. Regimen [WMS17]. Region [Bra14]. Regional [LQC^{+}12]. Regionally [KPRP^{+}19]. Regions [BM12, Som98, Som01]. Registration [BLY18]. Regression [AS02, AB17, ABS01, Bar18, BFRP13, BRC^{+}15, Bel11b, BV17, BdMM15, Bon18, BKL05, BRG20, BR17, Car13, CB21, CGC11, CRW05, CNZ10, DLN17, DSD13, DLC06, Dha10, DM18, DPSH18, DS15, Edl04, FS12, FW18, Ger14, GKZ16, Gou10, Gra07, Grö06, GZP14, GKZ12, HM18, Hel16b, Hil09a, HHV20, HGK^{+}20, Jam99, JLZJ05, JFDB15, KB16, KN05, KL14, KBMB19, KZ03, Koe08, KV13, Lei04, Len09a, LM14, LBC^{+}16, LHA^{+}15, LGF^{+}18, LWOW18, Mai09a, Mai09b, MHM20, MW10, MV17, MH05, MW07, Mül16, Pet97a, PRS18, RT14, RBHB15, SS19, SSM21, TFR16, TR14, TGJ17, UAK^{+}15, VKVC15, WMS17, WLK08, Wei02, WMR16, XWHL15, ZMG07, ZLHK02, ZKJ08, ZQS16, dL09a, Hed99, Grö16, Der10, Hel16b]. Regression-Based [WMS17]. Regression/Item [CGC11]. Regular [ZG05]. Regularization [FHT10, RT14, Sch11, SFHT11]. Regularized [FHH17, SHBZ14]. Reject [Hüs06]. relaimpo [Grö06]. Related [ASBMB17, BL18, DR11, MAK06]. Relational [But08a]. Relationships [ECW^{+}12]. Relative [DHM11, DM18, Grö06, PP18, Pet97a, Wan13a]. Relevant [CGBN14]. relevent [MB15]. Reliable [AGM07]. relsurv [PP18]. rEMM [HD10]. Remote [Gos11, SR07a, ADN15]. Renewal [KBMB19]. Repeated [Fri10, KAK05, LW03]. Replacement [Han07a]. Replicated [Cha03]. Reporting [FR15]. reporttools [Ruf09]. Representation [HHB^{+}08b]. Representative [Kon13]. Reproducibility [BBGL17]. Reproducible [LT16, SDDD12, SvdLN17, Wil14b]. repsample [Kon13]. Resampling [Kil16, KAK05]. Research [HBF17, LT16, LNY^{+}20, Rus15, SDDD12, Str10, Hel15, Wil14b, Chr11, Hil09b]. Researchers [Kha18c]. reshape [Wic07]. Reshaping [Wic07]. Residual [LSL20]. Residuals [FW18, LSL20]. Resistant [Can04b, Con03, FS12]. Resolution [BM07]. Resonance [CMS^{+}11, KCCG11, TW11, WS11]. Resource [Din06]. Response [AW16, Arc10, Bon18, CGK11, Cha12, Cha16, CGC11, CO16, Cur10, DBZ^{+}11, DP12, FEvdL07, Had10, HM18, HZSW10, Joh07, Len09b, MR12, MPM14, MH05, MTvdM15, NGVCFLS19, Riz06, RSW15, Whe14, WB08, WEH11]. Response-Adaptive [RSW15]. Response-Surface [Len09b]. Responses [FEvdL07, GSD12, GHR18, KR13, LBC^{+}16, MTvdM15, Tou15, ZC10]. Restoring [BK17]. Restricted [Hüs06, KH13]. Results [BK17, FR15, TWK03]. Retention [KSS^{+}07]. Retrieve [SSS^{+}20]. Review [Ach09, Arm19, Ban16, Bar18, BMR14, Bow09, Bow10, Boy05, Bul06, But15, Chr09, Chr11, CH11, Cox07, Dem11, Dem17, DN17, Dem18a, Dem18b, Der10, Dow16, Dow17, Dow18, Dow19, Edd09b, Edd09a, Edd09c, Edd11, Edd12b, Edd12a, Eks10, Eks13, Ell09, Eva11, Eva14, Fay18, Fox05a, Fri12, Gen06, Gle16, GL14, GR16b, GR17a, GR17b, GR18a, Gou05, Gou10, GvdV16, Grö10a, Grö11, Grö13, Grö15b, Grö15a, Grö16, Grö17, Grö18a, Grö19, Grö21, Gro18c, Gro08, Gum07, Ham10, Han15, Har08, Hel15, Hel16a, Hel16b, Hel17, Hel18, Hel20, Hil06, Hil07, Hil09a, Hil09b, Hil10b, Hil10a, Hil15a, Hil15b, Hof15, How11, How15a, How15b, How16a, How16b, Iac15, Iac17, JS13, Kar06, Kha15b, Kha15a, Kha16a, Kha16b, Kha16c, Kha17a]. Review [Kha17b, Kha18a, Kha18b, Kha18c, Kha20, Kil16, KN17, Kuk09, Lal16, Lal17, Lan17, Len09a, Leo10, Lor17, Lor18, Mac12, Mai06, Mai07, Mai09a, Mai09b, Mai08, Mai09c, Mal09, Mat13, Mat15, Mat16b, Mat16a, McN14, MT14, Mor18, Mül16, Mun14, Mur05, Mur06, Nun15, Oom10, Otn17, Pav16, Pav20, Pie05, Pos15, Ros09, Ros08, Ros17, Rui16, Rus15, Sab07, SL09, San06, San07, San09, San10b, San10a, San11, SCD07, Sar09, Sch08, She11, Soe10, Spa06, Sto12, Str10, Unw09, VM09b, VM09a, VM10, Web16, Wic08b, Wic08a, Wic10, WG10, Wil14a, Wil09, Wil14b, Xu06, Yaf07, Yal10, Zei15, Zei16, Zho10, dL05b, dL05c, dL05a, dL06, dL07, dL08, dL09a, dL09c, dL09b, dL12]. Reweighted [Pet97b]. rFerns [Kur14]. rft1d [Pat16]. Rgbp [TKM17]. rgenoud [MS11]. RgoogleMaps [LR15b]. RGtk2 [LL10]. Rich [AB12]. Richness [SP05, Wan11]. Riemannian [MRHA20]. RinRuby [DC09]. Risk [DHM11, DM18, MS19, SZ11, Wan13a, WB08, WEH11]. Risks [Pet97a, Put11, dWFP11]. RKWard [RFKM12]. Rmalschains [BMB16]. RMatlab [VFV13]. RMatlab-app2web [VFV13]. rmcfs [DK18]. Rmixmod [LIL^{+}15]. RNA [GD21]. RNA-Seq [GD21]. RNetLogo [Thi14]. RNGs [Mar03]. Robert [Pos15]. Robin [Gle16]. RobPer [TFR16]. Robust [AMYR16, BGH^{+}17, CCF19, Can04a, CYK^{+}09, Jam99, KSS^{+}07, Kol16, KSBZ16, LH12, Mil17, Str04, TFR16, TF09, VW13, WM18, Wie04, ZFZ10, dL06]. robustlmm [Kol16]. robustloggamma [AMYR16]. ROC [Sac17]. Roger [HH14]. ROI [TSH20]. Roll [PLLC11]. Root [Lup09]. ROPTLIB [MRHA20]. Rotations [LSvdVK09]. Routes [vE17]. Routine [CKY14, GR18b]. Routines [LM14, Mil04, Mül13]. rpanel [BCAB07]. rpartOrdinal [Arc10]. rpartScore [GSD12]. RProtoBuf [ESO16]. Rqc [dSdSCLC18]. RRreg [HM18]. rscala [Dah20]. RSKC [KSBZ16]. rsm [Len09b]. RSNNS [BB12]. rSoNIA [BdMM08]. Rssa [GKSU15]. RStudio [Grö15a]. Rtadf [Cas17]. rTensor [LBW18]. Rubba [Iac15]. Rubens [How15b]. Ruby [DC09]. Rule [Fok20, RBHB15]. Rule-Based [RBHB15]. Rules [HGH05]. Run [ECK04, EC06, LC13, Lec14, ZPC^{+}16]. runjags [Den16]. runmixregls [Lec14]. runmlwin [LC13]. Running [FK17, SLG05]. Ruxton [Han15]. Rweb [Ban99].

S [BLR97, How16b, dL05a, BH07, BH08a, BH08b, But05, Can04b, Con03, HC05, Kuo03, SP05, WLK08, Wie04, ZG11, Ziv11, Fox05a, SCD07]. S-PLUS [dL05a, But05, Can04b, Con03, HC05, Kuo03, SP05, ZG11, Wie04, SCD07, Fox05a]. S3 [ZG05]. S4 [KSHZ04, KR10a, RK14]. SAEM [CLL17]. saemix [CLL17]. Safety [LNY^{+}20, WEH11]. Salesperson [HH07a]. Sampaio [How15b]. Sample [BH08a, BH08b, Hüs06, KAK05, LQC^{+}12, MVS13, MLAN02, PLRC10, TH08, WMR16, XGY06, BH07]. Sampled [MGG^{+}04]. Sampler [MHJS16, VHM07]. SamplerCompare [Tho11]. Samples [iM16, Lum04]. sampleSelection [TH08]. Sampling [AH08, Bar14, BGSC10, BFM18, BH08a, BH08b, CC11, GKZ16, Han07a, Kie08, Kon13, MRT^{+}19, PG15, VSV09, Ban16]. SamplingStrata [Bar14]. Sandboxing [Oom13]. Sandwich [Zei06]. sanon [KK15]. Sarkar [Gro08]. SAS [GL14, Kha15a, AS02, BMGT15, Dha10, FS13, Fri06, GHR18, Gos14, Gre17, JMS^{+}09, Jur15, KCM08, KM16, KS10, LVK^{+}11, LS02a, Law08, LSC09, LH07, LG11, LV16, MH05, NO06, OC15, Pan09, PKC^{+}20, Rec10a, Rec10b, SED14, Sel11, SH09, SR07a, SB01, TR14, VDT14, VGDM^{+}18, WWG09, Wei12, XM14, Yua11, ZCI^{+}16, ZK12, Wil09, Der10, Grö10a, Har08, JS13, Kha18a, Len09a, Oom10, Spa06]. SAS-Based [FS13]. SAS-IML [SB01]. SAS/IML [AS02, JMS^{+}09, KCM08, NO06, Pan09, ZK12]. SAS(R)

[MT14]. SASweave [LH07]. Satellite [MCAP19]. Saturation [PRS18]. SAVE [PPGD15]. SC’11 [LCK11]. Scala [Dah20]. Scalable [KEW13, NL12, ZWH17]. Scale [AGZ19, CB17, CFHBK11, DCS20, Gra16, HN13, Lec14, LSL20, PCL13, SR07b, TKM07, vdA07, vdA12]. Scales [HR14, MHM20, PLR^{+}16]. Scaling [GvdV16, KM08, MdL10, PLLC11, ÜKD09, dL05c, dLM09a, dLM09b]. Scatterplot3d [LM03]. Schemes [TDRD15]. School [HJS16]. Science [Dow17, DGP08, Hel20, HHJCRB21, Iac17, Mai09c, San09, Str10, How15b, Lor17]. Sciences [Lan14, Mun14, SVMMRP17, Web16]. Scientific [DN17, HRBP15, Hil10b, How15a, Kar06]. Scientist [Grö18a]. Scilab [Das06]. Score [Alb16, ABC19, HP09, Sek11, dSJdSF14]. scoringRules [JKL19]. Scrambled [Vig16]. Scraping [Iac15]. Screening [PE19, SF18]. Scripting [dL97]. Scythe [PQM11]. sdcMicro [TKM15]. SDD [BDMP15]. Seamless [EF11]. Search [AEL10, BMB16, Esm14, RPC15]. Searching [GFC12]. Seasonal [KMTS14, SE18]. SEATS [SE18]. Seattle [LCK11]. Second [ECK04, RBN19, VM09b, dL05c]. Security [Oom13]. Segmentation [BLY18, HKL09]. Segmented [Mug10]. Segregation [Tiv19]. Select [CC08, Gos14]. Selecting [CDP20, LRRÁCSGS14]. Selection [AS02, CdM10, Cal07, CNLR19, DK18, HPWdL15, HGK^{+}20, HMS16, KR10b, LAF^{+}17, LBC^{+}16, MLMK12, ND12, Ntz02, Sch11, SR18, SH09, TMW18, TH08, WF12, WZQW20, WMR16, Wol15, ZA17]. Selective [SH14]. Self [WB07, WK18]. Self- [WB07]. Self-Organizing [WK18]. Semi [Cul11, Hua20, JHQ^{+}11, KL14, KSP15, LIL^{+}15, Nas05, NMB15, OH11, SP14, VGDM^{+}18, XHW20]. Semi- [JHQ^{+}11, SP14]. Semi-automatic [Nas05]. Semi-Confirmatory [Hua20]. Semi-Continuous [VGDM^{+}18]. Semi-Markov [KSP15]. Semi-Parametric [Cul11, KL14, NMB15, XHW20]. Semi-Supervised [Cul11, LIL^{+}15]. semiliterate [Lum98]. SemiMarkov [KSP15]. Semiparametric [BP19a, Bur12, GRS12, Gra07, LWOW18, MW10, MGH18]. SemiParSampleSel [WMR16]. semPLS [ML12]. Sense [Wic10]. Sensing [Gos11]. Sensitivity [FS20, GT10, HPK18, SP10]. Separation [FS12, MNT17, NMM^{+}21]. Separation-Resistant [FS12]. Seq [GD21]. seqHMM [HH19]. SeqNet [GD21]. Sequence [HNCI18, HH19, MB15, VV16]. Sequences [GRMS11, GR16a, Mel16, OH11]. Sequencing [dSdSCLC18]. Sequential [AH08, BG21, CGRvD15, CC11, GG16, Joh09, JMD08, Ros15, Was15, Zho15]. Sequentially [MMM12]. Serial [BDMP15]. Serialization [ESO16]. seriation [HHB08a]. Series [BFC02, Bra15, CN13, DS15, Edd09a, HK08, Kas16, Kha17a, KT10, LDHM21, LFF17, Lor18, MGG^{+}04, MCAP19, MYK07, MV14, NMM^{+}21, Pen08a, SSH16, San10b, San11, SCD07, VdL09, Yaf07, ZG05]. Server [dL97]. Server-side [dL97]. Service [SGK03]. Set [CGBN14, Góm15, HW07, Hüs06, LPLPD14, SBL04, dLHM09]. Sets [BL18, BDdM11, HGH05, MH09]. SFAMB [HB17]. sgmcmc [BFFN19].

Shaken [GKZ12]. Shape [LPLPD14, SR07b]. Shaped [PG15]. Shapes [Coo97]. Shared [BP19a, MGH18]. Sharpening [Wol12]. Shelf [DV20]. Short [ECK04, EC06]. Short-Run [EC06]. Shrinkage [DSH16]. Shrinking [CQZ^{+}10]. side [dL97]. Sided [BH08a, BH08b, SL11, BH07, MJ00]. Sightability [Fie12]. SightabilityModel [Fie12]. Significance [CFHBK11]. Significant [GFC12]. Signs [NO06]. Sim.DiffProc [GB20]. simcausal [SvdLN17]. simecol [PR07]. SIMEX [HXY12]. simFrame [ATF10]. Similarity [SWAF15]. SimInf [WBEE19]. simmer [USA19]. Simon [Iac15]. simPH [Gan15]. Simple [BV02, BCAB07, GC18, MN14, ZFZ10, dLM09c]. SimpleITK [BLY18]. Simplex [ZQS16]. simplexreg [ZQS16]. simPop [TMKD17]. Simreg [JLZJ05]. simsurv [BWMBC21]. Simulate [AH08, BMvS^{+}07, DC05, HHB^{+}08c, LGF^{+}18]. Simulated [LGF^{+}18]. Simulating [AS04, BS18, BVE^{+}15, BWMBC21, GRD13, GD21, Han07b, MCM12]. Simulation [ATF10, ABS01, AHvD09, BGH^{+}17, BM97a, BILL10, BFH^{+}14, Coe00, Dem11, DN17, DCS20, EWMR20, HHB^{+}08b, Hil10b, IMR17, KSPR16, Kuk09, LT16, MGFA^{+}20, MGH18, MN14, PK08, SMM^{+}15, SvdLN17, TMKD17, USA19]. Simulation-Based [KSPR16]. Simulations [BILL07, CGC11, HM16, WBEE19]. Simulator [MRC15, MLAN02, BB12]. SimultAn [ZG11]. Simultaneous [HH07b, JLZJ05, KPSH15, KR13, SH14, SB01, ZG11, MJ00]. Single [KWE^{+}17, SBMG06, Su07, YZZ^{+}20]. Single-Agent [YZZ^{+}20]. Single-Arm [KWE^{+}17]. Singular [GKSU15, LZ17, LS02b]. SIS [SF18]. Size [FS13, LQC^{+}12, MLAN02, Wol15]. Sizes [Kel07, MBGK18]. Skew [LS02a, LM13, LM18, PCL13]. Skew-Normal [PCL13]. Skew-symmetry [LS02a]. slfm [DM19]. sm [Scr01]. SMACOF [dLM09b]. Smale [GP12]. Small [HMM21, MJ00]. SMCTC [Joh09]. Smirnov [DKT20, SL11]. Smooth [Pat16]. SmoothHazard [TGJ17]. Smoothing [Bar18, Cam12, Can04b, Con03, CHML17, Gu14, LG11, MPM14, NW04, Pac07, PT07, PT09, PPT20, Scr01]. Smoothing-Spline-Based [LG11]. sms [Kav15]. sna [But08b]. Snippet [Aiz12, AD15, AW16, BDMP15, BK11, Bak20, BS18, Bel11a, CKE20, CG15, Car13, Car15, Cas17, CBAA19, CO16, Cur10, DMB18, DP12, Dha10, Dry09, EC16, Esm14, FS12, FKP17, GF15, Gos14, GC18, Gre17, GVM16, HF18, Han06a, HW07, Han07a, HZBB20, Har15, HXY12, Hla16a, Hof11, Hoh18, Hug07, Jal19, Jur15, KM16, Kam08, Kla18, Kon13, LSC09, LW16, Lec14, LR20, LQC^{+}12, LZ17, LG11, LRN18b, LV16, MF14, MB17, MHJS16, MTPL15, MBGK18, MF15, MP14, MHT07, NK06, Nag21, NAA17, OC15, Pan09, Pan18, Pap16, Pav15, PSY20, Pen08a, PZK^{+}12, PKC^{+}20, PV20, Phi10, Rec10a, Rec10b, RPC15, RBB18, RGK20, Ruf09, Sac17, SWAF15, SBMG06, SGS^{+}14, TKM07, TR14, VSV09, VDT14, WWG09, WZ16, WM18, Wei12, WEH11, Wol15, XMW10, YS12, dSJdSF14]. Snippet [dSdSCLC18]. SNP [BGM06]. SNP NLMM [VDT14]. SNPMClust [EC16]. sns [MHJS16]. SNSequate [Gon14]. Social [AT13, But08b, DCS20, Iac17, MRV21, Mai09c, Ost20, Str10, Web16]. SOCR [CCD09, Din06]. Software [AR14, AJ11, BFC02, BV02, BGM06, But05, CGK11, CKO11, DN99, FF14, FR07, Fri06, Gen06, GL14, HHB^{+}08b, HK20, IKP09, JLZJ05, JFDB15, JMD08, JMS^{+}09, KCM08, Kar06, Kha03, KPSH15, Kra07, KMM^{+}17, LC13, Lec14, LSM^{+}19, Mac12, Mat16c, Men11, MR05, NRG^{+}17, NV11, NW04, Ngu07, NGBK12, OO11, PBR15, Ros08, RSW15, Sek11, Sto12, SKZ05, Str04, TWK03, WKL11, WB08, Wie04, Yaf07, Yua11, Yuc11, ZMG07, dI17, Edl04, Boy05, Ell09, FL16]. Solar [Lam12, Lam12]. Solutions [Kha16c]. Solver [Tou15]. Solving [SPS10, VG09, Web16]. Some [BGRR15, JLZJ05, MT02, Mon06, Tak16, Tie05, Tiv19]. Somerville [Som07]. Sonja [Kha15a]. Sons [But15, Hel15, Iac15, Kha15a]. Sorted [Har15]. Source [DKMT11, MNT17, NMM^{+}21, PQM11]. Souza [How15b]. spa [Cul11]. Space [CDP20, CKO11, DLC06, Doa11, DG11, FF14, Góm15, Jal19, JWHS16, Luc11, Men11, Mur15, Pel11, PA11, PP11, PLR^{+}16, Sel11, SBMD14, ÜS10, Van11, VP18, Ziv11]. Space-Time [CDP20, FF14, JWHS16]. Spaces [ZPZW21]. spacetime [Peb12]. spam [FS10b]. sparr [DHM11]. Sparse [AMW14, Bra14, Bra17, FS10b, HD16, KN03, KSBZ16, LGF^{+}18, SCL^{+}18]. sparsebn [AGZ19]. sparseHessianFD [Bra17]. SparseM [KN03]. SpaSM [SCL^{+}18]. spate [SKS15]. Spatial [Alm10, BT05, Ban16, Bar96, BP15, BGRR15, BDdM11, BGSC10, DHM11, Eva14, FBC07, GGC^{+}15, GLC^{+}15, Gra16, LRGTA12, Lee13, LR15a, LRN18b, MP12, NO06, PPGRMB21, PBR15, PG15, PV14, Pir10, SYC08, TDRD13, TDRD15, TR17, VPCS14, ZMC21, ZHZ20, Gle16, GR16b]. Spatial-Temporal [LRN18b]. Spatially [RC17]. Spatially-Varying [RC17]. Spatio [BS15, FBG15, GRD13, HRBP15, LRN18a, Lei10, MHH17, Peb12, SKS15, TDRD13, ZMC21]. Spatio-Temporal [BS15, FBG15, GRD13, HRBP15, LRN18a, Lei10, MHH17, Peb12, SKS15, TDRD13, ZMC21]. Spatiotemporal [TDRD15]. spatstat [BT05]. spatsurv [TR17]. spBayes [FBC07, FBG15]. spBayesSurv [ZHZ20]. Special [KP07, MvS07a, Put11, TW11, dLM07, AFJZ11]. Species [GVM16, MBGK18, SP05, Wan11, Wan11]. Specific [MS19, PRS18, She10]. Specification [MHH08]. Spectra [Kra07]. Spectral [BMR14, JCPL19, Kle16, MGG^{+}04]. spectralGP [Pac07]. spectrino [Kra07]. Spectrometry [CS12]. Spectroscopic [MvS07b]. Spectroscopy [MvS07a]. Spectrum [BM07, GKSU15]. speedR [VYD^{+}12]. Spelling [HI20]. Spherical [FDGD16, HFKB12]. sphet [Pir10]. Spider [EL09]. Spider-Man [EL09]. spikeSlabGAM [Sch11]. Spline [CRW05, FS20, Gu14, LG11, WMR16]. Splines [Cam12, WLK08]. Split [Wic11]. Split-Apply-Combine [Wic11]. splittable [SLF14]. splm [MP12]. SPmlficmcm [NMB15]. SPODT [GGC^{+}15]. Spread [GW18, WBEE19]. Spreading [MGFA^{+}20]. Spreadsheets [ALO^{+}10, MY13]. Springer [Gro08, Hof15, How16b, Kha15b, Mat15, Mat16a, Nun15, Pos15, Rus15, Zei16]. Springer-Verlag [Gro08, Hof15, How16b, Kha15b, Mat15, Mat16a, Nun15, Pos15, Rus15, Zei16]. SPSS [GL14, BP12, San06]. spTest [Wel18]. spTimer [BS15]. Square [LS02a]. Squared [Rec10b]. SQUAREM [DV20]. Squares [Das06, FHH17, KS10, Len16, MW07, ML12, RS08, Pet97b]. SsfPack [Pel11, Ziv11]. SSMMATLAB [Góm15]. SSN [VPCS14]. SSpace [VP18]. Stability [TMW18]. Stable [dVSWAL17]. Stacked [HR14]. Stage [CKE20, ECK04, EC06, JPM19, MMB15]. STAMP [Men11]. stampr [LRN18b]. Stan [Bür17, CGH^{+}17]. Standard [BK11, Gon14, MW10, Mil17, SM12, WM18]. Standardized [Kel07]. Standards [Ngu07, WST11]. STARS [PV14]. Start [SH17]. STAT [dL05d, BM96, Fir00, Jos05, Lum96, Lum98, Nar05, Scr01, Tie05, VMU05]. Stata [KN17, Yaf07, CL13, DG11, FHH17, HX13, HHJCRB21, Kon13, KSPR16, LC13, Lec14, RGK20, RW11, VGDM^{+}18, Lal17]. StatCrunch [WWH04]. State [ASB11, AMMRP14, CP11, CKO11, DLC06, Doa11, DG11, GRMS11, GR16a, Góm15, Jac11, KSP15, Luc11, Men11, Mur15, Pel11, PA11, PP11, PPC15, Put11, SME^{+}09, Sel11, SBMD14, Van11, VP18, Yuc11, Ziv11, dWFP11]. State-Space [Mur15]. Stationary [CN13, EN11, TPE19, VV16]. STATISTICA [FS12]. Statistical [Ach09, ATF10, AJ11, AGM07, Bai04, Ban99, But09, CKO11, Dem18b, FBdlF12, FDGD16, FL16, Gen06, GL14, Góm15, Gon14, Grö15a, Har08, Hil06, Hil15a, HKL09, JvdBP07, KNI16, KMM^{+}17, Lei13, LSM^{+}19, Mai09a, Mat16c, Mil04, MBK18, MG09, NO11, OHD17, PPGD15, Pan18, PQM11, Pen08b, Per03, PV14, RFKM12, Ros17, SL09, SF18, SCL^{+}18, Str04, TP11, TKM15, TWK03, VPCS14, VS14, Wei05, WO97, Wic08a, Wil09, Xie13, Zha14, dL06, Chr11, Dem11, Hel17, Kha15b, Kha15a, Kha18c, Lal16, Lal17, Mai09b, Nun15, Spa06, dL09c]. Statistically [LAF^{+}17]. Statisticians [dL08]. Statistics [Bai04, Ber09, CCD09, Din06, Dry09, Edd09c, Fox05b, Han15, Hil15b, JS13, JWHS16, Kha16b, Kha16c, Kil16, KGP^{+}19, MB15, New05, PBR15, Ruf09, Sab07, SLGB14, VM09b, Web16, WF12, WG10, YFK^{+}15, Ylv16, Zei16, dL97, Hel15, Gou05, Hel15, Hil07, Kha16a, Mur05, Sar09, Web16, dL07]. StatLSSVM [DSD13]. statnet [GHH^{+}08, HHB^{+}08b, HGH13]. STEM [FF14]. Step [MT14, Som12]. Step-Down [Som12]. Step-Up [Som12]. stgenreg [CL13]. Stirred [GKZ12]. stm [RST19]. StMoMo [VKM18]. Stochastic [BFFN19, BFH^{+}14, CC08, CJM06, GB20, HB17, How15b, HPK18, Kas16, LP13, MHJS16, MGFA^{+}20, PHF10, PK08, SKS15, VKM18, WBEE19]. stochvol [Kas16]. Storage [LCK11]. stpp [GRD13]. Strategies [Hel16b, KEW13]. Strategy [Wic11]. Stratified [Bar14, KK15]. Stratonovich [GB20]. Stream [HD10, HBF17, PV14, VPCS14, HBF17]. StreamingBandit [KvEPK20]. Strings [RGK20]. strucchange [ZLHK02]. Strucplot [MZH06]. structSSI [SH14]. Structural [BKL05, DOV17, Hua20, MR18, ML12, Obe14, PT09, PRS18, RST19, Ros12, UNBK17, WH08, ZLHK02]. Structure [Gos14, JFDB15, Mai08, MW19, Scu17]. Structured [LGF^{+}18, SH14, UAK^{+}15, Jon07]. Student [Mat13, AHvD09]. Student- [AHvD09]. Studentized [Kla18]. Students [GL14]. Studies [ADH11, BPB09, CBDM07, GR18b, HSL11, JWHS16, KWE^{+}17, LT16, MF15, MLAN02, NAA17, PC11, SED14, SvdLN17, Wan13a, WZL20, ZFZ10, SL09, Unw09]. Study [ABS01, HM16, LVK^{+}11, MCM12, Som12, Coe00]. Studying [PG15]. Stuttgart [BB12]. Style [dL12]. Subgroup [HY21]. Subject [MvB14]. Subset [HGK^{+}20, MZ08, SFC20, Tri03, WZQW20, Day01]. Subsets [Day01, LAF^{+}17, Wol15]. Subspace [ZWH17]. Sufficient [AR14, CFT11]. Suffix [GR16a]. Suite [Alm10, LH14]. Sums [FDGD16, MKC21]. Sunflower [DP03]. Super [WB07]. Super-organizing [WB07]. Supervised [BD18, Cul11, LIL^{+}15, PE19]. Support [HP09, KMH06]. Supporting [Aiz12]. Sure [SF18]. Surface [Len09b]. Surfaces [FS20]. Surveillance [SSH16, MHH17]. Survey [CG15, GR17b, Kha18a, Kuo03, Lum04, Obe14]. Survey-Weighted [CG15]. Surveys [AT13, KAK05, Ban16]. survHE [Bai20]. Survival [Bai20, BWMBC21, CSNF18, CKY14, CYK^{+}09, CL13, DN99, FC12, Hel16b, Jac16, dUJ13, Lum98, MS19, MNA^{+}97, MMRP11, MIG12, MN14, OL17, PP18, RMG12, TM15, TR17, TR14, WZ16, XHW20, ZCI^{+}16, ZHZ20, Hil09b]. survivalBIV [MMM12]. survPresmooth [dUJ13]. survsim [MN14]. Sustainable [HWPG13]. SVAR [Pfa08]. svars [LDHM21]. SVEC [Pfa08]. svt [LZ17]. Swapping [SGK03]. Switching [ABB^{+}19, Pap16]. Switzerland [Rus15, Zei16]. Symbol [TKM06]. Symbolically [Phi10]. Symmetries [HL07]. symmetry [LS02a]. Synergism [BN10]. Synergism/Antagonism [BN10]. Synth [ADH11]. Synthetic [ADH11, NRD16, RD21, TMKD17]. synthpop [NRD16]. System [DKMT11, HKL09, Jos05, KOS20, NV11, VG09, vdHvB16]. systemfit [HH07b]. Systems [GB20, HH07b, How15a, Lam12, LCL^{+}08, OO11, RBHB15].

t [AWBM18, Eks10, Hug07, LM18]. t-Distributions [LM18]. t-to-z [Hug07]. Table [WGSL12]. TableMaker [Hla16a]. Tables [Hla16a, ILS11, LS02a, Lei13, MZH06, OK14, Per03, Ruf09, SSV14, WH08]. Taeger [Kha15a]. Tailed [Gil15]. Tandem [MDvdV19]. Targeted [GvdL12, LSPvdL17]. tcltk [BCAB07]. tclust [FGEMI12]. Teaching [BGSC10, DKMT11, Dry09, MP06]. Technical [Mor18]. Techniques [CA17, FR15, Mai09a, Mat16b, MP14, WMS17]. Technologies [Mun14]. teigen [AWBM18]. Teleconnections [ADN15]. Temperature [Mug10]. Tempering [GT10]. Template [HGH13]. Templates [Den16]. Temporal [BS15, BDdM11, FBG15, GRD13, HRBP15, LRN18a, Lei10, LCD18, LRN18b, MHH17, Peb12, PSS20, SKS15, TDRD13, ZMC21]. Tensor [Lei10, LBW18, PT09]. Term [PC11]. Terminal [KMM^{+}17]. Terms [MHH08]. termstrc [FH10]. Ternary [HF18]. Terry [Fir05, TF12]. Test [ABW13, Bat15, Cha16, Gon14, GR18b, Mar05, MAK06, NGVCFLS19, NAA17, Som12, Ver18, XGY06]. Testing [Cas17, CFHBK11, GFC12, JJJ14, Kha15a, LR20, Lup09, MR12, MB17, MAK06, Pav15, SB01, ZA17, ZLHK02, Gro18c]. Tests [BV02, BKvT^{+}14, Caf06, Cer17, FS10a, FO15, HH14, HHvdWZ08, JLH^{+}20, KBC17, LT16, LRRÁCSGS14, LX12, MT02, MVS13, PJSPC17, Pet97a, PKC^{+}20, SBL04, Wee10, Wel18, WH08, ZFZ10, dSJdSF14]. texreg [Lei13]. Text [FHM08, HMR^{+}13, Iac15, Ros09, TFH12, Fay18]. textcat [HMR^{+}13]. tgp [Gra07, GT10]. TGUI [DKMT11]. Their [BTMB13, Coo97]. Them [HZBB20]. Thematic [Ten18]. Theory [Cha12, Cha16, CGC11, Cur10, Der10, Han07b, Kel07, MPM14, MTvdM15, SBL04, ÜS10, Whe14, dL05c]. Things [HHB08a]. Third [Wil09]. Thomas [Hil15a]. Threading [Nag21]. Three [GKD14, LX12, MD18, NG07]. Three-dimensional [NG07]. Three-Mode [MD18]. Three-Way [GKD14, MD18]. ThreeWay [GKD14]. threg [XWHL15]. Threshold [HS18, MMB15, PJSPC17, XWHL15]. Thresholding [JS05, KO06, LZ17]. ThresholdROC [PJSPC17]. Throughput [PSS^{+}17, dSdSCLC18]. Tidy [Fay18, Wic14]. Tiles [LR15b]. Time [AVD21, BL14, BFC02, Bra15, CDP20, CN13, CKY14, CBHK21, CS12, DOV17, DS15, Edd09a, FF14, GHR18, Gio09, Har10, HW20, HXY12, Hüs06, HK08, JvdBP07, Jal19, JWHS16, Kas16, Kha17a, KSS^{+}07, KT10, LDHM21, LFF17, Lor18, MGG^{+}04, MCAP19, MYK07, MV14, NMM^{+}21, Pen08a, RG07, Riz10, Riz16, RSW15, SSH16, San10b, San11, SCD07, Sav16, TR14, VR14, VdL09, Wan13b, Yaf07, ZG05]. Time-Indexed [Sav16]. Time-Intensity [RG07]. Time-to-Event [CBHK21, GHR18, Riz10, Riz16, RSW15]. Time-Varying [HW20, TR14]. Time-Weighted [MCAP19]. timereg [SZ11]. Times [FEvdL07, GW11, MTvdM15]. TIMP [LMB^{+}07, MvS07b, SLS^{+}12]. Tissue [FT11]. tm [TFH12]. tmap [Ten18]. TMB [KNB^{+}16]. tmle [GvdL12]. Tolerance [You10, You10]. Tool [BM07, BV02, LT16, MR05, MAK06, NGVCFLS19, Sac17, SDJ20]. Toolbox [ÁBZ17, ÁBZ20, BRC^{+}15, Ber09, CGRvD15, CSY15, GHH20, LSvdV19, McI16, NGVCFLS19, PA11, RS08, SLGB14, SCL^{+}18, VP18, YR19, ZFH^{+}20]. Toolkit [CCD09, KT10, LL10]. Tools [AGM07, Bec18, BGSC10, BW01, Fuj17, HHB^{+}08b, Lum96, MTPL15, NOT08, PJSPC17, RG07, Iac17]. Toolset [PV14]. Topic [GH11, RST19]. topicmodels [GH11]. Total [SAR11]. Tour [HCW12, MLVMY05]. tourr [WCHB11]. tourrGui [HCW12]. Toxicokinetics [PSS^{+}17]. TP.idm [BdUÁ18]. TPmsm [AMMRP14]. trackeR [FK17]. Tracking [FK17]. Tractography [CMS^{+}11]. TractoR [CMS^{+}11]. Trait [BGM06]. Traits [She10]. TraMineR [GRMS11]. Transform [RS07]. Transformation [HSH07, XTL13]. Transformation-Based [XTL13]. Transformations [Hot20, Dem11]. Transforms [Hug07]. Transition [ASB11, AMMRP14, BdUÁ18]. Translation [RS07]. Translation-Invariant [RS07]. Transparent [SvdLN17]. Traveling [HH07a]. Traversal [Nas05]. Treatment [LQC^{+}12, TM15, VS14, WMS17, WLH^{+}18]. Tree [Arc10, BR17, DP12, HNCI18, Sch06]. Tree-Based [BR17, DP12]. Treed [Gra07, GT10]. Trees [CSNF18, GR16a, GSD12, GZP14, KB16, Lee18, Nas05, SSM21]. Trend [MAK06]. Trends [BL14]. Trial [CKE20, She11]. Trials [JPM19, KPH^{+}21, LQC^{+}12, RSW15, SMS13, USHH18, WWG09, WCLS20, YZZ^{+}20, Ham10]. Trigonometric [FDGD16]. Trimmed [BM12, Bel11a]. Trimming [FGEMI12]. Truncated [KL14, MdUÁC10, NK06]. truncSP [KL14]. Trust [Bra14]. trustOptim [Bra14]. TSclust [MV14]. tscount [LFF17]. TSP [HH07a]. Turning [Mat16c]. Tutorial [GHH^{+}08, KMM^{+}17, TR14]. Two [BH08b, CKE20, DHF15, EC06, Esm14, GFPB19, HSL11, MMB15, Meu13, MVS13, MBR11, MHT07, NG07, Rec10b, RGD12, SED14, SL11, WZ16, WH08, XGY06, MJ00]. Two- [NG07]. Two-Dimensional [GFPB19]. Two-Level [Esm14, Rec10b]. Two-Mode [Meu13]. Two-Phase [HSL11, SED14]. Two-Sample [MVS13, XGY06]. Two-Sided [BH08b, SL11, MJ00]. Two-Stage [CKE20, EC06, MMB15]. Two-Way [Meu13, WH08, MHT07]. Two-Zone [MBR11]. TWPack [Kro16]. Type [FO15, FM18]. Types [CGK11, RG07].

UCLA [Ylv16]. ukcatalogue.oup.com [Han15, Hil15a, Hil15b]. Ultrahigh [SF18]. Ultrahigh-Dimensional [SF18]. Unbiased [PG15]. Uncertain [BSVT12]. Uncertainty [How15b]. Unconstrained [LC10]. Uncovering [Mai08]. Under- [SF16]. Under-Dispersed [SF19]. Underlying [DKT20]. Unfolding [LBW18]. Unidimensional [She08b]. Unified [Han07b, Lal17, SYC08]. Unifying [Mil17, NV11]. UniLogistic [Dha10]. Unimodal [Wol12]. Unit [LRN18a, LBC^{+}16, Lup09]. Unit-Based [LBC^{+}16]. Units [Ban16]. Univariable [Dha10]. Univariate [DR11, FBC07, FBG15, Kha16a, MMM12, San06]. University [Han15, Hil15a, Hil15b, How15a]. Unknown [VV16]. Unleashing [LR15b]. unmarked [FC11]. Unobserved [Bos11]. Unreplicated [Law08]. Unstratified [Gre17]. Unsupervised [BER^{+}16, LIL^{+}15]. untb [Han07b]. Unwin [Zei15]. Upcrossing [Pat16]. Updates [MB17]. Upper [Kla18]. Urn [Han07a]. Usage [Mai08, dL12]. UScensus2000 [Alm10]. USD [But15, Gle16, Grö15a, Grö16, Gro08, Hel15, Hil15a, Hil15b, How15a, How15b, How16b, Kha15b, Kha15a, Kha16b, Mat15, Mat16a, Rus15, Zei15, Zei16]. Use [VYD^{+}12]. Used [PV14]. Useful [RG07]. User [AB12, CCH15, Dun99, Fox05b, HL09, LL10, LSvdV19, Mat16a, RFKM12, RGK20, SLS^{+}12, SGS^{+}14, VML12, WLH^{+}18]. User-Friendly [SGS^{+}14]. Users [Ber01, NV11, WDT^{+}12]. Using [AS02, Arm19, BM07, Bai04, BS15, Bak20, Bar18, BFRP13, BE13, BMBW15, BB12, BER^{+}16, BLM^{+}15, BL18, BPGC14, Bos11, BCAB07, BWMBC21, Bür17, But15, Can17, CG15, Car13, CN13, CBDM07, CP11, Cha03, CP12, CCH15, CGC11, CLL17, CRW05, CG10, Dah20, Das06, DHM11, DN17, DM18, DPSH18, EVBK19, Eva11, FDGD16, Fie12, FR07, GF15, GKD14, GLC^{+}15, GR17b, Grö10b, Grö15b, HF18, HD18, HSH07, Hel15, Hil06, Hil09a, Hil10b, Hil15a, HWPG13, HB17, How16b, HCW12, HY21, Jam99, JS13, JFDB15, JCPL19, KMC^{+}12, Kao09, Kas16, Kha17a, KS10, KCCG11, KY10, Kuh08, KGP^{+}19, LPLPD14, LW16, LRN18a, Lei10, LH07, Len09b, LSL20, LHA^{+}15, MMB15, MB15, MS11, MMR20, MHH17, MP06, MIG12, MTvdM15, ML12, MAK06, Mur15, NMM^{+}21, Ntz02, Oom13]. Using [Otn17, Pac07, PB15, Per03, Pet97a, PV20, PLPL17, QY19, Rec10a, Rec10b, RS05, Riz16, RMG12, Röv20, San10a, San11, SIR^{+}11, SZ11, Sel11, She11, SH09, Sim07, SP10, Spa06, Str10, SM07, TSN21, TV11, TKM15, TM05, TR14, TBE^{+}21, VR14, WWG09, Wee10, Whe14, Wie04, Wil14a, Wil09, XMW10, Yal10, YZZ^{+}20, YSH15, Yua11, ZQS16, ZHZ20, Ziv11, dLM09b, dLM09c, Day01, Fir03, Jos05, LS02a, SLMV17, The02, Dow16, Gou05, Grö15a]. Usual [PZK^{+}12]. Utilities [Den16]. Utility [Cro20].

Validating [CAA15]. Validation [BPDD08, BKvT^{+}14, KO06, TH21, vdLdJ21]. Value [GK16, HS18, LZ17, LS02b, MMB15, McB06]. Valued [dI17]. Values [CCH15, JH16, MF15, McB06, SB01, Som07, VS14, ZD17]. VAR [Pfa08]. Variable [AS02, AW16, CS12, HGK^{+}20, HMS16, LL11, LBC^{+}16, ND12, Ntz02, RJH14, Riz06, Sch11, SR18, TMW18, WF12, ZA17]. Variable-Subset [HGK^{+}20]. Variables [ALV07, Bon18, CC08, CKSLS12, FF14, GL08, KJ14, LS02a, MT98, MTW04, Mar06b, MT00]. Variance [KKH10, NO06, Pet97a, SF16, SF19]. Variance-Covariance [Pet97a]. Variances [ZKG20, Fir00]. Variety [Wei05]. Variogram [Bar96]. Various [ZKG20]. vars [Pfa08]. Varying [BR17, GL08, HW20, RC17, TR14]. vcd [MZH06]. vcrpart [BR17]. vdg [SlRC16]. vdmR [Fuj17]. Vector [KMH06, Mon06, Mur09, NL12]. Verlag [Gro08, Hof15, How16b, Kha15b, Mat15, Mat16a, Nun15, Pos15, Rus15, Zei16]. Versatile [ZKG20]. Version [GT10, GL08]. Vertex [HL07]. Very [RPC15]. VGAM [Yee10, YSH15]. via [AWBM18, AM14, BBGL17, DSD13, Dry09, DM19, FHT10, Gra16, Hua20, Jur15, KNI16, LM13, LWOW18, MW10, MR18, MP06, OWA20, SFHT11, SCS13, Su07, WMS17, Wan13b]. Video [MLVMY05]. Vignette [WKL11]. Vikram [Hof15]. VIM [KT16]. Vine [BS13]. Violators [dLHM09]. ViSta [MLVMY05]. Visual [Fuj17, Kam08, MHM20, PSW07, Kha03]. Visualisation [Grö17]. Visualization [Bec18, Fer11, GGK10, Grö21, Gro08, HHB^{+}08b, HRBP15, Kra07, LCL^{+}08, Mül16, OHD17, PU13, The02, Wic10, Wil14a]. Visualizations [ECW^{+}12]. Visualizing [BKT14, FW18, GRMS11, Gio09, IM05, Kie08, Lee18, LM03, MZH06, Mur06, Pen08a, RVCG19, VMY02, Dow18]. Visually [CCH15]. VNM [HWY18]. Volatility [Kas16]. Volume [AFJZ11, KP07, MvS07a, TW11, dLM07]. VOStat [BM07]. Votes [PLLC11]. vrmlgen [GGK10]. vsgoftest [LR20]. vSMC [Zho15].

WA [LCK11]. Warping [CS12, Gio09, MCAP19, SCS13]. WaveD [RS07].

Wavelet [AG19, ABS01, EN11, RS07, TPE19]. Wavelet-Based [AG19]. Way [Bro15, GKD14, Meu13, MD18, WH08, MZH06, MvS07b, MHT07]. Ways [AJ11]. Web [Iac15, Mai08, Mun14, Ban99, Fir00, Fir03, Fuj17, GGK10, MLAN02, SWAF15, SGK03, SGS^{+}14, SSS^{+}20, VFV13, Ver12, WO97]. Web-Based [Fuj17, SWAF15, Ban99, MLAN02]. WebBUGS [Zha14]. webchem [SSS^{+}20]. Webstat [WO97]. WebSwap [SGK03]. Weighted [BM12, CG15, DPSH18, FS10a, GLC^{+}15, MCAP19, MPI18, PT11, QY19, WS11, ZWH17]. Weighted-Mean [BM12]. Weighting [vdWG11]. Weights [PPT20]. Well [SH17]. wgaim [TV11]. WhatIf [SKZ05]. Whitney [FO15]. Whittle [AG19]. Wide [WO97]. Wilcoxon [JLH^{+}20]. Wildlife [Cal07, FC11]. Wiley [But15, Hel15, Iac15, Kha15a]. WileyCDA [But15, Kha15a]. WileyTitle [But15, Kha15a]. Williams [WWG09]. WinBUGS [CBDM07, CRW05, CG10, KCCG11, SR07a, SLG05, WEH11]. WinBUGSio [SR07a]. Windows [PG15, SGHY11]. Wise [BR17, PPT20]. Within [Mug10, Pfa08, LC13, Lec14, Ver12, WEH11, ZPC^{+}16]. Without [Han07a, BG18, KMG^{+}13]. wnominate [PLLC11]. Workflow [Bak20]. Working [Gos14, RGK20, WST11, GR17b]. Workplace [Kar06]. World [Tiv19, WO97]. Wrapper [CC08]. Wrappers [WLK08]. Write [Str04]. wsrf [ZWH17]. www.amherst.edu [Grö15a]. www.ceremade.dauphine.fr [Pos15]. www.crcpress.com [Gle16, Grö16, Kha16b, Zei15]. www.imperial.ac.uk [Hel15]. www.r-datacollection.com [Iac15]. www.sciencedirect.com [How15b]. www.springer.com [Hof15, How16b, Kha15b, Mat15, Mat16a, Nun15, Rus15, Zei16]. www.wiley.com [But15, Kha15a].

X [EC06, SE18]. X-13ARIMA-SEATS [SE18]. x12 [KMTS14]. x12GUI [KMTS14]. xian [Pos15]. xian/books.html [Pos15]. xiv [Kha15a]. XLISP [Lum96, Lum98, dL05d, BM96, Fir00, Scr01]. XML [Mun14]. Xorshift [Bre04, Mar03, PL05, Vig16]. xsample [VSV09]. xviii [Kha16b]. xx [Kha15b].

yaImpute [CF08]. Yield [FH10]. York [Gro08, Hil15a, Hil15b, How16b, Kha15b, Mat15, Mat16a, Nun15, Pos15]. YUIMA [BFH^{+}14, IMR17].

z [Hug07, Hug07]. Zelterman [Zei16]. zenplots [HO20]. Zero [FH10, LC10]. Zero-Coupon [FH10]. Zero-Inflated [LC10]. Zeros [WH08]. Ziggurat [LZL^{+}05, MT00]. Zigzag [HO20]. Zone [MBR11]. zoo [ZG05]. zt [BV02].

References

Ameijeiras-Alonso:2021:MRP

[AACRC21] Jose Ameijeiras-Alonso, Rosa M. Crujeiras, and Alberto Rodriguez-Casal. multimode: an R package for mode assessment. Journal of Statistical Software, 97(??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v097i09;
https://www.jstatsoft.org/index.php/jss/article/view/
v097i09/v97i09.pdf.

Austenfeld:2012:GUI

[AB12] Marcel Austenfeld and Wolfram Beyschlag. A graphical user interface for R in a rich client platform for ecological modeling. Journal of Statistical Software, 49(4):??, June 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v49/i04.

Anderson-Bergman:2017:IRM

[AB17] Clifford Anderson-Bergman. icenReg: Regression models for interval censored data in R. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v081i12;   https://www.jstatsoft.org/
index.php/jss/article/view/v081i12/v81i12.pdf.

Ardia:2019:MSG

[ABB^{+}19] David Ardia, Keven Bluteau, Kris Boudt, Leopoldo Catania, and Denis-Alexandre Trottier. Markov-switching GARCH models in R: The MSGARCH package. Journal of Statistical Software, 91(??):??, ???? 2019. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v091i04;   https://www.jstatsoft.
org/index.php/jss/article/view/v091i04/v91i04.pdf.

Ardia:2019:GAS

[ABC19] David Ardia, Kris Boudt, and Leopoldo Catania. Generalized autoregressive score models in R: The GAS package. Journal of Statistical Software, 88(??):??, ???? 2019. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v088i06;   https://www.
jstatsoft.org/index.php/jss/article/view/v088i06/v88i06.
pdf.

Antoniadis:2001:WEN

[ABS01] Anestis Antoniadis, Jérémie Bigot, and Theofanis Sapatinas. Wavelet estimators in nonparametric regression: a comparative simulation study. Journal of Statistical Software, 6(6):1–83, ???? 2001. CODEN JSSOBK. ISSN

1548-7660.  URL http://www.jstatsoft.org/v06/i06;
 http://www.jstatsoft.org/v06/i06/codes/;   http://
www.jstatsoft.org/v06/i06/html-help/;   http://www.
jstatsoft.org/v06/i06/jss057Rev.pdf;   http://www.
jstatsoft.org/v06/i06/updates.

Andersson:2013:PKM

[ABW13] Björn Andersson, Kenny Bränberg, and Marie Wiberg. Performing the kernel method of test equating with the package kequate. Journal of Statistical Software, 55(6):??, October 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v55/i06.

Alvarez:2017:PDT

[ÁBZ17] Inmaculada C. Álvarez, Javier Barbero, and José L. Zofío. A panel data toolbox for MATLAB. Journal of Statistical Software, 76(??):??, ???? 2017. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v076i06;   https://www.jstatsoft.
org/index.php/jss/article/view/v076i06/v76i06.pdf.

Alvarez:2020:DEA

[ÁBZ20] Inmaculada C. Álvarez, Javier Barbero, and José L. Zofío. A data envelopment analysis toolbox for MATLAB. Journal of Statistical Software, 95(??):??, ???? 2020. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v095i03;   https://www.
jstatsoft.org/index.php/jss/article/view/v095i03/v95i03.
pdf.

Achard:2009:BRB

[Ach09] Sophie Achard. Book review: The Statistical Analysis of Functional MRI Data. Journal of Statistical Software, 29(BR-3): ??, January 2009. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v29/b03.

Adragni:2012:GRP

[ACW12] Kofi Placid Adragni, R. Dennis Cook, and Seongho Wu. GrassmannOptim: An R package for Grassmann manifold optimization. Journal of Statistical Software, 50(5):??, July 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v50/i05.

Amatya:2015:CSO

[AD15] Anup Amatya and Hakan Demirtas. Code snippet: OrdNor: An R package for concurrent generation of correlated ordinal and normal data. Journal of Statistical Software, 68 (CS-2):??, ???? 2015. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/index.php/jss/
article/view/v068c02;   http://www.jstatsoft.org/index.
php/jss/article/view/v068c02/v68c02.pdf.

Abadie:2011:SRP

[ADH11] Alberto Abadie, Alexis Diamond, and Jens Hainmueller. Synth: An R package for synthetic control methods in comparative case studies. Journal of Statistical Software, 42(13): ??, June 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v42/i13.

Appelhans:2015:REO

[ADN15] Tim Appelhans, Florian Detsch, and Thomas Nauss. remote: Empirical orthogonal teleconnections in R. Journal of Statistical Software, 65(10):??, June 2015. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v65/i10.

Abreu:2010:HDG

[AEL10] Gabriel C. G. Abreu, David Edwards, and Rodrigo Labouriau. High-dimensional graphical model search with the gRapHD R package. Journal of Statistical Software, 37(1):??, November 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v37/i01.

Altman:2011:ISV

[AFJZ11] Micah Altman, John Fox, Simon Jackman, and Achim Zeileis. An introduction to the Special Volume on ’Political Methodology’. Journal of Statistical Software, 42(1):??, June 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v42/i01.

Achard:2019:WBF

[AG19] Sophie Achard and Irène Gannaz. Wavelet-based and Fourierbased multivariate Whittle estimation: multiwave. Journal of Statistical Software, 89(??):??, ???? 2019. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v089i06;   https://www.
jstatsoft.org/index.php/jss/article/view/v089i06/v89i06.
pdf.

Alfaro:2013:ARP

[AGG13] Esteban Alfaro, Matias Gamez, and Noelia García. adabag: An R package for classification with boosting and bagging. Journal of Statistical Software, 54(2):??, September 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v54/i02.

Atman:2007:ATA

[AGM07] Micah Atman, Jeff Gill, and Michael P. McDonald. accuracy: Tools for accurate and reliable statistical computing. Journal of Statistical Software, 21(1):1–30, July 2007. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v21/i01.

Aragam:2019:LLS

[AGZ19] Bryon Aragam, Jiaying Gu, and Qing Zhou. Learning largescale Bayesian networks with the sparsebn package. Journal of Statistical Software, 91(??):??, ???? 2019. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v091i11;   https://www.
jstatsoft.org/index.php/jss/article/view/v091i11/v91i11.
pdf.

Admiraal:2008:NPS

[AH08] Ryan Admiraal and Mark S. Handcock. networksis: a package to simulate bipartite graphs with fixed marginals through sequential importance sampling. Journal of Statistical Software, 24(8):1–21, May 2008. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v24/i08.

Altham:2012:MGM

[AH12] Pat M. E. Altham and Robin K. S. Hankin. Multivariate generalizations of the multiplicative binomial distribution: Introducing the MM package. Journal of Statistical Software, 46 (12):??, March 2012. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v46/i12.

Ardia:2009:AMS

[AHvD09] David Ardia, Lennart F. Hoogerheide, and Herman K. van Dijk. Adaptive mixture of Student-t distributions as a flexible candidate distribution for efficient simulation: The R package AdMit. Journal of Statistical Software, 29(3):??, January 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v29/i03.

Aizaki:2012:CSB

[Aiz12] Hideo Aizaki. Code snippet: Basic functions for supporting an implementation of choice experiments in R. Journal of Statistical Software, 50(CS-2):??, September 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v50/c02.

Altman:2011:NWL

[AJ11] Micah Altman and Simon Jackman. Nineteen ways of looking at statistical software. Journal of Statistical Software, 42(2): ??, June 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v42/i02.

Albano:2016:ERP

[Alb16] Anthony D. Albano. equate: An R package for observedscore linking and equating. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v074i08;   https://www.jstatsoft.org/
index.php/jss/article/view/v074i08/v74i08.pdf.

Almquist:2010:UCS

[Alm10] Zack W. Almquist. US census spatial and demographic data in R: The UScensus2000 suite of packages. Journal of Statistical Software, 37(6):??, November 2010. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v37/i06.

Almiron:2010:NAS

[ALO^{+}10] Marcelo G. Almiron, Bruno Lopes, Alyson L. C. Oliveira, Antonio C. Medeiros, and Alejandro C. Frery. On the numerical accuracy of spreadsheets. Journal of Statistical Software, 34 (4):??, April 2010. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v34/i04.

Anderson:2007:EMR

[ALV07] Carolyn J. Anderson, Zhushan Li, and Jeroen K. Vermunt. Estimation of models in a Rasch family for polytomous items and multiple latent variables. Journal of Statistical Software, 20(6):1–36, February 2007. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v20/i06.

Altman:2011:BBL

[AM11] Micah Altman and Michael P. McDonald. BARD: Better automated redistricting. Journal of Statistical Software, 42(4): ??, June 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v42/i04.

Azzalini:2014:CND

[AM14] Adelchi Azzalini and Giovanna Menardi. Clustering via nonparametric density estimation: The R package pdfCluster. Journal of Statistical Software, 57(11):??, May 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v57/i11.

Araujo:2014:PET

[AMMRP14] Artur Araújo, Luís Meira-Machado, and Javier Roca-Pardiñas. TPmsm: Estimation of the transition probabilities in 3-state models. Journal of Statistical Software, 62(4):??, December 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v62/i04.

Augugliaro:2014:PDR

[AMW14] Luigi Augugliaro, Angelo M. Mineo, and Ernst C. Wit. dglars: An R package to estimate sparse generalized linear models. Journal of Statistical Software, 59(8):??, September 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v59/i08.

Agostinelli:2016:REG

[AMYR16] Claudio Agostinelli, Alfio Marazzi, Víctor J. Yohai, and Alex Randriamiharisoa. Robust estimation of the generalized loggamma model: The R package robustloggamma. Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v070i07;   https://www.
jstatsoft.org/index.php/jss/article/view/v070i07/v70i07.
pdf.

Adragni:2014:PLR

[AR14] Kofi P. Adragni and Andrew M. Raim. ldr: An R software package for likelihood-based sufficient dimension reduction. Journal of Statistical Software, 61(3):??, November 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v61/i03.

Archer:2010:RRP

[Arc10] Kellie J. Archer. rpartOrdinal: An R package for deriving a classification tree for predicting an ordinal response. Journal of Statistical Software, 34(7):??, April 2010. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v34/i07.

Armero:2019:BRD

[Arm19] Carmen Armero. Book review: Data Analysis Using Hierarchical Generalized Linear Models with R. Journal of Statistical Software, 88(??):??, ???? 2019. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v088b01;   https://www.jstatsoft.
org/index.php/jss/article/view/v088b01/v88b01.pdf.

Al-Subaihi:2002:VSM

[AS02] Ali A. Al-Subaihi. Variable selection in multivariable regression using SAS/IML. Journal of Statistical Software, 7(12):1–20, ???? 2002. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v07/i12;   http:
//www.jstatsoft.org/v07/i12/mv.pdf.

Al-Subaihi:2004:SCM

[AS04] Ali A. Al-Subaihi. Simulating correlated multivariate pseudorandom numbers. Journal of Statistical Software, 9 (4):1–20, ???? 2004. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/counter.php?id=
85&url=v09/i04/paper.pdf&ct=1.

Allignol:2011:ETM

[ASB11] Arthur Allignol, Martin Schumacher, and Jan Beyersmann. Empirical transition matrix of multi-state models: The etm package. Journal of Statistical Software, 38(4):??, January 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v38/i04.

Anota:2017:QRP

[ASBMB17] Amélie Anota, Marion Savina, Caroline Bascoul-Mollevi, and Franck Bonnetain. QoLR: An R package for the longitudinal analysis of health-related quality of life in oncology. Journal of Statistical Software, 77(??):??, ???? 2017. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v077i12;   https://www.
jstatsoft.org/index.php/jss/article/view/v077i12/v77i12.
pdf.

Alfons:2013:ESE

[AT13] Andreas Alfons and Matthias Templ. Estimation of social exclusion indicators from complex surveys: The R package laeken. Journal of Statistical Software, 54(15):??, September 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v54/i15.

Alfons:2010:OOF

[ATF10] Andreas Alfons, Matthias Templ, and Peter Filzmoser. An object-oriented framework for statistical simulation: The R package simFrame. Journal of Statistical Software, 37(3):??, November 2010. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v37/i03.

Almutiry:2021:CTI

[AVD21] Waleed Almutiry, Vineetha Warriyar K V, and Rob Deardon. Continuous time individual-level models of infectious disease: Package EpiILMCT. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v098i10;   https://www.jstatsoft.org/
index.php/jss/article/view/v098i10/v98i10.pdf.

An:2016:CSL

[AW16] Weihua An and Xuefu Wang. Code snippet: LARF: Instrumental variable estimation of causal effects through local average response functions. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v071c01;   https://www.jstatsoft.org/
index.php/jss/article/view/v071c01/v71c01.pdf.

Andrews:2018:TRP

[AWBM18] Jeffrey L. Andrews, Jaymeson R. Wickins, Nicholas M. Boers, and Paul D. McNicholas. teigen: An R package for model-based clustering and classification via the multivariate t distribution. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v083i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v083i07/v83i07.pdf.

Badsberg:2001:GC

[Bad01] Jens Henrik Badsberg. A guide to CoCo. Journal of Statistical Software, 6(4):1–190, ???? 2001. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v06/i04;   http://www.jstatsoft.org/v06/i04/CoCo-1.
5/;   http://www.jstatsoft.org/v06/i04/guide-JSS-1.
pdf;   http://www.jstatsoft.org/v06/i04/PartII/;
http://www.jstatsoft.org/v06/i04/updates.

Baiocchi:2004:UPS

[Bai04] Giovanni Baiocchi. Using Perl for statistics: Data processing and statistical computing. Journal of Statistical Software, 11(1):1–81, ???? 2004. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/counter.php?id=
95&url=v11/i01/paper.pdf&ct=1.

Baio:2020:SSA

[Bai20] Gianluca Baio. survHE: Survival analysis for health economic evaluation and cost-effectiveness modeling. Journal of Statistical Software, 95(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v095i14;   https://www.jstatsoft.
org/index.php/jss/article/view/v095i14/v95i14.pdf.

Baker:2020:CSU

[Bak20] Peter Baker. Code snippet: Using GNU Make to manage the workflow of data analysis projects. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v094c01;   https://www.jstatsoft.
org/index.php/jss/article/view/v094c01/v94c01.pdf.

Banfield:1999:RWB

[Ban99] Jeff Banfield. Rweb: Web-based statistical analysis. Journal of Statistical Software, 4(1):1–15, ???? 1999. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v04/i01;
 http://www.jstatsoft.org/v04/i01/Rweb.pdf;   http:
//www.jstatsoft.org/v04/i01/Rweb.ps;   http://www.
jstatsoft.org/v04/i01/Rweb.tar.gz;   http://www.jstatsoft.
org/v04/i01/Rweb/Rweb.html;   http://www.jstatsoft.
org/v04/i01/updates.

Bandyopadhyay:2016:BRS

[Ban16] Dipankar Bandyopadhyay. Book review: Sampling Spatial Units for Agricultural Surveys. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v069b06;   https://www.jstatsoft.org/
index.php/jss/article/view/v069b06/v69b06.pdf.

Barry:1996:DAF

[Bar96] Ronald Barry. A diagnostic to assess the fit of a variogram model to spatial data. Journal of Statistical Software, 1(1):1–11, ???? 1996. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v01/i01;
http://www.jstatsoft.org/v01/i01/barry.pdf;   http:
//www.jstatsoft.org/v01/i01/code.txt;   http://www.
jstatsoft.org/v01/i01/updates.

Barcaroli:2014:PRP

[Bar14] Giulio Barcaroli. SamplingStrata: An R package for the optimization of stratified sampling. Journal of Statistical Software, 61(4):??, November 2014. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v61/i04.

Barber:2018:BRF

[Bar18] Xavier Barber. Book review: Flexible Regression and Smoothing: Using GAMLSS in R. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/

article/view/v085b02;   https://www.jstatsoft.org/
index.php/jss/article/view/v085b02/v85b02.pdf.

Battauz:2015:ERP

[Bat15] Michela Battauz. equateIRT: An R package for IRT test equating. Journal of Statistical Software, 68(7):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v068i07;
http://www.jstatsoft.org/index.php/jss/article/view/
v068i07/v68i07.pdf.

Bergmeir:2012:NNR

[BB12] Christoph Bergmeir and José M. Benítez. Neural networks in R using the Stuttgart Neural Network Simulator: RSNNS. Journal of Statistical Software, 46(7):??, January 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v46/i07.

Berge:2012:HRP

[BBG12] Laurent Bergé, Charles Bouveyron, and Stéphane Girard. HDclassif: An R package for model-based clustering and discriminant analysis of high-dimensional data. Journal of Statistical Software, 46(6):??, January 2012. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v46/i06.

Becker:2017:ERC

[BBGL17] Gabriel Becker, Cory Barr, Robert Gentleman, and Michael Lawrence. Enhancing reproducibility and collaboration via management of R package Cohorts. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v082i01;   https://www.jstatsoft.
org/index.php/jss/article/view/v082i01/v82i01.pdf.

Bowman:2007:RSI

[BCAB07] Adrian Bowman, Ewan Crawford, Gavin Alexander, and Richard W. Bowman. rpanel: Simple interactive controls for R functions using the tcltk package. Journal of Statistical Software, 17(9):1–18, January 2007. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v17/i09.

Benaglia:2009:MRP

[BCHY09] Tatiana Benaglia, Didier Chauveau, David R. Hunter, and Derek S. Young. mixtools: An R package for analyzing mixture models. Journal of Statistical Software, 32(6):??, October 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v32/i06.

Botcher:2003:PDP

[BD03] Susanne G. Bøtcher and Claus Dethlefsen. deal: a package for learning Bayesian networks. Journal of Statistical Software, 8 (20):1–40, December 28, 2003. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v08/i20/deal.
pdf;   http://www.jstatsoft.org/v08/i20/deal_1.2-4.
tar.gz;   http://www.jstatsoft.org/v08/i20/jcgs.tex.

Bougeard:2018:SMA

[BD18] Stéphanie Bougeard and Stéphane Dray. Supervised multiblock analysis in R with the ade4 package. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v086i01;   https://www.jstatsoft.
org/index.php/jss/article/view/v086i01/v86i01.pdf.

Bordier:2011:TSI

[BDdM11] Cécile Bordier, Michel Dojat, and Pierre Lafaye de Micheaux. Temporal and spatial independent component analysis for fMRI data sets embedded in the AnalyzeFMRI R package. Journal of Statistical Software, 44(9):??, October 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v44/i09.

Benham:2017:CCE

[BDKL17] Tim Benham, Qibin Duan, Dirk P. Kroese, and Benoît Liquet. CEoptim: Cross-entropy R package for optimization. Journal of Statistical Software, 76(??):??, ???? 2017. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v076i08;   https://www.
jstatsoft.org/index.php/jss/article/view/v076i08/v76i08.
pdf.

Bender-deMol:2008:PPM

[BdMM08] Skye Bender-deMoll, Martina Morris, and James Moody. Prototype packages for managing and animating longitudinal network data: dynamicnetwork and rSoNIA. Journal of Statistical Software, 24(7):1–36, May 2008. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v24/i07.

Bilodeau:2015:RPG

[BdMM15] Martin Bilodeau, Pierre Lafaye de Micheaux, and Smail Mahdi. The R package groc for generalized regression on orthogonal components. Journal of Statistical Software, 65(1): ??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v65/i01.

Bagnato:2015:CSP

[BDMP15] Luca Bagnato, Lucio De Capitani, Angelo Mazza, and Antonio Punzo. Code snippet: SDD: An R package for serial dependence diagrams. Journal of Statistical Software, 64(CS-2):??, March 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v64/c02.

Balboa:2018:ETP

[BdUÁ18] Vanesa Balboa and Jacobo de Uña-Álvarez. Estimation of transition probabilities for the illness-death model: Package TP.idm. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v083i10;
https://www.jstatsoft.org/index.php/jss/article/view/
v083i10/v83i10.pdf.

Bates:2013:FEN

[BE13] Douglas Bates and Dirk Eddelbuettel. Fast and elegant numerical linear algebra using the RcppEigen package. Journal of Statistical Software, 52(5):??, February 2013. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v52/i05.

Beath:2017:RRP

[Bea17] Ken J. Beath. randomLCA: An R package for latent class with random effects analysis. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v081i13;   https://www.jstatsoft.org/
index.php/jss/article/view/v081i13/v81i13.pdf.

Beck:2018:NVA

[Bec18] Marcus W. Beck. NeuralNetTools: Visualization and analysis tools for neural networks. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v085i11;   https://www.jstatsoft.org/
index.php/jss/article/view/v085i11/v85i11.pdf.

Burchett:2017:NIM

[BEHB17] Woodrow W. Burchett, Amanda R. Ellis, Solomon W. Harrar, and Arne C. Bathke. Nonparametric inference for multivariate data: The R package npmv. Journal of Statistical Software, 76(??):??, ???? 2017. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v076i04;   https://www.jstatsoft.
org/index.php/jss/article/view/v076i04/v76i04.pdf.

Beliakov:2011:CSF

[Bel11a] Gleb Beliakov. Code snippet: Fast computation of trimmed means. Journal of Statistical Software, 39(CS-2):??, March 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v39/c02.

Bell:2011:RFP

[Bel11b] William R. Bell. REGCMPNT — a Fortran program for regression models with ARIMA component errors. Journal of Statistical Software, 41(7):??, May 2011. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v41/i07.

Berchtold:2001:MCC

[Ber01] Andre Berchtold. Markov chain computation for homogeneous and non-homogeneous data: MARCH 1.1 users guide. Journal of Statistical Software, 6(3):1–81, ???? 2001. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v06/i03;   http://www.jstatsoft.org/
v06/i03/march-11-c-win.zip;   http://www.jstatsoft.
org/v06/i03/march-11-ug.ps;   http://www.jstatsoft.
org/v06/i03/march-ex1_0;   http://www.jstatsoft.org/
v06/i03/march-ex2_0;   http://www.jstatsoft.org/v06/
i03/march.tar.gz;   http://www.jstatsoft.org/v06/i03/
march_11_ug.pdf;   http://www.jstatsoft.org/v06/i03/
updates.

Berens:2009:CMT

[Ber09] Philipp Berens. CircStat: a MATLAB toolbox for circular statistics. Journal of Statistical Software, 31(10):??, September 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v31/i10.

Bilgrau:2016:GUC

[BER^{+}16] Anders Ellern Bilgrau, Poul Svante Eriksen, Jakob Gulddahl Rasmussen, Hans Erik Johnsen, Karen Dybkaer, and Martin Boegsted. GMCM: Unsupervised clustering and meta-analysis using Gaussian mixture copula models. Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v070i02;   https://www.jstatsoft.
org/index.php/jss/article/view/v070i02/v70i02.pdf.

Belaire-Franch:2002:RPN

[BFC02] Jorge Belaire-Franch and Dulce Contreras. Recurrence plots in nonlinear time series analysis: Free software. Journal of Statistical Software, 7(9):1–18, ???? 2002. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v07/i09;   http://www.jstatsoft.org/
v07/i09/JSS_055.pdf;   http://www.jstatsoft.org/v07/
i09/updates.

Baker:2019:SRP

[BFFN19] Jack Baker, Paul Fearnhead, Emily B. Fox, and Christopher Nemeth. sgmcmc: an R package for stochastic gradient Markov Chain Monte Carlo. Journal of Statistical Software, 91(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v091i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v091i03/v91i03.pdf.

Brouste:2014:YPC

[BFH^{+}14] Alexandre Brouste, Masaaki Fukasawa, Hideitsu Hino, Stefano M. Iacus, Kengo Kamatani, Yuta Koike, Hiroki Masuda, Ryosuke Nomura, Teppei Ogihara, Yasutaka Shimuzu, Masayuki Uchida, and Nakahiro Yoshida. The YUIMA Project: A computational framework for simulation and inference of stochastic differential equations. Journal of Statistical Software, 57(4):??, April 2014. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v57/i04.

Brewer:2018:DDN

[BFM18] Brendon J. Brewer and Daniel Foreman-Mackey. DNest4: Diffusive nested sampling in C++ and Python. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v086i07;   https://www.jstatsoft.
org/index.php/jss/article/view/v086i07/v86i07.pdf.

Barragan:2013:IRP

[BFRP13] Sandra Barragán, Miguel A. Fernández, Cristina Rueda, and Shyamal Das Peddada. isocir: An R package for constrained inference using isotonic regression for circular data, with an application to cell biology. Journal of Statistical Software, 54 (4):??, September 2013. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v54/i04.

Backlin:2018:DFC

[BG18] Christofer L. Bäcklin and Mats G. Gustafsson. Developerfriendly and computationally efficient predictive modeling without information leakage: The emil package for R. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v085i13;   https://www.
jstatsoft.org/index.php/jss/article/view/v085i13/v85i13.
pdf.

Binois:2021:HHG

[BG21] Mickaël Binois and Robert B. Gramacy. hetGP: Heteroskedastic Gaussian process modeling and sequential design in R. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v098i13;   https://
www.jstatsoft.org/index.php/jss/article/view/v098i13/
v98i13.pdf.

Basturk:2017:RPM

[BGH^{+}17] Nalan Bastürk, Stefano Grassi, Lennart Hoogerheide, Anne Opschoor, and Herman K. van Dijk. The R package MitISEM: Efficient and robust simulation procedures for Bayesian inference. Journal of Statistical Software, 79 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/

article/view/v079i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v079i01/v79i01.pdf.

Burkett:2006:HSL

[BGM06] Kelly Burkett, Jinko Graham, and Brad McNeney. hapassoc: Software for likelihood inference of trait associations with SNP haplotypes and other attributes. Journal of Statistical Software, 16(2):1–19, April 2006. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v16/i02.

Bivand:2015:SDA

[BGRR15] Roger S. Bivand, Virgilio Gómez-Rubio, and Håvard Rue. Spatial data analysis with r-INLA with some extensions. Journal of Statistical Software, 63(20):??, February 2015. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v63/i20.

Bowman:2010:ITT

[BGSC10] Adrian W. Bowman, Iain Gibson, Ewan Marian Scott, and Ewan Crawford. Interactive teaching tools for spatial sampling. Journal of Statistical Software, 36(13):??, October 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v36/i13.

Brown:2007:RAM

[BH07] J. Randall Brown and Milton E. Harvey. Rational arithmetic Mathematica functions to evaluate the one-sided one-sample K-S cumulative sample distribution. Journal of Statistical Software, 19(6):1–32, March 2007. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v19/i06.

Brown:2008:APM

[BH08a] J. Randall Brown and Milton E. Harvey. Arbitrary precision Mathematica functions to evaluate the one-sided one sample K-S cumulative sampling distribution. Journal of Statistical Software, 26(3):1–55, June 2008. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v26/i03.

Brown:2008:RAM

[BH08b] J. Randall Brown and Milton E. Harvey. Rational arithmetic Mathematica functions to evaluate the two-sided one sample K-S cumulative sampling distribution. Journal of Statistical Software, 26(2):1–40, June 2008. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v26/i02.

Bhatia:2017:BRP

[BIG17] Parmeet Singh Bhatia, Serge Iovleff, and Gérard Govaert. blockcluster: An R package for model-based co-clustering. Journal of Statistical Software, 76(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v076i09;   https://
www.jstatsoft.org/index.php/jss/article/view/v076i09/
v76i09.pdf.

Brouste:2007:FGR

[BILL07] Alexandre Brouste, Jacques Istas, and Sophie Lambert-Lacroix. On fractional Gaussian random fields simulations. Journal of Statistical Software, 23(1):1–23, November 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v23/i01.

Brouste:2010:SMI

[BILL10] Alexandre Brouste, Jacques Istas, and Sophie Lambert-Lacroix. On simulation of manifold indexed fractional Gaussian fields. Journal of Statistical Software, 36(4):??, August 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v36/i04.

Bailey:2011:CSI

[BK11] Delia Bailey and Jonathan N. Katz. Code snippet: Implementing panel-corrected standard errors in R: The pcse package. Journal of Statistical Software, 42(CS-1):??, June 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v42/c01.

Blazejowski:2015:BMA

[BK15] Marcin Błażejowski and Jacek Kwiatkowski. Bayesian model averaging and jointness measures for gretl. Journal of Statistical Software, 68(5):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v068i05;
http://www.jstatsoft.org/index.php/jss/article/view/
v068i05/v68i05.pdf.

Biecek:2017:ARP

[BK17] Przemyslaw Biecek and Marcin Kosiński. archivist: An R package for managing, recording and restoring data analysis results. Journal of Statistical Software, 82 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v082i11;   https://www.jstatsoft.org/
index.php/jss/article/view/v082i11/v82i11.pdf.

Brezger:2005:BAB

[BKL05] Andreas Brezger, Thomas Kneib, and Stefan Lang. BayesX: Analyzing Bayesian structural additive regression models. Journal of Statistical Software, 14(11):1–22, September 9, 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/counter.php?id=136&url=v14/i11&ct=2;
 http://www.jstatsoft.org/counter.php?id=136&url=v14/
i11/v14i11.pdf&ct=1.

Berentsen:2014:IPL

[BKT14] Geir Drage Berentsen, Tore Selland Kleppe, and Dag Tjøstheim. Introducing localgauss, an R package for estimating and visualizing local Gaussian correlation. Journal of Statistical Software, 56(12):??, February 2014. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v56/i12.

Budczies:2014:CRP

[BKvT^{+}14] Jan Budczies, Daniel Kosztyla, Christian von Törne, Albrecht Stenzinger, Silvia Darb-Esfahani, Manfred Dietel, and Carsten Denkert. cancerclass: An R package for development and validation of diagnostic tests from high-dimensional molecular data. Journal of Statistical Software, 59(1):??, August 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v59/i01.

Bada:2014:PPP

[BL14] Oualid Bada and Dominik Liebl. phtt: Panel data analysis with heterogeneous time trends in R. Journal of Statistical Software, 59(6):??, September 2014. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v59/i06.

Bolin:2018:CPE

[BL18] David Bolin and Finn Lindgren. Calculating probabilistic excursion sets and related quantities using excursions. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v086i05;   https://www.
jstatsoft.org/index.php/jss/article/view/v086i05/v86i05.
pdf.

Bischl:2015:PBP

[BLM^{+}15] Bernd Bischl, Michel Lang, Olaf Mersmann, Jörg Rahnenführer, and Claus Weihs. BatchJobs and BatchExperiments: Abstraction mechanisms for using R in batch environments. Journal of Statistical Software, 64(11):??, March 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v64/i11.

Brown:1997:APL

[BLR97] Barry W. Brown, James Lovato, and Kathy Russell. The Asypow S(plus) library for asymptotic power calculations. Journal of Statistical Software, 2(2):1–21, ???? 1997. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v02/i02;   http://www.jstatsoft.org/v02/i02/asypow-
2.1.tar.Z;   http://www.jstatsoft.org/v02/i02/dist;
 http://www.jstatsoft.org/v02/i02/docs/asypow.pdf;
 http://www.jstatsoft.org/v02/i02/docs/asypow.ps.gz;
 http://www.jstatsoft.org/v02/i02/docs/asypow/;   http:
//www.jstatsoft.org/v02/i02/docs/cheat.pdf;   http:
//www.jstatsoft.org/v02/i02/docs/cheat.ps.gz;   http:
//www.jstatsoft.org/v02/i02/docs/cheat/;   http://
www.jstatsoft.org/v02/i02/updates.

Beare:2018:ISR

[BLY18] Richard Beare, Bradley Lowekamp, and Ziv Yaniv. Image segmentation, registration and characterization in R with SimpleITK. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v086i08;
https://www.jstatsoft.org/index.php/jss/article/view/
v086i08/v86i08.pdf.

Bond:1996:HAX

[BM96] Jason Bond and George Michailides. Homogeneity analysis in Xlisp–Stat. Journal of Statistical Software, 1(2):??, ???? 1996. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v01/i02;   http://www.jstatsoft.org/
v01/i02/code_old.phtml;   http://www.jstatsoft.org/
v01/i02/data.phtml;   http://www.jstatsoft.org/v01/
i02/PAPER/paper.html;   http://www.jstatsoft.org/v01/
i02/updates.

Bois:1997:MMC

[BM97a] Freédéric Y. Bois and Don R. Maszle. MCSim: a Monte Carlo simulation program. Journal of Statistical Software, 2(9):1–60, ???? 1997. CODEN JSSOBK. ISSN

1548-7660.  URL http://www.jstatsoft.org/v02/i09;
 http://www.jstatsoft.org/v02/i09/mcsim/doc/mcsim.
html;   http://www.jstatsoft.org/v02/i09/mcsim/doc/
mcsim.pdf;   http://www.jstatsoft.org/v02/i09/mcsim/
doc/mcsim.ps;   http://www.jstatsoft.org/v02/i09/mcsim/
sim;   http://www.jstatsoft.org/v02/i09/updates.

Bond:1997:ICA

[BM97b] Jason Bond and George Michailides. Interactive correspondence analysis in a dynamic object-oriented environment. Journal of Statistical Software, 2(8):1–30, ???? 1997. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v02/i08;   http://www.jstatsoft.org/
v02/i08/jss.pdf;   http://www.jstatsoft.org/v02/i08/
jss.ps.gz;   http://www.jstatsoft.org/v02/i08/texstuff/
jss/;   http://www.jstatsoft.org/v02/i08/updates;
http://www.jstatsoft.org/v02/i08/xlisp-anacor.lsp;
 http://www.jstatsoft.org/v02/i08/xlisp-anacor.xli.

Babu:2007:URB

[BM07] G. Jogesh Babu and Ashish Mahabal. Using R-based VOStat as a low-resolution spectrum analysis tool. Journal of Statistical Software, 18(11):1–12, January 2007. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v18/i11.

Bazovkin:2012:EAW

[BM12] Pavel Bazovkin and Karl Mosler. An exact algorithm for weighted-mean trimmed regions in any dimension. Journal of Statistical Software, 47(13):??, May 2012. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v47/i13.

Bergmeir:2016:MAL

[BMB16] Christoph Bergmeir, Daniel Molina, and José M. Benítez. Memetic algorithms with local search chains in R: The Rmalschains package. Journal of Statistical Software, 75 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v075i04;   https://www.jstatsoft.org/
index.php/jss/article/view/v075i04/v75i04.pdf.

Bates:2015:FLM

[BMBW15] Douglas Bates, Martin Mächler, Ben Bolker, and Steve Walker. Fitting linear mixed-effects models using lme4. Journal of Statistical Software, 67(1):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067i01;
http://www.jstatsoft.org/index.php/jss/article/view/
v067i01/v67i01.pdf.

Bunouf:2015:SPC

[BMGT15] Pierre Bunouf, Geert Molenberghs, Jean-Marie Grouin, and Herbert Thijs. A SAS program combining R functionalities to implement pattern-mixture models. Journal of Statistical Software, 68(8):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v068i08;
http://www.jstatsoft.org/index.php/jss/article/view/
v068i08/v68i08.pdf.

Birgin:2014:SPG

[BMR14] Ernesto G. Birgin, Jose Mario Martínez, and Marcos Raydan. Spectral projected gradient methods: Review and perspectives. Journal of Statistical Software, 60(3):??, September 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v60/i03.

Binsl:2007:FRP

[BMvS^{+}07] Thomas W. Binsl, Katharine M. Mullen, Ivo H. M. van Stokkum, Jaap Heringa, and Johannes H. G. M. van Beek. FluxSimulator: An R package to simulate isotopomer distributions in metabolic networks. Journal of Statistical Software, 18(7):1–17, January 2007. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v18/i07.

Boik:2010:RPA

[BN10] John C. Boik and Balasubramanian Narasimhan. An R package for assessing drug synergism/antagonism. Journal of Statistical Software, 34(6):??, April 2010. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v34/i06.

Bonat:2018:MRV

[Bon18] Wagner Hugo Bonat. Multiple response variables regression models in R: The mcglm package. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v084i04;   https://www.jstatsoft.org/
index.php/jss/article/view/v084i04/v84i04.pdf.

Bos:2011:BAU

[Bos11] Charles S. Bos. A Bayesian analysis of unobserved component models using Ox. Journal of Statistical Software, 41(13):??, May 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v41/i13.

Bowman:2009:BRB

[Bow09] Adrian Bowman. Book review: Morphometrics with R. Journal of Statistical Software, 31(BR-1):??, September 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v31/b01.

Bowman:2010:BRB

[Bow10] Adrian Bowman. Book review: Functional Data Analysis with R and MATLAB. Journal of Statistical Software, 34(BR-3): ??, April 2010. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v34/b03.

Boyle:2005:BRA

[Boy05] Thomas Boyle. Book review: The Analysis of Gene Expression Data: Methods and Software. Journal of Statistical Software, 14(BR-1):1–2, April 2005. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v14/b01.

Basto:2012:SRM

[BP12] Mário Basto and José Manuel Pereira. An SPSS R-menu for ordinal factor analysis. Journal of Statistical Software, 46(4): ??, January 2012. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v46/i04.

Bivand:2015:CIE

[BP15] Roger Bivand and Gianfranco Piras. Comparing implementations of estimation methods for spatial econometrics. Journal of Statistical Software, 63(18):??, February 2015. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v63/i18.

Balan:2019:FRP

[BP19a] Theodor Adrian Balan and Hein Putter. frailtyEM: an R package for estimating semiparametric shared frailty models. Journal of Statistical Software, 90(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v090i07;   https://
www.jstatsoft.org/index.php/jss/article/view/v090i07/
v90i07.pdf.

Binois:2019:GRP

[BP19b] Mickaël Binois and Victor Picheny. GPareto: an R package for Gaussian-process-based multi-objective optimization and analysis. Journal of Statistical Software, 89 (??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v089i08;   https://www.jstatsoft.org/
index.php/jss/article/view/v089i08/v89i08.pdf.

Bornkamp:2009:MRP

[BPB09] Björn Bornkamp, José Pinheiro, and Frank Bretz. MCPMod: An R package for the design and analysis of dose-finding studies. Journal of Statistical Software, 29(7):??, February 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v29/i07.

Brock:2008:CRP

[BPDD08] Guy Brock, Vasyl Pihur, Susmita Datta, and Somnath Datta. clValid: An R package for cluster validation. Journal of Statistical Software, 25(4):1–22, March 2008. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v25/i04.

Bonhomme:2014:PMO

[BPGC14] Vincent Bonhomme, Sandrine Picq, Cédric Gaucherel, and Julien Claude. Momocs: Outline analysis using R. Journal of Statistical Software, 56(13):??, February 2014. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v56/i13.

Bartolucci:2017:LRP

[BPP17] Francesco Bartolucci, Silvia Pandolfi, and Fulvia Pennoni. LMest: An R package for latent Markov models for longitudinal categorical data. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v081i04;   https://www.jstatsoft.org/
index.php/jss/article/view/v081i04/v81i04.pdf.

Baillargeon:2007:PRL

[BR07] Sophie Baillargeon and Louis-Paul Rivest. Rcapture: Loglinear models for capture-recapture in R. Journal of Statistical Software, 19(5):1–31, April 2007. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v19/i05.

Burgin:2017:CWT

[BR17] Reto Bürgin and Gilbert Ritschard. Coefficient-wise tree-based varying coefficient regression with vcrpart. Journal of Statistical Software, 80(??):??, ???? 2017. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v080i06;   https://www.jstatsoft.
org/index.php/jss/article/view/v080i06/v80i06.pdf.

Braun:2014:PRP

[Bra14] Michael Braun. trustOptim: An R package for trust region optimization with sparse Hessians. Journal of Statistical Software, 60(4):??, September 2014. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v60/i04.

Brandmaier:2015:PPR

[Bra15] Andreas M. Brandmaier. pdc: An R package for complexitybased clustering of time series. Journal of Statistical Software, 67(5):??, ???? 2015. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/index.php/jss/
article/view/v067i05;   http://www.jstatsoft.org/index.
php/jss/article/view/v067i05/v067i05.pdf.

Braun:2017:SRP

[Bra17] Michael Braun. sparseHessianFD: An R package for estimating sparse Hessian matrices. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v082i10;   https://www.jstatsoft.org/
index.php/jss/article/view/v082i10/v82i10.pdf.

Boik:1998:DIB

[BRC98] Robert J. Boik and James F. Robison-Cox. Derivatives of the incomplete beta function. Journal of Statistical Software, 3(1):1–20, ???? 1998. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v03/i01;   http:
//www.jstatsoft.org/v03/i01/beta.der.pdf;   http://
www.jstatsoft.org/v03/i01/beta.der.ps.gz;   http://
www.jstatsoft.org/v03/i01/code/;   http://www.jstatsoft.
org/v03/i01/updates.

Baty:2015:TNR

[BRC^{+}15] Florent Baty, Christian Ritz, Sandrine Charles, Martin Brutsche, Jean-Pierre Flandrois, and Marie-Laure Delignette-Muller. A toolbox for nonlinear regression in R: The package nlstools. Journal of Statistical Software, 66(5):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066i05;
http://www.jstatsoft.org/index.php/jss/article/view/
v066i05/v66i05.pdf.

Brent:2004:NMX

[Bre04] Richard P. Brent. Note on Marsaglia’s xorshift random number generators. Journal of Statistical Software, 11(5):1–5, 2004. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/counter.php?id=101&url=v11/i05/

v11i05.pdf&ct=1. See [Mar03, PL05, Vig16]. This article shows the equivalence of xorshift generators and the wellunderstood linear feedback shift register generators.

Brockhaus:2020:BFR

[BRG20] Sarah Brockhaus, David Rügamer, and Sonja Greven. Boosting functional regression models with FDboost. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/

index.php/jss/article/view/v094i10;   https://www.
jstatsoft.org/index.php/jss/article/view/v094i10/v94i10.
pdf.

Brown:2015:MBG

[Bro15] Patrick E. Brown. Model-based geostatistics the easy way. Journal of Statistical Software, 63(12):??, February 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v63/i12.

Brechmann:2013:MDC

[BS13] Eike Christian Brechmann and Ulf Schepsmeier. Modeling dependence with C- and D-Vine copulas: The R package CDVine. Journal of Statistical Software, 52(3):??, February 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v52/i03.

Bakar:2015:SST

[BS15] Khandoker Shuvo Bakar and Sujit K. Sahu. spTimer: Spatiotemporal Bayesian modeling using R. Journal of Statistical Software, 63(15):??, February 2015. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v63/i15.

Barthelemy:2018:CSM

[BS18] Johan Barthélemy and Thomas Suesse. Code snippet: mipfp: An R package for multidimensional array fitting and simulating multivariate Bernoulli distributions. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v086c02;   https://www.jstatsoft.
org/index.php/jss/article/view/v086c02/v86c02.pdf.

Biecek:2012:RPB

[BSVT12] Przemysław Biecek, Ewa Szczurek, Martin Vingron, and Jerzy Tiuryn. The R package bgmm: Mixture modeling with uncertain knowledge. Journal of Statistical Software, 47(3):??, April 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v47/i03.

Baddeley:2005:PSR

[BT05] Adrian Baddeley and Rolf Turner. spatstat: An R package for analyzing spatial point patterns. Journal of Statistical Software, 12(6):1–42, ???? 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=113&url=v12/i06&ct=2;   http://www.
jstatsoft.org/counter.php?id=113&url=v12/i06/v12i06.
pdf&ct=1.

Baddeley:2013:HGP

[BTMB13] Adrian Baddeley, Rolf Turner, Jorge Mateu, and Andrew Bevan. Hybrids of Gibbs point process models and their implementation. Journal of Statistical Software, 55(11):??, November 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v55/i11.

Bullard:2006:BRB

[Bul06] James Bullard. Book review: R Graphics. Journal of Statistical Software, 16(BR-7):1–3, October 2006. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v16/b07.

Burr:2012:BRP

[Bur12] Deborah Burr. bspmma: An R package for Bayesian semiparametric models for meta-analysis. Journal of Statistical Software, 50(4):??, July 2012. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v50/i04.

Burkner:2017:BRP

[Bür17] Paul-Christian Bürkner. brms: An R package for Bayesian multilevel models using Stan. Journal of Statistical Software, 80(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v080i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v080i01/v80i01.pdf.

Buttrey:2005:CPL

[But05] Samuel E. Buttrey. Calling the lp solve linear program software from R, S-PLUS and Excel. Journal of Statistical Software, 14(4):1–13, May 1, 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=129&url=v14/i04&ct=2;   http://www.
jstatsoft.org/counter.php?id=129&url=v14/i04/v14i04.
pdf&ct=1.

Butts:2008:NPM

[But08a] Carter T. Butts. network: a package for managing relational data in R. Journal of Statistical Software, 24(2):1–36, May 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v24/i02.

Butts:2008:SNA

[But08b] Carter T. Butts. Social network analysis with sna. Journal of Statistical Software, 24(6):1–51, May 2008. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v24/i06.

Buttrey:2009:EAS

[But09] Samuel E. Buttrey. An Excel add-in for statistical process control charts. Journal of Statistical Software, 30(13):??, June 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v30/i13.

Buttrey:2015:BRD

[But15] Samuel E. Buttrey. Book review: Data Mining Algorithms Explained Using R, Paveł Cichosz, John Wiley & Sons, Chichester, 2015. ISBN 978-1-118-33258-0. 683 pp. USD 64.99.

http://www.wiley.com/WileyCDA/WileyTitle/productCd-111833258X.html.

Journal of Statistical Software, 66(BR-2):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066b02;
http://www.jstatsoft.org/index.php/jss/article/view/
v066b02/v66b02.pdf.

Bonnet:2002:ZST

[BV02] Eric Bonnet and Yves Van de Peer. zt: a software tool for simple and partial Mantel tests. Journal of Statistical Software, 7(10):1–12, ???? 2002. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v07/
i10;   http://www.jstatsoft.org/v07/i10/updates;
http://www.jstatsoft.org/v07/i10/zt.pdf;   http://
www.jstatsoft.org/v07/i10/zt_linux.tar.tgz;   http:
//www.jstatsoft.org/v07/i10/zt_solaris.tar.tgz.

Benoit:2017:BBA

[BV17] Dries F. Benoit and Dirk Van den Poel. bayesQR: A Bayesian approach to quantile regression. Journal of Statistical Software, 76(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v076i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v076i07/v76i07.pdf.

Blaser:2015:PGR

[BVE^{+}15] Nello Blaser, Luisa Salazar Vizcaya, Janne Estill, Cindy Zahnd, Bindu Kalesan, Matthias Egger, Thomas Gsponer, and Olivia Keiser. gems: An R package for simulating from disease progression models. Journal of Statistical Software, 64(10):??, March 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v64/i10.

Bradley:2001:IJT

[BW01] James W. Bradley and R. Webster West. Interactive Java tools for exploring high-dimensional data. Journal of Statistical Software, 6(1):??, ???? 2001. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v06/i01;   http:
//www.jstatsoft.org/v06/i01/bradley.tar;   http://
www.jstatsoft.org/v06/i01/bradley/index.html;   http:
//www.jstatsoft.org/v06/i01/updates.

Brilleman:2021:SSD

[BWMBC21] Samuel L. Brilleman, Rory Wolfe, Margarita Moreno-Betancur, and Michael J. Crowther. Simulating survival data using the simsurv R package. Journal of Statistical Software, 97(??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v097i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v097i03/v97i03.pdf.

Bove:2019:MBD

[BYPJ19] Daniel Sabanés Bové, Wai Yin Yeung, Giuseppe Palermo, and Thomas Jaki. Model-based dose escalation designs in R with crmPack. Journal of Statistical Software, 89 (??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v089i10;   https://www.jstatsoft.org/
index.php/jss/article/view/v089i10/v89i10.pdf.

Chiodi:2017:MNP

[CA17] Marcello Chiodi and Giada Adelfio. Mixed non-parametric and parametric estimation techniques in R package etasFLP for earthquakes’ description. Journal of Statistical Software, 76(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v076i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v076i03/v76i03.pdf.

Cebrian:2015:PNR

[CAA15] Ana C. Cebrián, Jesús Abaurrea, and Jesús Asín. NHPoisson: An R package for fitting and validating nonhomogeneous Poisson processes. Journal of Statistical Software, 64(6):??, March 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v64/i06.

Caffo:2006:EHT

[Caf06] Brian Caffo. Exact hypothesis tests for log-linear models with exactLoglinTest. Journal of Statistical Software, 17(7):1–17, November 2006. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v17/i07.

Calenge:2007:EHS

[Cal07] Clément Calenge. Exploring habitat selection by wildlife with adehabitat. Journal of Statistical Software, 22(6):1–19, September 2007. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v22/i06.

Camarda:2012:MRP

[Cam12] Carlo G. Camarda. MortalitySmooth: An R package for smoothing Poisson counts with P-splines. Journal of Statistical Software, 50(1):??, July 2012. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v50/i01.

Cantoni:2004:ARQ

[Can04a] Eva Cantoni. Analysis of robust quasi-deviances for generalized linear models. Journal of Statistical Software, 10 (4):1–9, ???? 2004. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/counter.php?id=
94&url=v10/i04/robGLMquasidev1.pdf&ct=1.

Cantoni:2004:RNS

[Can04b] Eva Cantoni. Resistant nonparametric smoothing with S-PLUS. Journal of Statistical Software, 10(2):1–6, ???? 2004. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/counter.php?id=93&url=v10/i02/
Mspline4.pdf&ct=1.

Canale:2017:MRP

[Can17] Antonio Canale. msBP: An R package to perform Bayesian nonparametric inference using multiscale Bernstein polynomials mixtures. Journal of Statistical Software, 78 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v078i06;   https://www.jstatsoft.org/
index.php/jss/article/view/v078i06/v78i06.pdf.

Carbon:2013:CSB

[Car13] Claus-Christian Carbon. Code snippet: BiDimRegression: Bidimensional regression modeling using R. Journal of Statistical Software, 52(CS-1):??, March 2013. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v52/c01.

Carvalho:2015:CSI

[Car15] Luis Carvalho. Code snippet: An improved evaluation of Kolmogorov’s distribution. Journal of Statistical Software, 65(CS-3):??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v65/c03.

Caspi:2017:CSR

[Cas17] Itamar Caspi. Code snippet: Rtadf: Testing for bubbles with EViews. Journal of Statistical Software, 81 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v081c01;   https://www.jstatsoft.org/
index.php/jss/article/view/v081c01/v81c01.pdf.

Caro:2017:IRP

[CB17] Daniel H. Caro and Przemyslaw Biecek. intsvy: An R package for analyzing international large-scale assessment data. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v081i07;   https://www.
jstatsoft.org/index.php/jss/article/view/v081i07/v81i07.
pdf.

Charvat:2021:MRP

[CB21] Hadrien Charvat and Aurélien Belot. mexhaz: An R package for fitting flexible hazard-based regression models for overall and excess mortality with a random effect. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v098i14;   https://www.jstatsoft.
org/index.php/jss/article/view/v098i14/v98i14.pdf.

Campelo:2020:MPC

[CBA20] Felipe Campelo, Lucas S. Batista, and Claus Aranha. The MOEADr package: a component-based framework for multiobjective evolutionary algorithms based on decomposition. Journal of Statistical Software, 92(??):??, ???? 2020. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v092i06;   https://www.
jstatsoft.org/index.php/jss/article/view/v092i06/v92i06.
pdf.

Coombes:2019:CSP

[CBAA19] Kevin R. Coombes, Guy Brock, Zachary B. Abrams, and Lynne V. Abruzzo. Code snippet: Polychrome: Creating and assessing qualitative palettes with many colors. Journal of Statistical Software, 90(??):??, ???? 2019. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v090c01;   https://www.
jstatsoft.org/index.php/jss/article/view/v090c01/v90c01.
pdf.

Carrigan:2007:CMD

[CBDM07] Gretchen Carrigan, Adrian G. Barnett, Annette J. Dobson, and Gita Mishra. Compensating for missing data from longitudinal studies using WinBUGS. Journal of Statistical Software, 19(7):1–17, June 2007. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v19/i07.

Choi:2021:FRP

[CBHK21] Yun-Hee Choi, Laurent Briollais, Wenqing He, and Karen Kopciuk. FamEvent: an R package for generating and modeling time-to-event data in family designs. Journal of Statistical Software, 97(??):??, ???? 2021. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v097i07;   https://www.jstatsoft.
org/index.php/jss/article/view/v097i07/v97i07.pdf.

Cao:2008:ORP

[CC08] Kim-Anh Lê Cao and Patrick Chabrier. ofw: An R package to select continuous variables for multiclass classification with a stochastic wrapper method. Journal of Statistical Software, 28(9):??, November 2008. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v28/i09.

Childs:2011:MFS

[CC11] Aaron Childs and Yalin Chen. Multilevel fixed and sequential acceptance sampling: The R package MFSAS. Journal of Statistical Software, 43(6):??, July 2011. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v43/i06.

Chu:2009:SAI

[CCD09] Annie Chu, Jenny Cui, and Ivo D. Dinov. SOCR analyses: Implementation and demonstration of a new graphical statistics educational toolkit. Journal of Statistical Software, 30(3): ??, April 2009. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v30/i03.

Calonico:2019:NNK

[CCF19] Sebastian Calonico, Matias D. Cattaneo, and Max H. Farrell. nprobust: Nonparametric kernel-based estimation and robust bias-corrected inference. Journal of Statistical Software, 91(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v091i08;   https://www.jstatsoft.org/
index.php/jss/article/view/v091i08/v91i08.pdf.

Cheng:2015:VEM

[CCH15] Xiaoyue Cheng, Dianne Cook, and Heike Hofmann. Visually exploring missing values in multivariable data using a graphical user interface. Journal of Statistical Software, 68(6):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v068i06;
http://www.jstatsoft.org/index.php/jss/article/view/
v068i06/v68i06.pdf.

Calcagno:2010:GRP

[CdM10] Vincent Calcagno and Claire de Mazancourt. glmulti: An R package for easy automated model selection with (generalized) linear models. Journal of Statistical Software, 34(12):??, May 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v34/i12.

Cappello:2020:CRP

[CDP20] Claudia Cappello, Sandra De Iaco, and Donato Posa. covatest: an R package for selecting a class of spacetime covariance functions. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v094i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v094i01/v94i01.pdf.

Cervantes:2017:DRP

[Cer17] Víctor H. Cervantes. DFIT: An R package for Raju’s differential functioning of items and tests framework. Journal of Statistical Software, 76(??):??, ???? 2017. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v076i05;   https://www.jstatsoft.
org/index.php/jss/article/view/v076i05/v76i05.pdf.

Crookston:2008:YRP

[CF08] Nicholas L. Crookston and Andrew O. Finley. yaImpute: An R package for kNN imputation. Journal of Statistical Software, 23(10):1–16, January 2008. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v23/i10.

Caimo:2014:PBB

[CF14] Alberto Caimo and Nial Friel. Bergm: Bayesian exponential random graphs in R. Journal of Statistical Software, 61(2): ??, October 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v61/i02.

Causeur:2011:FAM

[CFHBK11] David Causeur, Chloe Friguet, Magalie Houee-Bigot, and Maela Kloareg. Factor analysis for multiple testing (FAMT): An R package for large-scale significance testing under dependence. Journal of Statistical Software, 40(14):??, May 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v40/i14.

Conde:2015:DRP

[CFSR15] David Conde, Miguel Fernández, Bonifacio Salvador, and Cristina Rueda. dawai: An R package for discriminant analysis with additional information. Journal of Statistical Software, 66(10):??, ???? 2015. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/index.php/jss/
article/view/v066i10;   http://www.jstatsoft.org/index.
php/jss/article/view/v066i10/v66i10.pdf.

Cook:2011:LPL

[CFT11] R. Dennis Cook, Liliana M. Forzani, and Diego R. Tomassi. LDR: a package for likelihood-based sufficient dimension reduction. Journal of Statistical Software, 39(3):??, March 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v39/i03.

Crainiceanu:2010:BFD

[CG10] Ciprian M. Crainiceanu and A. Jeffrey Goldsmith. Bayesian functional data analysis using WinBUGS. Journal of Statistical Software, 32(11):??, January 2010. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v32/i11.

Capanu:2015:CSB

[CG15] Marinela Capanu and Mithat Gönen. Code snippet: Building a nomogram for survey-weighted Cox models using R. Journal of Statistical Software, 64(CS-1):??, March 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v64/c01.

Charrad:2014:PNR

[CGBN14] Malika Charrad, Nadia Ghazzali, Véronique Boiteau, and Azam Niknafs. NbClust: An R package for determining the relevant number of clusters in a data set. Journal of Statistical Software, 61(6):??, November 2014. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v61/i06.

Choi:2011:LRP

[CGC11] Seung W. Choi, Laura E. Gibbons, and Paul K. Crane. lordif: An R package for detecting differential item functioning using iterative hybrid ordinal logistic regression/item response theory and Monte Carlo simulations. Journal of Statistical Software, 39(8):??, March 2011. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v39/i08.

Colombi:2014:PHR

[CGC14] Roberto Colombi, Sabrina Giordano, and Manuela Cazzaro. hmmm: An R package for hierarchical multinomial marginal models. Journal of Statistical Software, 59(11):??, September 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v59/i11.

Carpenter:2017:SPP

[CGH^{+}17] Bob Carpenter, Andrew Gelman, Matthew D. Hoffman, Daniel Lee, Ben Goodrich, Michael Betancourt, Marcus Brubaker, Jiqiang Guo, Peter Li, and Allen Riddell. Stan: A probabilistic programming language. Journal of Statistical Software, 76(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v076i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v076i01/v76i01.pdf.

Carpenter:2011:RIS

[CGK11] James R. Carpenter, Harvey Goldstein, and Michael G. Kenward. REALCOM-IMPUTE software for multilevel multiple imputation with mixed response types. Journal of Statistical Software, 45(5):??, December 2011. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v45/i05.

Casarin:2015:PSM

[CGRvD15] Roberto Casarin, Stefano Grassi, Francesco Ravazzolo, and Herman K. van Dijk. Parallel sequential Monte Carlo for efficient density combination: The DeCo MATLAB toolbox. Journal of Statistical Software, 68(3):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v068i03;
http://www.jstatsoft.org/index.php/jss/article/view/
v068i03/v68i03.pdf.

Cule:2009:LRP

[CGS09] Madeleine Cule, Robert B. Gramacy, and Richard Samworth. LogConcDEAD: An R package for maximum likelihood estimation of a multivariate log-concave density. Journal of Statistical Software, 29(2):??, January 2009. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v29/i02.

Cook:2011:BRB

[CH11] Dianne Cook and Heike Hofmann. Book review: R Graphics (2nd Edition). Journal of Statistical Software, 43(BR-3):??, September 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v43/b03.

Chang:2003:MGP

[Cha03] Chwen-Ming Chang. MinPROMEP: Generation of partially replicated minimal orthogonal main-effect plans using a novel algorithm. Journal of Statistical Software, 8 (8):1–15, ???? 2003. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v08/i08/paper.
pdf;   http://www.jstatsoft.org/v08/i8.

Chausse:2010:CGM

[Cha10] Pierre Chaussé. Computing generalized method of moments and generalized empirical likelihood with R. Journal of Statistical Software, 34(11):??, May 2010. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v34/i11.

Chalmers:2012:MMI

[Cha12] R. Philip Chalmers. mirt: a multidimensional item response theory package for the R environment. Journal of Statistical Software, 48(6):??, May 2012. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v48/i06.

Chalmers:2016:GAN

[Cha16] R. Philip Chalmers. Generating adaptive and non-adaptive test interfaces for multidimensional item response theory applications. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v071i05;
https://www.jstatsoft.org/index.php/jss/article/view/
v071i05/v71i05.pdf.

Cornillon:2017:IBR

[CHML17] Pierre-André Cornillon, Nicolas Hengartner, and Eric Matzner-Løber. Iterative bias reduction multivariate smoothing in R: The ibr package. Journal of Statistical Software, 77 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v077i09;   https://www.jstatsoft.org/
index.php/jss/article/view/v077i09/v77i09.pdf.

Christ:2009:BRB

[Chr09] Aaron Christ. Book review: Mixed Effects Models and Extensions in Ecology with R. Journal of Statistical Software, 32 (BR-1):??, November 2009. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v32/b01.

Christou:2011:BRB

[Chr11] Nicolas Christou. Book review: Statistical Methods in e-Commerce Research. Journal of Statistical Software, 39(BR-1):??, March 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v39/b01.

Curini:2017:NOP

[CI17] Luigi Curini and Stefano M. Iacus. Nash optimal party positions: The nopp R package. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v081i11;   https://www.jstatsoft.org/
index.php/jss/article/view/v081i11/v81i11.pdf.

Culp:2006:ARP

[CJM06] Mark Culp, Kjell Johnson, and George Michailidis. ada: An R package for stochastic boosting. Journal of Statistical Software, 17(2):1–27, September 2006. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v17/i02.

Cameron:2020:CSP

[CKE20] Briana Cameron, Michael J. Kane, and Denise Esserman. Code snippet: preference: an R package for two-stage clinical trial design accounting for patient preference. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v094c02;   https://www.jstatsoft.
org/index.php/jss/article/view/v094c02/v94c02.pdf.

Commandeur:2011:SSS

[CKO11] Jacques J. F. Commandeur, Siem Jan Koopman, and Marius Ooms. Statistical software for state space methods. Journal of Statistical Software, 41(1):??, May 2011. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v41/i01.

Chavent:2012:CRP

[CKSLS12] Marie Chavent, Vanessa Kuentz-Simonet, Benoît Liquet, and Jérôme Saracco. ClustOfVar: An R package for the clustering of variables. Journal of Statistical Software, 50(13):??, September 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v50/i13.

Chiou:2014:FAF

[CKY14] Sy Han Chiou, Sangwook Kang, and Jun Yan. Fitting accelerated failure time models in routine survival analysis with R package aftgee. Journal of Statistical Software, 61(11):??, November 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v61/i11.

Crowther:2013:SSP

[CL13] Michael J. Crowther and Paul C. Lambert. stgenreg: A Stata package for general parametric survival analysis. Journal of Statistical Software, 53(12):??, May 2013. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v53/i12.

Comets:2017:PEN

[CLL17] Emmanuelle Comets, Audrey Lavenu, and Marc Lavielle. Parameter estimation in nonlinear mixed effect models using saemix, an R implementation of the SAEM algorithm. Journal of Statistical Software, 80(??):??, ???? 2017. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v080i03;   https://www.
jstatsoft.org/index.php/jss/article/view/v080i03/v80i03.
pdf.

Croissant:2008:PDE

[CM08] Yves Croissant and Giovanni Millo. Panel data econometrics in R: The plm package. Journal of Statistical Software, 27 (2):??, July 2008. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v27/i02.

Clayden:2011:TMR

[CMS^{+}11] Jonathan D. Clayden, Susana Muñoz Maniega, Amos J. Storkey, Martin D. King, Mark E. Bastin, and Chris A. Clark. TractoR: Magnetic resonance imaging and tractography with R. Journal of Statistical Software, 44(8):??, October 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v44/i08.

Cardinali:2013:CLS

[CN13] Alessandro Cardinali and Guy P. Nason. Costationarity of locally stationary time series using costat. Journal of Statistical Software, 55(1):??, October 2013. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v55/i01.

Catania:2018:DMA

[CN18] Leopoldo Catania and Nima Nonejad. Dynamic model averaging for practitioners in economics and finance: The eDMA package. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v084i11;
https://www.jstatsoft.org/index.php/jss/article/view/
v084i11/v84i11.pdf.

Conn:2019:FFE

[CNLR19] Daniel Conn, Tuck Ngun, Gang Li, and Christina M. Ramirez. Fuzzy forests: Extending random forest feature selection for correlated, high-dimensional data. Journal of Statistical Software, 91(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v091i09;   https://www.jstatsoft.org/
index.php/jss/article/view/v091i09/v91i09.pdf.

Cribari-Neto:2010:BRR

[CNZ10] Francisco Cribari-Neto and Achim Zeileis. Beta regression in R. Journal of Statistical Software, 34(2):??, April 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v34/i02.

Crippa:2016:CSM

[CO16] Alessio Crippa and Nicola Orsini. Code snippet: Multivariate dose-response meta-analysis: The dosresmeta R package. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v072c01;   https://
www.jstatsoft.org/index.php/jss/article/view/v072c01/
v72c01.pdf.

Coeurjolly:2000:SIF

[Coe00] Jean-François Coeurjolly. Simulation and identification of the fractional Brownian motion: a bibliographical and comparative study. Journal of Statistical Software, 5(7):1–53, ???? 2000. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v05/i07;   http://www.jstatsoft.
org/v05/i07/simest.pdf;   http://www.jstatsoft.org/
v05/i07/SimEstFBM;   http://www.jstatsoft.org/v05/
i07/updates.

Contoni:2003:RNS

[Con03] Eva Contoni. Resistant nonparametric smoothing with S-PLUS. Journal of Statistical Software, 8(3):??, ???? 2003. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v08/i03;   http://www.jstatsoft.org/
v08/i03/; http://www.jstatsoft.org/v08/i03/updates;
 http://www.jstatsoft.org/v08/i03/v8.txt.

Cook:1997:CYE

[Coo97] Dianne Cook. Calibrate your eyes to recognize highdimensional shapes from their low-dimensional projections. Journal of Statistical Software, 2(6):1–7, ???? 1997. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v02/i06;   http://www.jstatsoft.org/
v02/i06/paper.pdf;   http://www.jstatsoft.org/v02/
i06/paper.ps.gz;   http://www.jstatsoft.org/v02/i06/
paper/code.html;   http://www.jstatsoft.org/v02/i06/
paper/data;   http://www.jstatsoft.org/v02/i06/paper/
paper.html;   http://www.jstatsoft.org/v02/i06/updates.

Cox:2007:BRB

[Cox07] Nicholas Cox. Book review: The Grammar of Graphics. Journal of Statistical Software, 17(BR-3):1–7, January 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v17/b03.

Carstensen:2011:ULO

[CP11] Bendix Carstensen and Martyn Plummer. Using Lexis objects for multi-state models in R. Journal of Statistical Software, 38(6):??, January 2011. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v38/i06.

Cheng:2012:IDM

[CP12] Ya-Shan Cheng and Chien-Yu Peng. Integrated degradation models in R using iDEMO. Journal of Statistical Software, 49 (2):??, June 2012. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v49/i02.

Chang:2010:CRP

[CQZ^{+}10] Fang Chang, Weiliang Qiu, Ruben H. Zamar, Ross Lazarus, and Xiaogang Wang. clues: An R package for nonparametric clustering based on local shrinking. Journal of Statistical Software, 33(4):??, February 2010. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v33/i04.

Croissant:2020:ERU

[Cro20] Yves Croissant. Estimation of random utility models in R: The mlogit package. Journal of Statistical Software, 95(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v095i11;   https://www.jstatsoft.org/
index.php/jss/article/view/v095i11/v95i11.pdf.

Crainiceanu:2005:BAP

[CRW05] Ciprian M. Crainiceanu, David Ruppert, and Matthew P. Wand. Bayesian analysis for penalized spline regression using WinBUGS. Journal of Statistical Software, 14(14):1– 24, September 29, 2005. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/counter.php?id=
139&url=v14/i14&ct=2;   http://www.jstatsoft.org/counter.
php?id=139&url=v14/i14/v14i14.pdf&ct=1.

Clifford:2012:VPD

[CS12] David Clifford and Glenn Stone. Variable penalty dynamic time warping code for aligning mass spectrometry chromatograms in R. Journal of Statistical Software, 47(8):??, April 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v47/i08.

Calhoun:2018:CMS

[CSNF18] Peter Calhoun, Xiaogang Su, Martha Nunn, and Juanjuan Fan. Constructing multivariate survival trees: The MST package for R. Journal of Statistical Software, 83 (??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v083i12;   https://www.jstatsoft.org/
index.php/jss/article/view/v083i12/v83i12.pdf.

Cook:2015:PEM

[CSY15] Dennis Cook, Zhihua Su, and Yi Yang. envlp: A MAT-LAB toolbox for computing envelope estimators in multivariate analysis. Journal of Statistical Software, 62(8):??, January 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v62/i08.

Culp:2011:SSS

[Cul11] Mark Culp. spa: Semi-supervised semi-parametric graphbased estimation in R. Journal of Statistical Software, 40(10): ??, April 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v40/i10.

Curtis:2010:CSB

[Cur10] S. McKay Curtis. Code snippet: BUGS code for item response theory. Journal of Statistical Software, 36(CS-1):??, August 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v36/c01.

Cho:2009:RLB

[CYK^{+}09] HyungJun Cho, Ami Yu, Sukwoo Kim, Jaewoo Kang, and Seung-Mo Hong. Robust likelihood-based survival modeling with microarray data. Journal of Statistical Software, 29(1): ??, January 2009. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v29/i01.

Dahl:2020:IRS

[Dah20] David B. Dahl. Integration of R and Scala using rscala. Journal of Statistical Software, 92(??):??, ???? 2020. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v092i04;   https://www.
jstatsoft.org/index.php/jss/article/view/v092i04/v92i04.
pdf.

Dastidar:2006:GSP

[Das06] Surajit Ghosh Dastidar. Gompertz: a Scilab program for estimating Gompertz curve using Gauss–Newton method of least squares. Journal of Statistical Software, 15(12):1–12, April 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v15/i12.

Dayton:2001:SBS

[Day01] C. Mitchell Dayton. SUBSET: Best subsets using information criteria. Journal of Statistical Software, 6(2): 1–10, ???? 2001. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v06/i02;   http:
//www.jstatsoft.org/v06/i02/Code.rtf;   http://www.
jstatsoft.org/v06/i02/SUBSET.doc;   http://www.jstatsoft.
org/v06/i02/SUBSET.pdf;   http://www.jstatsoft.org/
v06/i02/updates.

Davies:2013:CEG

[DB13] Tilman M. Davies and David J. Bryant. On circulant embedding for Gaussian random fields in R. Journal of Statistical Software, 55(9):??, November 2013. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v55/i09.

Doran:2007:EMR

[DBBD07] Harold Doran, Douglas Bates, Paul Bliese, and Maritza Dowling. Estimating the multilevel Rasch model: With the lme4 package. Journal of Statistical Software, 20(2):1–18, February 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v20/i02.

DeBoeck:2011:EIR

[DBZ^{+}11] Paul De Boeck, Marjan Bakker, Robert Zwitser, Michel Nivard, Abe Hofman, Francis Tuerlinckx, and Ivailo Partchev. The estimation of item response models with the lmer function from the lme4 package in R. Journal of Statistical Software, 39(12):??, March 2011. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v39/i12.

Dejean:2005:FRP

[DC05] Sébastien Déjean and Serge Cohen. FracSim: An R package to simulate multifractional Lévy motions. Journal of Statistical Software, 14(18):1–19, December 2005. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v14/i18.

Dahl:2009:RAR

[DC09] David B. Dahl and Scott Crawford. RinRuby: Accessing the R interpreter from pure Ruby. Journal of Statistical Software, 29(4):??, January 2009. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v29/i04.

Dong:2020:FRP

[DCS20] Xu Dong, Luis Castro, and Nazrul Shaikh. fastnet: an R package for fast simulation and analysis of largescale social networks. Journal of Statistical Software, 96 (??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v096i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v096i07/v96i07.pdf.

Dray:2007:API

[DD07] Stéphane Dray and Anne-Béatrice Dufour. The ade4 package: Implementing the duality diagram for ecologists. Journal of Statistical Software, 22(4):1–20, September 2007. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v22/i04.

Demirtas:2011:BRB

[Dem11] Hakan Demirtas. Book review: Statistical Simulation: Power Method Polynomials and Other Transformations. Journal of Statistical Software, 43(BR-2):??, September 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v43/b02.

Demirtas:2017:BRB

[Dem17] Hakan Demirtas. Book review: Bayesian Computation with R (2nd Edition). Journal of Statistical Software, 79 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v079b01;   https://www.jstatsoft.org/
index.php/jss/article/view/v079b01/v79b01.pdf.

Demirtas:2018:BRF

[Dem18a] Hakan Demirtas. Book review: Flexible Imputation of Missing Data. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v085b04;
https://www.jstatsoft.org/index.php/jss/article/view/
v085b04/v85b04.pdf.

Demirtas:2018:BRH

[Dem18b] Hakan Demirtas. Book review: Handbook of Fitting Statistical Distributions with R. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v086b02;   https://www.jstatsoft.org/
index.php/jss/article/view/v086b02/v86b02.pdf.

Denwood:2016:RRP

[Den16] Matthew J. Denwood. runjags: An R package providing interface utilities, model templates, parallel computing methods and additional distributions for MCMC models in JAGS. Journal of Statistical Software, 71 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v071i09;   https://www.jstatsoft.org/
index.php/jss/article/view/v071i09/v71i09.pdf.

Derquenne:2010:BRB

[Der10] Christian Derquenne. Book review: Regression Modeling: Methods, Theory, and Computation with SAS. Journal of Statistical Software, 36(BR-5):??, September 2010. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v36/b05.

Drukker:2011:SSM

[DG11] David M. Drukker and Richard B. Gates. State space methods in Stata. Journal of Statistical Software, 41(10):??, May 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v41/i10.

Durrieu:2018:EPE

[DGJ^{+}18] Gilles Durrieu, Ion Grama, Kevin Jaunatre, Quang-Khoai Pham, and Jean-Marie Tricot. extremefit: A package for extreme quantiles. Journal of Statistical Software, 87 (??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v087i12;   https://www.jstatsoft.org/
index.php/jss/article/view/v087i12/v87i12.pdf.

Dutang:2008:ARP

[DGP08] Christophe Dutang, Vincent Goulet, and Mathieu Pigeon. actuar: An R package for actuarial science. Journal of Statistical Software, 25(7):1–37, March 2008. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v25/i07.

Dethlefsen:2005:CPG

[DH05] Claus Dethlefsen and Søren Højsgaard. A common platform for graphical models in R: The gRbase package. Journal of Statistical Software, 14(17):1–12, December 3, 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=142&url=v14/i17&ct=2;   http://www.
jstatsoft.org/counter.php?id=142&url=v14/i17/v14i17.
pdf&ct=1.

Dhand:2010:CSU

[Dha10] Navneet K. Dhand. Code snippet: UniLogistic: a SAS macro for descriptive and univariable logistic regression analyses. Journal of Statistical Software, 35(CS-1):??, July 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v35/c01.

Dupuy:2015:DDT

[DHF15] Delphine Dupuy, Céline Helbert, and Jessica Franco. DiceDesign and DiceEval: Two R packages for design and analysis of computer experiments. Journal of Statistical Software, 65(11):??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v65/i11.

Davies:2011:SAS

[DHM11] Tilman M. Davies, Martin L. Hazelton, and Jonathan C. Marshall. sparr: Analyzing spatial relative risk using fixed and adaptive kernel density estimation in R. Journal of Statistical Software, 39(1):??, March 2011. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v39/i01.

deIaco:2017:CSA

[dI17] Sandra de Iaco. The cgeostat software for analyzing complex-valued random fields. Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v079i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v079i05/v79i05.pdf.

Dinov:2006:SSO

[Din06] Ivo D. Dinov. SOCR: Statistics online computational resource. Journal of Statistical Software, 16(11):1–16, October 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v16/i11.

Draminski:2018:RRP

[DK18] Michal Dramiński and Jacek Koronacki. rmcfs: An R package for Monte Carlo feature selection and interdependency discovery. Journal of Statistical Software, 85 (??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v085i12;   https://www.jstatsoft.org/
index.php/jss/article/view/v085i12/v85i12.pdf.

Dinges:2011:OSA

[DKMT11] Gerlinde Dinges, Alexander Kowarik, Bernhard Meindl, and Matthias Templ. An open source approach for modern teaching methods: The interactive TGUI system. Journal of Statistical Software, 39(7):??, March 2011. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v39/i07.

Dimitrova:2020:CKS

[DKT20] Dimitrina S. Dimitrova, Vladimir K. Kaishev, and Senren Tan. Computing the Kolmogorov–Smirnov distribution when the underlying CDF is purely discrete, mixed, or continuous. Journal of Statistical Software, 95(??):??, ???? 2020. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v095i10;   https://www.
jstatsoft.org/index.php/jss/article/view/v095i10/v95i10.
pdf.

deLeeuw:1997:SSS

[dL97] Jan de Leeuw. Server-side statistics scripting in PHP. Journal of Statistical Software, 2(1):1–13, ???? 1997. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v02/i01;   http://www.jstatsoft.org/
v02/i01/code/;   http://www.jstatsoft.org/v02/i01/
Dstatistics/;   http://www.jstatsoft.org/v02/i01/statistics.
pdf;   http://www.jstatsoft.org/v02/i01/statistics.
ps;   http://www.jstatsoft.org/v02/i01/statistics/
statistics.html;   http://www.jstatsoft.org/v02/i01/
updates.

deLeeuw:2005:BRR

[dL05a] Jan de Leeuw. Book review: An R and S-PLUS Companion to Multivariate Analysis. Journal of Statistical Software, 14(BR-6):1–3, September 2005. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v14/b06.

deLeeuw:2005:BRC

[dL05b] Jan de Leeuw. Book review: Correspondence Analysis and Data Coding with Java and R. Journal of Statistical Software, 14(BR-5):1–2, September 2005. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v14/b05.

deLeeuw:2005:BRM

[dL05c] Jan de Leeuw. Book review: Modern Multidimensional Scaling: Theory and Applications (Second Edition). Journal of Statistical Software, 14(BR-4):1–2, September 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v14/b04.

deLeeuw:2005:AXS

[dL05d] Jan de Leeuw. On abandoning XLISP–STAT. Journal of Statistical Software, 13(7):1–5, ???? 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=121&url=v13/i07&ct=2;   http://www.
jstatsoft.org/counter.php?id=121&url=v13/i07/v13i07.
pdf&ct=1.

deLeeuw:2006:BRB

[dL06] Jan de Leeuw. Book review: Robust Statistical Methods with R. Journal of Statistical Software, 16(BR-2):1–2, July 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v16/b02.

deLeeuw:2007:BRB

[dL07] Jan de Leeuw. Book review: Computational Statistics. Journal of Statistical Software, 19(BR-1):1–3, March 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v19/b01.

deLeeuw:2008:BRB

[dL08] Jan de Leeuw. Book review: Matrix Algebra, A Matrix Handbook for Statisticians. Journal of Statistical Software, 25(BR-3):1–3, May 2008. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v25/b03.

deLeeuw:2009:BRBa

[dL09a] Jan de Leeuw. Book review: Nonlinear Regression with R. Journal of Statistical Software, 29(BR-6):??, January 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v29/b06.

deLeeuw:2009:BRBc

[dL09b] Jan de Leeuw. Book review: R Programming for Bioinformatics. Journal of Statistical Software, 29(BR-8):??, January 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v29/b08.

deLeeuw:2009:BRBb

[dL09c] Jan de Leeuw. Book review: Statistical Methods for Environmental Epidemiology with R. Journal of Statistical Software, 29(BR-7):??, January 2009. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v29/b07.

deLeeuw:2012:BRB

[dL12] Jan de Leeuw. Book review: Modern Fortran: Style and Usage. Journal of Statistical Software, 47(BR-1):??, April 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v47/b01.

Dethlefsen:2006:FSS

[DLC06] Claus Dethlefsen and Søren Lundbye-Christensen. Formulating state space models in R with focus on longitudinal regression models. Journal of Statistical Software, 16(1):1–15, April 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v16/i01.

deLeeuw:2009:IOR

[dLHM09] Jan de Leeuw, Kurt Hornik, and Patrick Mair. Isotone optimization in R: Pool-adjacent-violators algorithm (PAVA) and active set methods. Journal of Statistical Software, 32(5):??, October 2009. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v32/i05.

deLeeuw:2007:ISV

[dLM07] Jan de Leeuw and Patrick Mair. An introduction to the special volume on “Psychometrics in R”. Journal of Statistical Software, 20(1):1–5, May 2007. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v20/i01.

deLeeuw:2009:GMO

[dLM09a] Jan de Leeuw and Patrick Mair. Gifi methods for optimal scaling in R: The package homals. Journal of Statistical Software, 31(4):??, August 2009. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v31/i04.

deLeeuw:2009:MSU

[dLM09b] Jan de Leeuw and Patrick Mair. Multidimensional scaling using majorization: SMACOF in R. Journal of Statistical Software, 31(3):??, August 2009. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v31/i03.

deLeeuw:2009:SCC

[dLM09c] Jan de Leeuw and Patrick Mair. Simple and canonical correspondence analysis using the R package anacor. Journal of Statistical Software, 31(5):??, August 2009. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v31/i05.

Daouia:2017:NPN

[DLN17] Abdelaati Daouia, Thibault Laurent, and Hohsuk Noh. npbr: A package for nonparametric boundary regression in R. Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v079i09;   https://www.
jstatsoft.org/index.php/jss/article/view/v079i09/v79i09.
pdf.

Donoghoe:2018:LRP

[DM18] Mark W. Donoghoe and Ian C. Marschner. logbin: An R package for relative risk regression using the log-binomial model. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v086i09;
https://www.jstatsoft.org/index.php/jss/article/view/
v086i09/v86i09.pdf.

Duarte:2019:SRP

[DM19] João Daniel N. Duarte and Vinícius D. Mayrink. slfm: an R package to evaluate coherent patterns in microarray data via factor analysis. Journal of Statistical Software, 90(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v090i09;   https://www.jstatsoft.org/
index.php/jss/article/view/v090i09/v90i09.pdf.

DEnza:2018:CSI

[DMB18] Alfonso Iodice D’Enza, Angelos Markos, and Davide Buttarazzi. Code snippet: The idm package: Incremental decomposition methods in R. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v086c04;   https://www.jstatsoft.org/
index.php/jss/article/view/v086c04/v86c04.pdf.

Delignette-Muller:2015:PFR

[DMD15] Marie Laure Delignette-Muller and Christophe Dutang. fitdistrplus: An R package for fitting distributions. Journal of Statistical Software, 64(4):??, March 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v64/i04.

Doss:1999:DDC

[DN99] Hani J. Doss and Balasubramanian Narasimhan. Dynamic display of changing posterior in Bayesian survival analysis: The software. Journal of Statistical Software, 4 (3):1–72, ???? 1999. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v04/i03;   http:
//www.jstatsoft.org/v04/i03/bsa.pdf;   http://www.
jstatsoft.org/v04/i03/bsa.tgz;   http://www.jstatsoft.
org/v04/i03/ddg.pdf;   http://www.jstatsoft.org/v04/
i03/example/;   http://www.jstatsoft.org/v04/i03/html/
bsa.html;   http://www.jstatsoft.org/v04/i03/lisp/;
 http://www.jstatsoft.org/v04/i03/updates.

Demirtas:2017:BRI

[DN17] Hakan Demirtas and Rachel Nordgren. Book review: Introduction to Scientific Programming and Simulation Using R (2nd Edition). Journal of Statistical Software, 78 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v078b04;   https://www.jstatsoft.org/
index.php/jss/article/view/v078b04/v78b04.pdf.

Doan:2011:SSM

[Doa11] Thomas Doan. State space methods in RATS. Journal of Statistical Software, 41(9):??, May 2011. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v41/i09.

Driver:2017:CTS

[DOV17] Charles C. Driver, Johan H. L. Oud, and Manuel C. Voelkle. Continuous time structural equation modeling with R package ctsem. Journal of Statistical Software, 77(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v077i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v077i05/v77i05.pdf.

Downie:2016:BRU

[Dow16] Tim Downie. Book review: Using the R Commander: A Point-and-Click Interface for R. Journal of Statistical Software, 75(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v075b03;   https://www.jstatsoft.org/
index.php/jss/article/view/v075b03/v75b03.pdf.

Downie:2017:BRM

[Dow17] Tim Downie. Book review: Modern Data Science with R. Journal of Statistical Software, 80(??):??, ???? 2017. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v080b02;   https://www.
jstatsoft.org/index.php/jss/article/view/v080b02/v80b02.
pdf.

Downie:2018:BRV

[Dow18] Tim Downie. Book review: Visualizing Baseball. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v083b03;   https://www.
jstatsoft.org/index.php/jss/article/view/v083b03/v83b03.
pdf.

Downie:2019:BRA

[Dow19] Tim Downie. Book review: Analyzing Baseball Data with R (2nd Edition). Journal of Statistical Software, 90 (??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v090b01;   https://www.jstatsoft.org/
index.php/jss/article/view/v090b01/v90b01.pdf.

Dupont:2003:DDS

[DP03] William D. Dupont and W. Dale Plummer, Jr. Density distribution sunflower plots. Journal of Statistical Software, 8(3):1–5, ???? 2003. CODEN JSSOBK. ISSN

1548-7660.  URL http://www.jstatsoft.org/v08/i03/

SunflowerPlots.pdf;   http://www.jstatsoft.org/v08/
i3.

DeBoeck:2012:CSI

[DP12] Paul De Boeck and Ivailo Partchev. Code snippet: IRTrees: Tree-based item response models of the GLMM family. Journal of Statistical Software, 48(CS-1):??, May 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v48/c01.

Dunkler:2018:WCR

[DPSH18] Daniela Dunkler, Meinhard Ploner, Michael Schemper, and Georg Heinze. Weighted Cox regression using the R package coxphw. Journal of Statistical Software, 84 (??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v084i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v084i02/v84i02.pdf.

Dumbgen:2011:PLC

[DR11] Lutz Dümbgen and Kaspar Rufibach. logcondens: Computations related to univariate log-concave density estimation. Journal of Statistical Software, 39(6):??, March 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v39/i06.

Deldossi:2020:RPO

[DR20] Laura Deldossi and Marta Nai Ruscone. R package OBsMD for follow-up designs in an objective Bayesian framework. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v094i02;   https://www.
jstatsoft.org/index.php/jss/article/view/v094i02/v94i02.
pdf.

Quintela-del-Rio:2012:NKD

[dREP12] Alejandro Quintela del Río and Graciela Estévez-Pérez. Nonparametric kernel distribution function estimation with kerdiest: An R package for bandwidth choice and applications. Journal of Statistical Software, 50(8):??, August 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v50/i08.

Dryver:2009:CSE

[Dry09] Arthur Dryver. Code snippet: The enhancement of teaching materials for applied statistics courses by combining random number generation and portable document format files via LaTEX. Journal of Statistical Software, 31(CS-3):1–9, September 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v31/c03.

daSilva:2011:CPI

[dS11] Adelino R. Ferreira da Silva. cudaBayesreg: Parallel implementation of a Bayesian multilevel model for fMRI data analysis. Journal of Statistical Software, 44(4):??, October 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v44/i04.

Dunsmuir:2015:PGP

[DS15] William T. M. Dunsmuir and David J. Scott. The glarma package for observation-driven time series regression of counts. Journal of Statistical Software, 67(7):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067i07;
http://www.jstatsoft.org/index.php/jss/article/view/
v067i07/v67i07.pdf.

DeBrabanter:2013:NRP

[DSD13] Kris De Brabanter, Johan A. K. Suykens, and Bart De Moor. Nonparametric regression via StatLSSVM. Journal of Statistical Software, 55(2):??, October 2013. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v55/i02.

deSouza:2018:CSR

[dSdSCLC18] Wélliton de Souza, Benilton de Sá Carvalho, and Iscia Lopes-Cendes. Code snippet: Rqc: A bioconductor package for quality control of high-throughput sequencing data. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v087c02;   https://www.
jstatsoft.org/index.php/jss/article/view/v087c02/v87c02.
pdf.

Dunkler:2016:GPJ

[DSH16] Daniela Dunkler, Willi Sauerbrei, and Georg Heinze. Global, parameterwise and joint shrinkage factor estimation. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v069i08;   https://www.
jstatsoft.org/index.php/jss/article/view/v069i08/v69i08.
pdf.

daSilva-Junior:2014:CSP

[dSJdSF14] Antonio Hermes M. da Silva-Júnior, Damião Nóbrega da Silva, and Silvia L. P. Ferrari. Code snippet: mdscore: An R package to compute improved score tests in generalized linear models. Journal of Statistical Software, 61(CS-2):??, October 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v61/c02.

Lopez-de-Ullibarri:2013:SRP

[dUJ13] Ignacio López de Ullibarri and M. Amalia Jácome. survPresmooth: An R package for presmoothed estimation in survival analysis. Journal of Statistical Software, 54(11):??, September 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v54/i11.

Dunn:1999:GUI

[Dun99] Peter Dunn. A graphical user interface to generalized linear models in MATLAB. Journal of Statistical Software, 4(4):1–25, ???? 1999. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v04/i04;   http:/
/www.jstatsoft.org/v04/i04/guiglm.pdf;   http://www.
jstatsoft.org/v04/i04/updates.

Duong:2007:KKD

[Duo07] Tarn Duong. ks: Kernel density estimation and kernel discriminant analysis for multivariate data in R. Journal of Statistical Software, 21(7):1–16, October 2007. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v21/i07.

Du:2020:SRP

[DV20] Yu Du and Ravi Varadhan. SQUAREM: an R package for off-the-shelf acceleration of EM, MM and other EM-like monotone algorithms. Journal of Statistical Software, 92 (??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v092i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v092i07/v92i07.pdf.

Royuela-del-Val:2017:LFP

[dVSWAL17] Javier Royuela del Val, Federico Simmross-Wattenberg, and Carlos Alberola-López. libstable: Fast, parallel, and high-precision computation of α-stable distributions in R, C/C++, and MATLAB. Journal of Statistical Software, 78(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v078i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v078i01/v78i01.pdf.

Dong:2017:BRP

[DW17] Chen Dong and Michel Wedel. BANOVA: An R package for hierarchical Bayesian ANOVA. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v081i09;   https://www.jstatsoft.org/
index.php/jss/article/view/v081i09/v81i09.pdf.

deWreede:2011:MRP

[dWFP11] Liesbeth C. de Wreede, Marta Fiocco, and Hein Putter. mstate: An R package for the analysis of competing risks and multi-state models. Journal of Statistical Software, 38(7): ??, January 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v38/i07.

Esty:2003:BPP

[EB03] Warren W. Esty and Jeffrey D. Banfield. The box-percentile plot. Journal of Statistical Software, 8(17):1–14, ???? 2003. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v08/i17;   http://www.jstatsoft.
org/v08/i17/BoxPercentilePlot.pdf.

Erni:2013:MRP

[EBO^{+}13] Birgit Erni, Bo T. Bonnevie, Hans-Dieter Oschadleus, Res Altwegg, and Les G. Underhill. moult: An R package to analyze moult in birds. Journal of Statistical Software, 52(8): ??, February 2013. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v52/i08.

Elam:2006:CPC

[EC06] Matthew E. Elam and Kenneth E. Case. A computer program to calculate two-stage short-run control chart factors for (X, MR) charts. Journal of Statistical Software, 15(11):1–13, April 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v15/i11.

Erickson:2016:CSS

[EC16] Stephen W. Erickson and Joshua C. Callaway. Code snippet: SNPMClust: Bivariate Gaussian genotype clustering and calling for Illumina microarrays. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v071c02;   https://www.jstatsoft.org/
index.php/jss/article/view/v071c02/v71c02.pdf.

Elam:2004:SSS

[ECK04] Matthew E. Elam, Kenneth E. Case, and Jong-Min Kim. Second stage short run (X̄, v_{c}) and (X̄, s_{c}) control charts. Journal of Statistical Software, 9(7):1–31, ???? 2004. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=88&url=v09/i07/paper.pdf&ct=1.

Epskamp:2012:QNV

[ECW^{+}12] Sacha Epskamp, Angélique O. J. Cramer, Lourens J. Waldorp, Verena D. Schmittmann, and Denny Borsboom. qgraph: Network visualizations of relationships in psychometric data. Journal of Statistical Software, 48(4):??, May 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v48/i04.

Eddelbuettel:2009:BRBb

[Edd09a] Dirk Eddelbuettel. Book review: Analysis of Integrated and Cointegrated Time Series with R (2nd Edition). Journal of Statistical Software, 30(BR-5):??, April 2009. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v30/b05.

Eddelbuettel:2009:BRBa

[Edd09b] Dirk Eddelbuettel. Book review: Applied Econometrics with R. Journal of Statistical Software, 29(BR-14):??, February 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v29/b14.

Eddelbuettel:2009:BRBc

[Edd09c] Dirk Eddelbuettel. Book review: Computational Statistics: An Introduction to R. Journal of Statistical Software, 32(BR-2):

??, December 2009. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v32/b02.

Eddelbuettel:2011:BRB

[Edd11] Dirk Eddelbuettel. Book review: R in a Nutshell. Journal of Statistical Software, 40(BR-1):??, April 2011. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v40/b01.

Eddelbuettel:2012:BRBb

[Edd12a] Dirk Eddelbuettel. Book review: R for Dummies. Journal of Statistical Software, 50(BR-1):??, August 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v50/b01.

Eddelbuettel:2012:BRBa

[Edd12b] Dirk Eddelbuettel. Book review: R in Action. Journal of Statistical Software, 46(BR-2):??, February 2012. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v46/b02.

Edlund:2004:CMS

[Edl04] Ove Edlund. CMregr — a Matlab software package for finding CM-estimates for regression. Journal of Statistical Software, 10(3):1–11, ???? 2004. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/counter.php?id=
92&url=v10/i03/CMregr_rev2.pdf&ct=1.

Erdman:2007:BRP

[EE07] Chandra Erdman and John W. Emerson. bcp: An R package for performing a Bayesian analysis of change point problems. Journal of Statistical Software, 23(3):1–13, December 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v23/i03.

Eddelbuettel:2011:RSR

[EF11] Dirk Eddelbuettel and Romain François. Rcpp: Seamless R and C++ integration. Journal of Statistical Software, 40(8): 1–18, April 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v40/i08.

Ekstrom:2010:BRB

[Eks10] Joakim Ekström. Book review: Computation of Multivariate Normal and t Probabilities. Journal of Statistical Software, 33 (BR-2):??, February 2010. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v33/b02.

Ekstrom:2013:BRB

[Eks13] Joakim Ekström. Book review: The BUGS Book: A Practical Introduction to Bayesian Analysis. Journal of Statistical Software, 52(BR-1):??, February 2013. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v52/b01.

Eugster:2009:SMH

[EL09] Manuel J. A. Eugster and Friedrich Leisch. From Spider-Man to hero — archetypal analysis in R. Journal of Statistical Software, 30(8):??, April 2009. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v30/i08.

Ellis:2009:BRB

[Ell09] Byron Ellis. Book review: Software for Data Analysis. Journal of Statistical Software, 29(BR-4):??, January 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v29/b04.

Escobar:2020:NCA

[EMU20] Modesto Escobar and Luis Martinez-Uribe. Network coincidence analysis: The netCoin R package. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v093i11;   https://www.jstatsoft.
org/index.php/jss/article/view/v093i11/v93i11.pdf.

Eckley:2011:LIL

[EN11] Idris A. Eckley and Guy P. Nason. LS2W: Implementing the locally stationary 2D wavelet process approach in R. Journal of Statistical Software, 43(3):??, July 2011. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v43/i03.

Emily:2020:GBM

[ESKHB20] Mathieu Emily, Nicolas Sounac, Florian Kroell, and Magalie Houée-Bigot. Gene-based methods to detect gene-gene interaction in R: The GeneGeneInteR package. Journal of Statistical Software, 95(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v095i12;   https://www.jstatsoft.
org/index.php/jss/article/view/v095i12/v95i12.pdf.

Esmailzadeh:2014:CSM

[Esm14] Nabaz Esmailzadeh. Code snippet: A MATLAB package for computing two-level search design performance criteria. Journal of Statistical Software, 56(CS-1):??, January 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v56/c01.

Eddelbuettel:2016:REC

[ESO16] Dirk Eddelbuettel, Murray Stokely, and Jeroen Ooms. RProtoBuf: Efficient cross-language data serialization in R. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v071i02;   https://
www.jstatsoft.org/index.php/jss/article/view/v071i02/
v71i02.pdf.

Evans:2011:BRB

[Eva11] Gary Evans. Book review: Exploratory Multivariate Analysis by Example Using R. Journal of Statistical Software, 40(BR-2):??, April 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v40/b02.

Evans:2014:BRA

[Eva14] Gary Evans. Book review: Analyzing Spatial Models of Choice and Judgment with R. Journal of Statistical Software, 58(BR-1):??, July 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v58/b01.

Erichson:2019:RMD

[EVBK19] N. Benjamin Erichson, Sergey Voronin, Steven L. Brunton, and J. Nathan Kutz. Randomized matrix decompositions using R. Journal of Statistical Software, 89 (??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v089i11;   https://www.jstatsoft.org/
index.php/jss/article/view/v089i11/v89i11.pdf.

Ebert:2020:CES

[EWMR20] Anthony Ebert, Paul Wu, Kerrie Mengersen, and Fabrizio Ruggeri. Computationally efficient simulation of queues: The R package queuecomputer. Journal of Statistical Software, 95(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v095i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v095i05/v95i05.pdf.

Fay:2018:BRT

[Fay18] Colin Fay. Book review: Text Mining with R: A Tidy Approach. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v083b01;
https://www.jstatsoft.org/index.php/jss/article/view/
v083b01/v83b01.pdf.

Finley:2007:SRP

[FBC07] Andrew O. Finley, Sudipto Banerjee, and Bradley P. Carlin. spBayes: An R package for univariate and multivariate hierarchical point-referenced spatial models. Journal of Statistical Software, 19(4):1–24, April 2007. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v19/i04.

Febrero-Bande:2012:SCF

[FBdlF12] Manuel Febrero-Bande and Manuel Oviedo de la Fuente. Statistical computing in functional data analysis: The R package fda.usc. Journal of Statistical Software, 51(4):??, October 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v51/i04.

Fresno:2014:PLL

[FBF14] Cristóbal Fresno, Mónica G. Balzarini, and Elmer A. Fernández. lmdme: Linear models on designed multivariate experiments in R. Journal of Statistical Software, 56(7):??, January 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v56/i07.

Finley:2015:PLU

[FBG15] Andrew O. Finley, Sudipto Banerjee, and Alan E. Gelfand. spBayes for large univariate and multivariate point-referenced spatio-temporal data models. Journal of Statistical Software, 63(13):??, February 2015. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v63/i13.

Fiske:2011:URP

[FC11] Ian Fiske and Richard Chandler. unmarked: An R package for fitting hierarchical models of wildlife occurrence and abundance. Journal of Statistical Software, 43(10):??, August 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v43/i10.

Fox:2012:RSP

[FC12] John Fox and Marilia Sá Carvalho. The RcmdrPlugin.survival package: Extending the R Commander interface to survival analysis. Journal of Statistical Software, 49(7):??, June 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v49/i07.

Ferguson:2012:MRP

[FDB12] Nicole Ferguson, Somnath Datta, and Guy Brock. msSurv: An R package for nonparametric estimation of multistate models. Journal of Statistical Software, 50(14):??, September 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v50/i14.

Fernandez-Duran:2016:CRP

[FDGD16] Juan José Fernández-Durán and María Mercedes Gregorio-Domínguez. CircNNTSR: An R package for the statistical analysis of circular, multivariate circular, and spherical data using nonnegative trigonometric sums. Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v070i06;   https://www.jstatsoft.
org/index.php/jss/article/view/v070i06/v70i06.pdf.

Fellows:2012:DDA

[Fel12] Ian Fellows. Deducer: a data analysis GUI for R. Journal of Statistical Software, 49(8):??, June 2012. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v49/i08.

Ferl:2011:DRP

[Fer11] Gregory Z. Ferl. DATforDCEMRI: An R package for deconvolution analysis and visualization of DCE-MRI data. Journal of Statistical Software, 44(3):??, October 2011. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v44/i03.

Fox:2007:MRR

[FEvdL07] Jean-Paul Fox, Rinke Klein Entink, and Wim van der Linden. Modeling of responses and response times with the package cirt. Journal of Statistical Software, 20(7):1–14, February 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v20/i07.

Finazzi:2014:PDS

[FF14] Francesco Finazzi and Alessandro Fassò. D-STEM: A software for the analysis and mapping of environmental space-time variables. Journal of Statistical Software, 62(6):??, December 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v62/i06.

Fritz:2012:TRP

[FGEMI12] Heinrich Fritz, Luis A. García-Escudero, and Agustín Mayo-Iscar. tclust: An R package for a trimming approach to cluster analysis. Journal of Statistical Software, 47(12):??, May 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v47/i12.

Fox:2009:EDR

[FH09] John Fox and Jangman Hong. Effect displays in R for multinomial and proportional-odds logit models: Extensions to the effects package. Journal of Statistical Software, 32(1):??, October 2009. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v32/i01.

Ferstl:2010:ZCY

[FH10] Robert Ferstl and Josef Hayden. Zero-coupon yield curve estimation with the package termstrc. Journal of Statistical Software, 36(1):??, August 2010. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v36/i01.

Ferwerda:2017:KBR

[FHH17] Jeremy Ferwerda, Jens Hainmueller, and Chad J. Hazlett. Kernel-based regularized least squares in R (KRLS) and Stata (krls). Journal of Statistical Software, 79 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v079i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v079i03/v79i03.pdf.

Feinerer:2008:TMI

[FHM08] Ingo Feinerer, Kurt Hornik, and David Meyer. Text mining infrastructure in R. Journal of Statistical Software, 25(5):1– 54, March 2008. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v25/i05.

Friedman:2010:RPG

[FHT10] Jerome Friedman, Trevor Hastie, and Rob Tibshirani. Regularization paths for generalized linear models via coordinate descent. Journal of Statistical Software, 33(1):??, February 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v33/i01.

Fieberg:2012:EPA

[Fie12] John R. Fieberg. Estimating population abundance using sightability models: R SightabilityModel package. Journal of Statistical Software, 51(9):??, November 2012. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v51/i09.

Firth:2000:QVX

[Fir00] David Firth. Quasi-variances in Xlisp–Stat and on the Web. Journal of Statistical Software, 5(4):1–13, ???? 2000. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v05/i04;   http://www.jstatsoft.
org/v05/i04/code/;   http://www.jstatsoft.org/v05/
i04/QV_man.pdf;   http://www.jstatsoft.org/v05/i04/
updates.

Firth:2003:CFP

[Fir03] David Firth. CGIwithR: Facilities for processing Web forms using R. Journal of Statistical Software, 8(10):1–8, ???? 2003. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v08/i10;   http://www.jstatsoft.
org/v08/i10/CGIwithR-overview.pdf.

Firth:2005:BTM

[Fir05] David Firth. Bradley–Terry models in R. Journal of Statistical Software, 12(1):1–12, ???? 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=107&url=v12/i01&ct=2;   http://www.
jstatsoft.org/counter.php?id=107&url=v12/i01/v12i01.
pdf&ct=1.

Frick:2017:TIR

[FK17] Hannah Frick and Ioannis Kosmidis. trackeR: Infrastructure for running and cycling data from GPS-enabled tracking devices in R. Journal of Statistical Software, 82 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v082i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v082i07/v82i07.pdf.

Friedrich:2017:CSG

[FKP17] Sarah Friedrich, Frank Konietschke, and Markus Pauly. Code snippet: GFD: An R package for the analysis of general factorial designs. Journal of Statistical Software, 79 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v079c01;   https://www.jstatsoft.org/
index.php/jss/article/view/v079c01/v79c01.pdf.

Fox:2016:RJS

[FL16] John Fox and Allison Leanage. R and the Journal of Statistical Software. Journal of Statistical Software, 73 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v073i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v073i02/v73i02.pdf.

Freeman:2008:PRP

[FM08] Elizabeth A. Freeman and Gretchen Moisen. PresenceAbsence: An R package for presence absence analysis. Journal of Statistical Software, 23(11):1–31, January 2008. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v23/i11.

Foss:2018:KCM

[FM18] Alexander H. Foss and Marianthi Markatou. kamila: Clustering mixed-type data in R and Hadoop. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v083i13;   https://www.jstatsoft.
org/index.php/jss/article/view/v083i13/v83i13.pdf.

Fischer:2020:COM

[FMM^{+}20] Daniel Fischer, Karl Mosler, Jyrki Möttönen, Klaus Nordhausen, Oleksii Pokotylo, and Daniel Vogel. Computing the Oja median in R: The package OjaNP. Journal of Statistical Software, 92(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v092i08;   https://www.jstatsoft.
org/index.php/jss/article/view/v092i08/v92i08.pdf.

Fu:2020:CRP

[FNB20] Anqi Fu, Balasubramanian Narasimhan, and Stephen Boyd. CVXR: an R package for disciplined convex optimization. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v094i14;   https://www.
jstatsoft.org/index.php/jss/article/view/v094i14/v94i14.
pdf.

Fischer:2015:MWT

[FO15] Daniel Fischer and Hannu Oja. Mann–Whitney type tests for microarray experiments: The R package gMWT. Journal of Statistical Software, 65(9):??, June 2015. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v65/i09.

Fokkema:2020:FPR

[Fok20] Marjolein Fokkema. Fitting prediction rule ensembles with R package pre. Journal of Statistical Software, 92 (??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v092i12;   https://www.jstatsoft.org/
index.php/jss/article/view/v092i12/v92i12.pdf.

Fotouhi:2003:CEP

[Fot03] Ali Reza Fotouhi. Comparisons of estimation procedures for nonlinear multilevel models. Journal of Statistical Software, 8(9):1–39, ???? 2003. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v08/i09/Paper.
pdf;   http://www.jstatsoft.org/v08/i9.

Fox:2003:EDR

[Fox03] John Fox. Effect displays in R for generalised linear models. Journal of Statistical Software, 8(15):1–27, ???? 2003. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v08/i15;   http://www.jstatsoft.
org/v08/i15/effect-displays-revised.pdf.

Fox:2005:BRB

[Fox05a] John Fox. Book review: The Basics of S-Plus, Fourth Edition. Journal of Statistical Software, 14(BR-8):1–3, December 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v14/b08.

Fox:2005:RCB

[Fox05b] John Fox. The R Commander: a basic-statistics graphical user interface to R. Journal of Statistical Software, 14(9):1–42, August 19, 2005. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/counter.php?id=134&url=v14/i09&
ct=2;   http://www.jstatsoft.org/counter.php?id=134&
url=v14/i09/v14i09.pdf&ct=1.

Fox:2007:MIM

[Fox07] Jean-Paul Fox. Multilevel IRT modeling in practice with the package mlirt. Journal of Statistical Software, 20(5):1–16, February 2007. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v20/i05.

Fraley:2007:MBM

[FR07] Chris Fraley and Adrian Raftery. Model-based methods of classification: Using the mclust software in chemometrics. Journal of Statistical Software, 18(6):1–13, January 2007. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v18/i06.

Fortmann-Roe:2015:CCR

[FR15] Scott Fortmann-Roe. Consistent and clear reporting of results from diverse modeling techniques: The A3 method. Journal of Statistical Software, 66(7):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066i07;
http://www.jstatsoft.org/index.php/jss/article/view/
v066i07/v66i07.pdf.

Friendly:2006:DEP

[Fri06] Michael Friendly. Data ellipses, HE plots and reduced-rank displays for multivariate linear models: SAS software and examples. Journal of Statistical Software, 17(6):1–43, November 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v17/i06.

Friendly:2010:PRM

[Fri10] Michael Friendly. HE plots for repeated measures designs. Journal of Statistical Software, 37(4):??, November 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v37/i04.

Fricker:2012:BRB

[Fri12] Ronald D. Fricker. Book review: Analysis of Questionnaire Data with R. Journal of Statistical Software, 46(BR-1):??, January 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v46/b01.

Fay:2010:EAW

[FS10a] Michael P. Fay and Pamela A. Shaw. Exact and asymptotic weighted logrank tests for interval censored data: The interval R package. Journal of Statistical Software, 36(2): ??, August 2010. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v36/i02.

Furrer:2010:SSM

[FS10b] Reinhard Furrer and Stephan R. Sain. spam: a sparse matrix R package with emphasis on MCMC methods for Gaussian Markov random fields. Journal of Statistical Software, 36(10): ??, September 2010. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v36/i10.

Fijorek:2012:CSS

[FS12] Kamil Fijorek and Andrzej Sokolowski. Code snippet: Separation-resistant and bias-reduced logistic regression: STA-TISTICA macro. Journal of Statistical Software, 47(CS-2):??, April 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v47/c02.

Fontana:2013:MSM

[FS13] Roberto Fontana and Sabrina Sampò. Minimum-size mixedlevel orthogonal fractional factorial designs generation: A SASbased algorithm. Journal of Statistical Software, 53(10):??, May 2013. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v53/i10.

Francom:2020:BRP

[FS20] Devin Francom and Bruno Sansó. BASS: an R package for fitting and performing sensitivity analysis of Bayesian adaptive spline surfaces. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v094i08;   https://www.jstatsoft.org/
index.php/jss/article/view/v094i08/v94i08.pdf.

Frick:2012:FRM

[FSLZ12] Hannah Frick, Carolin Strobl, Friedrich Leisch, and Achim Zeileis. Flexible Rasch mixture models with package psychomix. Journal of Statistical Software, 48(7):??, May 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v48/i07.

Feng:2008:CDI

[FT08] Dai Feng and Luke Tierney. Computing and displaying isosurfaces in R. Journal of Statistical Software, 28(1):??, September 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v28/i01.

Feng:2011:MPM

[FT11] Dai Feng and Luke Tierney. mritc: a package for MRI tissue classification. Journal of Statistical Software, 44(7):??, October 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v44/i07.

Fujino:2017:VGW

[Fuj17] Tomokazu Fujino. vdmR: Generating Web-based visual data mining tools with R. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v082i06;   https://www.jstatsoft.org/
index.php/jss/article/view/v082i06/v82i06.pdf.

Fox:2018:VFL

[FW18] John Fox and Sanford Weisberg. Visualizing fit and lack of fit in complex regression models with predictor effect plots and partial residuals. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v087i09;   https://www.jstatsoft.org/
index.php/jss/article/view/v087i09/v87i09.pdf.

Gandrud:2015:SRP

[Gan15] Christopher Gandrud. simPH: An R package for illustrating estimates from Cox proportional hazard models including for interactive and nonlinear effects. Journal of Statistical Software, 65(3):??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v65/i03.

Gasparrini:2011:DLL

[Gas11] Antonio Gasparrini. Distributed lag linear and non-linear models in R: The package dlnm. Journal of Statistical Software, 43 (8):??, July 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v43/i08.

Genolini:2015:KKR

[GASA15] Christophe Genolini, Xavier Alacoque, Mariane Sentenac, and Catherine Arnaud. kml and kml3d: R packages to cluster longitudinal data. Journal of Statistical Software, 65(4):??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v65/i04.

Guidoum:2020:PPM

[GB20] Arsalane Chouaib Guidoum and Kamal Boukhetala. Performing parallel Monte Carlo and moment equations methods for Itô and Stratonovich stochastic differential systems: R package Sim.DiffProc. Journal of Statistical Software, 96(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v096i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v096i02/v96i02.pdf.

Graumann:2018:CSM

[GC18] Johannes Graumann and Richard Cotton. Code snippet: multipanelfigure: Simple assembly of multiple plots and images into a compound figure. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/

article/view/v084c03;   https://www.jstatsoft.org/
index.php/jss/article/view/v084c03/v84c03.pdf.

Goncalves:2012:RPB

[GCA12] M. Helena Gonçalves, M. Salomé Cabral, and Adelchi Azzalini. The R package bild for the analysis of binary longitudinal data. Journal of Statistical Software, 46(9):??, February 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v46/i09.

Grimes:2021:SRP

[GD21] Tyler Grimes and Somnath Datta. SeqNet: An R package for generating gene-gene networks and simulating RNA-seq data. Journal of Statistical Software, 98 (??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v098i12;   https://www.jstatsoft.org/
index.php/jss/article/view/v098i12/v98i12.pdf.

Gonzalez:2008:CRP

[GDMB08] Ignacio González, Sébastien Déjean, Pascal G. P. Martin, and Alain Baccini. CCA: An R package to extend canonical correlation analysis. Journal of Statistical Software, 23(12):1–14, January 2008. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v23/i12.

Gentleman:2006:BRB

[Gen06] Robert Gentleman. Book review: Developing Statistical Software in FORTRAN 95. Journal of Statistical Software, 17 (BR-2):1–2, December 2006. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v17/b02.

Geraci:2014:LQM

[Ger14] Marco Geraci. Linear quantile mixed models: The lqmm package for Laplace quantile regression. Journal of Statistical Software, 57(13):??, May 2014. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v57/i13.

Gerber:2015:CSP

[GF15] Florian Gerber and Reinhard Furrer. Code snippet: Pitfalls in the implementation of Bayesian hierarchical modeling of areal count data: An illustration using BYM and Leroux models. Journal of Statistical Software, 63(CS-1):??, February 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v63/c01.

Gopal:2012:BRP

[GFC12] Vikneswaran Gopal, Claudio Fuentes, and George Casella. bayesclust: An R package for testing and searching for significant clusters. Journal of Statistical Software, 47(14):??, May 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v47/i14.

Geitner:2019:CIT

[GFPB19] Robert Geitner, Robby Fritzsch, Jürgen Popp, and Thomas W. Bocklitz. corr2D: Implementation of two-dimensional correlation analysis in R. Journal of Statistical Software, 90(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v090i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v090i03/v90i03.pdf.

Gonzalez-Fernandez:2014:PCR

[GFS14] Yasser Gonzalez-Fernandez and Marta Soto. copulaedas: An R package for estimation of distribution algorithms based on copulas. Journal of Statistical Software, 58(9):??, June 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v58/i09.

Gerber:2016:GRP

[GG16] Florian Gerber and Thomas Gsponer. gsbDesign: An R package for evaluating the operating characteristics of a group sequential Bayesian design. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v069i11;   https://www.jstatsoft.org/
index.php/jss/article/view/v069i11/v69i11.pdf.

Gaudart:2015:PRP

[GGC^{+}15] Jean Gaudart, Nathalie Graffeo, Drissa Coulibaly, Guillaume Barbet, Stanislas Rebaudet, Nadine Dessay, Ogobara K. Doumbo, and Roch Giorgi. SPODT: An R package to perform spatial partitioning. Journal of Statistical Software, 63(16): ??, February 2015. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v63/i16.

Glaab:2010:VRP

[GGK10] Enrico Glaab, Jonathan M. Garibaldi, and Natalio Krasnogor. vrmlgen: An R package for 3D data visualization on the Web. Journal of Statistical Software, 36(8):??, August 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v36/i08.

Grun:2011:PTR

[GH11] Bettina Grün and Kurt Hornik. topicmodels: An R package for fitting topic models. Journal of Statistical Software, 40 (13):??, May 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v40/i13.

Goodreau:2008:ST

[GHH^{+}08] Steven M. Goodreau, Mark S. Handcock, David R. Hunter, Carter T. Butts, and Martina Morris. A statnet tutorial. Journal of Statistical Software, 24(9):1–26, May 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v24/i09.

Gorecki:2020:HAC

[GHH20] Jan Górecki, Marius Hofert, and Martin Holena. Hierarchical Archimedean copulas for MATLAB and Octave: The HACopula toolbox. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v093i10;   https://www.jstatsoft.org/
index.php/jss/article/view/v093i10/v93i10.pdf.

Garcia-Hernandez:2018:JSM

[GHR18] Alberto Garcia-Hernandez and Dimitris Rizopoulos. %JM: A SAS macro to fit jointly generalized mixed models for longitudinal data and time-to-event responses. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v084i12;   https://www.jstatsoft.
org/index.php/jss/article/view/v084i12/v84i12.pdf.

Gillespie:2015:FHT

[Gil15] Colin S. Gillespie. Fitting heavy tailed distributions: The poweRlaw package. Journal of Statistical Software, 64(2):??, March 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v64/i02.

Giorgino:2009:CVD

[Gio09] Toni Giorgino. Computing and visualizing dynamic time warping alignments in R: The dtw package. Journal of Statistical Software, 31(7):??, August 2009. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v31/i07.

Gilleland:2016:EEV

[GK16] Eric Gilleland and Richard W. Katz. extRemes 2.0: An extreme value analysis package in R. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v072i08;   https://www.jstatsoft.
org/index.php/jss/article/view/v072i08/v72i08.pdf.

Giordani:2014:TWC

[GKD14] Paolo Giordani, Henk A. L. Kiers, and Maria Antonietta Del Ferraro. Three-way component analysis using the R package ThreeWay. Journal of Statistical Software, 57(7):??, April 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v57/i07.

Golyandina:2015:MES

[GKSU15] Nina Golyandina, Anton Korobeynikov, Alex Shlemov, and Konstantin Usevich. Multivariate and 2D extensions of singular spectrum analysis with the Rssa package. Journal of Statistical Software, 67(2):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067i02;
http://www.jstatsoft.org/index.php/jss/article/view/
v067i02/v67i02.pdf.

Grun:2012:EBR

[GKZ12] Bettina Grün, Ioannis Kosmidis, and Achim Zeileis. Extended beta regression in R: Shaken, stirred, mixed, and partitioned. Journal of Statistical Software, 48(11):??, May 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v48/i11.

Ghysels:2016:MFD

[GKZ16] Eric Ghysels, Virmantas Kvedaras, and Vaidotas Zemlys. Mixed frequency data sampling regression models: The R package midasr. Journal of Statistical Software, 72 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v072i04;   https://www.jstatsoft.org/
index.php/jss/article/view/v072i04/v72i04.pdf.

Grun:2008:FVF

[GL08] Bettina Grün and Friedrich Leisch. FlexMix version 2: Finite mixtures with concomitant variables and varying and constant parameters. Journal of Statistical Software, 28(4):??, October 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v28/i04.

Godfrey:2014:SRS

[GL14] A. Jonathan R. Godfrey and M. Theodor Loots. Software review: Statistical Software (R, SAS, SPSS, and Minitab) for Blind Students and Practitioners. Journal of Statistical Software, 58(SR-1):??, July 2014. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v58/s01.

Gollini:2015:PGR

[GLC^{+}15] Isabella Gollini, Binbin Lu, Martin Charlton, Christopher Brunsdon, and Paul Harris. GWmodel: An R package for exploring spatial heterogeneity using geographically weighted models. Journal of Statistical Software, 63(17):??, February 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v63/i17.

Glenn:2016:BRS

[Gle16] Ezra Haber Glenn. Book review: Spatial Microsimulation with R, Robin Lovelace and Morgane Dumont, Chapman & Hall/CRC, Boca Raton, 2016. ISBN 978-1-4987-1154-8. 259 pp. USD 89.95 (P), 62.97 (eBook).

https://www.crcpress.com/9781498711548. Journal of

Statistical Software, 72(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v072b01;   https://www.
jstatsoft.org/index.php/jss/article/view/v072b01/v72b01.
pdf.

Gomez:2015:SSM

[Góm15] Victor Gómez. SSMMATLAB: A set of MATLAB programs for the statistical analysis of state space models. Journal of Statistical Software, 66(9):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066i09;
http://www.jstatsoft.org/index.php/jss/article/view/
v066i09/v66i09.pdf.

Gonzalez:2014:SSN

[Gon14] Jorge González. SNSequate: Standard and nonstandard statistical models and methods for test equating. Journal of Statistical Software, 59(7):??, September 2014. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v59/i07.

Goslee:2011:ARS

[Gos11] Sarah C. Goslee. Analyzing remote sensing data in R: The landsat package. Journal of Statistical Software, 43(4):??, July 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v43/i04.

Gosho:2014:CSC

[Gos14] Masahiko Gosho. Code snippet: Criteria to select a working correlation structure in SAS. Journal of Statistical Software, 57(CS-1):??, March 2014. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v57/c01.

Gould:2005:BRU

[Gou05] Robert Gould. Book review: Using R for Introductory Statistics. Journal of Statistical Software, 14(BR-7):1–3, November 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v14/b07.

Gould:2010:BRB

[Gou10] Robert Gould. Book review: A Modern Approach to Regression with R. Journal of Statistical Software, 33(BR-3): ??, March 2010. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v33/b03.

Gower:2016:BPO

[Gow16] John Gower. Biometrics and psychometrics: Origins, commonalities and differences. Journal of Statistical Software, 73(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v073i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v073i05/v73i05.pdf.

Gerber:2012:DAM

[GP12] Samuel Gerber and Kristin Potter. Data analysis with the Morse–Smale complex: The msr package for R. Journal of Statistical Software, 50(2):??, July 2012. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v50/i02.

Gabadinho:2016:ASS

[GR16a] Alexis Gabadinho and Gilbert Ritschard. Analyzing state sequences with probabilistic suffix trees: The PST R package. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v072i03;   https://
www.jstatsoft.org/index.php/jss/article/view/v072i03/
v72i03.pdf.

Gomez-Rubio:2016:BRS

[GR16b] Virgilio Gómez-Rubio. Book review: Spatial Point Patterns: Methodology and Applications with R. Journal of Statistical Software, 75(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v075b02;   https://www.jstatsoft.
org/index.php/jss/article/view/v075b02/v75b02.pdf.

Gomez-Rubio:2017:BRG

[GR17a] Virgilio Gómez-Rubio. Book review: ggplot2 — Elegant Graphics for Data Analysis (2nd Edition). Journal of Statistical Software, 77(??):??, ???? 2017. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v077b02;   https://www.jstatsoft.
org/index.php/jss/article/view/v077b02/v77b02.pdf.

Gomez-Rubio:2017:BRW

[GR17b] Virgilio Gómez-Rubio. Book review: Working with the American Community Survey in R: A Guide to Using the acs Package. Journal of Statistical Software, 76 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/

article/view/v076b01;   https://www.jstatsoft.org/
index.php/jss/article/view/v076b01/v76b01.pdf.

Gomez-Rubio:2018:BRG

[GR18a] Virgilio Gomez-Rubio. Book review: Generalized Additive Models: An Introduction with R (2nd Edition). Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v086b01;   https://www.
jstatsoft.org/index.php/jss/article/view/v086b01/v86b01.
pdf.

Guo:2018:MBB

[GR18b] Jingyi Guo and Andrea Riebler. meta4diag: Bayesian bivariate meta-analysis of diagnostic test studies for routine practice. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v083i01;   https://
www.jstatsoft.org/index.php/jss/article/view/v083i01/
v83i01.pdf.

Gramacy:2007:TRP

[Gra07] Robert B. Gramacy. tgp: An R package for Bayesian nonstationary, semiparametric nonlinear regression and design by treed Gaussian process models. Journal of Statistical Software, 19(9):1–46, June 2007. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v19/i09.

Graffelman:2015:EDG

[Gra15] Jan Graffelman. Exploring diallelic genetic markers: The HardyWeinberg package. Journal of Statistical Software, 64 (3):??, March 2015. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v64/i03.

Gramacy:2016:LLS

[Gra16] Robert B. Gramacy. laGP: Large-scale spatial modeling via local approximate Gaussian processes in R. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v072i01;   https://www.jstatsoft.
org/index.php/jss/article/view/v072i01/v72i01.pdf.

Gabriel:2013:SRP

[GRD13] Edith Gabriel, Barry S. Rowlingson, and Peter J. Diggle. stpp: An R package for plotting, simulating and analyzing spatiotemporal point patterns. Journal of Statistical Software, 53 (2):??, April 2013. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v53/i02.

Greene:2017:CSS

[Gre17] Erich J. Greene. Code snippet: A SAS macro for covariateconstrained randomization of general cluster-randomized and unstratified designs. Journal of Statistical Software, 77 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v077c01;   https://www.jstatsoft.org/
index.php/jss/article/view/v077c01/v77c01.pdf.

George:2016:RPC

[GRK^{+}16] Ann Cathrice George, Alexander Robitzsch, Thomas Kiefer, Jürgen Groß, and Ali Ünlü. The R package CDM for cognitive diagnosis models. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v074i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v074i02/v74i02.pdf.

Gomez-Rubio:2019:DMB

[GRMMR19] Virgilio Gómez-Rubio, Paula Moraga, John Molitor, and Barry Rowlingson. DClusterm: Model-based detection of disease clusters. Journal of Statistical Software, 90 (??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v090i14;   https://www.jstatsoft.org/
index.php/jss/article/view/v090i14/v90i14.pdf.

Gabadinho:2011:AVS

[GRMS11] Alexis Gabadinho, Gilbert Ritschard, Nicolas S. Müller, and Matthias Studer. Analyzing and visualizing state sequences in R with TraMineR. Journal of Statistical Software, 40(4):??, April 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v40/i04.

Gromping:2006:RIL

[Grö06] Ulrike Grömping. Relative importance for linear regression in R: The package relaimpo. Journal of Statistical Software, 17 (1):1–27, September 2006. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v17/i01.

Grothendieck:2008:BRB

[Gro08] Gabor Grothendieck. Book review: lattice: Multivariate Data Visualization with R, Deepayan Sarkar, Springer-Verlag, New York, NY, 2008. ISBN 978-0-387-75968-5. 268 pp. USD

54.95 (P). http://dsarkar.fhcrc.org/lattice/. Journal of

Statistical Software, 25(BR-2):1–3, May 2008. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v25/b02.

Gromping:2010:BRB

[Grö10a] Ulrike Grömping. Book review: Design and Analysis of Experiments with SAS. Journal of Statistical Software, 37(BR-2): ??, December 2010. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v37/b02.

Gromping:2010:ILE

[Grö10b] Ulrike Grömping. Inference with linear equality and inequality constraints using R: The package ic.infer. Journal of Statistical Software, 33(10):??, February 2010. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v33/i10.

Gromping:2011:BRB

[Grö11] Ulrike Grömping. Book review: Optimal Experimental Design with R. Journal of Statistical Software, 43(BR-5):??, October 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v43/b05.

Gromping:2013:BRB

[Grö13] Ulrike Grömping. Book review: The R Primer. Journal of Statistical Software, 52(BR-3):??, February 2013. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v52/b03.

Gromping:2014:RPP

[Grö14] Ulrike Grömping. R package FrF2 for creating and analyzing fractional factorial 2-level designs. Journal of Statistical Software, 56(1):??, January 2014. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v56/i01.

Gromping:2015:BRU

[Grö15a] Ulrike Grömping. Book review: Using R and RStudio for Data Management, Statistical Analysis and Graphics (2nd Edition), Nicholas J. Horton and Ken Kleinman, CRC Press, Boca Raton, 2015. ISBN 978-1-4822-3736-8. 313 pp.

USD 59.46 (H). http://www.amherst.edu/ nhorton/r2/.

Journal of Statistical Software, 68(BR-4):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v068b04;
http://www.jstatsoft.org/index.php/jss/article/view/
v068b04/v68b04.pdf.

Gromping:2015:BRM

[Grö15b] Ulrike Grömping. Book review: Multilevel Modeling Using R. Journal of Statistical Software, 62(BR-1):??, January 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v62/b01.

Gromping:2016:BRP

[Grö16] Ulrike Grömping. Book review: Practical Guide to Logistic Regression, Joseph M. Hilbe, Chapman & Hall/CRC, Boca Raton, 2016. ISBN 978-1-4987-0957-6. 174 pp. USD 49.95

(P). https://www.crcpress.com/9781498709576. Journal

of Statistical Software, 71(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v071b03;   https://www.
jstatsoft.org/index.php/jss/article/view/v071b03/v71b03.
pdf.

Gromping:2017:BRD

[Grö17] Ulrike Grömping. Book review: Data Visualisation with R: 100 Examples. Journal of Statistical Software, 81 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v081b02;   https://www.jstatsoft.org/
index.php/jss/article/view/v081b02/v81b02.pdf.

Gromping:2018:BRD

[Grö18a] Ulrike Grömping. Book review: A Data Scientist’s Guide to Acquiring, Cleaning and Managing Data in R. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v086b03;   https://www.
jstatsoft.org/index.php/jss/article/view/v086b03/v86b03.
pdf.

Gromping:2018:RPD

[Grö18b] Ulrike Grömping. R package DoE.base for factorial experiments. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v085i05;
https://www.jstatsoft.org/index.php/jss/article/view/
v085i05/v85i05.pdf.

Grose:2018:BRT

[Gro18c] Daniel Grose. Book review: Testing R Code. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v085b01;   https://www.jstatsoft.
org/index.php/jss/article/view/v085b01/v85b01.pdf.

Gromping:2019:BRP

[Grö19] Ulrike Grömping. Book review: Practical R for Mass Communication and Journalism. Journal of Statistical Software, 90(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v090b02;   https://www.jstatsoft.org/
index.php/jss/article/view/v090b02/v90b02.pdf.

Gromping:2021:BRD

[Grö21] Ulrike Grömping. Book review: Data Visualization: Charts, Maps, and Interactive Graphics. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v098b01;   https://www.jstatsoft.org/
index.php/jss/article/view/v098b01/v98b01.pdf.

Gorst-Rasmussen:2012:CDM

[GRS12] Anders Gorst-Rasmussen and Thomas H. Scheike. Coordinate descent methods for the penalized semiparametric additive hazards model. Journal of Statistical Software, 47(9):??, April 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v47/i09.

Galimberti:2012:CTO

[GSD12] Giuliano Galimberti, Gabriele Soffritti, and Matteo Di Maso. Classification trees for ordinal responses in R: The rpartScore package. Journal of Statistical Software, 47(10):??, May 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v47/i10.

Gronau:2020:BRP

[GSW20] Quentin F. Gronau, Henrik Singmann, and Eric-Jan Wagenmakers. bridgesampling: an R package for estimating normalizing constants. Journal of Statistical Software, 92(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v092i10;   https://www.jstatsoft.org/
index.php/jss/article/view/v092i10/v92i10.pdf.

Gramacy:2010:CIS

[GT10] Robert B. Gramacy and Matthew Alan Taddy. Categorical inputs, sensitivity analysis, optimization and importance tempering with tgp version 2, an R package for treed Gaussian process models. Journal of Statistical Software, 33(6):??, February 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v33/i06.

Goudie:2020:MPI

[GTDT20] Robert J. B. Goudie, Rebecca M. Turner, Daniela De Angelis, and Andrew Thomas. MultiBUGS: a parallel implementation of the BUGS modeling framework for faster Bayesian inference. Journal of Statistical Software, 95 (??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v095i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v095i07/v95i07.pdf.

Goslee:2007:EPD

[GU07] Sarah C. Goslee and Dean L. Urban. The ecodist package for dissimilarity-based analysis of ecological data. Journal of Statistical Software, 22(7):1–19, September 2007. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v22/i07.

Gu:2014:SSA

[Gu14] Chong Gu. Smoothing spline ANOVA models: R package gss. Journal of Statistical Software, 58(5):??, June 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v58/i05.

Guha:2007:CIF

[Guh07] Rajarshi Guha. Chemical informatics functionality in R. Journal of Statistical Software, 18(5):1–16, January 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v18/i05.

Gumprecht:2007:BRB

[Gum07] Daniela Gumprecht. Book review: Model-based Geostatistics. Journal of Statistical Software, 21(BR-3):1–2, September 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v21/b03.

Guolo:2012:RPM

[GV12] Annamaria Guolo and Cristiano Varin. The R package metaLik for likelihood inference in meta-analysis. Journal of Statistical Software, 50(7):??, August 2012. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v50/i07.

Gruber:2012:TRP

[GvdL12] Susan Gruber and Mark J. van der Laan. tmle: An R package for targeted maximum likelihood estimation. Journal of Statistical Software, 51(13):??, November 2012. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v51/i13.

Grasman:2009:FCC

[GvdMW09] Raoul P. P. P. Grasman, Han L. J. van der Maas, and Eric-Jan Wagenmakers. Fitting the cusp catastrophe in R: a cusp package primer. Journal of Statistical Software, 32(8):??, November 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v32/i08.

Groenen:2016:MSM

[GvdV16] Patrick J. F. Groenen and Michel van de Velden. Multidimensional scaling by majorization: A review. Journal of Statistical Software, 73(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v073i08;   https://www.jstatsoft.
org/index.php/jss/article/view/v073i08/v73i08.pdf.

Griffith:2016:CSC

[GVM16] Daniel M. Griffith, Joseph A. Veech, and Charles J. Marsh. Code snippet: cooccur: Probabilistic species cooccurrence analysis in R. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v069c02;   https://www.jstatsoft.org/
index.php/jss/article/view/v069c02/v69c02.pdf.

Grolemund:2011:DTM

[GW11] Garrett Grolemund and Hadley Wickham. Dates and times made easy with lubridate. Journal of Statistical Software, 40 (3):??, April 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v40/i03.

Groendyke:2018:ERP

[GW18] Chris Groendyke and David Welch. epinet: An R package to analyze epidemics spread across contact networks. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v083i11;   https://www.
jstatsoft.org/index.php/jss/article/view/v083i11/v83i11.
pdf.

Grun:2009:AGE

[GZ09] Bettina Grün and Achim Zeileis. Automatic generation of exams in R. Journal of Statistical Software, 29(10):??, February 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v29/i10.

Grubinger:2014:PEE

[GZP14] Thomas Grubinger, Achim Zeileis, and Karl-Peter Pfeiffer. evtree: Evolutionary learning of globally optimal classification and regression trees in R. Journal of Statistical Software, 61(1):??, October 2014. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v61/i01.

Hadfield:2010:MMM

[Had10] Jarrod D. Hadfield. MCMC methods for multi-response generalized linear mixed models: The MCMCglmm R package. Journal of Statistical Software, 33(2):??, February 2010. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v33/i02.

Hamilton:2010:BRB

[Ham10] Cody Hamilton. Book review: Bayesian Adaptive Methods for Clinical Trials. Journal of Statistical Software, 37(BR-1):??, November 2010. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v37/b01.

Hankin:2005:IBR

[Han05] Robin K. S. Hankin. Introducing BACCO, an R bundle for Bayesian analysis of computer code output. Journal of Statistical Software, 14(16):1–21, October 19, 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=141&url=v14/i16&ct=2;   http://www.
jstatsoft.org/counter.php?id=141&url=v14/i16/v14i16.
pdf&ct=1.

Hankin:2006:CSA

[Han06a] Robin K. S. Hankin. Code snippet: Additive integer partitions in R. Journal of Statistical Software, 16(CS-1):1–3, May 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v16/c01.

Hankin:2006:IPE

[Han06b] Robin K. S. Hankin. Introducing elliptic, an R package for elliptic and modular functions. Journal of Statistical Software, 15(7):1–22, January 2006. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v15/i07.

Hankin:2007:CSU

[Han07a] Robin K. S. Hankin. Code snippet: Urn sampling without replacement: Enumerative combinatorics in R. Journal of Statistical Software, 17(CS-1):1–6, January 2007. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v17/c01.

Hankin:2007:IUR

[Han07b] Robin K. S. Hankin. Introducing untb, an R package for simulating ecological drift under the unified neutral theory of biodiversity. Journal of Statistical Software, 22(12):1–15, September 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v22/i12.

Hankin:2010:GDD

[Han10] Robin K. S. Hankin. A generalization of the Dirichlet distribution. Journal of Statistical Software, 33(11):??, February 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v33/i11.

Hankin:2012:IMM

[Han12] Robin K. S. Hankin. Introducing multivator: a multivariate emulator. Journal of Statistical Software, 46(8):??, February 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v46/i08.

Hankin:2015:BRC

[Han15] Robin K. S. Hankin. Book review: Circular Statistics in R, Arthur Pewsey, Markus Neuhäuser, Graeme D. Ruxton, Oxford University Press, Oxford, 2013. ISBN 978-0-19-967113-7.

183 pp. GBP 24.99 (P). http://ukcatalogue.oup.com/product/9780199671137.do.

Journal of Statistical Software, 66(BR-5):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066b05;
http://www.jstatsoft.org/index.php/jss/article/view/
v066b05/v66b05.pdf.

Hartmann:2008:BRB

[Har08] Wolfgang M. Hartmann. Book review: SAS for Data Analysis: Intermediate Statistical Methods. Journal of Statistical Software, 28(BR-1):??, October 2008. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v28/b01.

Harte:2010:PRP

[Har10] David Harte. PtProcess: An R package for modelling marked point processes indexed by time. Journal of Statistical Software, 35(8):??, July 2010. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v35/i08.

Harris:2015:CSP

[Har15] Naftali Harris. Code snippet: LazySorted: A lazily, partially sorted Python list. Journal of Statistical Software, 65(CS-1):

??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v65/c01.

Holtkamp:2017:SFA

[HB17] Jonathan Holtkamp and Bernhard Brümmer. Stochastic frontier analysis using SFAMB for Ox. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v081i06;   https://www.jstatsoft.org/
index.php/jss/article/view/v081i06/v81i06.pdf.

Hahsler:2017:ISE

[HBF17] Michael Hahsler, Matthew Bolaños, and John Forrest. Introduction to stream: An extensible framework for data stream clustering research with R. Journal of Statistical Software, 76(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v076i14;   https://www.jstatsoft.org/
index.php/jss/article/view/v076i14/v76i14.pdf.

Hin:2005:ACD

[HC05] Lin Yee Hin and Vincent Carey. An algorithm for clustered data generalized additive modelling with S-PLUS. Journal of Statistical Software, 14(8):1–24, August 31, 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=133&url=v14/i08&ct=2;   http://www.
jstatsoft.org/counter.php?id=133&url=v14/i08/v14i08.
pdf&ct=1.

Huang:2012:TGG

[HCW12] Bei Huang, Dianne Cook, and Hadley Wickham. tourrGui: a gWidgets GUI for the tour to explore high-dimensional data using low-dimensional projections. Journal of Statistical Software, 49(6):??, June 2012. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v49/i06.

Hahsler:2010:REM

[HD10] Michael Hahsler and Margaret H. Dunham. rEMM: Extensible Markov model for data stream clustering in R. Journal of Statistical Software, 35(5):??, July 2010. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v35/i05.

Hatzinger:2012:PRP

[HD12] Reinhold Hatzinger and Regina Dittrich. prefmod: An R package for modeling preferences based on paired comparisons, rankings, or ratings. Journal of Statistical Software, 48(10): ??, May 2012. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v48/i10.

Hocevar:2016:CGO

[HD16] Tomaz Hocevar and Janez Demsar. Computation of graphlet orbits for nodes and edges in sparse graphs. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v071i10;   https://www.jstatsoft.
org/index.php/jss/article/view/v071i10/v71i10.pdf.

Han:2018:GRP

[HD18] Zifei Han and Victor De Oliveira. gcKrig: An R package for the analysis of geostatistical count data using Gaussian copulas. Journal of Statistical Software, 87 (??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v087i13;   https://www.jstatsoft.org/
index.php/jss/article/view/v087i13/v87i13.pdf.

Hedeker:1999:MCP

[Hed99] Donald Hedeker. MIXNO: a computer program for mixedeffects nominal logistic regression. Journal of Statistical Software, 4(5):1–92, ???? 1999. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v04/i05;   http:
//www.jstatsoft.org/v04/i05/mixno102.exe;   http://
www.jstatsoft.org/v04/i05/mixnocm.pdf;   http://www.
jstatsoft.org/v04/i05/updates.

Helmreich:2015:BRS

[Hel15] James E. Helmreich. Book review: Statistics: An Introduction Using R (2nd Edition), Michael J. Crawley, John Wiley & Sons, Chichester, 2015. ISBN 978-1-118-94109-6. 339 pp. USD

45.00 (P). http://www.imperial.ac.uk/bio/research/crawley/statistics/.

Journal of Statistical Software, 67(BR-5):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067b05;
http://www.jstatsoft.org/index.php/jss/article/view/
v067b05/v67b05.pdf.

Helmreich:2016:BRL

[Hel16a] James E. Helmreich. Book review: Learning Base R. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v069b04;   https://www.
jstatsoft.org/index.php/jss/article/view/v069b04/v69b04.
pdf.

Helmreich:2016:BRR

[Hel16b] James E. Helmreich. Book review: Regression Modeling Strategies with Applications to Linear Models, Logistic and Ordinal Regression and Survival Analysis (2nd Edition). Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v070b02;   https://www.
jstatsoft.org/index.php/jss/article/view/v070b02/v70b02.
pdf.

Helmreich:2017:BRS

[Hel17] James E. Helmreich. Book review: Statistical Analysis and Data Display: An Intermediate Course with Examples in R (2nd Edition). Journal of Statistical Software, 76(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v076b02;   https://www.jstatsoft.org/
index.php/jss/article/view/v076b02/v76b02.pdf.

Helmreich:2018:BRR

[Hel18] James E. Helmreich. Book review: The R Primer (2nd ed.). Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v083b02;   https://
www.jstatsoft.org/index.php/jss/article/view/v083b02/
v83b02.pdf.

Helmreich:2020:BRD

[Hel20] James E. Helmreich. Book review: Data Science with Julia. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v094b01;   https://
www.jstatsoft.org/index.php/jss/article/view/v094b01/
v94b01.pdf.

Hamilton:2018:CSG

[HF18] Nicholas E. Hamilton and Michael Ferry. Code snippet: ggtern: Ternary diagrams using ggplot2. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v087c03;   https://www.jstatsoft.
org/index.php/jss/article/view/v087c03/v87c03.pdf.

Hornik:2012:SMC

[HFKB12] Kurt Hornik, Ingo Feinerer, Martin Kober, and Christian Buchta. Spherical k-means clustering. Journal of Statistical Software, 50(10):??, September 2012. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v50/i10.

Hornik:2014:PMR

[HG14] Kurt Hornik and Bettina Grün. movMF: An R package for fitting mixtures of von Mises–Fisher distributions. Journal of Statistical Software, 58(10):??, July 2014. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v58/i10.

Hui:2008:LRP

[HGG08] Wallace Hui, Yulia R. Gel, and Joseph L. Gastwirth. lawstat: An R package for law, public policy and biostatistics. Journal of Statistical Software, 28(3):??, October 2008. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v28/i03.

Hahsler:2005:PAC

[HGH05] Michael Hahsler, Bettina Grün, and Kurt Hornik. arules — a computational environment for mining association rules and frequent item sets. Journal of Statistical Software, 14(15):1–25, September 29, 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=140&url=v14/i15&ct=2;   http://www.
jstatsoft.org/counter.php?id=140&url=v14/i15/v14i15.
pdf&ct=1.

Hunter:2013:EUT

[HGH13] David R. Hunter, Steven M. Goodreau, and Mark S. Handcock. ergm.userterms: a template package for extending statnet. Journal of Statistical Software, 52(2):??, February 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v52/i02.

Hofmann:2020:LEV

[HGK^{+}20] Marc Hofmann, Cristian Gatu, Erricos J. Kontoghiorghes, Ana Colubi, and Achim Zeileis. lmSubsets: Exact variable-subset selection in linear regression for R. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v093i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v093i03/v93i03.pdf.

Hahsler:2007:TIT

[HH07a] Michael Hahsler and Kurt Hornik. TSP — infrastructure for the traveling salesperson problem. Journal of Statistical Software, 23(2):1–21, December 2007. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v23/i02.

Henningsen:2007:SPE

[HH07b] Arne Henningsen and Jeff D. Hamann. systemfit: a package for estimating systems of simultaneous equations in R. Journal of Statistical Software, 23(4):1–40, December 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v23/i04.

Halekoh:2014:KRA

[HH14] Ulrich Halekoh and Søren Højsgaard. A Kenward–Roger approximation and parametric bootstrap methods for tests in linear mixed models — the R package pbkrtest. Journal of Statistical Software, 59(9):??, September 2014. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v59/i09.

Helske:2019:MHM

[HH19] Satu Helske and Jouni Helske. Mixture hidden Markov models for sequence data: The seqHMM package in R. Journal of Statistical Software, 88(??):??, ???? 2019. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v088i03;   https://www.jstatsoft.
org/index.php/jss/article/view/v088i03/v88i03.pdf.

Hahsler:2008:GTO

[HHB08a] Michael Hahsler, Kurt Hornik, and Christian Buchta. Getting things in order: An introduction to the R package seriation. Journal of Statistical Software, 25(3):1–34, March 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v25/i03.

Handcock:2008:SST

[HHB^{+}08b] Mark S. Handcock, David R. Hunter, Carter T. Butts, Steven M. Goodreau, and Martina Morris. statnet: Software tools for the representation, visualization, analysis and simulation of network data. Journal of Statistical Software, 24 (1):1–11, May 2008. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v24/i01.

Hunter:2008:EPF

[HHB^{+}08c] David R. Hunter, Mark S. Handcock, Carter T. Butts, Steven M. Goodreau, and Martina Morris. ergm: a package to fit, simulate and diagnose exponential-family models for networks. Journal of Statistical Software, 24(3):1–29, May 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v24/i03.

Ho:2021:DSS

[HHJCRB21] Anson T. Y. Ho, Kim P. Huynh, David T. Jacho-Chávez, and Diego Rojas-Baez. Data science in Stata 16: Frames, lasso, and Python integration. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v098s01;   https://www.jstatsoft.org/
index.php/jss/article/view/v098s01/v98s01.pdf.

Hirk:2020:MRP

[HHV20] Rainer Hirk, Kurt Hornik, and Laura Vana. mvord: an R package for fitting multivariate ordinal regression models. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v093i04;   https://www.
jstatsoft.org/index.php/jss/article/view/v093i04/v93i04.
pdf.

Hothorn:2008:ICP

[HHvdWZ08] Torsten Hothorn, Kurt Hornik, Mark A. van de Wiel, and Achim Zeileis. Implementing a class of permutation tests: The coin package. Journal of Statistical Software, 28(8):??, November 2008. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v28/i08.

Halekoh:2006:RPP

[HHY06] Ulrich Halekoh, Søren Højsgaard, and Jun Yan. The R package geepack for generalized estimating equations. Journal of Statistical Software, 15(2):1–11, January 2006. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v15/i02.

Howard:2020:PSA

[HI20] James P. Howard and II. Phonetic spelling algorithm implementations for R. Journal of Statistical Software, 95 (??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v095i08;   https://www.jstatsoft.org/
index.php/jss/article/view/v095i08/v95i08.pdf.

Ho:2011:MNP

[HIKS11] Daniel E. Ho, Kosuke Imai, Gary King, and Elizabeth A. Stuart. MatchIt: Nonparametric preprocessing for parametric causal inference. Journal of Statistical Software, 42(8):??, June 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v42/i08.

Hilbe:2006:BRB

[Hil06] Joseph Hilbe. Book review: A Handbook of Statistical Analyses Using R. Journal of Statistical Software, 16(BR-6):1–6, August 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v16/b06.

Hilbe:2007:BRB

[Hil07] Joseph Hilbe. Book review: Bayesian Core: A Practical Approach to Bayesian Computational Statistics. Journal of Statistical Software, 21(BR-1):1–5, July 2007. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v21/b01.

Hilbe:2009:BRBa

[Hil09a] Joseph Hilbe. Book review: Data Analysis Using Regression and Multilevel/Hierarchical Models. Journal of Statistical Software, 30(BR-3):??, April 2009. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v30/b03.

Hilbe:2009:BRBb

[Hil09b] Joseph Hilbe. Book review: Survival Analysis for Epidemiologic and Medical Research. Journal of Statistical Software, 30 (BR-4):??, April 2009. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v30/b04.

Hilbe:2010:BRBb

[Hil10a] Joseph Hilbe. Book review: Bayesian Analysis for Population Ecology. Journal of Statistical Software, 36(BR-1):??, August 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v36/b01.

Hilbe:2010:BRBa

[Hil10b] Joseph Hilbe. Book review: Introduction to Scientific Programming and Scientific Simulation Using R. Journal of Statistical Software, 36(BR-4):??, September 2010. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v36/b04.

Hilbe:2015:BRH

[Hil15a] Joseph M. Hilbe. Book review: Hurricane Climatology: A Modern Statistical Guide Using R, James B. Elsner and Thomas H. Jagger. Oxford University Press, New York, 2013. ISBN 978-0-19-982763-3. 390 pp. USD 105.00 (P).

http://ukcatalogue.oup.com/product/9780199827633.do.

Journal of Statistical Software, 67(BR-6):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067b06;
http://www.jstatsoft.org/index.php/jss/article/view/
v067b06/v67b06.pdf.

Hilbe:2015:BRN

[Hil15b] Joseph M. Hilbe. Book review: The New Statistics with R: An Introduction for Biologists, Andy Hector, Oxford University Press, New York, 2015. ISBN 978-0-19-872906-8. 208 pp. USD

49.95 (P). http://ukcatalogue.oup.com/product/9780198729068.do.

Journal of Statistical Software, 68(BR-1):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v068b01;
http://www.jstatsoft.org/index.php/jss/article/view/
v068b01/v68b01.pdf.

Husson:2016:JLF

[HJS16] François Husson, Julie Josse, and Gilbert Saporta. Jan de Leeuw and the French school of data analysis. Journal of Statistical Software, 73(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v073i06;   https://www.
jstatsoft.org/index.php/jss/article/view/v073i06/v73i06.
pdf.

Hyndman:2008:ATS

[HK08] Rob J. Hyndman and Yeasmin Khandakar. Automatic time series forecasting: The forecast package for R. Journal of Statistical Software, 27(3):??, July 2008. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v27/i03.

Happ-Kurz:2020:OOS

[HK20] Clara Happ-Kurz. Object-oriented software for functional data. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v093i05;
https://www.jstatsoft.org/index.php/jss/article/view/
v093i05/v93i05.pdf.

Honaker:2011:AIP

[HKB11] James Honaker, Gary King, and Matthew Blackwell. Amelia II: a program for missing data. Journal of Statistical Software, 45(7):??, December 2011. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v45/i07.

Holmes:2009:IJS

[HKL09] Susan Holmes, Adam Kapelner, and Peter P. Lee. An interactive Java statistical image segmentation system: GemIdent. Journal of Statistical Software, 30(10):??, June 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v30/i10.

Hojsgaard:2007:IGG

[HL07] Søren Højsgaard and Steffen L. Lauritzen. Inference in graphical Gaussian models with edge and vertex symmetries with the gRc package for R. Journal of Statistical Software, 23(6): 1–26, December 2007. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v23/i06.

Hoffmann:2009:FMA

[HL09] Thomas J. Hoffmann and Nan M. Laird. fgui: a method for automatically creating graphical user interfaces for commandline R packages. Journal of Statistical Software, 30(2):??, April 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v30/i02.

Hlavac:2016:CST

[Hla16a] Marek Hlavac. Code snippet: TableMaker: An Excel macro for publication-quality tables. Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v070c03;   https://www.jstatsoft.org/
index.php/jss/article/view/v070c03/v70c03.pdf.

Hlavac:2016:EEB

[Hla16b] Marek Hlavac. ExtremeBounds: Extreme bounds analysis in R. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v072i09;   https://
www.jstatsoft.org/index.php/jss/article/view/v072i09/
v72i09.pdf.

Hofert:2011:NAC

[HM11] Marius Hofert and Martin Mächler. Nested Archimedean copulas meet R: The nacopula package. Journal of Statistical Software, 39(9):??, March 2011. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v39/i09.

Hofert:2016:POS

[HM16] Marius Hofert and Martin Mächler. Parallel and other simulations in R made easy: An end-to-end study. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v069i04;   https://www.jstatsoft.
org/index.php/jss/article/view/v069i04/v69i04.pdf.

Heck:2018:RRP

[HM18] Daniel W. Heck and Morten Moshagen. RRreg: An R package for correlation and regression analyses of randomized response data. Journal of Statistical Software, 85 (??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v085i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v085i02/v85i02.pdf.

Hill:2021:RPF

[HMM21] Andreas Hill, Alexander Massey, and Daniel Mandallaz. The R package forestinventory: Design-based global and small area estimations for multiphase forest inventories. Journal of Statistical Software, 97(??):??, ???? 2021. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v097i04;   https://www.
jstatsoft.org/index.php/jss/article/view/v097i04/v97i04.
pdf.

Hornik:2013:TPG

[HMR^{+}13] Kurt Hornik, Patrick Mair, Johannes Rauch, Wilhelm Geiger, Christian Buchta, and Ingo Feinerer. The textcat package for n-gram based text categorization in R. Journal of Statistical Software, 52(6):??, February 2013. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v52/i06.

Hofner:2016:GRP

[HMS16] Benjamin Hofner, Andreas Mayr, and Matthias Schmid. gamboostLSS: An R package for model building and variable selection in the GAMLSS framework. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v074i01;   https://www.jstatsoft.
org/index.php/jss/article/view/v074i01/v74i01.pdf.

Hedeker:2013:MPM

[HN13] Donald Hedeker and Rachel Nordgren. MIXREGLS: a program for mixed-effects location scale analysis. Journal of Statistical Software, 52(12):??, March 2013. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v52/i12.

Hamuro:2018:MAP

[HNCI18] Yukinobu Hamuro, Masakazu Nakamoto, Stephane Cheung, and Edward H. Ip. mbonsai: Application package for sequence classification by tree methodology. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v086i06;   https://www.jstatsoft.
org/index.php/jss/article/view/v086i06/v86i06.pdf.

Hofert:2020:ZEN

[HO20] Marius Hofert and Wayne Oldford. Zigzag expanded navigation plots in R: The R package zenplots. Journal of Statistical Software, 95(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v095i04;   https://www.jstatsoft.
org/index.php/jss/article/view/v095i04/v95i04.pdf.

Hoffmann:2011:CSP

[Hof11] Thomas J. Hoffmann. Code snippet: Passing in command line arguments and parallel cluster/multicore batching in R with batch. Journal of Statistical Software, 39(CS-1):??, March 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v39/c01.

Hofmarcher:2015:BRI

[Hof15] Paul Hofmarcher. Book review: An Introduction to R for Quantitative Economics, Vikram Dayal. Springer-Verlag, New Dehli, 2015. ISBN 978-81-322-2339-9. 109 pp. EUR 51.99.

http://www.springer.com/book/9788132223399. Journal

of Statistical Software, 67(BR-4):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067b04;
http://www.jstatsoft.org/index.php/jss/article/view/
v067b04/v67b04.pdf.

Hohensinn:2018:CSP

[Hoh18] Christine Hohensinn. Code snippet: pcIRT: An R package for polytomous and continuous Rasch models. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v084c02;   https://www.jstatsoft.
org/index.php/jss/article/view/v084c02/v84c02.pdf.

Hojsgaard:2004:PMP

[Høj04] Søren Højsgaard. The mimR package for graphical modelling in R. Journal of Statistical Software, 11(6):1–13, ???? 2004. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/counter.php?id=102&url=v11/i06/
v11i06.pdf&ct=1.

Hojsgaard:2012:GIN

[Høj12] Søren Højsgaard. Graphical independence networks with the gRain package for R. Journal of Statistical Software, 46(10): ??, February 2012. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v46/i10.

Hornik:2005:CCE

[Hor05] Kurt Hornik. A CLUE for CLUster Ensembles. Journal of Statistical Software, 14(12):1–25, September 20, 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/counter.php?id=137&url=v14/i12&ct=2;   http://
www.jstatsoft.org/counter.php?id=137&url=v14/i12/v14i12.
pdf&ct=1.

Hothorn:2020:MLT

[Hot20] Torsten Hothorn. Most likely transformations: The mlt package. Journal of Statistical Software, 92(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v092i01;
https://www.jstatsoft.org/index.php/jss/article/view/
v092i01/v92i01.pdf.

Howard:2011:BRB

[How11] James P. Howard. Book review: R Cookbook. Journal of Statistical Software, 40(BR-3):??, April 2011. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v40/b03.

Howard:2015:BRD

[How15a] James P. Howard II. Book review: Data-Driven Modeling & Scientific Computation: Methods for Complex Systems & Big Data, J. Nathan Kutz, Oxford University Press, Oxford, 2013. ISBN 978-0-19-966034-6. 656 pp. USD 49.95 (P).

https://global.oup.com/academic/product/9780199660346.

Journal of Statistical Software, 67(BR-1):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067b01;
http://www.jstatsoft.org/index.php/jss/article/view/
v067b01/v67b01.pdf.

Howard:2015:BRU

[How15b] James P. Howard II. Book review: Uncertainty Quantification and Stochastic Modeling with MATLAB, Eduardo Souza de Cursi and Rubens Sampaio, ISTE Press, London, 2015. ISBN 978-1-78548-004-1. 442 pp. USD 180 (H).

http://www.sciencedirect.com/science/book/9781785480058.

Journal of Statistical Software, 67(BR-7):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067b07;
http://www.jstatsoft.org/index.php/jss/article/view/
v067b07/v67b07.pdf.

Howard:2016:BRM

[How16a] James P. Howard II. Book review: Meta-Analysis with R. Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v070b01;   https://
www.jstatsoft.org/index.php/jss/article/view/v070b01/
v70b01.pdf.

Howard:2016:BRP

[How16b] James P. Howard II. Book review: Political Analysis Using R, James E. Monogan III, Springer-Verlag, New York, 2015. ISBN 978-3-319-23446-5. 242 pp. USD 59.99

(S). http://www.springer.com/9783319234458. Journal of

Statistical Software, 74(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v074b01;   https://www.
jstatsoft.org/index.php/jss/article/view/v074b01/v74b01.
pdf.

Helmreich:2009:PRP

[HP09] James E. Helmreich and Robert M. Pruzek. PSAgraphics: An R package to support propensity score analysis. Journal of Statistical Software, 29(6):??, February 2009. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v29/i06.

Hahsler:2019:DFD

[HPD19] Michael Hahsler, Matthew Piekenbrock, and Derek Doran. dbscan: Fast density-based clustering with R. Journal of Statistical Software, 91(??):??, ???? 2019. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v091i01;   https://www.
jstatsoft.org/index.php/jss/article/view/v091i01/v91i01.
pdf.

Hu:2018:IMP

[HPK18] Weilong Hu, Yannis Pantazis, and Markos A. Katsoulakis. ISAP — MATLAB package for sensitivity analysis of highdimensional stochastic chemical networks. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v085i03;   https://www.jstatsoft.
org/index.php/jss/article/view/v085i03/v85i03.pdf.

Haggstrom:2015:CRP

[HPWdL15] Jenny Häggström, Emma Persson, Ingeborg Waernbaum, and Xavier de Luna. CovSel: An R package for covariate selection when estimating average causal effects. Journal of Statistical Software, 68(1):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v068i01;
http://www.jstatsoft.org/index.php/jss/article/view/
v068i01/v68i01.pdf.

Hayfield:2008:NEN

[HR08] Tristen Hayfield and Jeffrey S. Racine. Nonparametric econometrics: The np package. Journal of Statistical Software, 27 (5):??, July 2008. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v27/i05.

Heiberger:2014:DDS

[HR14] Richard M. Heiberger and Naomi B. Robbins. Design of diverging stacked bar charts for Likert scales and other applications. Journal of Statistical Software, 57(5):??, April 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v57/i05.

Hengl:2015:PSV

[HRBP15] Tomislav Hengl, Pierre Roudier, Dylan Beaudette, and Edzer Pebesma. plotKML: Scientific visualization of spatio-temporal data. Journal of Statistical Software, 63(5):??, February 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v63/i05.

Hooker:2016:CCI

[HRX16] Giles Hooker, James O. Ramsay, and Luo Xiao. CollocInfer: Collocation inference in differential equation models. Journal of Statistical Software, 75(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v075i02;   https://www.
jstatsoft.org/index.php/jss/article/view/v075i02/v75i02.
pdf.

Hu:2018:ERP

[HS18] Yang Hu and Carl Scarrott. evmix: An R package for extreme value mixture modeling, threshold estimation and boundary corrected kernel density estimation. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v084i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v084i05/v84i05.pdf.

Huang:2012:PDR

[HSG12] B. Emma Huang, Rohan Shah, and Andrew W. George. dlmap: An R package for mixed model QTL and association analysis. Journal of Statistical Software, 50(6):??, August 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v50/i06.

Headrick:2007:NCG

[HSH07] Todd C. Headrick, Yanyan Sheng, and Flaviu-Adrian Hodis. Numerical computing and graphics for the power method transformation using Mathematica. Journal of Statistical Software, 19(3):1–17, April 2007. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v19/i03.

Haneuse:2011:ORP

[HSL11] Sebastien Haneuse, Takumi Saegusa, and Thomas Lumley. osDesign: An R package for the analysis, evaluation, and design of two-phase and case-control studies. Journal of Statistical Software, 43(11):??, August 2011. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v43/i11.

Huang:2020:LSC

[Hua20] Po-Hsien Huang. lslx: Semi-confirmatory structural equation modeling via penalized likelihood. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v093i07;   https://www.jstatsoft.
org/index.php/jss/article/view/v093i07/v93i07.pdf.

Hughett:2007:CSA

[Hug07] Paul Hughett. Code snippet: Accurate computation of the F-to-z and t-to-z transforms for large arguments. Journal of Statistical Software, 23(CS-1):1–5, November 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v23/c01.

Husing:2006:ARA

[Hüs06] Johannes Hüsing. An accept-and-reject algorithm to sample from a set of permutations restricted by a time constraint. Journal of Statistical Software, 16(6):1–17, May 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v16/i06.

Hankin:2007:CSS

[HW07] Robin K. S. Hankin and Luke J. West. Code snippet: Set partitions in R. Journal of Statistical Software, 23(CS-2):1– 12, December 2007. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v23/c02.

Haslbeck:2020:MET

[HW20] Jonas M. B. Haslbeck and Lourens J. Waldorp. mgm: Estimating time-varying mixed graphical models in highdimensional data. Journal of Statistical Software, 93 (??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v093i08;   https://www.jstatsoft.org/
index.php/jss/article/view/v093i08/v93i08.pdf.

Hasan:2016:FEM

[HWM16] Asad Hasan, Zhiyu Wang, and Alireza S. Mahani. Fast estimation of multinomial logit models: R package mnlogit. Journal of Statistical Software, 75(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v075i03;   https://www.
jstatsoft.org/index.php/jss/article/view/v075i03/v75i03.
pdf.

Hocking:2013:SED

[HWPG13] Toby Dylan Hocking, Thomas Wutzler, Keith Ponting, and Philippe Grosjean. Sustainable, extensible documentation generation using inlinedocs. Journal of Statistical Software, 54 (6):??, September 2013. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v54/i06.

Hyun:2018:VRP

[HWY18] Seung Won Hyun, Weng Kee Wong, and Yarong Yang. VNM: An R package for finding multiple-objective optimal designs for the 4-parameter logistic model. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v083i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v083i05/v83i05.pdf.

Hainmueller:2013:ESP

[HX13] Jens Hainmueller and Yiqing Xu. ebalance: A Stata package for entropy balancing. Journal of Statistical Software, 54(7): ??, September 2013. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v54/i07.

He:2012:CSS

[HXY12] Wenqing He, Juan Xiong, and Grace Y. Yi. Code snippet: SIMEX R package for accelerated failure time models with covariate measurement error. Journal of Statistical Software, 46(CS-1):??, January 2012. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v46/c01.

Huling:2021:SIU

[HY21] Jared D. Huling and Menggang Yu. Subgroup identification using the personalized package. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v098i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v098i05/v98i05.pdf.

Happ:2020:CSP

[HZBB20] Martin Happ, Georg Zimmermann, Edgar Brunner, and Arne C. Bathke. Code snippet: Pseudo-ranks: How to calculate them efficiently in R. Journal of Statistical Software, 95(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v095c01;   https://www.jstatsoft.org/
index.php/jss/article/view/v095c01/v95c01.pdf.

Hocevar:2021:CPO

[HZS21] Tomaz Hocevar, Blaz Zupan, and Jonna Stålring. Conformal prediction with Orange. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v098i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v098i07/v98i07.pdf.

Hu:2010:COD

[HZSW10] Jiaqiao Hu, Wei Zhu, Yi Su, and Weng Kee Wong. Controlled optimal design program for the logit dose response model. Journal of Statistical Software, 35(6):??, July 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v35/i06.

Iacus:2015:BRA

[Iac15] Stefano M. Iacus. Book review: Automated Data Collection with R — A Practical Guide to Web Scraping and Text Mining, Simon Munzert, Christian Rubba, Peter Meißner, Dominic Nyhuis, John Wiley & Sons, Chichester, 2014. ISBN 978-1-118-83481-7. 480 pp. GBP 50.00 (H).

http://www.r-datacollection.com/. Journal of Statistical

Software, 68(BR-3):??, ???? 2015. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/index.php/
jss/article/view/v068b03;   http://www.jstatsoft.org/
index.php/jss/article/view/v068b03/v68b03.pdf.

Iacus:2017:BRB

[Iac17] Stefano M. Iacus. Book review: Big Data and Social Science — A Practical Guide to Methods and Tools. Journal of Statistical Software, 78(??):??, ???? 2017. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v078b02;   https://www.jstatsoft.
org/index.php/jss/article/view/v078b02/v78b02.pdf.

Irurozki:2016:PRP

[ICL16] Ekhine Irurozki, Borja Calvo, and Jose A. Lozano. PerMallows: An R package for Mallows and generalized Mallows models. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v071i12;   https://
www.jstatsoft.org/index.php/jss/article/view/v071i12/
v71i12.pdf.

Iacus:2009:CSC

[IKP09] Stefano M. Iacus, Gary King, and Giuseppe Porro. cem: Software for coarsened exact matching. Journal of Statistical Software, 30(9):??, June 2009. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v30/i09.

Imai:2011:ERP

[ILS11] Kosuke Imai, Ying Lu, and Aaron Strauss. eco: R package for ecological inference in 2 × 2 tables. Journal of Statistical Software, 42(5):??, June 2011. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v42/i05.

Iversen:2005:VED

[IM05] Philip W. Iversen and Mervyn G. Marasinghe. Visualizing experimental designs for balanced ANOVA models. Journal of Statistical Software, 13(3):1–18, ???? 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=117&url=v13/i03&ct=2;   http://www.
jstatsoft.org/counter.php?id=117&url=v13/i03/v13i03.
pdf&ct=1.

Fernandez-i-Marin:2016:GAM

[iM16] Xavier Fernández i Marín. ggmcmc: Analysis of MCMC samples and Bayesian inference. Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v070i09;   https://www.jstatsoft.org/
index.php/jss/article/view/v070i09/v70i09.pdf.

Iacus:2017:CPQ

[IMR17] Stefano M. Iacus, Lorenzo Mercuri, and Edit Rroji. COGARCH(p, q): Simulation and inference with the yuima package. Journal of Statistical Software, 80(??):??, ???? 2017. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v080i04;   https://www.
jstatsoft.org/index.php/jss/article/view/v080i04/v80i04.
pdf.

Iam-on:2010:LMP

[IoG10] Natthakan Iam-on and Simon Garrett. LinkCluE: a MATLAB package for link-based cluster ensembles. Journal of Statistical Software, 36(9):??, August 2010. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v36/i09.

Iacus:2008:IMF

[IP08] Stefano M. Iacus and Giuseppe Porro. Invariant and metric free proximities for data matching: An R package. Journal of Statistical Software, 25(11):1–22, April 2008. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v25/i11.

Imai:2005:PMR

[IV05] Kosuke Imai and David A. Van Dyk. MNP: R package for fitting the Multinomial Probit Model. Journal of Statistical Software, 14(3):1–32, May 8, 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=128&url=v14/i03&ct=2;   http://www.
jstatsoft.org/counter.php?id=128&url=v14/i03/v14i03.
pdf&ct=1.

Jackson:2011:MSM

[Jac11] Christopher H. Jackson. Multi-state models for panel data: The msm package for R. Journal of Statistical Software, 38(8): ??, January 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v38/i08.

Jackson:2016:FPP

[Jac16] Christopher Jackson. flexsurv: A platform for parametric survival modeling in R. Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v070i08;   https://www.jstatsoft.org/
index.php/jss/article/view/v070i08/v70i08.pdf.

Jalilian:2019:CSE

[Jal19] Abdollah Jalilian. Code snippet: ETAS: An R package for fitting the space–time ETAS model to earthquake data. Journal of Statistical Software, 88(??):??, ???? 2019. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v088c01;   https://www.
jstatsoft.org/index.php/jss/article/view/v088c01/v88c01.
pdf.

Jamshidian:1999:ARR

[Jam99] Mortaza Jamshidian. Adaptive robust regression by using a nonlinear regression program. Journal of Statistical Software, 4(6):1–25, ???? 1999. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v04/i06;   http://www.jstatsoft.org/v04/i06/code;
http://www.jstatsoft.org/v04/i06/tdistIIr1.pdf;   http:
//www.jstatsoft.org/v04/i06/updates.

Jo:2019:BRP

[JCPL19] Seongil Jo, Taeryon Choi, Beomjo Park, and Peter Lenk. bsamGP: an R package for Bayesian spectral analysis models using Gaussian process priors. Journal of Statistical Software, 90(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v090i10;   https://www.jstatsoft.org/
index.php/jss/article/view/v090i10/v90i10.pdf.

Jing:2015:PGR

[JD15] Liang Jing and Victor De Oliveira. geoCount: An R package for the analysis of geostatistical count data. Journal of Statistical Software, 63(11):??, February 2015. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v63/i11.

Jay:2015:PSP

[JFDB15] Flora Jay, Olivier François, Eric Y. Durand, and Michael G. B. Blum. POPS: [a] software for prediction of population genetic structure using latent regression models. Journal of Statistical Software, 68(9):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v068i09;
http://www.jstatsoft.org/index.php/jss/article/view/
v068i09/v68i09.pdf.

Jenness:2018:ERP

[JGM18] Samuel M. Jenness, Steven M. Goodreau, and Martina Morris. EpiModel: An R package for mathematical modeling of infectious disease over networks. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v084i08;   https://www.jstatsoft.org/
index.php/jss/article/view/v084i08/v84i08.pdf.

Josse:2016:MPH

[JH16] Julie Josse and François Husson. missMDA: A package for handling missing values in multivariate data analysis. Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v070i01;   https://www.
jstatsoft.org/index.php/jss/article/view/v070i01/v70i01.
pdf.

Jara:2011:DBS

[JHQ^{+}11] Alejandro Jara, Timothy E. Hanson, Fernando A. Quintana, Peter Müller, and Gary L. Rosner. DPpackage: Bayesian semiand nonparametric modeling in R. Journal of Statistical Software, 40(5):??, April 2011. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v40/i05.

Jamshidian:2014:PMR

[JJJ14] Mortaza Jamshidian, Siavash Jalal, and Camden Jansen. MissMech: An R package for testing homoscedasticity, multivariate normality, and missing completely at random (MCAR). Journal of Statistical Software, 56(6):??, January 2014. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v56/i06.

Jordan:2019:EPF

[JKL19] Alexander Jordan, Fabian Krüger, and Sebastian Lerch. Evaluating probabilistic forecasts with scoringRules. Journal of Statistical Software, 90(??):??, ???? 2019. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v090i12;   https://www.
jstatsoft.org/index.php/jss/article/view/v090i12/v90i12.
pdf.

Jiang:2020:WRB

[JLH^{+}20] Yujing Jiang, Mei-Ling Ting Lee, Xin He, Bernard Rosner, and Jun Yan. Wilcoxon rank-based tests for clustered data with R package clusrank. Journal of Statistical Software, 96(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v096i06;   https://www.jstatsoft.org/
index.php/jss/article/view/v096i06/v96i06.pdf.

Jamshidian:2005:SSI

[JLZJ05] Mortaza Jamshidian, Wei Liu, Ying Zhang, and Farid Jamshidian. Simreg: a software including some new developments in multiple comparison and simultaneous confidence bands for linear regression models. Journal of Statistical Software, 12(2):1–22, ???? 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=109&url=v12/i02&ct=2;   http://www.
jstatsoft.org/counter.php?id=109&url=v12/i02/v12i02.
pdf&ct=1.

James:2015:PER

[JM15] Nicholas A. James and David S. Matteson. ecp: An R package for nonparametric multiple change point analysis of multivariate data. Journal of Statistical Software, 62(7):??, January 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v62/i07.

Johnson:2008:SIS

[JMD08] Kjell Johnson, Abhyuday Mandal, and Tan Ding. Software for implementing the sequential elimination of level combinations algorithm. Journal of Statistical Software, 25(6):1–13, March 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v25/i06.

Johnson:2009:PSI

[JMS^{+}09] Jacqueline L. Johnson, Keith E. Muller, James C. Slaughter, Matthew J. Gurka, Matthew J. Gribbin, and Sean L. Simpson. POWERLIB: SAS/IML software for computing power in multivariate linear models. Journal of Statistical Software, 30 (5):??, April 2009. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v30/i05.

Johnson:2007:MML

[Joh07] Matthew S. Johnson. Marginal maximum likelihood estimation of item response models in R. Journal of Statistical Software, 20(10):1–24, February 2007. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v20/i10.

Johansen:2009:SSM

[Joh09] Adam M. Johansen. SMCTC: Sequential Monte Carlo in C++. Journal of Statistical Software, 30(6):??, April 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v30/i06.

Jones:2007:DPC

[Jon07] James Holland Jones. demogR: a package for the construction and analysis of age-structured demographic models in R. Journal of Statistical Software, 22(10):1–28, September 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v22/i10.

Josselin:2005:IGI

[Jos05] Didier Josselin. Interactive geographical information system using Lisp–Stat: Prototypes and applications. Journal of Statistical Software, 13(6):1–34, ???? 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=122&url=v13/i06&ct=2;   http://www.
jstatsoft.org/counter.php?id=122&url=v13/i06/v13i06.
pdf&ct=1.

Jelsema:2016:CRP

[JP16] Casey M. Jelsema and Shyamal D. Peddada. CLME: An R package for linear mixed effects models under inequality constraints. Journal of Statistical Software, 75 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v075i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v075i01/v75i01.pdf.

Jaki:2019:RPM

[JPM19] Thomas Jaki, Philip Pallmann, and Dominic Magirr. The R package MAMS for designing multi-arm multi-stage clinical trials. Journal of Statistical Software, 88(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v088i04;
https://www.jstatsoft.org/index.php/jss/article/view/
v088i04/v88i04.pdf.

Johnstone:2005:ERP

[JS05] Iain M. Johnstone and Bernard W. Silverman. EbayesThresh: R programs for empirical Bayes thresholding. Journal of Statistical Software, 12(8):1–38, ???? 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=125&url=v12/i08&ct=2;   http://www.
jstatsoft.org/counter.php?id=125&url=v12/i08/v12i08.
pdf&ct=1.

Jasen:2013:BRB

[JS13] Kelcey Jasen and Luigi Salmaso. Book review: Applied Medical Statistics Using SAS. Journal of Statistical Software, 52(BR-2):??, February 2013. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v52/b02.

Juretig:2015:CSP

[Jur15] Francisco Juretig. Code snippet: MLGA: A SAS macro to compute maximum likelihood estimators via genetic algorithms. Journal of Statistical Software, 66(CS-2):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066c02;
http://www.jstatsoft.org/index.php/jss/article/view/
v066c02/v66c02.pdf.

Jachner:2007:SMQ

[JvdBP07] Stefanie Jachner, K. Gerald van den Boogaart, and Thomas Petzoldt. Statistical methods for the qualitative assessment of dynamic models with time delay (R package qualV). Journal of Statistical Software, 22(8):1–30, September 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v22/i08.

Jirjies:2016:PPI

[JWHS16] Saman Jirjies, Garrick Wallstrom, Rolf U. Halden, and Matthew Scotch. pyJacqQ: Python implementation of Jacquez’s Q-statistics for space-time clustering of disease exposure in case-control studies. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v074i06;   https://www.jstatsoft.org/
index.php/jss/article/view/v074i06/v74i06.pdf.

Knoke:2005:ARM

[KAK05] James D. Knoke, Christy M. Anderson, and Gary G. Koch. Analyzing repeated measures marginal models on sample surveys with resampling methods. Journal of Statistical Software, 15(8):1–13, December 2005. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v15/i08.

Kampstra:2008:CSB

[Kam08] Peter Kampstra. Code snippet: Beanplot: a boxplot alternative for visual comparison of distributions. Journal of Statistical Software, 28(CS-1):??, November 2008. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v28/c01.

Kao:2009:MOO

[Kao09] Ming-Hung Kao. Multi-objective optimal experimental designs for ER-fMRI using MATLAB. Journal of Statistical Software, 30(11):??, June 2009. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v30/i11.

Karlsson:2006:SRS

[Kar06] Andreas Karlsson. Software review: Scientific Workplace 5.5 and LyX 1.4.2. Journal of Statistical Software, 17(SR-1):1– 11, October 2006. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v17/s01.

Kastner:2016:DSV

[Kas16] Gregor Kastner. Dealing with stochastic volatility in time series using the R package stochvol. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v069i05;   https://www.jstatsoft.
org/index.php/jss/article/view/v069i05/v69i05.pdf.

Kavroudakis:2015:SRP

[Kav15] Dimitris Kavroudakis. sms: An R package for the construction of microdata for geographical analysis. Journal of Statistical Software, 68(2):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v068i02;
http://www.jstatsoft.org/index.php/jss/article/view/
v068i02/v68i02.pdf.

Kapelner:2016:BML

[KB16] Adam Kapelner and Justin Bleich. bartMachine: Machine learning with Bayesian additive regression trees. Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v070i04;   https://www.
jstatsoft.org/index.php/jss/article/view/v070i04/v70i04.
pdf.

Kuznetsova:2017:LPT

[KBC17] Alexandra Kuznetsova, Per B. Brockhoff, and Rune H. B. Christensen. lmerTest package: Tests in linear mixed effects models. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v082i13;
https://www.jstatsoft.org/index.php/jss/article/view/
v082i13/v82i13.pdf.

Kharrat:2019:FRM

[KBMB19] Tarak Kharrat, Georgi N. Boshnakov, Ian McHale, and Rose Baker. Flexible regression models for count data based on renewal processes: The Countr package. Journal of Statistical Software, 90(??):??, ???? 2019. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v090i13;   https://www.jstatsoft.
org/index.php/jss/article/view/v090i13/v90i13.pdf.

King:2011:MCM

[KCCG11] Martin D. King, Fernando Calamante, Chris A. Clark, and David G. Gadian. Markov Chain Monte Carlo random effects modeling in magnetic resonance image processing using the BRugs interface to WinBUGS. Journal of Statistical Software, 44(2):??, October 2011. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v44/i02.

Kairalla:2008:GSI

[KCM08] John A. Kairalla, Christopher S. Coffey, and Keith E. Muller. GLUMIP 2.0: SAS/IML software for planning internal pilots. Journal of Statistical Software, 28(7):??, November 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v28/i07.

Killick:2014:PCR

[KE14] Rebecca Killick and Idris A. Eckley. changepoint: An R package for changepoint analysis. Journal of Statistical Software, 58(3):??, June 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v58/i03.

Kelley:2007:CIS

[Kel07] Ken Kelley. Confidence intervals for standardized effect sizes: Theory, application, and implementation. Journal of Statistical Software, 20(8):1–24, February 2007. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v20/i08.

Kane:2013:SSC

[KEW13] Michael J. Kane, John W. Emerson, and Stephen Weston. Scalable strategies for computing with massive data. Journal of Statistical Software, 55(14):??, November 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v55/i14.

Koenker:2017:RRP

[KG17] Roger Koenker and Jiaying Gu. REBayes: An R package for empirical Bayes mixture methods. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v082i08;   https://www.jstatsoft.
org/index.php/jss/article/view/v082i08/v82i08.pdf.

Kurz:2019:DSF

[KGP^{+}19] Gerhard Kurz, Igor Gilitschenski, Florian Pfaff, Lukas Drude, Uwe D. Hanebeck, Reinhold Haeb-Umbach, and Roland Y. Siegwart. Directional statistics and filtering using libDirectional. Journal of Statistical Software, 89 (??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v089i04;   https://www.jstatsoft.org/
index.php/jss/article/view/v089i04/v89i04.pdf.

Krivitsky:2008:FLC

[KH08] Pavel N. Krivitsky and Mark S. Handcock. Fitting latent cluster models for networks with latentnet. Journal of Statistical Software, 24(5):1–23, May 2008. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v24/i05.

Kuiper:2013:FPG

[KH13] Rebecca M. Kuiper and Herbert Hoijtink. A Fortran 90 program for the generalized order-restricted information criterion. Journal of Statistical Software, 54(8):??, September 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v54/i08.

Khan:2003:VBS

[Kha03] Haseeb Ahmad Khan. A Visual Basic software for computing Fisher’s exact probability. Journal of Statistical Software, 8 (21):1–7, December 28, 2003. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v08/i21/Article-
JSS.PDF.

Khademi:2015:BRSb

[Kha15a] Abdolvahab Khademi. Book review: Statistical Hypothesis Testing with SAS and R, Dirk Taeger and Sonja Kuhnt, John Wiley & Sons, Chichester, 2014. ISBN 978-1-119-95021-9. xiv

+ 292 pp. USD 97.95 (H). http://www.wiley.com/WileyCDA/WileyTitle/productCd-1118762614.html.

Journal of Statistical Software, 68(BR-2):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v068b02;
http://www.jstatsoft.org/index.php/jss/article/view/
v068b02/v68b02.pdf.

Khademi:2015:BRSa

[Kha15b] Abdolvahab Khademi. Book review: Statistical Modeling and Computation, Dirk P. Kroese and Joshua C. C. Chan, Springer-Verlag, New York, 2014. ISBN 978-1-4614-8774-6. xx

+ 400 pp. USD 79.99 (H). http://www.springer.com/book/9781461487746.

Journal of Statistical Software, 66(BR-3):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066b03;
http://www.jstatsoft.org/index.php/jss/article/view/
v066b03/v66b03.pdf.

Khademi:2016:BRA

[Kha16a] Abdolvahab Khademi. Book review: Applied Univariate, Bivariate, and Multivariate Statistics. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v072b02;   https://www.jstatsoft.org/
index.php/jss/article/view/v072b02/v72b02.pdf.

Khademi:2016:BRB

[Kha16b] Abdolvahab Khademi. Book review: Basics of Matrix Algebra for Statistics with R, Nick Fieller, Chapman & Hall/CRC, Boca Raton, 2016. ISBN 978-1-4987-1236-1. xviii + 226 pp.

USD 59.95 (H). https://www.crcpress.com/9781498712361.

Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v071b01;   https://
www.jstatsoft.org/index.php/jss/article/view/v071b01/
v71b01.pdf.

Khademi:2016:BRM

[Kha16c] Abdolvahab Khademi. Book review: Multivariate Statistics: Exercises and Solutions, 2nd Edition. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v069b01;   https://www.jstatsoft.
org/index.php/jss/article/view/v069b01/v69b01.pdf.

Khademi:2017:BRH

[Kha17a] Abdolvahab Khademi. Book review: Hidden Markov Models for Time Series: An Introduction Using R (2nd Edition). Journal of Statistical Software, 80(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v080b01;   https://
www.jstatsoft.org/index.php/jss/article/view/v080b01/
v80b01.pdf.

Khademi:2017:BRM

[Kha17b] Abdolvahab Khademi. Book review: Multivariate Analysis with LISREL. Journal of Statistical Software, 78 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v078b01;   https://www.jstatsoft.org/
index.php/jss/article/view/v078b01/v78b01.pdf.

Khademi:2018:BRCa

[Kha18a] Abdolvahab Khademi. Book review: Complex Survey Data Analysis with SAS. Journal of Statistical Software, 84 (??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v084b02;   https://www.jstatsoft.org/
index.php/jss/article/view/v084b02/v84b02.pdf.

Khademi:2018:BRCb

[Kha18b] Abdolvahab Khademi. Book review: Computational Methods for Numerical Analysis with R. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v087b02;   https://www.jstatsoft.org/
index.php/jss/article/view/v087b02/v87b02.pdf.

Khademi:2018:BRS

[Kha18c] Abdolvahab Khademi. Book review: Statistical Intervals: A Guide for Practitioners and Researchers (2nd Edition). Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v084b01;   https://www.
jstatsoft.org/index.php/jss/article/view/v084b01/v84b01.
pdf.

Khademi:2020:BRF

[Kha20] Abdolvahab Khademi. Book review: Flexible Imputation of Missing Data (2nd edition). Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v093b01;   https://www.jstatsoft.org/
index.php/jss/article/view/v093b01/v93b01.pdf.

Kahm:2010:GFB

[KHLF^{+}10] Matthias Kahm, Guido Hasenbrink, Hella Lichtenberg-Fraté, Jost Ludwig, and Maik Kschischo. grofit: Fitting biological growth curves with R. Journal of Statistical Software, 33(7): ??, February 2010. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v33/i07.

Kiermeier:2008:VAA

[Kie08] Andreas Kiermeier. Visualizing and assessing acceptance sampling plans: The R package AcceptanceSampling. Journal of Statistical Software, 26(6):??, July 2008. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v26/i06.

Killick:2016:BRI

[Kil16] Rebecca Killick. Book review: Introductory Statistics and Analytics: A Resampling Perspective. Journal of Statistical Software, 75(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v075b01;   https://www.jstatsoft.
org/index.php/jss/article/view/v075b01/v75b01.pdf.

King:2004:TPE

[Kin04] Gary King. EI: a program for ecological inference. Journal of Statistical Software, 11(7):1–41, ???? 2004. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/counter.
php?id=103&url=v11/i07/v11i07.pdf&ct=1.

Korzen:2014:PPP

[KJ14] Marcin Korzeń and Szymon Jaroszewicz. PaCAL: A Python package for arithmetic computations with random variables. Journal of Statistical Software, 57(10):??, May 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v57/i10.

Komarek:2014:CRP

[KK14] Arnost Komárek and Lenka Komárková. Capabilities of R package mixAK for clustering based on multivariate continuous and discrete longitudinal data. Journal of Statistical Software, 59(12):??, September 2014. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v59/i12.

Kawaguchi:2015:PSR

[KK15] Atsushi Kawaguchi and Gary G. Koch. sanon: An R package for stratified analysis with nonparametric covariable adjustment. Journal of Statistical Software, 67(9):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067i09;
http://www.jstatsoft.org/index.php/jss/article/view/
v067i09/v67i09.pdf.

Kuiper:2010:FPC

[KKH10] Rebecca M. Kuiper, Irene Klugkist, and Herbert Hoijtink. A Fortran 90 program for confirmatory analysis of variance. Journal of Statistical Software, 34(8):??, May 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v34/i08.

Karlsson:2014:PRP

[KL14] Maria Karlsson and Anita Lindmark. truncSP: An R package for estimation of semi-parametric truncated linear regression models. Journal of Statistical Software, 57(14):??, May 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v57/i14.

Klasson:2018:CSQ

[Kla18] K. Thomas Klasson. Code snippet: QXLA: Adding upper quantiles for the studentized range to Excel for multiple comparison procedures. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v085c01;   https://www.jstatsoft.org/
index.php/jss/article/view/v085c01/v85c01.pdf.

Kley:2016:QBS

[Kle16] Tobias Kley. Quantile-based spectral analysis in an objectoriented framework and a reference implementation in R: The quantspec package. Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v070i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v070i03/v70i03.pdf.

Knight:2020:GNA

[KLNN20] Marina Knight, Kathryn Leeming, Guy Nason, and Matthew Nunes. Generalized network autoregressive processes and the GNAR package. Journal of Statistical Software, 96 (??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v096i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v096i05/v96i05.pdf.

Knoblauch:2008:MML

[KM08] Kenneth Knoblauch and Laurence T. Maloney. MLDS: Maximum likelihood difference scaling in R. Journal of Statistical Software, 25(2):1–26, March 2008. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v25/i02.

Koenker:2014:COR

[KM14] Roger Koenker and Ivan Mizera. Convex optimization in R. Journal of Statistical Software, 60(5):??, September 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v60/i05.

Kalema:2016:CSG

[KM16] George Kalema and Geert Molenberghs. Code snippet: Generating correlated and/or overdispersed count data: A SAS implementation. Journal of Statistical Software, 70 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v070c01;   https://www.jstatsoft.org/
index.php/jss/article/view/v070c01/v70c01.pdf.

Kalisch:2012:CIU

[KMC^{+}12] Markus Kalisch, Martin Mächler, Diego Colombo, Marloes H. Maathuis, and Peter Bühlmann. Causal inference using graphical models with the R package pcalg. Journal of Statistical Software, 47(11):??, May 2012. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v47/i11.

Kreidler:2013:GOP

[KMG^{+}13] Sarah M. Kreidler, Keith E. Muller, Gary K. Grunwald, Brandy M. Ringham, Zacchary T. Coker-Dukowitz, Uttara R. Sakhadeo, Anna E. Barón, and Deborah H. Glueck. GLIMMPSE: Online power computation for linear models with and without a baseline covariate. Journal of Statistical Software, 54(10):??, September 2013. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v54/i10.

Karatzoglou:2006:SVM

[KMH06] Alexandros Karatzoglou, David Meyer, and Kurt Hornik. Support vector machines in R. Journal of Statistical Software, 15 (9):1–28, April 2006. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v15/i09.

Krol:2017:TJM

[KMM^{+}17] Agnieszka Król, Audrey Mauguen, Yassin Mazroui, Alexandre Laurent, Stefan Michiels, and Virginie Rondeau. Tutorial in joint modeling and prediction: A statistical software for correlated longitudinal outcomes, recurrent events and a terminal event. Journal of Statistical Software, 81 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v081i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v081i03/v81i03.pdf.

Kowarik:2014:SAR

[KMTS14] Alexander Kowarik, Angelika Meraner, Matthias Templ, and Daniel Schopfhauser. Seasonal adjustment with the R packages x12 and x12GUI. Journal of Statistical Software, 62(2):

??, December 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v62/i02.

Koenker:2003:SSM

[KN03] Roger Koenker and Pin Ng. SparseM: a sparse matrix package for R. Journal of Statistical Software, 8(6):1–9, ???? 2003. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v08/i06/SparseM.pdf;   http://
www.jstatsoft.org/v08/i6.

Karlis:2005:BPD

[KN05] Dimitris Karlis and Ioannis Ntzoufras. Bivariate Poisson and diagonal inflated bivariate Poisson regression models in R. Journal of Statistical Software, 14(10):1–36, September 5, 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/counter.php?id=135&url=v14/i10&ct=2;
 http://www.jstatsoft.org/counter.php?id=135&url=v14/
i10/v14i10.pdf&ct=1.

Kirchkamp:2017:BRI

[KN17] Oliver Kirchkamp and Hiltrud Niggemann. Book review: An Introduction to Stata Programming (2nd Edition). Journal of Statistical Software, 77(??):??, ???? 2017. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v077b03;   https://www.
jstatsoft.org/index.php/jss/article/view/v077b03/v77b03.
pdf.

Kristensen:2016:TAD

[KNB^{+}16] Kasper Kristensen, Anders Nielsen, Casper W. Berg, Hans Skaug, and Bradley M. Bell. TMB: Automatic differentiation and Laplace approximation. Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v070i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v070i05/v70i05.pdf.

King:2016:SIP

[KNI16] Aaron A. King, Dao Nguyen, and Edward L. Ionides. Statistical inference for partially observed Markov processes via the R package pomp. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/

article/view/v069i12;   https://www.jstatsoft.org/
index.php/jss/article/view/v069i12/v69i12.pdf.

Kim:2006:CRP

[KO06] Donghoh Kim and Hee-Seok Oh. CVTresh: R package for level-dependent cross-validation thresholding. Journal of Statistical Software, 15(10):1–13, April 2006. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v15/i10.

Koenker:2008:CQR

[Koe08] Roger Koenker. Censored quantile regression redux. Journal of Statistical Software, 27(6):??, July 2008. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v27/i06.

Koller:2016:RRP

[Kol16] Manuel Koller. robustlmm: An R package for robust estimation of linear mixed-effects models. Journal of Statistical Software, 75(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v075i06;   https://www.jstatsoft.
org/index.php/jss/article/view/v075i06/v75i06.pdf.

Kontopantelis:2013:CSG

[Kon13] Evangelos Kontopantelis. Code snippet: A greedy algorithm for representative sampling: repsample in Stata. Journal of Statistical Software, 55(CS-1):??, November 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v55/c01.

Kahle:2020:CAS

[KOS20] David Kahle, Christopher O’Neill, and Jeff Sommars. A computer algebra system for R: Macaulay2 and the m2r package. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v093i09;   https://
www.jstatsoft.org/index.php/jss/article/view/v093i09/
v93i09.pdf.

Kneib:2007:ISV

[KP07] Thomas Kneib and Thomas Petzoldt. Introduction to the special volume on “Ecology and Ecological Modeling in R”. Journal of Statistical Software, 22(1):1–7, September 2007.

CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v22/i01.

Kunzmann:2021:APA

[KPH^{+}21] Kevin Kunzmann, Maximilian Pilz, Carolin Herrmann, Geraldine Rauch, and Meinhard Kieser. The adoptr package: Adaptive optimal designs for clinical trials in R. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v098i09;   https://www.jstatsoft.
org/index.php/jss/article/view/v098i09/v98i09.pdf.

Kreutzmann:2019:RPE

[KPRP^{+}19] Ann-Kristin Kreutzmann, Sören Pannier, Natalia Rojas-Perilla, Timo Schmid, Matthias Templ, and Nikos Tzavidis. The R package emdi for estimating and mapping regionally disaggregated indicators. Journal of Statistical Software, 91(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v091i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v091i07/v91i07.pdf.

Konietschke:2015:PNR

[KPSH15] Frank Konietschke, Marius Placzek, Frank Schaarschmidt, and Ludwig A. Hothorn. nparcomp: An R software package for nonparametric multiple comparisons and simultaneous confidence intervals. Journal of Statistical Software, 64(9):??, March 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v64/i09.

Kontopantelis:2009:MMA

[KR09] Evangelos Kontopantelis and David Reeves. MetaEasy: a meta-analysis add-in for Microsoft Excel. Journal of Statistical Software, 30(7):??, April 2009. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v30/i07.

Kohl:2010:RPD

[KR10a] Matthias Kohl and Peter Ruckdeschel. R package distrMod: S4 classes and methods for probability models. Journal of Statistical Software, 35(10):??, July 2010. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v35/i10.

Kursa:2010:FSB

[KR10b] Miron B. Kursa and Witold R. Rudnicki. Feature selection with the Boruta package. Journal of Statistical Software, 36 (11):??, September 2010. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v36/i11.

Kuhnt:2013:SOM

[KR13] Sonja Kuhnt and Nikolaus Rudak. Simultaneous optimization of multiple responses with the R package JOP. Journal of Statistical Software, 54(9):??, September 2013. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v54/i09.

Krastev:2007:SSS

[Kra07] Teodor Krastev. spectrino software: Spectra visualization and preparation for R. Journal of Statistical Software, 18(10): 1–16, January 2007. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v18/i10.

Kroonenberg:2016:MMA

[Kro16] Pieter M. Kroonenberg. My multiway analysis: From Jan de Leeuw to TWPack and back. Journal of Statistical Software, 73(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v073i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v073i03/v73i03.pdf.

Kim:2010:PQS

[KS10] Hanjoo Kim and Justine Shults. %QLS SAS macro: a SAS macro for analysis of correlated data using quasi-least squares. Journal of Statistical Software, 35(2):??, July 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v35/i02.

Kenkel:2014:EEF

[KS14] Brenton Kenkel and Curtis S. Signorino. Estimating extensive form games in R. Journal of Statistical Software, 56(8):??, January 2014. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v56/i08.

Kondo:2016:RRP

[KSBZ16] Yumi Kondo, Matias Salibian-Barrera, and Ruben Zamar. RSKC: An R package for a robust and sparse K-means clustering algorithm. Journal of Statistical Software, 72 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v072i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v072i05/v72i05.pdf.

Karatzoglou:2004:PKS

[KSHZ04] Alexandros Karatzoglou, Alex Smola, Kurt Hornik, and Achim Zeileis. kernlab — an S4 package for kernel methods in R. Journal of Statistical Software, 11(9):1–20, ???? 2004. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/counter.php?id=105&url=v11/i09/
v11i09.pdf&ct=1.

Krl:2015:SRP

[KSP15] Agnieszka Król and Philippe Saint-Pierre. SemiMarkov: An R package for parametric estimation in multi-state semi-Markov models. Journal of Statistical Software, 66(6):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066i06;
http://www.jstatsoft.org/index.php/jss/article/view/
v066i06/v66i06.pdf.

Kontopantelis:2016:SBP

[KSPR16] Evangelos Kontopantelis, David A. Springate, Rosa Parisi, and David Reeves. Simulation-based power calculations for mixed effects modeling: ipdpower in Stata. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v074i12;   https://www.jstatsoft.
org/index.php/jss/article/view/v074i12/v74i12.pdf.

Kirchner:2007:ARP

[KSS^{+}07] Marc Kirchner, Benjamin Saussen, Hanno Steen, Judith A. J. Steen, and Fred A. Hamprecht. amsrpm: Robust point matching for retention time alignment of LC/MS data with R. Journal of Statistical Software, 18(4):1–12, January 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v18/i04.

Kugiumtzis:2010:MAT

[KT10] Dimitris Kugiumtzis and Alkiviadis Tsimpiris. Measures of analysis of time series (MATS): a MATLAB toolkit for computation of multiple measures on time series data bases. Journal of Statistical Software, 33(5):??, February 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v33/i05.

Kowarik:2016:IRP

[KT16] Alexander Kowarik and Matthias Templ. Imputation with the R package VIM. Journal of Statistical Software, 74 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v074i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v074i07/v74i07.pdf.

Kuhn:2008:BPM

[Kuh08] Max Kuhn. Building predictive models in R using the caret package. Journal of Statistical Software, 28(5):??, November 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v28/i05.

Kukuyeva:2009:BRB

[Kuk09] Irina Kukuyeva. Book review: Principles of Modeling and Simulation: A Multidisciplinary Approach. Journal of Statistical Software, 31(BR-3):??, September 2009. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v31/b03.

Kuonen:2003:NIP

[Kuo03] Diego Kuonen. Numerical integration in S-PLUS or R: a survey. Journal of Statistical Software, 8(13):1–14, ???? 2003. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v08/i13;   http://www.jstatsoft.
org/v08/i13/article.pdf.

Kursa:2014:PIR

[Kur14] Miron B. Kursa. rFerns: An implementation of the random ferns method for general-purpose machine learning. Journal of Statistical Software, 61(10):??, November 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v61/i10.

Kovalchik:2013:FAB

[KV13] Stephanie Kovalchik and Ravi Varadhan. Fitting additive binomial regression models with the R package blm. Journal of Statistical Software, 54(1):??, September 2013. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v54/i01.

Kruijswijk:2020:SEB

[KvEPK20] Jules Kruijswijk, Robin van Emden, Petri Parvinen, and Maurits Kaptein. StreamingBandit: Experimenting with bandit policies. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v094i09;
https://www.jstatsoft.org/index.php/jss/article/view/
v094i09/v94i09.pdf.

Kieser:2017:ORP

[KWE^{+}17] Meinhard Kieser, Marius Wirths, Stefan Englert, Cornelia Ursula Kunz, and Geraldine Rauch. OneArmPhaseTwoStudy: An R package for planning, conducting, and analysing single-arm phase II studies. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v081i08;   https://www.jstatsoft.org/
index.php/jss/article/view/v081i08/v81i08.pdf.

Kojadinovic:2010:MMD

[KY10] Ivan Kojadinovic and Jun Yan. Modeling multivariate distributions with continuous margins using the copula R package. Journal of Statistical Software, 34(9):??, May 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v34/i09.

King:2003:LRR

[KZ03] Gary King and Langche Zeng. Logistic regression in rare events data. Journal of Statistical Software, 8(2):1–30, ???? 2003. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v08/i02/0s.pdf;   http://
www.jstatsoft.org/v08/i2.

Lagani:2017:FSR

[LAF^{+}17] Vincenzo Lagani, Giorgos Athineou, Alessio Farcomeni, Michail Tsagris, and Ioannis Tsamardinos. Feature selection with the R package MXM: Discovering statistically equivalent feature subsets. Journal of Statistical Software, 80(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v080i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v080i07/v80i07.pdf.

Lalanne:2016:BRS

[Lal16] Christophe Lalanne. Book review: Statistical Data Analytics. Foundations for Data Mining, Informatics, and Knowledge Discovery. Journal of Statistical Software, 69 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v069b03;   https://www.jstatsoft.org/
index.php/jss/article/view/v069b03/v69b03.pdf.

Lalanne:2017:BRS

[Lal17] Christophe Lalanne. Book review: Statistical Analysis of Questionnaires: A Unified Approach Based on R and Stata. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v082b01;   https://
www.jstatsoft.org/index.php/jss/article/view/v082b01/
v82b01.pdf.

Lamigueiro:2012:SSR

[Lam12] Oscar Perpiñán Lamigueiro. solaR: Solar radiation and photovoltaic systems with R. Journal of Statistical Software, 50 (9):??, August 2012. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v50/i09.

Lang:2014:ADA

[Lan14] Albert-Georg Lang. ART: A data aggregation program for the behavioral sciences. Journal of Statistical Software, 59(3): ??, August 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v59/i03.

Lang:2017:BRE

[Lan17] Michel Lang. Book review: Efficient R Programming. Journal of Statistical Software, 80(??):??, ???? 2017. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v080b03;   https://www.
jstatsoft.org/index.php/jss/article/view/v080b03/v80b03.
pdf.

Lawson:2008:SMA

[Law08] John Lawson. SAS macros for analysis of unreplicated 2^{k} and 2^{k−p} designs with a possible outlier. Journal of Statistical Software, 25(13):1–17, May 2008. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v25/i13.

Liquet:2016:RGP

[LBC^{+}16] Benoît Liquet, Leonardo Bottolo, Gianluca Campanella, Sylvia Richardson, and Marc Chadeau-Hyam. R2GUESS: A graphics processing unit-based R package for Bayesian variable selection regression of multivariate responses. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v069i02;   https://www.
jstatsoft.org/index.php/jss/article/view/v069i02/v69i02.
pdf.

Li:2018:RRP

[LBW18] James Li, Jacob Bien, and Martin T. Wells. rTensor: An R package for multidimensional array (tensor) unfolding, multiplication, and decomposition. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v087i10;   https://www.jstatsoft.org/
index.php/jss/article/view/v087i10/v87i10.pdf.

Liu:2010:ICR

[LC10] Hai Liu and Kung-Sik Chan. Introducing COZIGAM: An R package for unconstrained and constrained zero-inflated generalized additive model analysis. Journal of Statistical Software, 35(11):??, July 2010. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v35/i11.

Leckie:2013:PRP

[LC13] George Leckie and Chris Charlton. runmlwin: a program to run the MLwiN multilevel modeling software from within Stata. Journal of Statistical Software, 52(11):??, March 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v52/i11.

Leifeld:2018:TER

[LCD18] Philip Leifeld, Skyler J. Cranmer, and Bruce A. Desmarais. Temporal exponential random graph models with btergm: Estimation and bootstrap confidence intervals. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v083i06;   https://www.jstatsoft.
org/index.php/jss/article/view/v083i06/v83i06.pdf.

Lathrop:2011:SPI

[LCK11] Scott Lathrop, Jim Costa, and William Kramer, editors. SC’11: Proceedings of 2011 International Conference for High Performance Computing, Networking, Storage and Analysis, Seattle, WA, November 12–18 2011. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4503-0771-X. LCCN ????

Lawrence:2008:EME

[LCL^{+}08] Michael Lawrence, Dianne Cook, Eun-Kyung Lee, Heather Babka, and Eve Syrkin Wurtele. explorase: Multivariate exploratory analysis and visualization for systems biology. Journal of Statistical Software, 25(9):1–23, April 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v25/i09.

Luo:2014:PMR

[LCSC14] Shen Luo, Yong Chen, Xiao Su, and Haitao Chu. mmeta: An R package for multivariate meta-analysis. Journal of Statistical Software, 56(11):??, January 2014. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v56/i11.

Lange:2021:SRP

[LDHM21] Alexander Lange, Bernhard Dalheimer, Helmut Herwartz, and Simone Maxand. svars: an R package for data-driven identification in multivariate time series analysis. Journal of Statistical Software, 97(??):??, ???? 2021. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v097i05;   https://www.jstatsoft.
org/index.php/jss/article/view/v097i05/v97i05.pdf.

Leckie:2014:CSP

[Lec14] George Leckie. Code snippet: runmixregls: A program to run the MIXREGLS mixed-effects location scale software from within Stata. Journal of Statistical Software, 59(CS-2):??, August 2014. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v59/c02.

Lee:2013:CRP

[Lee13] Duncan Lee. CARBayes: An R package for Bayesian spatial modeling with conditional autoregressive priors. Journal of Statistical Software, 55(13):??, November 2013. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v55/i13.

Lee:2018:PRP

[Lee18] Eun-Kyung Lee. PPtreeViz: An R package for visualizing projection pursuit classification trees. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v083i08;   https://www.jstatsoft.
org/index.php/jss/article/view/v083i08/v83i08.pdf.

Leisch:2004:FGF

[Lei04] Friedrich Leisch. FlexMix: a general framework for finite mixture models and latent class regression in R. Journal of Statistical Software, 11(8):1–18, ???? 2004. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/counter.php?
id=104&url=v11/i08/v11i08.pdf&ct=1.

Leibovici:2010:STM

[Lei10] Didier G. Leibovici. Spatio-temporal multiway data decomposition using principal tensor analysis on k-modes: The R package PTAk. Journal of Statistical Software, 34(10):??, May 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v34/i10.

Leifeld:2013:TCS

[Lei13] Philip Leifeld. texreg: Conversion of statistical model output in R to LaTEX and HTML tables. Journal of Statistical Software, 55(8):??, November 2013. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v55/i08.

Lenth:2009:BRB

[Len09a] Russell V. Lenth. Book review: Design and Analysis of Experiments: Classical and Regression Approaches with SAS. Journal of Statistical Software, 29(BR-5):??, January 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v29/b05.

Lenth:2009:RSM

[Len09b] Russell V. Lenth. Response-surface methods in R, using rsm. Journal of Statistical Software, 32(7):??, October 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v32/i07.

Lenth:2016:LSM

[Len16] Russell V. Lenth. Least-squares means: The R package lsmeans. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v069i01;
https://www.jstatsoft.org/index.php/jss/article/view/
v069i01/v69i01.pdf.

Leong:2010:BRB

[Leo10] Chee Kian Leong. Book review: R in a Nutshell. Journal of Statistical Software, 36(BR-2):??, August 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v36/b02.

Liboschik:2017:TRP

[LFF17] Tobias Liboschik, Konstantinos Fokianos, and Roland Fried. tscount: An R package for analysis of count time series following generalized linear models. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v082i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v082i05/v82i05.pdf.

Liu:2011:CSF

[LG11] Ziyue Liu and Wensheng Guo. Code snippet: fmixed: a SAS macro for smoothing-spline-based functional mixed effects models. Journal of Statistical Software, 43(CS-1):??, July 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v43/c01.

Lofstedt:2018:SDL

[LGF^{+}18] Tommy Löfstedt, Vincent Guillemot, Vincent Frouin, Edouard Duchesnay, and Fouad Hadj-Selem. Simulated data for linear regression with structured and sparse penalties: Introducing pylearn-simulate. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v087i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v087i03/v87i03.pdf.

laGrange:2009:BIB

[lGlRGL09] Anthony la Grange, Niël le Roux, and Sugnet Gardner-Lubbe. BiplotGUI: Interactive biplots in R. Journal of Statistical Software, 30(12):??, June 2009. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v30/i12.

Lenth:2007:SLP

[LH07] Russell V. Lenth and Søren Højsgaard. SASweave: Literate programming using SAS. Journal of Statistical Software, 19 (8):1–20, May 2007. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v19/i08.

Langfelder:2012:FRF

[LH12] Peter Langfelder and Steve Horvath. Fast R functions for robust correlations and hierarchical clustering. Journal of Statistical Software, 46(11):??, March 2012. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v46/i11.

Loy:2014:PHS

[LH14] Adam Loy and Heike Hofmann. HLMdiag: A suite of diagnostics for hierarchical linear models in R. Journal of Statistical Software, 56(5):??, January 2014. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v56/i05.

Liverani:2015:PRP

[LHA^{+}15] Silvia Liverani, David I. Hastie, Lamiae Azizi, Michail Papathomas, and Sylvia Richardson. PReMiuM: An R package for profile regression mixture models using Dirichlet processes. Journal of Statistical Software, 64(7):??, March 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v64/i07.

Leiva:2008:RPG

[LHS08] Víctor Leiva, Hugo Hernández, and Antonio Sanhueza. An R package for a general class of inverse Gaussian distributions. Journal of Statistical Software, 26(4):1–21, June 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v26/i04.

Lebret:2015:PRR

[LIL^{+}15] Rémi Lebret, Serge Iovleff, Florent Langrognet, Christophe Biernacki, Gilles Celeux, and Gérard Govaert. Rmixmod: The R package of the model-based unsupervised, supervised, and semi-supervised classification mixmod library. Journal of Statistical Software, 67(6):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067i06;
http://www.jstatsoft.org/index.php/jss/article/view/
v067i06/v67i06.pdf.

Le:2008:FRP

[LJH08] Sébastien Lê, Julie Josse, and François Husson. FactoMineR: An R package for multivariate analysis. Journal of Statistical Software, 25(1):1–18, March 2008. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v25/i01.

Lawrence:2010:RGU

[LL10] Michael Lawrence and Duncan Temple Lang. RGtk2: a graphical user interface toolkit for R. Journal of Statistical Software, 37(8):??, December 2010. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v37/i08.

Linzer:2011:PRP

[LL11] Drew A. Linzer and Jeffrey B. Lewis. poLCA: An R package for polytomous variable latent class analysis. Journal of Statistical Software, 42(10):??, June 2011. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v42/i10.

Linn:2015:PII

[LLS15] Kristin A. Linn, Eric B. Laber, and Leonard A. Stefanski. iqLearn: Interactive Q-learning in R. Journal of Statistical Software, 64(1):??, March 2015. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v64/i01.

Ligges:2003:SRP

[LM03] Uwe Ligges and Martin Mächler. Scatterplot3d — an R package for visualizing multivariate data. Journal of Statistical Software, 8(11):1–20, ???? 2003. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v08/i11;
 http://www.jstatsoft.org/v08/i11/JSSs3d.pdf.

Lee:2013:ERP

[LM13] Sharon X. Lee and Geoffrey J. McLachlan. EMMIXuskew: An R package for fitting mixtures of multivariate skew t distributions via the EM algorithm. Journal of Statistical Software, 55 (12):??, November 2013. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v55/i12.

Liao:2014:PCR

[LM14] Xiyue Liao and Mary C. Meyer. coneproj: An R package for the primal or dual cone projections with routines for constrained regression. Journal of Statistical Software, 61(12):??, November 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v61/i12.

Lee:2018:ERP

[LM18] Sharon X. Lee and Geoffrey J. McLachlan. EMMIXcskew: An R package for the fitting of a mixture of canonical fundamental skew t-distributions. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v083i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v083i03/v83i03.pdf.

Liao:2019:CRP

[LM19] Xiyue Liao and Mary C. Meyer. cgam: an R package for the constrained generalized additive model. Journal of Statistical Software, 89(??):??, ???? 2019. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v089i05;   https://www.jstatsoft.
org/index.php/jss/article/view/v089i05/v89i05.pdf.

Mullen:2007:FLI

[LMB^{+}07] Sergey Laptenok, Katharine M. Mullen, Jan Willem Borst, Ivo H. M. van Stokkum, Vladimir V. Apanasovich, and Antonie J. W. G. Visser. Fluorescence lifetime imaging microscopy (FLIM) data analysis with TIMP. Journal of Statistical Software, 18(8):1–20, January 2007. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v18/i08.

Lehnert:2019:HDA

[LMO^{+}19] Lukas W. Lehnert, Hanna Meyer, Wolfgang A. Obermeier, Brenner Silva, Bianca Regeling, and Jörg Bendix. Hyperspectral data analysis in R: The hsdar package. Journal of Statistical Software, 89(??):??, ???? 2019. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v089i12;   https://www.jstatsoft.
org/index.php/jss/article/view/v089i12/v89i12.pdf.

Lauve:2020:LRP

[LNY^{+}20] Nicolas R. Lauve, Stuart J. Nelson, S. Stanley Young, Robert L. Obenchain, and Christophe G. Lambert. LocalControl: an R package for comparative safety and effectiveness research. Journal of Statistical Software, 96(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v096i04;   https://
www.jstatsoft.org/index.php/jss/article/view/v096i04/
v96i04.pdf.

Lortie:2017:BRR

[Lor17] Christopher Lortie. Book review: R for Data Science. Journal of Statistical Software, 77(??):??, ???? 2017. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v077b01;   https://www.
jstatsoft.org/index.php/jss/article/view/v077b01/v77b01.
pdf.

Lortie:2018:BRA

[Lor18] Christopher Lortie. Book review: Applied Time Series Analysis with R (2nd Edition). Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v085b03;   https://www.jstatsoft.org/
index.php/jss/article/view/v085b03/v85b03.pdf.

Ladroue:2013:DPC

[LP13] Christophe Ladroue and Anastasia Papavasiliou. A distributed procedure for computing stochastic expansions with Mathematica. Journal of Statistical Software, 53(11):??, May 2013.

CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v53/i11.

Lucchetti:2017:DDP

[LP17] Riccardo Lucchetti and Claudia Pigini. DPB: Dynamic panel binary data models in gretl. Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v079i08;   https://www.jstatsoft.org/
index.php/jss/article/view/v079i08/v79i08.pdf.

Lafarge:2014:RIP

[LPLPD14] Thomas Lafarge, Beatriz Pateiro-López, Antonio Possolo, and Joy P. Dunkers. R implementation of a polyhedral approximation to a 3D set of points using the α-shape. Journal of Statistical Software, 56(4):??, January 2014. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v56/i04.

Li:2012:CSR

[LQC^{+}12] Mingyu Li, Hui Quan, Joshua Chen, Yoko Tanaka, Peter Ouyang, Xiaolong Luo, and Gang Li. Code snippet: R functions for sample size and probability calculations for assessing consistency of treatment effects in multi-regional clinical trials. Journal of Statistical Software, 47(CS-1):??, April 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v47/c01.

Lindgren:2015:BSM

[LR15a] Finn Lindgren and Håvard Rue. Bayesian spatial modelling with R-INLA. Journal of Statistical Software, 63(19):??, February 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v63/i19.

Loecher:2015:PPU

[LR15b] Markus Loecher and Karl Ropkins. RgoogleMaps and loa: Unleashing R graphics power on map tiles. Journal of Statistical Software, 63(4):??, February 2015. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v63/i04.

Lequesne:2020:CSV

[LR20] Justine Lequesne and Philippe Regnault. Code snippet: vsgoftest: an R package for goodness-of-fit testing based on Kullback–Leibler divergence. Journal of Statistical Software, 96(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v096c01;   https://www.jstatsoft.org/
index.php/jss/article/view/v096c01/v96c01.pdf.

Laurent:2012:GRP

[LRGTA12] Thibault Laurent, Anne Ruiz-Gazen, and Christine Thomas-Agnan. GeoXp: An R package for exploratory spatial data analysis. Journal of Statistical Software, 47(2):??, April 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v47/i02.

Lee:2018:STA

[LRN18a] Duncan Lee, Alastair Rushworth, and Gary Napier. Spatiotemporal areal unit modeling in R with conditional autoregressive priors using the CARBayesST package. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v084i09;   https://www.jstatsoft.
org/index.php/jss/article/view/v084i09/v84i09.pdf.

Long:2018:CSS

[LRN18b] Jed Long, Colin Robertson, and Trisalyn Nelson. Code snippet: stampr: Spatial-temporal analysis of moving polygons in R. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v084c01;
https://www.jstatsoft.org/index.php/jss/article/view/
v084c01/v84c01.pdf.

Lopez-Raton:2014:POR

[LRRÁCSGS14] Mónica López-Ratón, María Xosé Rodríguez-Álvarez, Carmen Cadarso-Suárez, and Francisco Gude-Sampedro. OptimalCutpoints: An R package for selecting optimal cutpoints in diagnostic tests. Journal of Statistical Software, 61(8):??, November 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v61/i08.

Lawal:2002:GFV

[LS02a] H. Bayo Lawal and Richard A. Sundheim. Generating factor variables for asymmetry, non-independence and skewsymmetry models in square contingency tables using SAS.

Journal of Statistical Software, 7(8):1–23, ???? 2002. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v07/i08;   http://www.jstatsoft.org/
v07/i08/generate.pdf;   http://www.jstatsoft.org/v07/
i08/GENERATE.txt;   http://www.jstatsoft.org/v07/i08/
updates.

Lipkovich:2002:BSV

[LS02b] Ilya Lipkovich and Eric P. Smith. Biplot and singular value decomposition macros for Excel\begin{matrix}c \\ ◯ \end{matrix}. Journal of Statistical Software, 7(5):1–15, ???? 2002. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v07/i05;   http:
//www.jstatsoft.org/v07/i05/biplot01.xla;   http://
www.jstatsoft.org/v07/i05/BIPLOT_paper_6_6_02.doc;
 http://www.jstatsoft.org/v07/i05/BIPLOT_paper_6_6_
02.pdf;   http://www.jstatsoft.org/v07/i05/BIPLOThelp.
zip;   http://www.jstatsoft.org/v07/i05/updates.

Lawson:2009:CSB

[LSC09] John Lawson, G. Bruce Schaalje, and Bruce Jay Collings. Code snippet: Blocking mixed-level factorials with SAS. Journal of Statistical Software, 32(CS-1):??, November 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v32/c01.

Liu:2020:PRP

[LSL20] Qi Liu, Bryan Shepherd, and Chun Li. PResiduals: an R package for residual analysis using probability-scale residuals. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v094i12;   https://
www.jstatsoft.org/index.php/jss/article/view/v094i12/
v94i12.pdf.

Love:2019:JGS

[LSM^{+}19] Jonathon Love, Ravi Selker, Maarten Marsman, Tahira Jamil, Damian Dropmann, Josine Verhagen, Alexander Ly, Quentin F. Gronau, Martin Smíra, Sacha Epskamp, Dora Matzke, Anneliese Wild, Patrick Knight, Jeffrey N. Rouder, Richard D. Morey, and Eric-Jan Wagenmakers. JASP: Graphical statistical software for common statistical designs. Journal of Statistical Software, 88 (??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v088i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v088i02/v88i02.pdf.

Lendle:2017:LRP

[LSPvdL17] Samuel D. Lendle, Joshua Schwab, Maya L. Petersen, and Mark J. van der Laan. ltmle: An R package implementing targeted minimum loss-based estimation for longitudinal data. Journal of Statistical Software, 81 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v081i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v081i01/v81i01.pdf.

Lorenzo-Seva:2019:MGU

[LSvdV19] Urbano Lorenzo-Seva and Michel van de Velden. MultipleCar: a graphical user interface MATLAB toolbox to compute multiple correspondence analysis. Journal of Statistical Software, 90(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v090i04;   https://www.jstatsoft.org/
index.php/jss/article/view/v090i04/v90i04.pdf.

Lorenzo-Seva:2009:CMP

[LSvdVK09] Urbano Lorenzo-Seva, Michel van de Velden, and Henk A. L. Kiers. CAR: a MATLAB package to compute correspondence analysis with rotations. Journal of Statistical Software, 31(8): ??, September 2009. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v31/i08.

LafayedeMicheaux:2016:PPR

[LT16] Pierre Lafaye de Micheaux and Viet Anh Tran. PoweR: A reproducible research tool to ease Monte Carlo power simulation studies for goodness-of-fit tests in R. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v069i03;   https://www.jstatsoft.
org/index.php/jss/article/view/v069i03/v69i03.pdf.

Lucchetti:2011:SSM

[Luc11] Riccardo Lucchetti. State space methods in gretl. Journal of Statistical Software, 41(11):??, May 2011. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v41/i11.

Lumley:1996:XST

[Lum96] Thomas Lumley. XLISP–Stat tools for building generalised estimating equation models. Journal of Statistical Software, 1(3):??, ???? 1996. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v01/i03;   http://www.jstatsoft.org/v01/i03/code/;
 http://www.jstatsoft.org/v01/i03/data/;   http://
www.jstatsoft.org/v01/i03/paper/geejss/geejss.html;
 http://www.jstatsoft.org/v01/i03/updates.

Lumley:1998:SAX

[Lum98] Thomas Lumley. Survival analysis in XLISP–Stat. A semiliterate program. Journal of Statistical Software, 3(2):1–90, ???? 1998. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v03/i02;   http://www.jstatsoft.org/
v03/i02/code/;   http://www.jstatsoft.org/v03/i02/
coxregdoc.pdf;   http://www.jstatsoft.org/v03/i02/
coxregdoc.ps.gz;   http://www.jstatsoft.org/v03/i02/
survival.tar.gz;   http://www.jstatsoft.org/v03/i02/
updates.

Lumley:2004:ACS

[Lum04] Thomas Lumley. Analysis of complex survey samples. Journal of Statistical Software, 9(8):1–19, ???? 2004. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=89&url=v09/i08/paper.pdf&ct=1.

Lupi:2009:URC

[Lup09] Claudio Lupi. Unit root CADF testing with R. Journal of Statistical Software, 32(2):??, October 2009. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v32/i02.

Lovaglio:2016:CSE

[LV16] Pietro Giorgio Lovaglio and Gianmarco Vacca. Code snippet: %ERA: A SAS macro for extended redundancy analysis. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/

index.php/jss/article/view/v074c01;   https://www.
jstatsoft.org/index.php/jss/article/view/v074c01/v74c01.
pdf.

Laaksonen:2011:SMC

[LVK^{+}11] Maarit A. Laaksonen, Esa Virtala, Paul Knekt, Hannu Oja, and Tommi Härkänen. SAS macros for calculation of population attributable fraction in a cohort study design. Journal of Statistical Software, 43(7):??, July 2011. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v43/i07.

Liu:2003:EAU

[LW03] Honghu Liu and Tongtong Wu. Estimating the area under a receiver operating characteristic curve for repeated measures design. Journal of Statistical Software, 8(12):1–18, ???? 2003. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v08/i12;   http://www.jstatsoft.
org/v08/i12/roc.pdf.

Lawson:2016:CSM

[LW16] John Lawson and Cameron Willden. Code snippet: Mixture experiments in R using mixexp. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v072c02;   https://www.jstatsoft.org/
index.php/jss/article/view/v072c02/v72c02.pdf.

Luts:2018:SRA

[LWOW18] Jan Luts, Shen S. J. Wang, John T. Ormerod, and Matt P. Wand. Semiparametric regression analysis via Infer.NET. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v087i02;   https://
www.jstatsoft.org/index.php/jss/article/view/v087i02/
v87i02.pdf.

Luo:2012:DRP

[LX12] Jingqin Luo and Chengjie Xiong. DiagTest3Grp: An R package for analyzing diagnostic tests with three ordinal groups. Journal of Statistical Software, 51(3):??, October 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v51/i03.

Liu:2015:CMP

[LZ15] Xiaohui Liu and Yijun Zuo. CompPD: A MATLAB package for computing projection depth. Journal of Statistical Software, 65(2):??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v65/i02.

Li:2017:CSS

[LZ17] Cai Li and Hua Zhou. Code snippet: svt: Singular value thresholding in MATLAB. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v081c02;   https://www.jstatsoft.org/
index.php/jss/article/view/v081c02/v81c02.pdf.

Lin:2017:PAB

[LZHC17] Lifeng Lin, Jing Zhang, James S. Hodges, and Haitao Chu. Performing arm-based network meta-analysis in R with the pcnetmeta package. Journal of Statistical Software, 80(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v080i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v080i05/v80i05.pdf.

Leong:2005:CIZ

[LZL^{+}05] Philip Leong, Guanglie Zhang, Dong-U Lee, Wayne Luk, and John Villasenor. A comment on the implementation of the ziggurat method. Journal of Statistical Software, 12(7):1–44, ???? 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/counter.php?id=114&url=v12/i07&ct=2;
 http://www.jstatsoft.org/counter.php?id=114&url=v12/
i07/v12i07.pdf&ct=1. See [MT00].

MacAskill:2012:SRB

[Mac12] Michael R. MacAskill. Software review: DataGraph 3.0. Journal of Statistical Software, 47(SR-2):??, May 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v47/s02.

Mullen:2011:DRP

[MAG^{+}11] Katharine M. Mullen, David Ardia, David L. Gil, Donald Windover, and James Cline. DEoptim: An R package for global optimization by differential evolution. Journal of Statistical Software, 40(6):??, April 2011. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v40/i06.

Maindonald:2006:BRB

[Mai06] John Maindonald. Book review: Generalized Additive Models: An Introduction with R. Journal of Statistical Software, 16 (BR-3):1–2, July 2006. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v16/b03.

Maindonald:2007:BRB

[Mai07] John Maindonald. Book review: Pattern Recognition and Machine Learning. Journal of Statistical Software, 17(BR-5):1–3, January 2007. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v17/b05.

Mair:2008:BRB

[Mai08] Patrick Mair. Book review: Data Mining the Web: Uncovering Patterns in Web Content, Structure, and Usage. Journal of Statistical Software, 25(BR-1):1–3, May 2008. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v25/b01.

Maindonald:2009:BRBa

[Mai09a] John Maindonald. Book review: Modern Multivariate Statistical Techniques: Regression, Classification and Manifold Learning. Journal of Statistical Software, 29(BR-11):??, February 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v29/b11.

Maindonald:2009:BRBb

[Mai09b] John Maindonald. Book review: Statistical Learning from a Regression Perspective. Journal of Statistical Software, 29(BR-12):??, February 2009. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v29/b12.

Mair:2009:BRB

[Mai09c] Patrick Mair. Book review: Analysis of Multivariate Social Science Data (2nd Edition). Journal of Statistical Software, 29(BR-13):??, February 2009. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v29/b13.

Moon:2006:CTT

[MAK06] Hojin Moon, Hongshik Ahn, and Ralph L. Kodell. A computational tool for testing dose-related trend using an age-adjusted bootstrap-based poly-k test. Journal of Statistical Software, 16(7):1–14, August 2006. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v16/i07.

Maloney:2009:BRB

[Mal09] Laurence T. Maloney. Book review: Introduction to Probability with R. Journal of Statistical Software, 30(BR-1):??, April 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v30/b01.

Marsaglia:2003:XR

[Mar03] George Marsaglia. Xorshift RNGs. Journal of Statistical Software, 8(14):1–6, 2003. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v08/i14;   http:
//www.jstatsoft.org/v08/i14/xorshift.pdf. See [Bre04]

for corrections and the equivalence of xorshift generators and the well-understood linear feedback shift register generators. See also [SMDS11, SM12, SLF14] for the failure of Marsaglia’s xorwow() generator from this paper. See [PL05, Vig16] for detailed analysis.

Marsaglia:2004:END

[Mar04] George Marsaglia. Evaluating the normal distribution. Journal of Statistical Software, 11(4):1–7, ???? 2004. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/counter.php?id=100&url=v11/i04/cphi.pdf&ct=1.

Marsaglia:2005:MGF

[Mar05] George Marsaglia. Monkeying with the goodness-of-fit test. Journal of Statistical Software, 14(13):1–4, September 20, 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/counter.php?id=138&url=v14/i13&ct=2;
 http://www.jstatsoft.org/counter.php?id=138&url=v14/
i13/v14i13.pdf&ct=1.

Marchetti:2006:IIG

[Mar06a] Giovanni M. Marchetti. Independencies induced from a graphical Markov model after marginalization and conditioning: The R package ggm. Journal of Statistical Software, 15(6):1–15, February 2006. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v15/i06.

Marsaglia:2006:RNV

[Mar06b] George Marsaglia. Ratios of normal variables. Journal of Statistical Software, 16(4):1–10, May 2006. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v16/i04.

Matloff:2013:BRB

[Mat13] Norman S. Matloff. Book review: The R Student Companion. Journal of Statistical Software, 52(BR-4):??, February 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v52/b04.

Matloff:2015:BRR

[Mat15] Norman S. Matloff. Book review: R for Cloud Computing, Ajay Ohri, Springer-Verlag, New York, 2014. ISBN 978-1-4939-

1701-3. 267 pp. USD 99.00 (P). http://www.springer.com/book/9781493917013.

Journal of Statistical Software, 66(BR-4):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066b04;
http://www.jstatsoft.org/index.php/jss/article/view/
v066b04/v66b04.pdf.

Matloff:2016:BRU

[Mat16a] Norm Matloff. Book review: A User’s Guide to Network Analysis in R, Douglas A. Luke, Springer-Verlag, New York, 2015. ISBN 978-3-319-23882-1. 238 pp. USD 59.99

(P). http://www.springer.com/9783319238821. Journal

of Statistical Software, 72(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v072b03;   https://www.
jstatsoft.org/index.php/jss/article/view/v072b03/v72b03.
pdf.

Matloff:2016:BRN

[Mat16b] Norman Matloff. Book review: Numerical Python: A Practical Techniques Approach for Industry. Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v070b04;   https://www.jstatsoft.
org/index.php/jss/article/view/v070b04/v70b04.pdf.

Matloff:2016:SAT

[Mat16c] Norman Matloff. Software alchemy: Turning complex statistical computations into embarrassingly-parallel ones. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v071i04;   https://www.
jstatsoft.org/index.php/jss/article/view/v071i04/v71i04.
pdf.

Marcum:2015:CMS

[MB15] Christopher Steven Marcum and Carter T. Butts. Constructing and modifying sequence statistics for relevent: Using informR in R. Journal of Statistical Software, 64(5):??, March 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v64/i05.

Magis:2017:CSC

[MB17] David Magis and Juan Ramon Barrada. Code snippet: Computerized adaptive testing with R: Recent updates of the package catR. Journal of Statistical Software, 76 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v076c01;   https://www.jstatsoft.org/
index.php/jss/article/view/v076c01/v76c01.pdf.

Marsh:2018:CSD

[MBGK18] Charles J. Marsh, Louise J. Barwell, Yoni Gavish, and William E. Kunin. Code snippet: downscale: An R package for downscaling species occupancy from coarse-grain data to predict occupancy at fine-grain sizes. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v086c03;   https://www.jstatsoft.
org/index.php/jss/article/view/v086c03/v86c03.pdf.

Muhleisen:2018:DIO

[MBK18] Hannes Mühleisen, Alexander Bertram, and Maarten-Jan Kallen. Database-inspired optimizations for statistical analysis. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v087i04;
https://www.jstatsoft.org/index.php/jss/article/view/
v087i04/v87i04.pdf.

Monteiro:2011:BRP

[MBR11] João Vitor Dias Monteiro, Sudipto Banerjee, and Gurumurthy Ramachandran. B2Z: R package for Bayesian two-zone models. Journal of Statistical Software, 43(2):??, July 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v43/i02.

Maus:2019:DTW

[MCAP19] Victor Maus, Gilberto Câmara, Marius Appel, and Edzer Pebesma. dtwSat: Time-weighted dynamic time warping for satellite image time series analysis in R. Journal of Statistical Software, 88(??):??, ???? 2019. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v088i05;   https://www.jstatsoft.
org/index.php/jss/article/view/v088i05/v88i05.pdf.

McBane:2006:PCD

[McB06] George C. McBane. Programs to compute distribution functions and critical values for extreme value ratios for outlier detection. Journal of Statistical Software, 16(3):1–9, May 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v16/i03.

McIlhagga:2016:PMT

[McI16] William McIlhagga. penalized: A MATLAB toolbox for fitting generalized linear models with penalties. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v072i06;   https://www.jstatsoft.
org/index.php/jss/article/view/v072i06/v72i06.pdf.

Melnykov:2012:MRP

[MCM12] Volodymyr Melnykov, Wei-Chen Chen, and Ranjan Maitra. MixSim: An R package for simulating data to study performance of clustering algorithms. Journal of Statistical Software, 51(12):??, November 2012. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v51/i12.

McNamara:2014:BRD

[McN14] Amelia McNamara. Book review: Dynamic Documents with R and knitr. Journal of Statistical Software, 56(BR-2):??, January 2014. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v56/b02.

Meulders:2018:LCP

[MD18] Michel Meulders and Philippe De Bruecker. Latent class probabilistic latent feature analysis of three-way threemode binary data. Journal of Statistical Software, 87 (??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v087i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v087i01/v87i01.pdf.

Mair:2010:GFM

[MdL10] Patrick Mair and Jan de Leeuw. A general framework for multivariate analysis with optimal scaling: The R package aspect. Journal of Statistical Software, 32(9):??, January 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v32/i09.

Ma:2020:GRP

[MdlT20] Wenchao Ma and Jimmy de la Torre. GDINA: an R package for cognitive diagnosis modeling. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v093i14;   https://www.jstatsoft.
org/index.php/jss/article/view/v093i14/v93i14.pdf.

Moreira:2010:DRP

[MdUÁC10] Carla Moreira, Jacobo de Uña-Álvarez, and Rosa M. Crujeiras. DTDA: An R package to analyze randomly truncated data. Journal of Statistical Software, 37(7):??, November 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v37/i07.

Markos:2019:BTA

[MDvdV19] Angelos Markos, Alfonso Iodice D’Enza, and Michel van de Velden. Beyond tandem analysis: Joint dimension reduction and clustering in R. Journal of Statistical Software, 91(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v091i10;   https://www.jstatsoft.org/
index.php/jss/article/view/v091i10/v91i10.pdf.

Melnykov:2016:CRP

[Mel16] Volodymyr Melnykov. ClickClust: An R package for modelbased clustering of categorical sequences. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v074i09;   https://www.jstatsoft.
org/index.php/jss/article/view/v074i09/v74i09.pdf.

Mendelssohn:2011:SSS

[Men11] Roy Mendelssohn. The STAMP software for state space models. Journal of Statistical Software, 41(2):??, May 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v41/i02.

Meulders:2013:RPP

[Meu13] Michel Meulders. An R package for probabilistic latent feature analysis of two-way two-mode frequencies. Journal of Statistical Software, 54(14):??, September 2013. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v54/i14.

Meyer:2013:RNE

[Mey13] Christian Meyer. Recursive numerical evaluation of the cumulative bivariate normal distribution. Journal of Statistical Software, 52(10):??, March 2013. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v52/i10.

Magis:2014:CSP

[MF14] David Magis and Bruno Facon. Code snippet: deltaPlotR: An R package for differential item functioning analysis with Angoff’s delta plot. Journal of Statistical Software, 59(CS-1): ??, August 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v59/c01.

Matthews:2015:CSP

[MF15] Gregory J. Matthews and Andrea S. Foulkes. Code snippet: MixMAP: An R package for mixed modeling of metaanalysis p values in genetic association studies. Journal of Statistical Software, 66(3):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066c03;
http://www.jstatsoft.org/index.php/jss/article/view/
v066c03/v66c03.pdf.

Murrell:2009:QCS

[MG09] Paul Murrell and Stephen Gardiner. Quality control for statistical graphics: The graphicsQC package for R. Journal of Statistical Software, 30(1):??, April 2009. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v30/i01.

Marques:2020:HRP

[MGFA^{+}20] Fernando S. Marques, José H. H. Grisi-Filho, Marcos Amaku, Jean C. R. Silva, Erivânia C. Almeida, and José L. Silva Júnior. hybridModels: an R package for the stochastic simulation of disease spreading in dynamic networks. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v094i06;   https://www.
jstatsoft.org/index.php/jss/article/view/v094i06/v94i06.
pdf.

Mathias:2004:ASA

[MGG^{+}04] Adolf Mathias, Florian Grond, Ramon Guardans, Detlef Seese, Miguel Canela, and Hans H. Diebner. Algorithms for spectral analysis of irregularly sampled time series. Journal of Statistical Software, 11(2):1–30, ???? 2004. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/counter.php?
id=96&url=v11/i02/paper.pdf&ct=1.

Monaco:2018:GSS

[MGH18] John V. Monaco, Malka Gorfine, and Li Hsu. General semiparametric shared frailty model: Estimation and simulation with frailtySurv. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v086i04;   https://www.jstatsoft.org/
index.php/jss/article/view/v086i04/v86i04.pdf.

McMillan:2005:SMB

[MH05] Garnett McMillan and Timothy Hanson. SAS macro BDM for fitting the Dale regression model to bivariate ordinal response data. Journal of Statistical Software, 14(2):1–12, May 1, 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/counter.php?id=127&url=v14/i02&ct=2;
 http://www.jstatsoft.org/counter.php?id=127&url=v14/
i02/v14i02.pdf&ct=1.

Mair:2007:ERM

[MH07] Patrick Mair and Reinhold Hatzinger. Extended Rasch modeling: The eRm package for the application of IRT models in R. Journal of Statistical Software, 20(9):1–20, February 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v20/i09.

Meyer:2009:GCS

[MH09] David Meyer and Kurt Hornik. Generalized and customizable sets in R. Journal of Statistical Software, 31(2):??, August 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v31/i02.

Marcon:2015:PER

[MH15] Eric Marcon and Bruno Hérault. entropart: An R package to measure and partition diversity. Journal of Statistical Software, 67(8):??, ???? 2015. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/index.php/jss/
article/view/v067i08;   http://www.jstatsoft.org/index.
php/jss/article/view/v067i08/v67i08.pdf.

Moral:2017:HNP

[MHD17] Rafael A. Moral, John Hinde, and Clarice G. B. Demétrio. Half-normal plots and overdispersed models in R: The hnp package. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v081i10;
https://www.jstatsoft.org/index.php/jss/article/view/
v081i10/v81i10.pdf.

Mulder:2012:BFP

[MHdL12] Joris Mulder, Herbert Hoijtink, and Christiaan de Leeuw. BIEMS: a Fortran 90 program for calculating Bayes factors for inequality and equality constrained models. Journal of Statistical Software, 46(2):??, January 2012. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v46/i02.

Manolopoulou:2020:BRP

[MHE20] Ioanna Manolopoulou, Axel Hille, and Brent Emerson. BPEC: an R package for Bayesian phylogeographic and ecological clustering. Journal of Statistical Software, 92 (??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v092i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v092i05/v92i05.pdf.

Morris:2008:SEF

[MHH08] Martina Morris, Mark S. Handcock, and David R. Hunter. Specification of exponential-family random graph models: Terms and computational aspects. Journal of Statistical Software, 24(4):1–24, May 2008. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v24/i04.

Meyer:2017:STA

[MHH17] Sebastian Meyer, Leonhard Held, and Michael Höhle. Spatiotemporal analysis of epidemic phenomena using the R package surveillance. Journal of Statistical Software, 77 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v077i11;   https://www.jstatsoft.org/
index.php/jss/article/view/v077i11/v77i11.pdf.

Mahani:2016:CSS

[MHJS16] Alireza S. Mahani, Asad Hasan, Marshall Jiang, and Mansour T. A. Sharabiani. Code snippet: Stochastic Newton sampler: The R package sns. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v074c02;   https://www.jstatsoft.org/
index.php/jss/article/view/v074c02/v74c02.pdf.

Manuguerra:2020:COR

[MHM20] Maurizio Manuguerra, Gillian Z. Heller, and Jun Ma. Continuous ordinal regression for analysis of visual analogue scales: The R package ordinalCont. Journal of Statistical Software, 96(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v096i08;   https://www.jstatsoft.org/
index.php/jss/article/view/v096i08/v96i08.pdf.

Myers:2007:CSE

[MHT07] James Myers, Shih-Feng Huang, and Jhishen Tsay. Code snippet: Exact conditional inference for two-way randomized Bernoulli experiments. Journal of Statistical Software, 21(CS-1):1–7, September 2007. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v21/c01.

Mogensen:2012:ERF

[MIG12] Ulla B. Mogensen, Hemant Ishwaran, and Thomas A. Gerds. Evaluating random forests for survival analysis using prediction error curves. Journal of Statistical Software, 50(11):??, September 2012. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v50/i11.

Miller:2004:CMS

[Mil04] Alan J. Miller. A collection of mathematical and statistical routines in FORTRAN 90. Journal of Statistical Software, 9 (3):1–3, February 25, 2004. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v09/i03/F90lib.
pdf.

Millo:2017:RSE

[Mil17] Giovanni Millo. Robust standard error estimators for panel models: A unifying approach. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v082i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v082i03/v82i03.pdf.

May:2000:CTS

[MJ00] Warren L. May and William D. Johnson. Constructing twosided simultaneous confidence intervals for multinomial proportions for small counts in a large number of cells. Journal of Statistical Software, 5(6):1–24, ???? 2000. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v05/i06;   http://www.jstatsoft.org/v05/i06/sison5.
sas;   http://www.jstatsoft.org/v05/i06/Sisonfinal.
pdf;   http://www.jstatsoft.org/v05/i06/updates.

Merl:2010:ARP

[MJGM10] Daniel Merl, Leah R. Johnson, Robert B. Gramacy, and Marc Mangel. amei: An R package for the adaptive management of epidemiological interventions. Journal of Statistical Software, 36(6):??, August 2010. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v36/i06.

Meier:2021:MPM

[MKC21] Alexander Meier, Claudia Kirch, and Haeran Cho. mosum: a package for moving sums in change-point analysis. Journal of Statistical Software, 97(??):??, ???? 2021. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v097i08;   https://www.
jstatsoft.org/index.php/jss/article/view/v097i08/v97i08.
pdf.

Molas:2011:HGL

[ML11] Marek Molas and Emmanuel Lesaffre. Hierarchical generalized linear models: The R package HGLMMM. Journal of Statistical Software, 39(13):??, March 2011. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v39/i13.

Monecke:2012:SSE

[ML12] Armin Monecke and Friedrich Leisch. semPLS: Structural equation modeling using partial least squares. Journal of Statistical Software, 48(3):??, May 2012. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v48/i03.

Moon:2002:WBS

[MLAN02] Hojin Moon, J. Jack Lee, Hongshik Ahn, and Rumiana G. Nikolova. A Web-based simulator for sample size and power estimation in animal carcinogenicity studies. Journal of Statistical Software, 7(13):1–36, ???? 2002. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v07/i13;
 http://www.jstatsoft.org/v07/i13/ACSSreview.pdf.

Mussel:2012:MOP

[MLMK12] Christoph Müssel, Ludwig Lausser, Markus Maucher, and Hans A. Kestler. Multi-objective parameter selection for classifiers. Journal of Statistical Software, 46(5):??, January 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v46/i05.

Molina:2005:VTT

[MLVMY05] J. Gabriel Molina, Rubén Ledesma, Pedro M. Valero-Mora, and Forrest W. Young. A video tour through ViSta 6.4. Journal of Statistical Software, 13(8):1–13, ???? 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=122&url=v13/i08&ct=2;   http://www.

jstatsoft.org/counter.php?id=122&url=v13/i08/v13i08.
pdf&ct=1.

Marsaglia:2004:EAD

[MM04] George Marsaglia and John Marsaglia. Evaluating the Anderson–Darling distribution. Journal of Statistical Software, 9(2):1–5, February 25, 2004. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v09/i02/ad.pdf;
 http://www.jstatsoft.org/v09/i02/ADinf.c;   http:/
/www.jstatsoft.org/v09/i02/AnDarl.c.

Mair:2016:HLF

[MM16] Patrick Mair and Katharine Mullen. Honoring the lion: A Festschrift for Jan de Leeuw. Journal of Statistical Software, 73(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v073i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v073i01/v73i01.pdf.

Mankad:2015:TVE

[MMB15] Shawn Mankad, George Michailidis, and Moulinath Banerjee. Threshold value estimation using adaptive two-stage plans in R. Journal of Statistical Software, 67(3):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067i03;
http://www.jstatsoft.org/index.php/jss/article/view/
v067i03/v67i03.pdf.

Moreira:2012:SEB

[MMM12] Ana Moreira and Luís Meira-Machado. survivalBIV: Estimation of the bivariate distribution function for sequentially ordered events under univariate censoring. Journal of Statistical Software, 46(13):??, March 2012. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v46/i13.

Meehan:2020:EAA

[MMR20] Timothy D. Meehan, Nicole L. Michel, and Håvard Rue. Estimating animal abundance with N-mixture models using the R-INLA package for R. Journal of Statistical Software, 95(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v095i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v095i02/v95i02.pdf.

Meira-Machado:2011:PPM

[MMRP11] Luís Meira-Machado and Javier Roca-Pardiñas. p3state.msm: Analyzing survival data from an illness-death model. Journal of Statistical Software, 38(3):??, January 2011. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v38/i03.

Morina:2014:RPP

[MN14] David Moriña and Albert Navarro. The R package survsim for the simulation of simple and complex survival data. Journal of Statistical Software, 59(2):??, August 2014. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v59/i02.

McLachlan:1997:AFM

[MNA^{+}97] Geoff McLachlan, Angus Ng, Peter Adams, David McGiffin, and Andrew Gailbraith. An algorithm for fitting mixtures of Gompertz distributions to censored survival data. Journal of Statistical Software, 2(7):1–23, ???? 1997. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v02/i07;
 http://www.jstatsoft.org/v02/i07/0010.pdf;   http:
//www.jstatsoft.org/v02/i07/0010.ps.gz;   http://www.
jstatsoft.org/v02/i07/code/;   http://www.jstatsoft.
org/v02/i07/updates.

Miettinen:2017:BSS

[MNT17] Jari Miettinen, Klaus Nordhausen, and Sara Taskinen. Blind source separation based on joint diagonalization in R: The packages JADE and BSSasymp. Journal of Statistical Software, 76(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v076i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v076i02/v76i02.pdf.

Monahan:2006:SAC

[Mon06] John F. Monahan. Some algorithms for the conditional mean vector and covariance matrix. Journal of Statistical Software, 16(8):1–9, August 2006. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v16/i08.

Mora:2018:BRB

[Mor18] Pedro M. Valero Mora. Book review: bookdown: Authoring Books and Technical Documents with R Markdown. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v087b01;   https://www.
jstatsoft.org/index.php/jss/article/view/v087b01/v87b01.
pdf.

McLachlan:1999:EAF

[MP99] Geoff J. McLachlan and David Peel. The EMMIX algorithm for the fitting of normal and t-components. Journal of Statistical Software, 4(2):1–14, ???? 1999. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v04/
i02;   http://www.jstatsoft.org/v04/i02/emmix.tar.
gz;   http://www.jstatsoft.org/v04/i02/guide.pdf;
http://www.jstatsoft.org/v04/i02/paper.pdf;   http:
//www.jstatsoft.org/v04/i02/updates;   http://www.
maths.uq.edu.au/~gjm/emmix/emmix.html.

Mineo:2006:URP

[MP06] Angelo M. Mineo and Alfredo Pontillo. Using R via PHP for teaching purposes: R-php. Journal of Statistical Software, 17 (4):1–20, October 2006. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v17/i04.

Millo:2012:SSP

[MP12] Giovanni Millo and Gianfranco Piras. splm: Spatial panel data models in R. Journal of Statistical Software, 47(1):??, April 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v47/i01.

Mazza:2014:CSD

[MP14] Angelo Mazza and Antonio Punzo. Code snippet: DBKGrad: An R package for mortality rates graduation by discrete beta kernel techniques. Journal of Statistical Software, 57(CS-2): ??, April 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v57/c02.

Mazza:2018:FFF

[MPI18] Angelo Mazza, Antonio Punzo, and Salvatore Ingrassia. flexCWM: A flexible framework for cluster-weighted models. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v086i02;   https://

www.jstatsoft.org/index.php/jss/article/view/v086i02/
v86i02.pdf.

Mazza:2014:PKR

[MPM14] Angelo Mazza, Antonio Punzo, and Brian McGuire. KernSmoothIRT: An R package for kernel smoothing in item response theory. Journal of Statistical Software, 58(6):??, June 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v58/i06.

Martin:2011:MMC

[MQP11] Andrew D. Martin, Kevin M. Quinn, and Jong Hee Park. MCMCpack: Markov chain Monte Carlo in R. Journal of Statistical Software, 42(9):??, June 2011. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v42/i09.

Mineo:2005:STE

[MR05] Angelo M. Mineo and Mariantonietta Ruggieri. A software tool for the exponential power distribution: The normalp package. Journal of Statistical Software, 12(4):1–24, ???? 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/counter.php?id=111&url=v12/i04&ct=2;
 http://www.jstatsoft.org/counter.php?id=111&url=v12/
i04/v12i04.pdf&ct=1.

Magis:2012:RGR

[MR12] David Magis and Gilles Raîche. Random generation of response patterns under computerized adaptive testing with the R package catR. Journal of Statistical Software, 48(8):??, May 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v48/i08.

Merkle:2018:BBS

[MR18] Edgar C. Merkle and Yves Rosseel. blavaan: Bayesian structural equation models via parameter expansion. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v085i04;   https://www.
jstatsoft.org/index.php/jss/article/view/v085i04/v85i04.
pdf.

MacDonald:2015:GRP

[MRC15] Blake MacDonald, Pritam Ranjan, and Hugh Chipman. GPfit: An R package for fitting a Gaussian process model to deterministic simulator outputs. Journal of Statistical Software, 64(12):??, April 2015. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v64/i12.

Martin:2020:MRI

[MRHA20] Sean Martin, Andrew M. Raim, Wen Huang, and Kofi P. Adragni. ManifoldOptim: an R interface to the ROPTLIB library for Riemannian manifold optimization. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v093i01;   https://www.
jstatsoft.org/index.php/jss/article/view/v093i01/v93i01.
pdf.

Munda:2012:PPF

[MRL12] Marco Munda, Federico Rotolo, and Catherine Legrand. parfm: Parametric frailty models in R. Journal of Statistical Software, 51(11):??, November 2012. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v51/i11.

Miller:2019:DSR

[MRT^{+}19] David L. Miller, Eric Rexstad, Len Thomas, Laura Marshall, and Jeffrey L. Laake. Distance sampling in R. Journal of Statistical Software, 89(??):??, ???? 2019. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v089i01;   https://www.
jstatsoft.org/index.php/jss/article/view/v089i01/v89i01.
pdf.

Magnani:2021:AMS

[MRV21] Matteo Magnani, Luca Rossi, and Davide Vega. Analysis of multiplex social networks with R. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v098i08;   https://www.jstatsoft.org/
index.php/jss/article/view/v098i08/v98i08.pdf.

Mebane:2011:GOU

[MS11] Walter R. Mebane, Jr. and Jasjeet S. Sekhon. Genetic optimization using derivatives: The rgenoud package for R. Journal of Statistical Software, 42(11):??, June 2011. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v42/i11.

Mahani:2019:BNB

[MS19] Alireza S. Mahani and Mansour T. A. Sharabiani. Bayesian, and non-Bayesian, cause-specific competing-risk analysis for parametric and nonparametric survival functions: The R package CFC. Journal of Statistical Software, 89(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v089i09;
https://www.jstatsoft.org/index.php/jss/article/view/
v089i09/v89i09.pdf.

Medrano-Soto:2005:BBA

[MSCCV05] Arturo Medrano-Soto, J. Andrés Christen, and Julio Collado-Vides. BClass: a Bayesian approach based on mixture models for clustering and classification of heterogeneous biological data. Journal of Statistical Software, 13(2):1–18, ???? 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/counter.php?id=116&url=v13/i02&ct=2;
 http://www.jstatsoft.org/counter.php?id=116&url=v13/
i02/v13i02.pdf&ct=1.

Marsaglia:1998:MPM

[MT98] George Marsaglia and Wai Wan Tsang. The Monty Python method for generating gamma variables. Journal of Statistical Software, 3(3):1–8, ???? 1998. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v03/i03;   http:
//www.jstatsoft.org/v03/i03/GERMGAM.PDF;   http://
www.jstatsoft.org/v03/i03/GERMGAM.PS;   http://www.
jstatsoft.org/v03/i03/updates.

Marsaglia:2000:ZMG

[MT00] George Marsaglia and Wai Wan Tsang. The ziggurat method for generating random variables. Journal of Statistical Software, 5(8):1–7, ???? 2000. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v05/
i08;   http://www.jstatsoft.org/v05/i08/rnorrexp.c;
 http://www.jstatsoft.org/v05/i08/updates;   http:/
/www.jstatsoft.org/v05/i08/ziggurat.pdf. See comment

[LZL^{+}05].

Marsaglia:2002:SDP

[MT02] George Marsaglia and Wai Wan Tsang. Some difficult-to-pass tests of randomness. Journal of Statistical Software, 7(3):1–8, ???? 2002. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v07/i03;   http://www.jstatsoft.
org/v07/i03/tuftests.c;   http://www.jstatsoft.org/
v07/i03/tuftests.pdf;   http://www.jstatsoft.org/v07/
i03/updates.

Meng:2014:BRH

[MT14] Xiangxiang Meng and Wayne Thompson. Book review: Handbook of SAS(R) DATA Step Programming. Journal of Statistical Software, 56(BR-1):??, January 2014. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v56/b01.

Marcon:2015:CST

[MTPL15] Eric Marcon, Stéphane Traissac, Florence Puech, and Gabriel Lang. Code snippet: Tools to characterize point patterns: dbmss for R. Journal of Statistical Software, 67(CS-3):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067c03;
http://www.jstatsoft.org/index.php/jss/article/view/
v067c03/v67c03.pdf.

Molenaar:2015:FDI

[MTvdM15] Dylan Molenaar, Francis Tuerlinckx, and Han L. J. van der Maas. Fitting diffusion item response theory models for responses and response times using the R package diffIRT. Journal of Statistical Software, 66(4):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066i04;
http://www.jstatsoft.org/index.php/jss/article/view/
v066i04/v66i04.pdf.

Marsaglia:2003:EKD

[MTW03] George Marsaglia, Wai Wan Tsang, and Jingbo Wang. Evaluating Kolmogorov’s distribution. Journal of Statistical Software, 8(18):1–4, ???? 2003. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v08/i18;   http:
//www.jstatsoft.org/v08/i18/k.pdf.

Marsaglia:2004:FGD

[MTW04] George Marsaglia, Wai Wan Tsang, and Jingbo Wang. Fast generation of discrete random variables. Journal of Statistical Software, 11(3):1–8, ???? 2004. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/counter.php?
id=99&url=v11/i03/discrete.pdf&ct=1.

Muggeo:2010:ATE

[Mug10] Vito M. R. Muggeo. Analyzing temperature effects on mortality within the R environment: The constrained segmented distributed lag parameterization. Journal of Statistical Software, 32(12):??, January 2010. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v32/i12.

Mullner:2013:PFF

[Mül13] Daniel Müllner. fastcluster: Fast hierarchical, agglomerative clustering routines for R and Python. Journal of Statistical Software, 53(9):??, May 2013. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v53/i09.

Mullen:2014:CGO

[Mul14] Katharine M. Mullen. Continuous global optimization in R. Journal of Statistical Software, 60(6):??, September 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v60/i06.

Muller:2016:BRM

[Mül16] Marlene Müller. Book review: Multivariate Nonparametric Regression and Visualization With R and Applications to Finance. Journal of Statistical Software, 70 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v070b05;   https://www.jstatsoft.org/
index.php/jss/article/view/v070b05/v70b05.pdf.

Munzert:2014:BRB

[Mun14] Simon Munzert. Book review: XML and Web Technologies for Data Sciences with R. Journal of Statistical Software, 61 (BR-1):??, October 2014. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v61/b01.

Murdoch:2003:POR

[Mur03] Duncan Murdoch. Orientlib: An R package for orientation data. Journal of Statistical Software, 8(19):1–11, ???? 2003. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v08/i19;   http://www.jstatsoft.
org/v08/i19/orientlib_paper.pdf.

Murdoch:2005:BRC

[Mur05] Duncan Murdoch. Book review: Computational Statistics. Journal of Statistical Software, 14(BR-2):1–4, April 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v14/b02.

Murrell:2006:BRB

[Mur06] Paul Murrell. Book review: Graphics of Large Datasets: Visualizing a Million. Journal of Statistical Software, 17(BR-1): 1–3, December 2006. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v17/b01.

Murrell:2009:IVG

[Mur09] Paul Murrell. Importing vector graphics: The grImport package for R. Journal of Statistical Software, 30(4):??, April 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v30/i04.

Murray:2015:BSS

[Mur15] Lawrence M. Murray. Bayesian state-space modelling on highperformance hardware using LibBi. Journal of Statistical Software, 67(10):??, ???? 2015. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/index.php/jss/
article/view/v067i10;   http://www.jstatsoft.org/index.
php/jss/article/view/v067i10/v67i10.pdf.

Montero:2014:PRP

[MV14] Pablo Montero and José A. Vilar. TSclust: An R package for time series clustering. Journal of Statistical Software, 62(1): ??, December 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v62/i01.

Masarotto:2017:GCR

[MV17] Guido Masarotto and Cristiano Varin. Gaussian copula regression in R. Journal of Statistical Software, 77 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v077i08;   https://www.jstatsoft.org/
index.php/jss/article/view/v077i08/v77i08.pdf.

Maus:2014:PCP

[MvB14] Bärbel Maus and Gerard J. P. van Breukelen. POBE: A computer program for optimal design of multi-subject blocked fMRI experiments. Journal of Statistical Software, 56(9):??, January 2014. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v56/i09.

Mullen:2007:ISV

[MvS07a] Katharine M. Mullen and Ivo H. M. van Stokkum. An introduction to the “Special Volume Spectroscopy and Chemometrics in R”. Journal of Statistical Software, 18(1):1–5, January 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v18/i01.

Mullen:2007:TRP

[MvS07b] Katharine M. Mullen and Ivo H. M. van Stokkum. TIMP: An R package for modeling multi-way spectroscopic measurements. Journal of Statistical Software, 18(3):1–46, January 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v18/i03.

Miecznikowski:2013:DRP

[MVS13] Jeffrey C. Miecznikowski, Albert Vexler, and Lori Shepherd. dbEmpLikeGOF: An R package for nonparametric likelihood ratio tests for goodness-of-fit and two-sample comparisons based on sample entropy. Journal of Statistical Software, 54(3):??, September 2013. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v54/i03.

Mevik:2007:PPP

[MW07] Björn-Helge Mevik and Ron Wehrens. The pls package: Principal component and partial least squares regression in R. Journal of Statistical Software, 18(2):1–24, January 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v18/i02.

Marley:2010:NSS

[MW10] Jennifer K. Marley and Matthew P. Wand. Non-standard semiparametric regression via BRugs. Journal of Statistical Software, 37(5):??, November 2010. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v37/i05.

Mildenberger:2012:BPB

[MW12] Thoralf Mildenberger and Henrike Weinert. The benchden package: Benchmark densities for nonparametric density estimation. Journal of Statistical Software, 46(14):??, March 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v46/i14.

Mohammadi:2019:BRP

[MW19] Reza Mohammadi and Ernst C. Wit. BDgraph: an R package for Bayesian structure learning in graphical models. Journal of Statistical Software, 89(??):??, ???? 2019. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v089i03;   https://www.
jstatsoft.org/index.php/jss/article/view/v089i03/v89i03.
pdf.

McCullough:2013:SCR

[MY13] Bruce D. McCullough and A. Talha Yalta. Spreadsheets in the cloud — not ready yet. Journal of Statistical Software, 52(7): ??, February 2013. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v52/i07.

McLeod:2007:ALT

[MYK07] A. Ian McLeod, Hao Yu, and Zinovi L. Krougly. Algorithms for linear time series analysis: With R package. Journal of Statistical Software, 23(5):1–26, December 2007. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v23/i05.

McLeod:2008:ISA

[MZ08] A. Ian McLeod and Ying Zhang. Improved subset autoregression: With R package. Journal of Statistical Software, 28(2): ??, October 2008. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v28/i02.

Meyer:2006:SFV

[MZH06] David Meyer, Achim Zeileis, and Kurt Hornik. The Strucplot framework: Visualizing multi-way contingency tables with vcd. Journal of Statistical Software, 17(3):1–48, October 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v17/i03.

Nyaga:2017:CSC

[NAA17] Victoria N. Nyaga, Marc Arbyn, and Marc Aerts. Code snippet: CopulaDTA: An R package for copula-based bivariate beta-binomial models for diagnostic test accuracy studies in a Bayesian framework. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v082c01;   https://www.jstatsoft.org/
index.php/jss/article/view/v082c01/v82c01.pdf.

Nagler:2018:KRP

[Nag18] Thomas Nagler. kdecopula: An R package for the kernel estimation of bivariate copula densities. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v084i07;   https://www.jstatsoft.
org/index.php/jss/article/view/v084i07/v84i07.pdf.

Nagler:2021:CSR

[Nag21] Thomas Nagler. Code snippet: R-friendly multi-threading in C++. Journal of Statistical Software, 97(??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v097c01;
https://www.jstatsoft.org/index.php/jss/article/view/
v097c01/v97c01.pdf.

Narasimhan:2005:LSJ

[Nar05] Balasubramanian Narasimhan. Lisp–Stat to Java to R. Journal of Statistical Software, 13(4):1–10, ???? 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=118&url=v13/i04&ct=2;   http://www.
jstatsoft.org/counter.php?id=118&url=v13/i04/v13i04.
pdf&ct=1.

Nason:2005:PPS

[Nas05] Guy Nason. pinktoe: Semi-automatic traversal of trees. Journal of Statistical Software, 14(1):1–11, April 29, 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/counter.php?id=127&url=v14/i01&ct=2;   http://

www.jstatsoft.org/counter.php?id=127&url=v14/i01/v14i01.
pdf&ct=1.

Nash:2014:BPO

[Nas14] John C. Nash. On best practice optimization methods in R. Journal of Statistical Software, 60(2):??, September 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v60/i02.

Nia:2012:HDB

[ND12] Vahid Partovi Nia and Anthony C. Davison. High-dimensional Bayesian clustering with variable selection: The R package bclust. Journal of Statistical Software, 47(5):??, April 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v47/i05.

Noorian:2016:GGE

[NdSL16] Farzad Noorian, Anthony M. de Silva, and Philip H. W. Leong. gramEvol: Grammatical evolution in R. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v071i01;   https://www.jstatsoft.
org/index.php/jss/article/view/v071i01/v71i01.pdf.

Narasimhan:2020:DGM

[NE20] Balasubramanian Narasimhan and Bradley Efron. deconvolveR: a G-modeling program for deconvolution and empirical Bayes estimation. Journal of Statistical Software, 94 (??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v094i11;   https://www.jstatsoft.org/
index.php/jss/article/view/v094i11/v94i11.pdf.

Newson:2005:ECJ

[New05] Roger Newson. Efficient calculation of jackknife confidence intervals for rank statistics. Journal of Statistical Software, 15 (1):1–10, December 2005. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v15/i01.

Nenadic:2007:RS

[NG07] Oleg Nenadić and Michael Greenacre. Correspondence analysis in R, with two- and three-dimensional graphics: The ca package. Journal of Statistical Software, 20(3):1–13, February 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v20/i03.

Noguchi:2012:NRS

[NGBK12] Kimihiro Noguchi, Yulia R. Gel, Edgar Brunner, and Frank Konietschke. nparLD: An R software package for the nonparametric analysis of longitudinal data in factorial experiments. Journal of Statistical Software, 50(12):??, September 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v50/i12.

Nguyen:2007:NSE

[Ngu07] Paul Nguyen. nonbinROC: Software for evaluating diagnostic accuracies with non-binary gold standards. Journal of Statistical Software, 21(10):1–10, November 2007. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v21/i10.

Navarro-Gonzalez:2019:PTT

[NGVCFLS19] David Navarro-Gonzalez, Andreu Vigil-Colet, Pere Joan Ferrando, and Urbano Lorenzo-Seva. Psychological test toolbox: a new tool to compute factor analysis controlling response bias. Journal of Statistical Software, 91 (??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v091i06;   https://www.jstatsoft.org/
index.php/jss/article/view/v091i06/v91i06.pdf.

Nadarajah:2006:CSR

[NK06] Saralees Nadarajah and Samuel Kotz. Code snippet: R programs for truncated distributions. Journal of Statistical Software, 16(CS-2):1–8, August 2006. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v16/c02.

Nolan:2012:IAS

[NL12] Deborah Nolan and Duncan Temple Lang. Interactive and animated scalable vector graphics and R data displays. Journal of Statistical Software, 46(1):??, January 2012. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v46/i01.

Nguile-Makao:2015:SPM

[NMB15] Moliere Nguile-Makao and Alexandre Bureau. Semi-parametric maximum likelihood method for interaction in case-mother control-mother designs: Package SPmlficmcm. Journal of Statistical Software, 68(10):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v068i10;
http://www.jstatsoft.org/index.php/jss/article/view/
v068i10/v68i10.pdf.

Nordhausen:2021:DRT

[NMM^{+}21] Klaus Nordhausen, Markus Matilainen, Jari Miettinen, Joni Virta, and Sara Taskinen. Dimension reduction for time series in a blind source separation context using R. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v098i15;   https://www.jstatsoft.
org/index.php/jss/article/view/v098i15/v98i15.pdf.

Nevalainen:2006:SIM

[NO06] Jaakko Nevalainen and Hannu Oja. SAS/IML macros for a multivariate analysis of variance based on spatial signs. Journal of Statistical Software, 16(5):1–17, May 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v16/i05.

Nordhausen:2011:MLS

[NO11] Klaus Nordhausen and Hannu Oja. Multivariate L1 statistical methods: The package MNM. Journal of Statistical Software, 43(5):??, July 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v43/i05.

Nordh:2017:PPF

[Nor17] Jerker Nordh. pyParticleEst: A Python framework for particle-based estimation methods. Journal of Statistical Software, 78(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v078i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v078i03/v78i03.pdf.

Nordhausen:2008:TEM

[NOT08] Klaus Nordhausen, Hannu Oja, and David E. Tyler. Tools for exploring multivariate data: The package ICS. Journal of Statistical Software, 28(6):??, November 2008. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v28/i06.

Nadarajah:2016:NRP

[NR16] Saralees Nadarajah and Ricardo Rocha. Newdistns: An R package for new families of distributions. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v069i10;   https://www.jstatsoft.
org/index.php/jss/article/view/v069i10/v69i10.pdf.

Nowok:2016:SBC

[NRD16] Beata Nowok, Gillian M. Raab, and Chris Dibben. synthpop: Bespoke creation of synthetic data in R. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v074i11;   https://www.jstatsoft.
org/index.php/jss/article/view/v074i11/v74i11.pdf.

Narasimhan:2017:SDC

[NRG^{+}17] Balasubramanian Narasimhan, Daniel L. Rubin, Samuel M. Gross, Marina Bendersky, and Philip W. Lavori. Software for distributed computation on medical databases: A demonstration project. Journal of Statistical Software, 77 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v077i13;   https://www.jstatsoft.org/
index.php/jss/article/view/v077i13/v77i13.pdf.

Ntzoufras:2002:GVS

[Ntz02] Ioannis Ntzoufras. Gibbs variable selection using BUGS. Journal of Statistical Software, 7(7):1–19, ???? 2002. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v07/i07;   http://www.jstatsoft.org/
v07/i07/gvs-bugsfiles.zip;   http://www.jstatsoft.
org/v07/i07/GVSusingBUGS.pdf;   http://www.jstatsoft.
org/v07/i07/updates.

Nunes:2015:BRS

[Nun15] Matthew Nunes. Book review: Statistical Analysis of Network Data with R, Eric D. Kolaczyk and Gábor Csárdi, Springer-Verlag, New York, 2014. ISBN 978-1-4939-0983-4. 207 pp.

GBP 39.99 (P). http://www.springer.com/book/9781493909827.

Journal of Statistical Software, 66(BR-1):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066b01;

http://www.jstatsoft.org/index.php/jss/article/view/
v066b01/v66b01.pdf.

Nash:2011:UOA

[NV11] John C. Nash and Ravi Varadhan. Unifying optimization algorithms to aid software system users: optimx for R. Journal of Statistical Software, 43(9):??, August 2011. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v43/i09.

Ngo:2004:SMM

[NW04] Long Ngo and Matthew P. Wand. Smoothing with mixed model software. Journal of Statistical Software, 9(1):1–54, January 1, 2004. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v09/i01/Abstract%20for%20JCGS.
txt;   http://www.jstatsoft.org/v09/i01/code.zip;
http://www.jstatsoft.org/v09/i01/softpaper_010804.
pdf.

Oberski:2014:PLS

[Obe14] Daniel Oberski. lavaan.survey: An R package for complex survey analysis of structural equation models. Journal of Statistical Software, 57(1):??, March 2014. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v57/i01.

Olsbjerg:2015:CSP

[OC15] Maja Olsbjerg and Karl Bang Christensen. Code snippet: %lrasch mml: A SAS macro for marginal maximum likelihood estimation in longitudinal polytomous Rasch models. Journal of Statistical Software, 67(CS-2):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067c02;
http://www.jstatsoft.org/index.php/jss/article/view/
v067c02/v67c02.pdf.

Oliveira:2014:NRP

[OCRC14] María Oliveira, Rosa M. Crujeiras, and Alberto Rodríguez-Casal. NPCirc: An R package for nonparametric circular methods. Journal of Statistical Software, 61(9):??, November 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v61/i09.

OConnell:2011:HSM

[OH11] Jared O’Connell and Søren Højsgaard. Hidden semi Markov models for multiple observation sequences: The mhsmm package for R. Journal of Statistical Software, 39(4):??, March 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v39/i04.

OConnell:2017:CVS

[OHD17] Mark O’Connell, Catherine B. Hurley, and Katarina Domijan. Conditional visualization for statistical models: An introduction to the condvis package in R. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v081i05;   https://www.jstatsoft.
org/index.php/jss/article/view/v081i05/v81i05.pdf.

Overstall:2014:PCR

[OK14] Antony M. Overstall and Ruth King. conting: An R package for Bayesian analysis of complete and incomplete contingency tables. Journal of Statistical Software, 58(7):??, June 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v58/i07.

Oller:2017:FRP

[OL17] Ramon Oller and Klaus Langohr. FHtest: An R package for the comparison of survival curves with censored data. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v081i15;   https://www.
jstatsoft.org/index.php/jss/article/view/v081i15/v81i15.
pdf.

Ocana:2011:ISA

[OO11] Francisco A. Ocaña and Pablo Oñate. IndElec: a software for analyzing party systems and electoral systems. Journal of Statistical Software, 42(6):??, June 2011. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v42/i06.

Ooms:2010:BRB

[Oom10] Jeroen Ooms. Book review: SAS and R. Journal of Statistical Software, 37(BR-3):??, December 2010. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v37/b03.

Ooms:2013:RPE

[Oom13] Jeroen Ooms. The RAppArmor package: Enforcing security policies in R using dynamic sandboxing on Linux. Journal of Statistical Software, 55(7):??, November 2013. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v55/i07.

Okhrin:2014:HAC

[OR14] Ostap Okhrin and Alexander Ristig. Hierarchical Archimedean copulae: The HAC package. Journal of Statistical Software, 58 (4):??, June 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v58/i04.

Ostoic:2020:AAM

[Ost20] J. Antonio Rivero Ostoic. Algebraic analysis of multiple social networks with multiplex. Journal of Statistical Software, 92(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v092i11;   https://www.jstatsoft.org/
index.php/jss/article/view/v092i11/v92i11.pdf.

Otneim:2017:BRB

[Otn17] Håkon Otneim. Book review: Business Analytics Using R — A Practical Approach. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v081b01;   https://www.jstatsoft.org/
index.php/jss/article/view/v081b01/v81b01.pdf.

Overstall:2020:ARP

[OWA20] Antony M. Overstall, David C. Woods, and Maria Adamou. acebayes: an R package for Bayesian optimal design of experiments via approximate coordinate exchange. Journal of Statistical Software, 95(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v095i13;   https://www.jstatsoft.
org/index.php/jss/article/view/v095i13/v95i13.pdf.

Peng:2011:SSM

[PA11] Jyh-Ying Peng and John A. D. Aston. The state space models toolbox for MATLAB. Journal of Statistical Software, 41(6): ??, May 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v41/i06.

Paciorek:2007:BSG

[Pac07] Christopher J. Paciorek. Bayesian smoothing with Gaussian processes using Fourier basis functions in the spectralGP package. Journal of Statistical Software, 19(2):1–38, April 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v19/i02.

Pan:2009:CSS

[Pan09] Wei Pan. Code snippet: a SAS/IML macro for computing percentage points of Pearson distributions. Journal of Statistical Software, 31(CS-2):??, September 2009. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v31/c02.

Panse:2018:CSR

[Pan18] Christian Panse. Code snippet: Rectangular statistical cartograms in R: The recmap package. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v086c01;   https://www.jstatsoft.
org/index.php/jss/article/view/v086c01/v86c01.pdf.

Papastamoulis:2016:CSL

[Pap16] Panagiotis Papastamoulis. Code snippet: label.switching: An R package for dealing with the label switching problem in MCMC outputs. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v069c01;   https://www.jstatsoft.org/
index.php/jss/article/view/v069c01/v69c01.pdf.

Patefield:2000:FAC

[Pat00] Mike Patefield. Fast and accurate calculation of Owen’s T function. Journal of Statistical Software, 5(5):1–25, ???? 2000. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v05/i05;   http://www.jstatsoft.
org/v05/i05/t.f; http://www.jstatsoft.org/v05/i05/
t.pdf;   http://www.jstatsoft.org/v05/i05/updates.

Pataky:2016:RSO

[Pat16] Todd C. Pataky. rft1d: Smooth one-dimensional random field upcrossing probabilities in Python. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v071i07;   https://www.jstatsoft.
org/index.php/jss/article/view/v071i07/v71i07.pdf.

Pavia:2015:CST

[Pav15] Jose M. Pavia. Code snippet: Testing goodness-of-fit with the kernel density estimator: GoFKernel. Journal of Statistical Software, 66(CS-1):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066c01;
http://www.jstatsoft.org/index.php/jss/article/view/
v066c01/v66c01.pdf.

Pavia:2016:BRH

[Pav16] Jose M. Pavía. Book review: Humanities Data in R. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v069b02;   https://www.
jstatsoft.org/index.php/jss/article/view/v069b02/v69b02.
pdf.

Pavia:2020:BRR

[Pav20] Jose M. Pavía. Book review: R Graphics (3rd Edition). Journal of Statistical Software, 92(??):??, ???? 2020. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v092b01;   https://www.
jstatsoft.org/index.php/jss/article/view/v092b01/v92b01.
pdf.

Padoan:2015:ARF

[PB15] Simone A. Padoan and Moreno Bevilacqua. Analysis of random fields using CompRandFld. Journal of Statistical Software, 63(9):??, February 2015. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v63/i09.

Pebesma:2015:SSS

[PBR15] Edzer Pebesma, Roger Bivand, and Paulo Justiano Ribeiro. Software for spatial statistics. Journal of Statistical Software, 63(1):??, February 2015. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v63/i01.

Plummer:2011:LRC

[PC11] Martyn Plummer and Bendix Carstensen. Lexis: An R class for epidemiological studies with long-term follow-up. Journal of Statistical Software, 38(5):??, January 2011. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v38/i05.

Prates:2013:MFF

[PCL13] Marcos Oliveira Prates, Celso Rômulo Barbosa Cabral, and Víctor Hugo Lachos. mixsmsn: Fitting finite mixture of scale mixture of skew-normal distributions. Journal of Statistical Software, 54(12):??, September 2013. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v54/i12.

Petersen:2019:DYA

[PE19] Anne Helby Petersen and Claus Thorn Ekstrøm. dataMaid: Your assistant for documenting supervised data quality screening in R. Journal of Statistical Software, 90(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v090i06;
https://www.jstatsoft.org/index.php/jss/article/view/
v090i06/v90i06.pdf.

Pebesma:2012:SST

[Peb12] Edzer Pebesma. spacetime: Spatio-temporal data in R. Journal of Statistical Software, 51(7):??, November 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v51/i07.

Pelagatti:2011:SSM

[Pel11] Matteo M. Pelagatti. State space methods in Ox/SsfPack. Journal of Statistical Software, 41(3):??, May 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v41/i03.

Peng:2003:MDP

[Pen03] Roger D. Peng. Multi-dimensional point process models in R. Journal of Statistical Software, 8(16):1–27, ???? 2003. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v08/i16;   http://www.jstatsoft.
org/v08/i16/ptprocR102.pdf.

Peng:2008:CSM

[Pen08a] Roger Peng. Code snippet: a method for visualizing multivariate time series data. Journal of Statistical Software, 25(CS-1): 1–17, March 2008. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v25/c01.

Peng:2008:CDS

[Pen08b] Roger D. Peng. Caching and distributing statistical analyses in R. Journal of Statistical Software, 26(7):??, July 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v26/i07.

PerezSainzdeRozas:2003:UMB

[Per03] Gloria Perez Sainz de Rozas. Using Mathematica to build non-parametric statistical tables. Journal of Statistical Software, 8(4):1–26, ???? 2003. CODEN JSSOBK. ISSN

1548-7660.  URL http://www.jstatsoft.org/v08/i04/
tablasjss2.pdf;   http://www.jstatsoft.org/v08/i4.

Peterson:1997:CCP

[Pet97a] Leif E. Peterson. COVAR: Computer program for multifactor relative risks and tests of hypotheses using a variancecovariance matrix from linear and log-linear regression. Journal of Statistical Software, 2(4):1–21, ???? 1997. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v02/i04;   http://www.jstatsoft.org/v02/i04/covar.
pdf;   http://www.jstatsoft.org/v02/i04/updates.

Peterson:1997:PPI

[Pet97b] Leif E. Peterson. PIRLS: Poisson Iteratively Reweighted Least Squares computer program for additive, multiplicative, power, and non-linear models. Journal of Statistical Software, 2(5):1–28, ???? 1997. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v02/i05;
 http://www.jstatsoft.org/v02/i05/code/;   http://
www.jstatsoft.org/v02/i05/pirls.pdf;   http://www.
jstatsoft.org/v02/i05/pirls.ps.gz;   http://www.jstatsoft.
org/v02/i05/pirls/;   http://www.jstatsoft.org/v02/
i05/updates.

Petris:2010:RPD

[Pet10] Giovanni Petris. An R package for dynamic linear models. Journal of Statistical Software, 36(12):??, October 2010.

CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v36/i12.

Pfaff:2008:VSS

[Pfa08] Bernhard Pfaff. VAR, SVAR and SVEC models: Implementation within R package vars. Journal of Statistical Software, 27(4):??, July 2008. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v27/i04.

Pelissier:2015:PAP

[PG15] Raphaël Pélissier and François Goreaud. ads package for R: A fast unbiased implementation of the K-function family for studying spatial point patterns in irregular-shaped sampling windows. Journal of Statistical Software, 63(6):??, February 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v63/i06.

Patil:2010:PBS

[PHF10] Anand Patil, David Huard, and Christopher J. Fonnesbeck. PyMC: Bayesian stochastic modelling in Python. Journal of Statistical Software, 35(4):??, July 2010. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v35/i04.

Phillips:2010:CSR

[Phi10] Kem Phillips. Code snippet: R functions to symbolically compute the central moments of the multivariate normal distribution. Journal of Statistical Software, 33(CS-1):??, February 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v33/c01.

Pierce:2005:BRI

[Pie05] John G. Pierce. Book review: Introduction to MATLAB with Numerical Preliminaries. Journal of Statistical Software, 14 (BR-3):1–3, August 2005. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v14/b03.

Piras:2010:SSM

[Pir10] Gianfranco Piras. sphet: Spatial models with heteroskedastic innovations in R. Journal of Statistical Software, 35(1):??, June 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v35/i01.

Perez-Jaume:2017:TOT

[PJSPC17] Sara Perez-Jaume, Konstantina Skaltsa, Natàlia Pallarès, and Josep L. Carrasco. ThresholdROC: Optimum threshold estimation tools for continuous diagnostic tests in R. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v082i04;   https://www.
jstatsoft.org/index.php/jss/article/view/v082i04/v82i04.
pdf.

Pineda-Krch:2008:GIG

[PK08] Mario Pineda-Krch. GillespieSSA: Implementing the Gillespie stochastic simulation algorithm in R. Journal of Statistical Software, 25(12):1–18, April 2008. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v25/i12.

Pham:2020:CSA

[PKC^{+}20] Thanh V. Pham, Jeffrey D. Kromrey, Yi-Hsin Chen, Eun Sook Kim, Diep T. Nguyen, and Yan Wang. Code snippet: ANOVA robust: a SAS macro for parametric tests of mean differences in one-factor ANOVA models. Journal of Statistical Software, 95(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v095c02;   https://www.jstatsoft.
org/index.php/jss/article/view/v095c02/v95c02.pdf.

Panneton:2005:XRN

[PL05] François Panneton and Pierre L’Ecuyer. On the xorshift random number generators. ACM Transactions on Modeling and Computer Simulation, 15(4):346–361, October 2005. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See [Mar03, Bre04, Vig16].

Poole:2011:SRC

[PLLC11] Keith Poole, Jeffrey Lewis, James Lo, and Royce Carroll. Scaling roll call votes with wnominate in R. Journal of Statistical Software, 42(14):??, June 2011. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v42/i14.

Proust-Lima:2017:EEM

[PLPL17] Cécile Proust-Lima, Viviane Philipps, and Benoit Liquet. Estimation of extended mixed models using latent classes and latent processes: The R package lcmm. Journal of Statistical Software, 78(??):??, ???? 2017. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v078i02;   https://www.jstatsoft.
org/index.php/jss/article/view/v078i02/v78i02.pdf.

Poole:2016:RBS

[PLR^{+}16] Keith T. Poole, Jeffrey B. Lewis, Howard Rosenthal, James Lo, and Royce Carroll. Recovering a basic space from issue scales in R. Journal of Statistical Software, 69 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v069i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v069i07/v69i07.pdf.

Pateiro-Lopez:2010:GCH

[PLRC10] Beatriz Pateiro-López and Alberto Rodríguez-Casal. Generalizing the convex hull of a sample: The R package alphahull. Journal of Statistical Software, 34(5):??, April 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v34/i05.

Paciorek:2015:PGP

[PLZ^{+}15] Christopher J. Paciorek, Benjamin Lipshitz, Wei Zhuo, Prabhat, Cari G. G. Kaufman, and Rollin C. Thomas. Parallelizing Gaussian process calculations in R. Journal of Statistical Software, 63(10):??, February 2015. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v63/i10.

Pokotylo:2019:DDB

[PMD19] Oleksii Pokotylo, Pavlo Mozharovskyi, and Rainer Dyckerhoff. Depth and depth-based classification with R package ddalpha. Journal of Statistical Software, 91 (??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v091i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v091i05/v91i05.pdf.

Punzo:2018:CRP

[PMM18] Antonio Punzo, Angelo Mazza, and Paul D. McNicholas. ContaminatedMixt: An R package for fitting parsimonious mixtures of multivariate contaminated normal distributions. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v085i10;   https://
www.jstatsoft.org/index.php/jss/article/view/v085i10/
v85i10.pdf.

Payton:2015:PMP

[PMW^{+}15] Quinn C. Payton, Michael G. McManus, Marc H. Weber, Anthony R. Olsen, and Thomas M. Kincaid. micromap: A package for linked micromaps. Journal of Statistical Software, 63 (2):??, February 2015. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v63/i02.

Pikounis:2013:CPC

[PO13] Bill Pikounis and John Oleynick. The cg package for comparison of groups. Journal of Statistical Software, 52(1):??, February 2013. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v52/i01.

Posekany:2015:BRB

[Pos15] Alexandra Posekany. Book review: Bayesian Essentials with R, Jean Michel Marin and Christian Robert, Springer-Verlag, New York, 2014. ISBN 978-1-4614-8686-2. 296 pp. $69.99 (P).

https://www.ceremade.dauphine.fr/ xian/books.html. Jour-

nal of Statistical Software, 67(BR-3):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067b03;
http://www.jstatsoft.org/index.php/jss/article/view/
v067b03/v67b03.pdf.

Petris:2011:SSM

[PP11] Giovanni Petris and Sonia Petrone. State space models in R. Journal of Statistical Software, 41(4):??, May 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v41/i04.

Pan:2017:JRP

[PP17] Jianxin Pan and Yi Pan. jmcm: An R package for joint meancovariance modeling of longitudinal data. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v082i09;   https://www.jstatsoft.
org/index.php/jss/article/view/v082i09/v82i09.pdf.

Perme:2018:NRS

[PP18] Maja Pohar Perme and Klemen Pavlic. Nonparametric relative survival analysis with the R package relsurv. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v087i08;   https://www.
jstatsoft.org/index.php/jss/article/view/v087i08/v87i08.
pdf.

Pickle:2015:MEC

[PPC15] Linda Williams Pickle, James B. Pearson, and Daniel B. Carr. micromapST: Exploring and communicating geospatial patterns in US state data. Journal of Statistical Software, 63(3): ??, February 2015. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v63/i03.

Palomo:2015:SRP

[PPGD15] Jesús Palomo, Rui Paulo, and Gonzalo García-Donato. SAVE: An R package for the statistical analysis of computer models. Journal of Statistical Software, 64(13):??, April 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v64/i13.

Palmi-Perales:2021:BMS

[PPGRMB21] Francisco Palmí-Perales, Virgilio Gómez-Rubio, and Miguel A. Martinez-Beneito. Bayesian multivariate spatial models for lattice data with INLA. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v098i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v098i02/v98i02.pdf.

Polzehl:2020:PWA

[PPT20] Jörg Polzehl, Kostas Papafitsoros, and Karsten Tabelow. Patch-wise adaptive weights smoothing in R. Journal of Statistical Software, 95(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v095i06;   https://www.jstatsoft.
org/index.php/jss/article/view/v095i06/v95i06.pdf.

Pemstein:2011:SSL

[PQM11] Daniel Pemstein, Kevin M. Quinn, and Andrew D. Martin. The Scythe Statistical Library: An open source C++ library for statistical computation. Journal of Statistical Software, 42 (12):??, June 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v42/i12.

Petzoldt:2007:SOO

[PR07] Thomas Petzoldt and Karsten Rinke. simecol: An objectoriented framework for ecological modeling in R. Journal of Statistical Software, 22(9):1–31, September 2007. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v22/i09.

Pretis:2018:AGS

[PRS18] Felix Pretis, J. James Reade, and Genaro Sucarrat. Automated general-to-specific (GETS) regression modeling and indicator saturation for outliers and structural breaks. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v086i03;   https://www.jstatsoft.
org/index.php/jss/article/view/v086i03/v86i03.pdf.

Pearce:2017:HRP

[PSS^{+}17] Robert G. Pearce, R. Woodrow Setzer, Cory L. Strope, Nisha S. Sipes, and John F. Wambaugh. httk: R package for high-throughput toxicokinetics. Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v079i04;   https://www.jstatsoft.org/
index.php/jss/article/view/v079i04/v79i04.pdf.

Pham:2020:PRP

[PSS20] Thong Pham, Paul Sheridan, and Hidetoshi Shimodaira. PAFit: an R package for the non-parametric estimation of preferential attachment and node fitness in temporal complex networks. Journal of Statistical Software, 92 (??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v092i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v092i03/v92i03.pdf.

Perez:2007:RJI

[PSW07] Carlos Javier Pérez, Hansgeorg Schwibbe, and Petra Weidner. RAGE: a Java-implemented visual random generator. Journal of Statistical Software, 17(10):1–10, January 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v17/i10.

Pena:2020:CSG

[PSY20] Daniel Peña, Ezequiel Smucler, and Victor J. Yohai. Code snippet: gdpc: an R package for generalized dynamic principal components. Journal of Statistical Software, 92 (??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v092c02;   https://www.jstatsoft.org/
index.php/jss/article/view/v092c02/v92c02.pdf.

Polzehl:2007:ASD

[PT07] Jörg Polzehl and Karsten Tabelow. Adaptive smoothing of digital images: The R package adimpro. Journal of Statistical Software, 19(1):1–17, March 2007. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v19/i01.

Polzehl:2009:SAS

[PT09] Jörg Polzehl and Karsten Tabelow. Structural adaptive smoothing in diffusion tensor imaging: The R package dti. Journal of Statistical Software, 31(9):??, September 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v31/i09.

Polzehl:2011:BGM

[PT11] Jörg Polzehl and Karsten Tabelow. Beyond the Gaussian model in diffusion-weighted imaging: The package dti. Journal of Statistical Software, 44(12):??, October 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v44/i12.

Pilhofer:2013:NAV

[PU13] Alexander Pilhöfer and Antony Unwin. New approaches in visualization of categorical data: R package extracat. Journal of Statistical Software, 53(7):??, May 2013. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v53/i07.

Putter:2011:SIA

[Put11] Hein Putter. Special issue about competing risks and multistate models. Journal of Statistical Software, 38(1):??, January 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v38/i01.

Peterson:2014:SAT

[PV14] Erin E. Peterson and Jay M. Ver Hoef. STARS: An ArcGIS toolset used to calculate the spatial information needed to fit spatial statistical models to stream network data. Journal of Statistical Software, 56(2):??, January 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v56/i02.

Philippe:2020:CSA

[PV20] Anne Philippe and Marie-Anne Vibet. Code snippet: Analysis of archaeological phases using the R package ArchaeoPhases. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v093c01;   https://
www.jstatsoft.org/index.php/jss/article/view/v093c01/
v93c01.pdf.

Perez:2012:CSI

[PZK^{+}12] Adriana Pérez, Saijuan Zhang, Victor Kipniss, Laurence S. Freedman, and Raymond J. Carroll. Code snippet: Intake epis food(): An R function for fitting a bivariate nonlinear measurement error model to estimate usual and energy intake for episodically consumed foods. Journal of Statistical Software, 46(CS-3): ??, March 2012. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v46/c03.

Qian:2019:WDB

[QY19] Zhaozhi Qian and Philip L. H. Yu. Weighted distance-based models for ranking data using the R package rankdist. Journal of Statistical Software, 90(??):??, ???? 2019. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v090i05;   https://www.
jstatsoft.org/index.php/jss/article/view/v090i05/v90i05.
pdf.

Ramirez:2000:GD

[Ram00] Donald E. Ramirez. The generalized F distribution. Journal of Statistical Software, 5(1):1–14, ???? 2000. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v05/i01;   http://www.jstatsoft.org/v05/i01/GENF.f;
 http://www.jstatsoft.org/v05/i01/GENF.pdf;   http:
//www.jstatsoft.org/v05/i01/updates.

Rigdon:2018:CSN

[RBB18] Joseph Rigdon, Michael Baiocchi, and Sanjay Basu. Code snippet: Near-far matching in R: The nearfar package. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v086c05;   https://www.
jstatsoft.org/index.php/jss/article/view/v086c05/v86c05.
pdf.

Riza:2015:FFR

[RBHB15] Lala Septem Riza, Christoph Bergmeir, Francisco Herrera, and José Manuel Benítez. frbs: Fuzzy rule-based systems for classification and regression in R. Journal of Statistical Software, 65(6):??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v65/i06.

Rakshit:2019:ECS

[RBN19] Suman Rakshit, Adrian Baddeley, and Gopalan Nair. Efficient code for second order analysis of events on a linear network. Journal of Statistical Software, 90 (??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v090i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v090i01/v90i01.pdf.

Risser:2017:LLE

[RC17] Mark D. Risser and Catherine A. Calder. Local likelihood estimation for covariance functions with spatially-varying parameters: The convoSPAT package for R. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v081i14;   https://www.jstatsoft.
org/index.php/jss/article/view/v081i14/v81i14.pdf.

Robbins:2021:MSC

[RD21] Michael W. Robbins and Steven Davenport. microsynth: Synthetic control methods for disaggregated and microlevel data in R. Journal of Statistical Software, 97 (??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v097i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v097i02/v97i02.pdf.

Recchia:2010:CSC

[Rec10a] Anthony Recchia. Code snippet: Contiguity-constrained hierarchical agglomerative clustering using SAS. Journal of Statistical Software, 33(CS-2):??, February 2010. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v33/c02.

Recchia:2010:CSR

[Rec10b] Anthony Recchia. Code snippet: R-squared measures for twolevel hierarchical linear models using SAS. Journal of Statistical Software, 32(CS-2):??, January 2010. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v32/c02.

Rodiger:2012:RCG

[RFKM12] Stefan Rödiger, Thomas Friedrichsmeier, Prasenjit Kapat, and Meik Michalke. RKWard: a comprehensive graphical user interface and integrated development environment for statistical analysis with R. Journal of Statistical Software, 49(9):??, June 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v49/i09.

Ritter:2007:ICD

[RG07] Christian Ritter and Jean Gilliard. ImpuR: a collection of diagnostic tools developed in R in the context of peak impurity detection in HPLC-DAD but potentially useful with other types of time-intensity matrices. Journal of Statistical Software, 18 (9):1–18, January 2007. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v18/i09.

Roustant:2012:DDT

[RGD12] Olivier Roustant, David Ginsbourger, and Yves Deville. DiceKriging, DiceOptim: Two R packages for the analysis of computer experiments by kriging-based metamodeling and optimization. Journal of Statistical Software, 51(1):??, October 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v51/i01.

Rossmann:2020:CSW

[RGK20] Joss Roßmann, Tobias Gummer, and Lars Kaczmirek. Code snippet: Working with user agent strings in Stata: The parseuas command. Journal of Statistical Software, 92 (??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/

article/view/v092c01;   https://www.jstatsoft.org/
index.php/jss/article/view/v092c01/v92c01.pdf.

Rizopoulos:2006:LRP

[Riz06] Dimitris Rizopoulos. ltm: An R package for latent variable modeling and item response analysis. Journal of Statistical Software, 17(5):1–25, November 2006. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v17/i05.

Rizopoulos:2010:JRP

[Riz10] Dimitris Rizopoulos. JM: An R package for the joint modelling of longitudinal and time-to-event data. Journal of Statistical Software, 35(9):??, July 2010. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v35/i09.

Rizopoulos:2016:RPJ

[Riz16] Dimitris Rizopoulos. The R package JMbayes for fitting joint models for longitudinal and time-to-event data using MCMC. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v072i07;   https://
www.jstatsoft.org/index.php/jss/article/view/v072i07/
v72i07.pdf.

Ritter:2014:RPM

[RJH14] Stephan J. Ritter, Nicholas P. Jewell, and Alan E. Hubbard. R package multiPIM: A causal inference approach to variable importance analysis. Journal of Statistical Software, 57(8):??, April 2014. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v57/i08.

Ruckdeschel:2014:GPC

[RK14] Peter Ruckdeschel and Matthias Kohl. General purpose convolution algorithm in S4 classes by means of FFT. Journal of Statistical Software, 59(4):??, August 2014. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v59/i04.

Rondeau:2012:FRP

[RMG12] Virginie Rondeau, Yassin Mazroui, and Juan R. Gonzalez. frailtypack: An R package for the analysis of correlated survival data with frailty models using penalized likelihood estimation or parametrical estimation. Journal of Statistical Software, 47(4):??, April 2012. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v47/i04.

Rosenblad:2008:SRG

[Ros08] Andreas Rosenblad. Software review: gret1 1.7.3. Journal of Statistical Software, 25(SR-1):1–14, March 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v25/s01.

Rosario:2009:BRB

[Ros09] Ryan Rosario. Book review: Practical Text Mining with Perl. Journal of Statistical Software, 29(BR-9):??, January 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v29/b09.

Rosseel:2012:LRP

[Ros12] Yves Rosseel. lavaan: An R package for structural equation modeling. Journal of Statistical Software, 48(2):??, May 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v48/i02.

Ross:2015:PNS

[Ros15] Gordon J. Ross. Parametric and nonparametric sequential change detection in R: The cpm package. Journal of Statistical Software, 66(3):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066i03;
http://www.jstatsoft.org/index.php/jss/article/view/
v066i03/v66i03.pdf.

Rosenblad:2017:BRM

[Ros17] Andreas Rosenblad. Book review: Multivariate Statistical Methods: A Primer (4th Edition). Journal of Statistical Software, 78(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v078b03;   https://www.jstatsoft.org/
index.php/jss/article/view/v078b03/v78b03.pdf.

Rover:2020:BRE

[Röv20] Christian Röver. Bayesian random-effects meta-analysis using the bayesmeta R package. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/

article/view/v093i06;   https://www.jstatsoft.org/
index.php/jss/article/view/v093i06/v93i06.pdf.

Riani:2015:CSF

[RPC15] Marco Riani, Domenico Perrotta, and Andrea Cerioli. Code snippet: The forward search for very large datasets. Journal of Statistical Software, 67(CS-1):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067c01;
http://www.jstatsoft.org/index.php/jss/article/view/
v067c01/v67c01.pdf.

Roelstraete:2011:FRP

[RR11] Bjorn Roelstraete and Yves Rosseel. FIAR: An R package for analyzing functional integration in the brain. Journal of Statistical Software, 44(13):??, October 2011. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v44/i13.

Ritz:2005:BAU

[RS05] Christian Ritz and Jens C. Streibig. Bioassay analysis using R. Journal of Statistical Software, 12(5):1–22, ???? 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/counter.php?id=112&url=v12/i05&ct=2;
 http://www.jstatsoft.org/counter.php?id=112&url=v12/
i05/v12i05.pdf&ct=1.

Raimondo:2007:WTR

[RS07] Marc Raimondo and Michael Stewart. The WaveD transform in R: Performs fast translation-invariant wavelet deconvolution. Journal of Statistical Software, 21(2):1–28, July 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v21/i02.

Ratcliffe:2008:GMT

[RS08] Sarah J. Ratcliffe and Justine Shults. GEEQBOX: a MAT-LAB toolbox for generalized estimating equations and quasileast squares. Journal of Statistical Software, 25(14):1–14, May 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v25/i14.

Roberts:2019:SRP

[RST19] Margaret E. Roberts, Brandon M. Stewart, and Dustin Tingley. stm: an R package for structural topic models. Journal of Statistical Software, 91(??):??, ???? 2019. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v091i02;   https://www.
jstatsoft.org/index.php/jss/article/view/v091i02/v91i02.
pdf.

Ryeznik:2015:RMS

[RSW15] Yevgen Ryeznik, Oleksandr Sverdlov, and Weng Kee Wong. RARtool: A MATLAB software package for designing response-adaptive randomized clinical trials with time-to-event outcomes. Journal of Statistical Software, 66(1):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066i01;
http://www.jstatsoft.org/index.php/jss/article/view/
v066i01/v66i01.pdf.

Reid:2014:RPC

[RT14] Stephen Reid and Rob Tibshirani. Regularization paths for conditional logistic regression: The clogitL1 package. Journal of Statistical Software, 58(12):??, July 2014. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v58/i12.

Rufibach:2009:CSR

[Ruf09] Kaspar Rufibach. Code snippet: reporttools: R functions to generate LaTeX tables of descriptive statistics. Journal of Statistical Software, 31(CS-1):??, August 2009. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v31/c01.

Ruiz:2016:BRM

[Rui16] Diego Andres Perez Ruiz. Book review: Modern Optimization with R. Journal of Statistical Software, 70 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v070b03;   https://www.jstatsoft.org/
index.php/jss/article/view/v070b03/v70b03.pdf.

Rusch:2015:BRR

[Rus15] Thomas Rusch. Book review: R for Marketing Research and Analytics, Chris Chapman, Elea McDonnell Feit, Springer-Verlag, Switzerland, 2015. ISBN 978-3-319-14435-1. 454 pp.

USD 64.99 (P). http://www.springer.com/book/9783319144351.

Journal of Statistical Software, 67(BR-2):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067b02;
http://www.jstatsoft.org/index.php/jss/article/view/
v067b02/v67b02.pdf.

Rutter:2019:GRP

[RVCG19] Lindsay Rutter, Susan VanderPlas, Dianne Cook, and Michelle A. Graham. ggenealogy: an R package for visualizing genealogical data. Journal of Statistical Software, 89(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v089i13;   https://www.jstatsoft.org/
index.php/jss/article/view/v089i13/v89i13.pdf.

Royston:2011:MIC

[RW11] Patrick Royston and Ian R. White. Multiple Imputation by Chained Equations (MICE): Implementation in Stata. Journal of Statistical Software, 45(4):??, December 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v45/i04.

Sabatti:2007:BRB

[Sab07] Chiara Sabatti. Book review: The Statistics of Gene Mapping. Journal of Statistical Software, 21(BR-2):1–2, September 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v21/b02.

Sachs:2017:CSP

[Sac17] Michael C. Sachs. Code snippet: plotROC: A tool for plotting ROC curves. Journal of Statistical Software, 79 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v079c02;   https://www.jstatsoft.org/
index.php/jss/article/view/v079c02/v79c02.pdf.

Sanchez:2006:BRB

[San06] Juana Sanchez. Book review: Handbook of Univariate and Multivariate Data Analysis and Interpretation with SPSS. Journal of Statistical Software, 16(BR-4):1–3, July 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v16/b04.

Sanchez:2007:BRB

[San07] Juana Sanchez. Book review: “Linear Models with R” and “Extending the Linear Model with R”. Journal of Statistical Software, 17(BR-4):1–3, January 2007. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v17/b04.

Sanchez:2009:BRB

[San09] Juana Sanchez. Book review: Probability with R — an Introduction with Computer Science Applications. Journal of Statistical Software, 29(BR-2):??, January 2009. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v29/b02.

Sanchez:2010:BRBb

[San10a] Juana Sanchez. Book review: Data Analysis and Graphics Using R — An Example-Based Approach (3rd Edition). Journal of Statistical Software, 36(BR-6):??, September 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v36/b06.

Sanchez:2010:BRBa

[San10b] Juana Sanchez. Book review: Introductory Time Series with R. Journal of Statistical Software, 32(BR-4):??, January 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v32/b04.

Sanchez:2011:BRB

[San11] Juana Sanchez. Book review: Hidden Markov Models for Time Series — An Introduction Using R. Journal of Statistical Software, 43(BR-4):??, October 2011. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v43/b04.

Sargin:2009:BRB

[Sar09] Anatol Sargin. Book review: Statistics and Data with R: An Applied Approach Through Examples. Journal of Statistical Software, 30(BR-6):??, April 2009. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v30/b06.

Sevcikova:2011:BRP

[SAR11] Hana Sevcíková, Leontine Alkema, and Adrian Raftery. bayesTFR: An R package for probabilistic projections of the total fertility rate. Journal of Statistical Software, 43(1):??, July 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v43/i01.

Sarrias:2016:DCM

[Sar16] Mauricio Sarrias. Discrete choice models with random parameters in R: The Rchoice package. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v074i10;   https://www.jstatsoft.
org/index.php/jss/article/view/v074i10/v74i10.pdf.

Savitsky:2016:BNM

[Sav16] Terrance D. Savitsky. Bayesian nonparametric mixture estimation for time-indexed functional data in R. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v072i02;   https://www.jstatsoft.
org/index.php/jss/article/view/v072i02/v72i02.pdf.

Somerville:2001:FSI

[SB01] Paul N. Somerville and Frank Bretz. FORTRAN 90 and SAS-IML programs for computation of critical values for multiple testing and simultaneous confidence intervals. Journal of Statistical Software, 6(5):1–14, ???? 2001. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v06/
i05;   http://www.jstatsoft.org/v06/i05/EXAMPLE/;
http://www.jstatsoft.org/v06/i05/FORTRAN/;   http:/
/www.jstatsoft.org/v06/i05/jss2r.pdf;   http://www.
jstatsoft.org/v06/i05/SAS/;   http://www.jstatsoft.
org/v06/i05/updates.

Strand:2004:SNT

[SBL04] Matthew Strand, Shalender Bhasin, and Hsiao-Pin Liu. A set of nonparametric tests for experiments with lattice-ordered means: Theory, programs, and applications. Journal of Statistical Software, 9(6):1–16, ???? 2004. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/counter.php?
id=87&url=v09/i06/paper.pdf&ct=1.

Santana:2010:MMP

[SBL^{+}10] Roberto Santana, Concha Bielza, Pedro Larrañaga, Jose A. Lozano, Carlos Echegoyen, Alexander Mendiburu, Rubén Armañanzas, and Siddartha Shakya. Mateda-2.0: a MAT-LAB package for the implementation and analysis of estimation of distribution algorithms. Journal of Statistical Software, 35(7):??, July 2010. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v35/i07.

Strickland:2014:PPM

[SBMD14] Christopher M. Strickland, Robert L. Burdett, Kerrie L. Mengersen, and Robert J. Denham. PySSM: A Python module for Bayesian inference of linear Gaussian state space models. Journal of Statistical Software, 57(6):??, April 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v57/i06.

Shin:2006:CSL

[SBMG06] Ji-Hyung Shin, Sigal Blay, Brad McNeney, and Jinko Graham. Code snippet: LDheatmap: An R function for graphical display of pairwise linkage disequilibria between single nucleotide polymorphisms. Journal of Statistical Software, 16(CS-3):1–9, September 2006. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v16/c03.

Sardy:2007:BRB

[SCD07] Sylvain Sardy and Valerie Chavez-Demoulin. Book review: Modeling Financial Time Series with S-PLUS (2nd Edition). Journal of Statistical Software, 17(BR-6):1–2, February 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v17/b06.

Schrepp:2006:IPC

[Sch06] Martin Schrepp. ITA 2.0: a program for classical and inductive item tree analysis. Journal of Statistical Software, 16(10):1–14, September 2006. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v16/i10.

Schmidberger:2008:BRB

[Sch08] Markus Schmidberger. Book review: Introduction to Machine Learning and Bioinformatics. Journal of Statistical Software, 28(BR-2):??, November 2008. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v28/b02.

Scheipl:2011:SBV

[Sch11] Fabian Scheipl. spikeSlabGAM: Bayesian variable selection, model choice and regularization for generalized additive mixed models in R. Journal of Statistical Software, 43(14):??, September 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v43/i14.

Scholz:2016:RPC

[Sch16] Michael Scholz. R package clickstream: Analyzing clickstream data with Markov chains. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v074i04;   https://www.jstatsoft.org/
index.php/jss/article/view/v074i04/v74i04.pdf.

Sjostrand:2018:SMT

[SCL^{+}18] Karl Sjöstrand, Line Harder Clemmensen, Rasmus Larsen, Gudmundur Einarsson, and Bjarne Ersbøll. SpaSM: A MAT-LAB toolbox for sparse statistical modeling. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v084i10;   https://www.jstatsoft.
org/index.php/jss/article/view/v084i10/v84i10.pdf.

Scrucca:2001:NKS

[Scr01] Luca Scrucca. Nonparametric kernel smoothing methods. the sm library in Xlisp–Stat. Journal of Statistical Software, 6(7):1–49, ???? 2001. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v06/i07;   http:
//www.jstatsoft.org/v06/i07/sm-doc_r.pdf;   http:/
/www.jstatsoft.org/v06/i07/sm_r.zip;   http://www.
jstatsoft.org/v06/i07/updates.

Scrucca:2013:GPG

[Scr13] Luca Scrucca. GA: a package for genetic algorithms in R. Journal of Statistical Software, 53(4):??, April 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v53/i04.

Slaets:2013:WFD

[SCS13] Leen Slaets, Gerda Claeskens, and Bernard W. Silverman. Warping functional data in R and C via a Bayesian multiresolution approach. Journal of Statistical Software, 55(3):??, October 2013. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v55/i03.

Scutari:2010:LBN

[Scu10] Marco Scutari. Learning Bayesian networks with the bnlearn R package. Journal of Statistical Software, 35(3):??, July 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v35/i03.

Scutari:2017:BNC

[Scu17] Marco Scutari. Bayesian network constraint-based structure learning algorithms: Parallel and optimized implementations in the bnlearn R package. Journal of Statistical Software, 77(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v077i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v077i02/v77i02.pdf.

Stajdohar:2013:INE

[SD13] Miha Stajdohar and Janez Demsar. Interactive network exploration with Orange. Journal of Statistical Software, 53(6): ??, April 2013. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v53/i06.

Sarrias:2017:MLM

[SD17] Mauricio Sarrias and Ricardo Daziano. Multinomial logit models with continuous and discrete individual heterogeneity in R: The gmnl package. Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v079i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v079i02/v79i02.pdf.

Schulte:2012:MLC

[SDDD12] Eric Schulte, Dan Davison, Thomas Dye, and Carsten Dominik. A multi-language computing environment for literate programming and reproducible research. Journal of Statistical Software, 46(3):??, January 2012. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v46/i03.

Speidel:2020:RPH

[SDJ20] Matthias Speidel, Jörg Drechsler, and Shahab Jolani. The R package hmi: a convenient tool for hierarchical multiple imputation and beyond. Journal of Statistical Software, 95(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v095i09;   https://www.jstatsoft.org/
index.php/jss/article/view/v095i09/v95i09.pdf.

Sax:2018:SAX

[SE18] Christoph Sax and Dirk Eddelbuettel. Seasonal adjustment by X-13ARIMA-SEATS in R. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v087i11;   https://www.jstatsoft.org/
index.php/jss/article/view/v087i11/v87i11.pdf.

Schill:2014:SPL

[SED14] Walter Schill, Dirk Enders, and Karsten Drescher. A SAS package for logistic two-phase studies. Journal of Statistical Software, 57(9):??, April 2014. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v57/i09.

Sekhon:2011:MPS

[Sek11] Jasjeet S. Sekhon. Multivariate and propensity score matching software with automated balance optimization: The matching package for R. Journal of Statistical Software, 42(7):??, June 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v42/i07.

Selukar:2011:SSM

[Sel11] Rajesh Selukar. State space modeling using SAS. Journal of Statistical Software, 41(12):??, May 2011. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v41/i12.

Smith:2016:MVM

[SF16] David M. Smith and Malcolm J. Faddy. Mean and variance modeling of under- and overdispersed count data. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v069i06;   https://www.
jstatsoft.org/index.php/jss/article/view/v069i06/v69i06.
pdf.

Saldana:2018:SRP

[SF18] Diego Franco Saldana and Yang Feng. SIS: An R package for sure independence screening in ultrahigh-dimensional statistical models. Journal of Statistical Software, 83 (??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v083i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v083i02/v83i02.pdf.

Smith:2019:MVM

[SF19] David M. Smith and Malcolm J. Faddy. Mean and variance modeling of under-dispersed and over-dispersed grouped binary data. Journal of Statistical Software, 90 (??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v090i08;   https://www.jstatsoft.org/
index.php/jss/article/view/v090i08/v90i08.pdf.

Schnell:2020:CMA

[SFC20] Patrick M. Schnell, Mark Fiecas, and Bradley P. Carlin. credsubs: Multiplicity-adjusted subset identification. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v094i07;   https://www.
jstatsoft.org/index.php/jss/article/view/v094i07/v94i07.
pdf.

Simon:2011:RPC

[SFHT11] Noah Simon, Jerome H. Friedman, Trevor Hastie, and Rob Tibshirani. Regularization paths for Cox’s Proportional Hazards Model via coordinate descent. Journal of Statistical Software, 39(5):??, March 2011. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v39/i05.

Santos-Fernandez:2012:MRP

[SFS12] Edgar Santos-Fernández and Michele Scagliarini. MPCI: An R package for computing multivariate process capability indices. Journal of Statistical Software, 47(7):??, April 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v47/i07.

Su:2011:MID

[SGHY11] Yu-Sung Su, Andrew Gelman, Jennifer Hill, and Masanao Yajima. Multiple imputation with diagnostics (mi) in R: Opening windows into the black box. Journal of Statistical Software, 45(2):??, December 2011. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v45/i02.

Sanil:2003:NWW

[SGK03] Ashish Sanil, Shanti Gomatam, and Alan F. Karr. NISS Web-Swap: a Web service for data swapping. Journal of Statistical Software, 8(7):1–12, ???? 2003. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v08/i07/
nisswebswap200206.pdf;   http://www.jstatsoft.org/
v08/i7.

Strona:2014:CSN

[SGS^{+}14] Giovanni Strona, Paolo Galli, Davide Seveso, Simone Montano, and Simone Fattorini. Code snippet: Nestedness for dummies (NeD): A user-friendly Web interface for exploratory nestedness analysis. Journal of Statistical Software, 59(CS-3): ??, August 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v59/c03.

Schmid:2007:BAP

[SH07] Volker J. Schmid and Leonhard Held. Bayesian age-period-cohort modeling and prediction — BAMP. Journal of Statistical Software, 21(8):1–15, October 2007. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v21/i08.

Siddique:2009:MSM

[SH09] Juned Siddique and Ofer Harel. MIDAS: a SAS macro for multiple imputation using distance-aided selection of donors. Journal of Statistical Software, 29(9):??, February 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v29/i09.

Sankaran:2014:PSS

[SH14] Kris Sankaran and Susan Holmes. structSSI: Simultaneous and selective inference for grouped or hierarchically structured data. Journal of Statistical Software, 59(13):??, September 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v59/i13.

Saul:2017:RIS

[SH17] Bradley C. Saul and Michael G. Hudgens. A recipe for inferference: Start with causal inference. Add interference. Mix well with R. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v082i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v082i02/v82i02.pdf.

Saul:2020:CME

[SH20] Bradley C. Saul and Michael G. Hudgens. The calculus of m-estimation in R with geex. Journal of Statistical Software, 92(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v092i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v092i02/v92i02.pdf.

Sill:2014:PCE

[SHBZ14] Martin Sill, Thomas Hielscher, Natalia Becker, and Manuela Zucknick. c060: Extended inference with lasso and elastic-net regularized Cox and generalized linear models. Journal of Statistical Software, 62(5):??, December 2014. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v62/i05.

Sheng:2008:MPMa

[She08a] Yanyan Sheng. Markov chain Monte Carlo estimation of normal Ogive IRT models in MATLAB. Journal of Statistical Software, 25(8):1–15, April 2008. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v25/i08.

Sheng:2008:MPMb

[She08b] Yanyan Sheng. A MATLAB package for Markov chain Monte Carlo with a multi-unidimensional IRT model. Journal of Statistical Software, 28(10):??, November 2008. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v28/i10.

Sheng:2010:BEM

[She10] Yanyan Sheng. Bayesian estimation of MIRT models with general and specific latent traits in MATLAB. Journal of Statistical Software, 34(3):??, April 2010. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v34/i03.

Shentu:2011:BRB

[She11] Yue Shentu. Book review: Clinical Trial Data Analysis Using R. Journal of Statistical Software, 43(BR-1):??, July 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v43/b01.

Sieger:2017:IDR

[SHFB17] Tomás Sieger, Catherine B. Hurley, Karel Fiser, and Claudia Beleites. Interactive dendrograms: The R packages idendro and idendr0. Journal of Statistical Software, 76(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v076i10;   https://www.jstatsoft.org/
index.php/jss/article/view/v076i10/v76i10.pdf.

Shotwell:2013:PRP

[Sho13] Matthew S. Shotwell. profdpm: An R package for MAP estimation in a class of conjugate product partition models. Journal of Statistical Software, 53(8):??, May 2013. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v53/i08.

Struyf:1996:COO

[SHR96] Anja Struyf, Mia Hubert, and Peter Rousseeuw. Clustering in an object-oriented environment. Journal of Statistical Software, 1(4):1–30, ???? 1996. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v01/i04;   http://www.jstatsoft.org/v01/i04/code/;
 http://www.jstatsoft.org/v01/i04/data/;   http://
www.jstatsoft.org/v01/i04/paper/clus.pdf;   http://
www.jstatsoft.org/v01/i04/updates.

Simpson:2007:AMP

[Sim07] Gavin L. Simpson. Analogue methods in palaeoecology: Using the analogue package. Journal of Statistical Software, 22(2): 1–29, September 2007. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v22/i02.

Scharpf:2011:URP

[SIR^{+}11] Robert B. Scharpf, Rafael A. Irizarry, Matthew E. Ritchie, Benilton Carvalho, and Ingo Ruczinski. Using the R package crlmm for genotyping and copy number estimation. Journal of Statistical Software, 40(12):??, May 2011. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v40/i12.

Sigrist:2015:PRP

[SKS15] Fabio Sigrist, Hans R. Künsch, and Werner A. Stahel. spate: An R package for spatio-temporal modeling with a stochastic advection-diffusion process. Journal of Statistical Software, 63 (14):??, February 2015. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v63/i14.

Stoll:2005:WRS

[SKZ05] Heather Stoll, Gary King, and Langche Zeng. WhatIf: R software for evaluating counterfactuals. Journal of Statistical Software, 15(4):1–18, December 2005. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v15/i04.

Sabatti:2009:BRB

[SL09] Chiara Sabatti and Christophe Lalanne. Book review: Applied Statistical Genetics with R for Population-Based Association Studies. Journal of Statistical Software, 31(BR-2):??, September 2009. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v31/b02.

Simard:2011:CTS

[SL11] Richard Simard and Pierre L’Ecuyer. Computing the twosided Kolmogorov–Smirnov distribution. Journal of Statistical Software, 39(11):1–18, March 2011. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v39/i11.

Schweinberger:2018:HHE

[SL18] Michael Schweinberger and Pamela Luna. hergm: Hierarchical exponential-family random graph models. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v085i01;   https://www.jstatsoft.
org/index.php/jss/article/view/v085i01/v85i01.pdf.

Steele:2014:FSP

[SLF14] Guy L. Steele, Jr., Doug Lea, and Christine H. Flood. Fast splittable pseudorandom number generators. ACM SIGPLAN Notices, 49(10):453–472, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sturtz:2005:PRP

[SLG05] Sibylle Sturtz, Uwe Ligges, and Andrew Gelman. R2WinBUGS: a package for running WinBUGS from R. Journal of Statistical Software, 12(3):1–16, ???? 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=110&url=v12/i03&ct=2;   http://www.
jstatsoft.org/counter.php?id=110&url=v12/i03/v12i03.
pdf&ct=1.

Sinz:2014:PNP

[SLGB14] Fabian H. Sinz, Jörn-Philipp Lies, Sebastian Gerwinn, and Matthias Bethge. Natter: A Python natural image statistics toolbox. Journal of Statistical Software, 61(5):??, November 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v61/i05.

Steen:2017:MRP

[SLMV17] Johan Steen, Tom Loeys, Beatrijs Moerkerke, and Stijn Vansteelandt. medflex: An R package for flexible mediation analysis using natural effect models. Journal of Statistical Software, 76(??):??, ???? 2017. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v076i11;   https://www.jstatsoft.
org/index.php/jss/article/view/v076i11/v76i11.pdf.

Schoonees:2016:FGA

[SlRC16] Pieter Schoonees, Niël le Roux, and Roelof Coetzer. Flexible graphical assessment of experimental designs in R: The vdg package. Journal of Statistical Software, 74 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v074i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v074i03/v74i03.pdf.

Snellenburg:2012:GJB

[SLS^{+}12] Joris J. Snellenburg, Sergey P. Laptenok, Ralf Seger, Katharine M. Mullen, and Ivo H. M. van Stokkum. Glotaran: a Java-based graphical user interface for the R package TIMP. Journal of Statistical Software, 49(3):??, June 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v49/i03.

Stubben:2007:EAD

[SM07] Chris Stubben and Brook Milligan. Estimating and analyzing demographic models using the popbio package in R. Journal of Statistical Software, 22(11):1–23, September 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v22/i11.

Saito:2012:DCS

[SM12] Mutsuo Saito and Makoto Matsumoto. A deviation of CURAND: Standard pseudorandom number generator in CUDA for GPGPU. Slides presented at the Tenth International Conference on Monte Carlo and Quasi-Monte Carlo Methods in Scientific Computing, February 2012. URL

http://www.mcqmc2012.unsw.edu.au/slides/MCQMC2012_
Matsumoto.pdf.

Salmon:2011:PRN

[SMDS11] John K. Salmon, Mark A. Moraes, Ron O. Dror, and David E. Shaw. Parallel random numbers: as easy as 1, 2, 3. In Lathrop et al. [LCK11], pages 16:1–16:12. ISBN 1-4503-0771-X. LCCN ????

Schmidberger:2009:SAP

[SME^{+}09] Markus Schmidberger, Martin Morgan, Dirk Eddelbuettel, Hao Yu, Luke Tierney, and Ulrich Mansmann. State of the art in parallel computing with R. Journal of Statistical Software, 31(1):??, August 2009. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v31/i01.

Smith:2007:PBR

[Smi07] Brian J. Smith. boa: An R package for MCMC output convergence assessment and posterior inference. Journal of Statistical Software, 21(11):1–37, November 2007. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v21/i11.

Schlather:2015:ASP

[SMM^{+}15] Martin Schlather, Alexander Malinowski, Peter J. Menck, Marco Oesting, and Kirstin Strokorb. Analysis, simulation and prediction of multivariate random fields with package RandomFields. Journal of Statistical Software, 63(8):??, February 2015. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v63/i08.

Sweeting:2013:BBC

[SMS13] Michael Sweeting, Adrian Mander, and Tony Sabin. bcrm: Bayesian continual reassessment method designs for Phase I dose-finding trials. Journal of Statistical Software, 54(13):??, September 2013. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v54/i13.

Soetaert:2010:BRB

[Soe10] Karline Soetaert. Book review: A Primer of Ecology with R. Journal of Statistical Software, 32(BR-3):??, January 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v32/b03.

Solymos:2009:PED

[Sól09] Péter Sólymos. Processing ecological data in R with the mefa package. Journal of Statistical Software, 29(8):??, February 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v29/i08.

Somerville:1998:FPE

[Som98] Paul N. Somerville. A Fortran 90 program for evaluation of multivariate normal and multivariate t integrals over convex regions. Journal of Statistical Software, 3(4):1–10, ???? 1998. CODEN JSSOBK. ISSN

1548-7660.  URL http://www.jstatsoft.org/v03/i04;
 http://www.jstatsoft.org/v03/i04/mvi3.for;   http:
//www.jstatsoft.org/v03/i04/somerville.pdf;   http:
//www.jstatsoft.org/v03/i04/somerville.ps;   http:
//www.jstatsoft.org/v03/i04/updates;   http://www.
jstatsoft.org/v03/i04/winmvi3.exe.

Somerville:2001:NCM

[Som01] Paul N. Somerville. Numerical computation of multivariate normal and multivariate t probabilities over ellipsoidal regions. Journal of Statistical Software, 6(8): 1–10, ???? 2001. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v06/i08;   http:
//www.jstatsoft.org/v06/i08/EXAMPLE/;   http://www.
jstatsoft.org/v06/i08/FORTRAN/;   http://www.jstatsoft.
org/v06/i08/jss2r.pdf;   http://www.jstatsoft.org/
v06/i08/updates.

Somerville:2007:CCV

[Som07] Paul N. Somerville. Calculation of critical values for Somerville’s FDR procedures. Journal of Statistical Software, 21(6):1–14, October 2007. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v21/i06.

Somerville:2012:PSP

[Som12] Paul N. Somerville. A package to study the performance of step-up and step-down test procedures. Journal of Statistical Software, 51(6):??, November 2012. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v51/i06.

Smith:2005:JES

[SP05] Christina D. Smith and Jeffrey S. Pontius. Jackknife estimator of species richness with S-PLUS. Journal of Statistical Software, 15(3):1–12, December 2005. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v15/i03.

Soetaert:2010:IMS

[SP10] Karline Soetaert and Thomas Petzoldt. Inverse modelling, sensitivity and Monte Carlo analysis in R using package FME. Journal of Statistical Software, 33(3):??, February 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v33/i03.

Savitsky:2014:BSN

[SP14] Terrance D. Savitsky and Susan M. Paddock. Bayesian semiand non-parametric models for longitudinal data with multiple membership effects in R. Journal of Statistical Software, 57 (3):??, March 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v57/i03.

Sparapani:2006:BRB

[Spa06] Rodney Sparapani. Book review: Statistical Analysis of Medical Data Using SAS. Journal of Statistical Software, 16(BR-5):1–5, July 2006. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v16/b05.

Spedicato:2013:LPP

[Spe13] Giorgio Alfredo Spedicato. The lifecontingencies package: Performing financial and actuarial mathematics calculations in R. Journal of Statistical Software, 55(10):??, November 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v55/i10.

Soetaert:2010:SDE

[SPS10] Karline Soetaert, Thomas Petzoldt, and R. Woodrow Setzer. Solving differential equations in R: Package deSolve. Journal of Statistical Software, 33(9):??, February 2010. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v33/i09.

Smith:2007:WSM

[SR07a] Michael K. Smith and Helen Richardson. WinBUGSio: a SAS macro for the remote execution of WinBUGS. Journal of Statistical Software, 23(9):1–10, December 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v23/i09.

Stasinopoulos:2007:GAM

[SR07b] D. Mikis Stasinopoulos and Robert A. Rigby. Generalized additive models for location scale and shape (GAMLSS) in R. Journal of Statistical Software, 23(7):1–46, December 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v23/i07.

Sevcikova:2016:BPP

[SR16] Hana Sevcíková and Adrian E. Raftery. bayesPop: Probabilistic population projections. Journal of Statistical Software, 75(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v075i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v075i05/v75i05.pdf.

Scrucca:2018:CPI

[SR18] Luca Scrucca and Adrian E. Raftery. clustvarsel: A package implementing variable selection for Gaussian modelbased clustering in R. Journal of Statistical Software, 84 (??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v084i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v084i01/v84i01.pdf.

Shou:2019:CRP

[SS19] Yiyun Shou and Michael Smithson. cdfquantreg: An R package for CDF-quantile regression. Journal of Statistical Software, 88(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v088i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v088i01/v88i01.pdf.

Salmon:2016:MCT

[SSH16] Maëlle Salmon, Dirk Schumacher, and Michael Höhle. Monitoring count time series in R: Aberration detection in public health surveillance. Journal of Statistical Software, 70 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v070i10;   https://www.jstatsoft.org/
index.php/jss/article/view/v070i10/v70i10.pdf.

Sparapani:2021:NML

[SSM21] Rodney Sparapani, Charles Spanbauer, and Robert Mc-Culloch. Nonparametric machine learning and efficient computation with Bayesian additive regression trees: The BART R package. Journal of Statistical Software, 97 (??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v097i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v097i01/v97i01.pdf.

Szocs:2020:WRP

[SSS^{+}20] Eduard Szöcs, Tamás Stirling, Eric R. Scott, Andreas Scharmüller, and Ralf B. Schäfer. webchem: an R package to retrieve chemical information from the Web. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v093i13;   https://www.jstatsoft.
org/index.php/jss/article/view/v093i13/v93i13.pdf.

Subirana:2014:BBT

[SSV14] Isaac Subirana, Héctor Sanz, and Joan Vila. Building bivariate tables: The compareGroups package for R. Journal of Statistical Software, 57(12):??, May 2014. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v57/i12.

Stokes:2012:SRB

[Sto12] Barrie Stokes. Software review: mathStatica 2.5. Journal of Statistical Software, 47(SR-1):??, April 2012. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v47/s01.

Stromberg:2004:WWS

[Str04] Arnold J. Stromberg. Why write statistical software? the case of robust statistical methods. Journal of Statistical Software, 10(5):1–8, ???? 2004. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/counter.php?id=
90&url=v10/i05/finalintro.pdf&ct=1.

Strobl:2010:BRB

[Str10] Carolin Strobl. Book review: Advances in Social Science Research Using R. Journal of Statistical Software, 34(BR-2):??, April 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v34/b02.

Su:2007:FSM

[Su07] Steve Su. Fitting single and mixture of generalized lambda distributions to data via discretized and maximum likelihood methods: GLDEX in R. Journal of Statistical Software, 21(9): 1–17, October 2007. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v21/i09.

Sofrygin:2017:SRP

[SvdLN17] Oleg Sofrygin, Mark J. van der Laan, and Romain Neugebauer. simcausal R package: Conducting transparent and reproducible simulation studies of causal effect estimation with complex longitudinal data. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v081i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v081i02/v81i02.pdf.

Sestelo:2017:NRP

[SVMMRP17] Marta Sestelo, Nora M. Villanueva, Luis Meira-Machado, and Javier Roca-Pardiñas. npregfast: An R package for nonparametric estimation and inference in life sciences. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/

index.php/jss/article/view/v082i12;   https://www.
jstatsoft.org/index.php/jss/article/view/v082i12/v82i12.
pdf.

Sales:2015:CSM

[SWAF15] Célia Maria Dias Sales, Peter P. Wakker, Paula Cristina Gomes Alves, and Luís Faísca. Code snippet: MF calculator: A Webbased application for analyzing similarity. Journal of Statistical Software, 65(CS-2):??, June 2015. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v65/c02.

Smith:2008:UGM

[SYC08] Brian J. Smith, Jun Yan, and Mary Kathryn Cowles. Unified geostatistical modeling for data fusion and spatial heteroskedasticity with R package ramps. Journal of Statistical Software, 25(10):1–21, April 2008. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v25/i10.

Scheike:2011:ACR

[SZ11] Thomas H. Scheike and Mei-Jie Zhang. Analyzing competing risk data using the R timereg package. Journal of Statistical Software, 38(2):??, January 2011. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v38/i02.

Takane:2016:MEI

[Tak16] Yoshio Takane. My early interactions with Jan and some of his lost papers. Journal of Statistical Software, 73 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v073i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v073i07/v73i07.pdf.

Taylor:2017:RPA

[TB17] Julian Taylor and David Butler. R package ASMap: Efficient genetic linkage map construction and diagnosis. Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v079i06;   https://www.
jstatsoft.org/index.php/jss/article/view/v079i06/v79i06.
pdf.

Tortora:2021:MBC

[TBE^{+}21] Cristina Tortora, Ryan P. Browne, Aisha ElSherbiny, Brian C. Franczak, and Paul D. McNicholas. Model-based clustering, classification, and discriminant analysis using the generalized hyperbolic distribution: MixGHD R package. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v098i03;   https://www.jstatsoft.
org/index.php/jss/article/view/v098i03/v98i03.pdf.

Turner:2014:OAP

[TBS14] Rolf Turner, Pradeep Banerjee, and Rayomand Shahlori. Optimal asset pricing. Journal of Statistical Software, 58(11):??, July 2014. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v58/i11.

Thioulouse:2007:IMD

[TD07] Jean Thioulouse and Stéphane Dray. Interactive multivariate data analysis in R with the ade4 and ade4TkGUI packages. Journal of Statistical Software, 22(5):1–14, September 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v22/i05.

Taylor:2013:LRP

[TDRD13] Benjamin M. Taylor, Tilman M. Davies, Barry S. Rowlingson, and Peter J. Diggle. lgcp: An R package for inference with spatial and spatio-temporal log-gaussian Cox processes. Journal of Statistical Software, 52(4):??, February 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v52/i04.

Taylor:2015:BID

[TDRD15] Benjamin M. Taylor, Tilman M. Davies, Barry S. Rowlingson, and Peter J. Diggle. Bayesian inference and data augmentation schemes for spatial, spatiotemporal and multivariate log-Gaussian Cox processes in R. Journal of Statistical Software, 63(7):??, February 2015. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v63/i07.

Tennekes:2018:TTM

[Ten18] Martijn Tennekes. tmap: Thematic maps in R. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v084i06;   https://www.
jstatsoft.org/index.php/jss/article/view/v084i06/v84i06.
pdf.

Todorov:2009:OOF

[TF09] Valentin Todorov and Peter Filzmoser. An object-oriented framework for robust multivariate analysis. Journal of Statistical Software, 32(3):??, October 2009. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v32/i03.

Turner:2012:BTM

[TF12] Heather Turner and David Firth. Bradley–Terry models in R: The BradleyTerry2 package. Journal of Statistical Software, 48(9):??, May 2012. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v48/i09.

Theussl:2012:TPD

[TFH12] Stefan Theußl, Ingo Feinerer, and Kurt Hornik. A tm plug-in for distributed text mining in R. Journal of Statistical Software, 51(5):??, November 2012. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v51/i05.

Thieler:2016:RRP

[TFR16] Anita M. Thieler, Roland Fried, and Jonathan Rathjens. RobPer: An R package to calculate periodograms for light curves based on robust regression. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v069i09;   https://www.jstatsoft.org/
index.php/jss/article/view/v069i09/v69i09.pdf.

Touraine:2017:SRP

[TGJ17] Célia Touraine, Thomas A. Gerds, and Pierre Joly. Smooth-Hazard: An R package for fitting regression models to intervalcensored observations of illness-death models. Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v079i07;   https://www.
jstatsoft.org/index.php/jss/article/view/v079i07/v79i07.
pdf.

Toomet:2008:SSM

[TH08] Ott Toomet and Arne Henningsen. Sample selection models in R: Package sampleSelection. Journal of Statistical Software, 27(7):??, July 2008. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v27/i07.

Thiele:2021:CIE

[TH21] Christian Thiele and Gerrit Hirschfeld. cutpointr: Improved estimation and validation of optimal cutpoints in R. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v098i11;   https://www.
jstatsoft.org/index.php/jss/article/view/v098i11/v98i11.
pdf.

Theus:2002:IDV

[The02] Martin Theus. Interactive data visualization using Mondrian. Journal of Statistical Software, 7(11):1–9, ???? 2002. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v07/i11;   http://www.jstatsoft.
org/v07/i11/MondrianJSSV2.pdf.

Thiele:2014:RMN

[Thi14] Jan C. Thiele. R marries NetLogo: Introduction to the RNetLogo package. Journal of Statistical Software, 58(2):??, June 2014. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v58/i02.

Thompson:2011:IS

[Tho11] Madeleine B. Thompson. Introduction to SamplerCompare. Journal of Statistical Software, 43(12):??, August 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v43/i12.

Tierney:2005:SNP

[Tie05] Luke Tierney. Some notes on the past and future of Lisp– Stat. Journal of Statistical Software, 13(9):1–15, ???? 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/counter.php?id=123&url=v13/i09&ct=2;
 http://www.jstatsoft.org/counter.php?id=123&url=v13/
i09/v13i09.pdf&ct=1.

Tivadar:2019:ORP

[Tiv19] Mihai Tivadar. OasisR: an R package to bring some order to the world of segregation measurement. Journal of Statistical Software, 89(??):??, ???? 2019. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v089i07;   https://www.jstatsoft.
org/index.php/jss/article/view/v089i07/v89i07.pdf.

Tikka:2017:ICE

[TK17] Santtu Tikka and Juha Karvanen. Identifying causal effects with the R package causaleffect. Journal of Statistical Software, 76(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v076i12;   https://www.jstatsoft.org/
index.php/jss/article/view/v076i12/v76i12.pdf.

Tanimura:2006:PSM

[TKM06] Susumu Tanimura, Chusi Kuroiwa, and Tsutomu Mizota. Proportional symbol mapping in R. Journal of Statistical Software, 15(5):1–7, January 2006. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v15/i05.

Tanimura:2007:CSA

[TKM07] Susumu Tanimura, Chusi Kuroiwa, and Tsutomu Mizota. Code snippet: Auxiliary cartographic functions in R: North arrow, scale bar, and label with leader arrow. Journal of Statistical Software, 19(CS-1):1–8, April 2007. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v19/c01.

Templ:2015:SDC

[TKM15] Matthias Templ, Alexander Kowarik, and Bernhard Meindl. Statistical disclosure control for micro-data using the R package sdcMicro. Journal of Statistical Software, 67(4):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v067i04;
http://www.jstatsoft.org/index.php/jss/article/view/
v067i04/v67i04.pdf.

Tak:2017:RRP

[TKM17] Hyungsuk Tak, Joseph Kelly, and Carl Morris. Rgbp: An R package for Gaussian, Poisson, and binomial random effects models with frequency coverage evaluations. Journal of Statistical Software, 78(??):??, ???? 2017. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v078i05;   https://www.jstatsoft.
org/index.php/jss/article/view/v078i05/v78i05.pdf.

Terpstra:2005:RBA

[TM05] Jeff T. Terpstra and Joseph W. McKean. Rank-based analysis of linear models using R. Journal of Statistical Software, 14(7):1–26, July 1, 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=132&url=v14/i07&ct=2;   http://www.
jstatsoft.org/counter.php?id=132&url=v14/i07/v14i07.
pdf&ct=1.

Tang:2015:DRP

[TM15] Xinyu Tang and Maria Melguizo. DTR: An R package for estimation and comparison of survival outcomes of dynamic treatment. Journal of Statistical Software, 65(7):??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v65/i07.

Templ:2017:SSC

[TMKD17] Matthias Templ, Bernhard Meindl, Alexander Kowarik, and Olivier Dupriez. Simulation of synthetic complex data: The R package simPop. Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v079i10;   https://www.jstatsoft.org/
index.php/jss/article/view/v079i10/v79i10.pdf.

Tendeiro:2016:PRP

[TMN16] Jorge N. Tendeiro, Rob R. Meijer, and A. Susan M. Niessen. PerFit: An R package for person-fit analysis in IRT. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v074i05;   https://www.
jstatsoft.org/index.php/jss/article/view/v074i05/v74i05.
pdf.

Tarr:2018:MRP

[TMW18] Garth Tarr, Samuel Müller, and Alan H. Welsh. mplot: An R package for graphical model stability and variable selection procedures. Journal of Statistical Software, 83 (??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v083i09;   https://www.jstatsoft.org/
index.php/jss/article/view/v083i09/v83i09.pdf.

Touloumis:2015:RPP

[Tou15] Anestis Touloumis. R package multgee: A generalized estimating equations solver for multinomial responses. Journal of Statistical Software, 64(8):??, March 2015. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v64/i08.

Tabelow:2011:SPM

[TP11] Karsten Tabelow and Jörg Polzehl. Statistical parametric maps for functional MRI experiments in R: The package fmri. Journal of Statistical Software, 44(11):??, October 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v44/i11.

Taylor:2019:MLS

[TPE19] Simon A. C. Taylor, Timothy Park, and Idris A. Eckley. Multivariate locally stationary wavelet analysis with the mvLSW R package. Journal of Statistical Software, 90 (??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v090i11;   https://www.jstatsoft.org/
index.php/jss/article/view/v090i11/v90i11.pdf.

Thomas:2014:CST

[TR14] Laine Thomas and Eric M. Reyes. Code snippet: Tutorial: Survival estimation for Cox regression models with timevarying coefficients using SAS and R. Journal of Statistical Software, 61(CS-1):??, October 2014. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v61/c01.

Taylor:2017:SRP

[TR17] Benjamin M. Taylor and Barry S. Rowlingson. spatsurv: An R package for Bayesian inference with spatial survival models. Journal of Statistical Software, 77(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v077i04;   https://
www.jstatsoft.org/index.php/jss/article/view/v077i04/
v77i04.pdf.

Trindade:2003:IMB

[Tri03] A. Alexandre Trindade. Implementing modifed Burg algorithms in multivariate subset autoregressive modeling. Journal of Statistical Software, 8(5):1–68, ???? 2003. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v08/i05/pap1r2.pdf;   http://www.jstatsoft.org/
v08/i5.

Theussl:2020:RER

[TSH20] Stefan Theußl, Florian Schwendinger, and Kurt Hornik. ROI: an extensible R optimization infrastructure. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v094i15;   https://www.jstatsoft.
org/index.php/jss/article/view/v094i15/v94i15.pdf.

Tanaka:2021:NRP

[TSN21] Ushio Tanaka, Masami Saga, and Junji Nakano. NScluster: An R package for maximum palm likelihood estimation for cluster point process models using OpenMP. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v098i06;   https://www.jstatsoft.
org/index.php/jss/article/view/v098i06/v98i06.pdf.

Traets:2020:GOD

[TSV20] Frits Traets, Daniel Gil Sanchez, and Martina Vandebroek. Generating optimal designs for discrete choice experiments in R: The idefix package. Journal of Statistical Software, 96(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v096i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v096i03/v96i03.pdf.

Tusell:2011:KFR

[Tus11] Fernando Tusell. Kalman filtering in R. Journal of Statistical Software, 39(2):??, March 2011. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v39/i02.

Taylor:2011:RPW

[TV11] Julian Taylor and Arunas Verbyla. R package wgaim: QTL analysis in bi-parental populations using linear mixed models.

Journal of Statistical Software, 40(7):??, April 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v40/i07.

Tabelow:2011:SVM

[TW11] Karsten Tabelow and Brandon Whitcher. Special volume on magnetic resonance imaging in R. Journal of Statistical Software, 44(1):??, October 2011. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v44/i01.

Tomz:2003:PCS

[TWK03] Michael Tomz, Jason Wittenberg, and Gary King. Clarify: Software for interpreting and presenting statistical results. Journal of Statistical Software, 8(1):1–30, ???? 2003. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v08/i01/clarify.pdf;   http://
www.jstatsoft.org/v08/i1.

Tingley:2014:PMR

[TYH^{+}14] Dustin Tingley, Teppei Yamamoto, Kentaro Hirose, Luke Keele, and Kosuke Imai. mediation: R package for causal mediation analysis. Journal of Statistical Software, 59(5):??, September 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v59/i05.

Umlauf:2015:SAR

[UAK^{+}15] Nikolaus Umlauf, Daniel Adler, Thomas Kneib, Stefan Lang, and Achim Zeileis. Structured additive regression models: An R interface to BayesX. Journal of Statistical Software, 63(21): ??, February 2015. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v63/i21.

Udina:2000:IIC

[Udi00] Frederic Udina. Implementing interactive computing in an object-oriented environment. Journal of Statistical Software, 5(3):1–20, ???? 2000. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v05/
i03;   http://www.jstatsoft.org/v05/i03/lisp-code/;
 http://www.jstatsoft.org/v05/i03/object-computing.
pdf;   http://www.jstatsoft.org/v05/i03/updates.

Udina:2005:IBC

[Udi05] Frederic Udina. Interactive biplot construction. Journal of Statistical Software, 13(5):1–16, ???? 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=119&url=v13/i05&ct=2;   http://www.
jstatsoft.org/counter.php?id=119&url=v13/i05/v13i05.
pdf&ct=1.

Unlu:2009:FSR

[ÜKD09] Ali Ünlü, Thomas Kiefer, and Ehtibar N. Dzhafarov. Fechnerian scaling in R: The package fechner. Journal of Statistical Software, 31(6):??, August 2009. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v31/i06.

Umbach:2017:FNS

[UNBK17] Nora Umbach, Katharina Naumann, Holger Brandt, and Augustin Kelava. Fitting nonlinear structural equation models in R with package nlsem. Journal of Statistical Software, 77(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v077i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v077i07/v77i07.pdf.

Unwin:2009:BRB

[Unw09] Antony Unwin. Book review: Bioconductor Case Studies. Journal of Statistical Software, 29(BR-1):??, January 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v29/b01.

Unwin:2012:OI

[Unw12] Antony Unwin. Oscars and interfaces. Journal of Statistical Software, 49(11):??, June 2012. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v49/i11.

Unlu:2010:PDR

[ÜS10] Ali Ünlü and Anatol Sargin. DAKS: An R package for data analysis methods in knowledge space theory. Journal of Statistical Software, 37(2):??, November 2010. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v37/i02.

Ucar:2019:SDE

[USA19] Iñaki Ucar, Bart Smeets, and Arturo Azcorra. simmer: Discrete-event simulation for R. Journal of Statistical Software, 90(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v090i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v090i02/v90i02.pdf.

Uschner:2018:RRP

[USHH18] Diane Uschner, David Schindler, Ralf-Dieter Hilgers, and Nicole Heussen. randomizeR: An R package for the assessment and implementation of randomization in clinical trials. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v085i08;   https://www.
jstatsoft.org/index.php/jss/article/view/v085i08/v85i08.
pdf.

VandenBossche:2011:FSS

[Van11] Filip A. M. Van den Bossche. Fitting state space models with EViews. Journal of Statistical Software, 41(8):??, May 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v41/i08.

Varadhan:2014:NOR

[Var14] Ravi Varadhan. Numerical optimization in R: Beyond optim. Journal of Statistical Software, 60(1):??, September 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v60/i01.

vanBuuren:2011:MMI

[vBGO11] Stef van Buuren and Karin Groothuis-Oudshoorn. mice: Multivariate Imputation by Chained Equations in R. Journal of Statistical Software, 45(3):??, December 2011. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v45/i03.

vanderArk:2007:MSA

[vdA07] L. Andries van der Ark. Mokken scale analysis in R. Journal of Statistical Software, 20(11):1–19, February 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v20/i11.

vanderArk:2012:NDM

[vdA12] L. Andries van der Ark. New developments in Mokken scale analysis in R. Journal of Statistical Software, 48(5):??, May 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v48/i05.

vanderHeijden:2016:LBG

[vdHvB16] Peter G. M. van der Heijden and Stef van Buuren. Looking back at the Gifi system of nonlinear multivariate analysis. Journal of Statistical Software, 73(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v073i04;   https://
www.jstatsoft.org/index.php/jss/article/view/v073i04/
v73i04.pdf.

Vinod:2009:MEB

[VdL09] Hrishikesh D. Vinod and Javier López de Lacalle. Maximum entropy bootstrap for time series: The meboot R package. Journal of Statistical Software, 29(5):??, January 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v29/i05.

vanderLoo:2021:MDR

[vdL21] Mark P. J. van der Loo. Monitoring data in R with the lumberjack package. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v098i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v098i01/v98i01.pdf.

vanderLoo:2021:DVI

[vdLdJ21] Mark P. J. van der Loo and Edwin de Jonge. Data validation infrastructure for R. Journal of Statistical Software, 97(??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v097i10;   https://www.jstatsoft.org/
index.php/jss/article/view/v097i10/v97i10.pdf.

Vock:2014:CSS

[VDT14] David Vock, Marie Davidian, and Anastasios Tsiatis. Code snippet: SNP NLMM: A SAS macro to implement a flexible random effects density for generalized linear and nonlinear mixed models. Journal of Statistical Software, 56(CS-2):

??, January 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v56/c02.

vanderWal:2011:IRP

[vdWG11] Willem M. van der Wal and Ronald B. Geskus. ipw: An R package for inverse probability weighting. Journal of Statistical Software, 43(13):??, September 2011. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v43/i13.

vanEtten:2017:RPG

[vE17] Jacob van Etten. R package gdistance: Distances and routes on geographical grids. Journal of Statistical Software, 76(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v076i13;   https://www.jstatsoft.org/
index.php/jss/article/view/v076i13/v76i13.pdf.

Verzani:2012:GCI

[Ver12] John Verzani. gWidgetsWWW: Creating interactive Web pages within R. Journal of Statistical Software, 49(10):??, June 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v49/i10.

Verde:2018:BRP

[Ver18] Pablo Emilio Verde. bamdit: An R package for Bayesian meta-analysis of diagnostic test data. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v086i10;   https://www.jstatsoft.
org/index.php/jss/article/view/v086i10/v86i10.pdf.

Varmaz:2013:RAW

[VFV13] Armin Varmaz, Christian Fieberg, and Andreas Varwig. RMatlab-app2web: Web deployment of R/MATLAB applications. Journal of Statistical Software, 54(5):??, September 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v54/i05.

Varadhan:2009:BRP

[VG09] Ravi Varadhan and Paul Gilbert. BB: An R package for solving a large system of nonlinear equations and for optimizing a highdimensional nonlinear objective function. Journal of Statistical Software, 32(4):??, October 2009. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v32/i04.

Voronca:2018:MSM

[VGDM^{+}18] Delia C. Voronca, Mulugeta Gebregziabher, Valerie Durkalski-Mauldin, Lei Liu, and Leonard E. Egede. MTPmle: A SAS macro and Stata programs for marginalized inference in semi-continuous data. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v087i06;   https://www.jstatsoft.org/
index.php/jss/article/view/v087i06/v87i06.pdf.

Verhelst:2007:RS

[VHM07] Norman D. Verhelst, Reinhold Hatzinger, and Patrick Mair. The Rasch sampler. Journal of Statistical Software, 20(4):1– 14, February 2007. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v20/i04.

Viechtbauer:2010:CMA

[Vie10] Wolfgang Viechtbauer. Conducting meta-analyses in R with the metafor package. Journal of Statistical Software, 36(3): ??, August 2010. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v36/i03.

Vigna:2016:EEM

[Vig16] Sebastiano Vigna. An experimental exploration of Marsaglia’s xorshift generators, scrambled. ACM Transactions on Mathematical Software, 42(4):30:1–30:23, July 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://dl.acm.org/citation.cfm?id=2845077.

Vinue:2017:ARP

[Vin17] Guillermo Vinué. Anthropometry: An R package for analysis of anthropometric data. Journal of Statistical Software, 77(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v077i06;   https://www.jstatsoft.org/
index.php/jss/article/view/v077i06/v77i06.pdf.

Villegas:2018:SRP

[VKM18] Andrés M. Villegas, Vladimir K. Kaishev, and Pietro Millossovich. StMoMo: An R package for stochastic mortality modeling. Journal of Statistical Software, 84 (??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v084i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v084i03/v84i03.pdf.

Vervloet:2015:PRP

[VKVC15] Marlies Vervloet, Henk A. L. Kiers, Wim Van den Noortgate, and Eva Ceulemans. PCovR: An R package for principal covariates regression. Journal of Statistical Software, 65(8):??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v65/i08.

Valero-Mora:2009:BRBb

[VM09a] Pedro M. Valero-Mora. Book review: Interactive and Dynamic Graphics for Data Analysis: With R and GGobi. Journal of Statistical Software, 30(BR-7):??, May 2009. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v30/b07.

Valero-Mora:2009:BRBa

[VM09b] Pedro M. Valero-Mora. Book review: Introductory Statistics with R (Second Edition). Journal of Statistical Software, 29 (BR-10):??, January 2009. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v29/b10.

Valero-Mora:2010:BRB

[VM10] Pedro M. Valero-Mora. Book review: ggplot2: Elegant Graphics for Data Analysis. Journal of Statistical Software, 35(BR-1):??, July 2010. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v35/b01.

Valero-Mora:2012:GUI

[VML12] Pedro M. Valero-Mora and Rubén D. Ledesma. Graphical user interfaces for R. Journal of Statistical Software, 49(1):??, June 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v49/i01.

Valero-Mora:2005:HLS

[VMU05] Pedro Valero-Mora and Frederic Udina. The health of Lisp– Stat. Journal of Statistical Software, 13(10):1–5, ???? 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/counter.php?id=124&url=v13/i10&ct=2;

 http://www.jstatsoft.org/counter.php?id=124&url=v13/
i10/v13i10.pdf&ct=1.

Valero-Mora:2002:CVL

[VMY02] Pedro M. Valero-Mora and Forrest W. Young. Computing and visualizing log-linear analysis interactively. Journal of Statistical Software, 7(6):1–9, ???? 2002. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v07/i06;
http://www.jstatsoft.org/v07/i06/lglin.zip;   http:
//www.jstatsoft.org/v07/i06/updates;   http://www.
jstatsoft.org/v07/i06/vislog6.pdf.

Villegas:2018:STS

[VP18] Marco A. Villegas and Diego J. Pedregal. SSpace: A toolbox for state space modeling. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v087i05;   https://www.jstatsoft.org/
index.php/jss/article/view/v087i05/v87i05.pdf.

VerHoef:2014:PSR

[VPCS14] Jay M. Ver Hoef, Erin E. Peterson, David Clifford, and Rohan Shah. SSN: An R package for spatial statistical modeling on stream networks. Journal of Statistical Software, 56(3): ??, January 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v56/i03.

Villa:2014:LCT

[VR14] Simone Villa and Marco Rossetti. Learning continuous time Bayesian network classifiers using MapReduce. Journal of Statistical Software, 62(3):??, December 2014. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v62/i03.

Visser:2010:DRP

[VS10] Ingmar Visser and Maarten Speekenbrink. depmixS4: An R package for hidden Markov models. Journal of Statistical Software, 36(7):??, August 2010. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v36/i07.

Verma:2014:DCC

[VS14] Mahendra P. Verma and Mario Cesar Suarez A. DixonTest.CriticalValues: A computer code to calculate critical values for the Dixon statistical data treatment approach. Journal of Statistical Software, 57(2):??, March 2014. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v57/i02.

VandenMeersche:2009:CSF

[VSV09] Karel Van den Meersche, Karline Soetaert, and Dick Van Oevelen. Code snippet: xsample(): An R function for sampling linear inverse problems. Journal of Statistical Software, 30(CS-1):??, April 2009. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v30/c01.

Villacorta:2016:FRP

[VV16] Pablo J. Villacorta and José L. Verdegay. FuzzyStatProb: An R package for the estimation of fuzzy stationary probabilities from a sequence of observations of an unknown Markov chain. Journal of Statistical Software, 71 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v071i08;   https://www.jstatsoft.org/
index.php/jss/article/view/v071i08/v71i08.pdf.

VanAelst:2013:FRB

[VW13] Stefan Van Aelst and Gert Willems. Fast and robust bootstrap for multivariate inference: The R package FRB. Journal of Statistical Software, 53(3):??, April 2013. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v53/i03.

Visne:2012:SRP

[VYD^{+}12] Ilhami Visne, Ahmet Yildiz, Erkan Dilaveroglu, Klemens Vierlinger, Christa Nöhammer, Friedrich Leisch, and Albert Kriegner. speedR: An R package for interactive data import, filtering and ready-to-use code generation. Journal of Statistical Software, 51(2):??, October 2012. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v51/i02.

Wang:2011:SRP

[Wan11] Ji-Ping Wang. SPECIES: An R package for species richness estimation. Journal of Statistical Software, 40(9):??, April 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v40/i09.

Wang:2013:COR

[Wan13a] Zhu Wang. Converting odds ratio to relative risk in cohort studies with partial data information. Journal of Statistical Software, 55(5):??, October 2013. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v55/i05.

Wang:2013:CRP

[Wan13b] Zhu Wang. cts: An R package for continuous time autoregressive models via Kalman filter. Journal of Statistical Software, 53(5):??, April 2013. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v53/i05.

Warnes:2002:HJL

[War02] Gregory R. Warnes. HYDRA: a Java library for Markov chain Monte Carlo. Journal of Statistical Software, 7 (4):1–32, ???? 2002. CODEN JSSOBK. ISSN 1548-

7660. URL http://software.biostat.washington.edu/
statsoft/MCMC/Hydra;   http://software.biostat.washington.
edu/statsoft/MCMC/Hydra/INSTALL.html;   http://www.
jstatsoft.org/v07/i04;   http://www.jstatsoft.org/
v07/i04/Description.pdf;   http://www.jstatsoft.org/
v07/i04/Hydra_1.0.0.jar;   http://www.jstatsoft.org/
v07/i04/Hydra_1.0.0.tgz;   http://www.jstatsoft.org/
v07/i04/updates;   http://www.jstatsoft.org/v07/i04/
UserGuide.pdf.

Wason:2015:ORP

[Was15] James Wason. OptGS: An R package for finding nearoptimal group-sequential designs. Journal of Statistical Software, 66(2):??, ???? 2015. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/index.php/jss/
article/view/v066i02;   http://www.jstatsoft.org/index.
php/jss/article/view/v066i02/v66i02.pdf.

Wehrens:2007:SSO

[WB07] Ron Wehrens and Lutgarde M. C. Buydens. Self- and superorganizing maps in R: The kohonen package. Journal of Statistical Software, 21(5):1–19, October 2007. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v21/i05.

Wheeler:2008:MAS

[WB08] Matthew W. Wheeler and A. John Bailer. Model averaging software for dichotomous dose response risk estimation. Journal of Statistical Software, 26(5):1–15, June 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v26/i05.

Widgren:2019:SRP

[WBEE19] Stefan Widgren, Pavol Bauer, Robin Eriksson, and Stefan Engblom. SimInf: an R package for data-driven stochastic disease spread simulations. Journal of Statistical Software, 91(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v091i12;   https://www.jstatsoft.org/
index.php/jss/article/view/v091i12/v91i12.pdf.

Wickham:2011:TRP

[WCHB11] Hadley Wickham, Dianne Cook, Heike Hofmann, and Andreas Buja. tourr: An R package for exploring multivariate data with projections. Journal of Statistical Software, 40(2):??, April 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v40/i02.

Wang:2020:IRP

[WCLS20] Chenguang Wang, Elizabeth Colantuoni, Andrew Leroux, and Daniel O. Scharfstein. idem: an R package for inferences in clinical trials with death and missingness. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v093i12;   https://www.jstatsoft.
org/index.php/jss/article/view/v093i12/v93i12.pdf.

Welvaert:2011:NRP

[WDM^{+}11] Marijke Welvaert, Joke Durnez, Beatrijs Moerkerke, Geert Verdoolaege, and Yves Rosseel. neuRosim: An R package for generating fMRI data. Journal of Statistical Software, 44(10): ??, October 2011. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v44/i10.

Wallace:2012:CGB

[WDT^{+}12] Byron C. Wallace, Issa J. Dahabreh, Thomas A. Trikalinos, Joseph Lau, Paul Trow, and Christopher H. Schmid. Closing the gap between methodologists and end-users: R as a computational back-end. Journal of Statistical Software, 49(5):??, June 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v49/i05.

Weeda:2011:AIF

[WdVW^{+}11] Wouter D. Weeda, Frank de Vos, Lourens J. Waldorp, Raoul P. P. P. Grasman, and Hilde M. Huizenga. arf3DS4: An integrated framework for localization and connectivity analysis of fMRI data. Journal of Statistical Software, 44(14):??, October 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v44/i14.

Webb:2016:BRE

[Web16] Paul Webb. Book review: Excel 2013 for Social Sciences Statistics: A Guide to Solving Practical Problems, Excel for Statistics. Journal of Statistical Software, 69 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v069b05;   https://www.jstatsoft.org/
index.php/jss/article/view/v069b05/v69b05.pdf.

Weeks:2010:PRP

[Wee10] Jonathan P. Weeks. plink: An R package for linking mixedformat tests using IRT-based methods. Journal of Statistical Software, 35(12):??, July 2010. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v35/i12.

Williams:2011:CSB

[WEH11] Michael S. Williams, Eric D. Ebel, and Jennifer A. Hoeting. Code snippet: Bayesian analysis for food-safety risk assessment: Evaluation of dose-response functions within Win-BUGS. Journal of Statistical Software, 43(CS-2):??, July 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v43/c02.

Weisberg:2002:DRR

[Wei02] Sanford Weisberg. Dimension reduction regression in R. Journal of Statistical Software, 7(1):1–22, ???? 2002. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v07/i01;   http://www.jstatsoft.org/
v07/i01/dr_1.0.tar.gz;   http://www.jstatsoft.org/
v07/i01/drdoc.pdf;   http://www.jstatsoft.org/v07/
i01/updates.

Weisberg:2005:LOW

[Wei05] Sanford Weisberg. Lost opportunities: Why we need a variety of statistical languages. Journal of Statistical Software, 13(1):1–12, ???? 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=115&url=v13/i01&ct=2;   http://www.
jstatsoft.org/counter.php?id=115&url=v13/i01/v13i01.
pdf&ct=1.

Wei:2012:CSP

[Wei12] Xin Wei. Code snippet: %PROC R: a SAS macro that enables native R programming in the base SAS environment. Journal of Statistical Software, 46(CS-2):??, January 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v46/c02.

Weller:2018:SRP

[Wel18] Zachary D. Weller. spTest: An R package implementing nonparametric tests of isotropy. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v083i04;   https://www.jstatsoft.org/
index.php/jss/article/view/v083i04/v83i04.pdf.

Wehrens:2012:MSV

[WF12] Ron Wehrens and Pietro Franceschi. Meta-statistics for variable selection: The R package BioMark. Journal of Statistical Software, 51(10):??, November 2012. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v51/i10.

Wooff:2000:BLP

[WG00] David Wooff and Michael Goldstein. The Bayes linear programming language [B/D]. Journal of Statistical Software, 5(2):1–19, ???? 2000. CODEN JSSOBK. ISSN 1548-

7660. URL http://maths.dur.ac.uk/stats/bd/;   http:
//www.jstatsoft.org/v05/i02; http://www.jstatsoft.
org/v05/i02/blm1.pdf;   http://www.jstatsoft.org/v05/
i02/blm2.pdf;   http://www.jstatsoft.org/v05/i02/blm3.
pdf;   http://www.jstatsoft.org/v05/i02/man.pdf;
http://www.jstatsoft.org/v05/i02/updates.

Wickham:2010:BRBb

[WG10] Hadley Wickham and Garrett Grolemund. Book review: Graphics for Statistics and Data Analysis with R. Journal of Statistical Software, 36(BR-3):??, September 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v36/b03.

Walker:2012:MTD

[WGSL12] Steven C. Walker, Guillaume Guénard, Péter Sólymos, and Pierre Legendre. Multiple-table data in R with the multitable package. Journal of Statistical Software, 51(8):??, November 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v51/i08.

West:2008:ETT

[WH08] Luke J. West and Robin K. S. Hankin. Exact tests for two-way contingency tables with structural zeros. Journal of Statistical Software, 28(11):??, November 2008. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v28/i11.

Wheadon:2014:CAC

[Whe14] Christopher Wheadon. Classification accuracy and consistency under item response theory models using the package classify. Journal of Statistical Software, 56(10):??, January 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v56/i10.

Wickham:2007:RDR

[Wic07] Hadley Wickham. Reshaping data with the reshape package. Journal of Statistical Software, 21(12):1–20, November 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v21/i12.

Wickham:2008:BRBb

[Wic08a] Hadley Wickham. Book review: A First Course in Statistical Programming with R. Journal of Statistical Software, 28(BR-3):??, December 2008. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v28/b03.

Wickham:2008:BRBa

[Wic08b] Hadley Wickham. Book review: Data Manipulation with R. Journal of Statistical Software, 26(BR-1):1–2, June 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v26/b01.

Wickham:2010:BRBa

[Wic10] Hadley Wickham. Book review: Making Sense of Data II: A Practical Guide to Data Visualization, Advanced Data Mining Methods, and Applications. Journal of Statistical Software, 34 (BR-1):??, April 2010. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v34/b01.

Wickham:2011:SAC

[Wic11] Hadley Wickham. The split-apply-combine strategy for data analysis. Journal of Statistical Software, 40(1):??, April 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v40/i01.

Wickham:2014:TD

[Wic14] Hadley Wickham. Tidy data. Journal of Statistical Software, 59(10):??, September 2014. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v59/i10.

Wiens:2004:MSP

[Wie04] Douglas P. Wiens. MATCH — a software package for robust profile matching using S-Plus. Journal of Statistical Software, 10(1):1–35, ???? 2004. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/counter.php?
id=91&url=v10/i01/manual2.pdf&ct=1.

Wilson:2009:BRB

[Wil09] Brigid Wilson. Book review: A Handbook of Statistical Analyses Using SAS (Third Edition). Journal of Statistical Software, 30(BR-2):??, April 2009. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v30/b02.

Wiley:2014:BRG

[Wil14a] Joshua F. Wiley. Book review: Growth Curve Analysis and Visualization Using R. Journal of Statistical Software, 58(BR-2):??, July 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v58/b02.

Wilson:2014:BRI

[Wil14b] Brigid Wilson. Book review: Implementing Reproducible Research. Journal of Statistical Software, 61(BR-2):??, October 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v61/b02.

Wehrens:2018:FSO

[WK18] Ron Wehrens and Johannes Kruisselbrink. Flexible selforganizing maps in kohonen 3.0. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v087i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v087i07/v87i07.pdf.

Wand:2011:ASA

[WKL11] Jonathan Wand, Gary King, and Olivia Lau. anchors: Software for anchoring vignette data. Journal of Statistical Software, 42(3):??, June 2011. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v42/i03.

Wang:2018:BRP

[WLH^{+}18] Chenguang Wang, Thomas A. Louis, Nicholas C. Henderson, Carlos O. Weiss, and Ravi Varadhan. beanz: An R package for Bayesian analysis of heterogeneous treatment effects with a graphical user interface. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v085i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v085i07/v85i07.pdf.

Wallstrom:2008:IBA

[WLK08] Garrick Wallstrom, Jeffrey Liebner, and Robert E. Kass. An implementation of Bayesian adaptive regression splines (BARS) in C with S and R wrappers. Journal of Statistical Software, 26(1):1–21, June 2008. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v26/i01.

White:2014:PBR

[WM14] Arthur White and Thomas Brendan Murphy. BayesLCA: An R package for Bayesian latent class analysis. Journal of Statistical Software, 61(13):??, November 2014. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v61/i13.

Wang:2018:CSM

[WM18] Ting Wang and Edgar C. Merkle. Code snippet: merDeriv: Derivative computations for linear mixed effects models with application to robust standard errors. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v087c01;   https://www.jstatsoft.
org/index.php/jss/article/view/v087c01/v87c01.pdf.

Wojtys:2016:CRS

[WMR16] Magorzata Wojtyś, Giampiero Marra, and Rosalba Radice. Copula regression spline sample selection models: The R package SemiParSampleSel. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v071i06;   https://www.jstatsoft.org/
index.php/jss/article/view/v071i06/v71i06.pdf.

Wallace:2017:DTR

[WMS17] Michael P. Wallace, Erica E. M. Moodie, and David A. Stephens. Dynamic treatment regimen estimation via regression-based techniques: Introducing R package DTRreg. Journal of Statistical Software, 80(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v080i02;   https://
www.jstatsoft.org/index.php/jss/article/view/v080i02/
v80i02.pdf.

West:1997:SAW

[WO97] R. Webster West and R. Todd Ogden. Statistical analysis with Webstat, Java applet for the World Wide Web. Journal of Statistical Software, 2(3):1–7, ???? 1997. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v02/i03;   http://www.jstatsoft.org/v02/i03/HTML/;
 http://www.jstatsoft.org/v02/i03/updates;   http:
//www.jstatsoft.org/v02/i03/webstat.pdf;   http://
www.jstatsoft.org/v02/i03/webstat.ps.

Wolters:2012:GAU

[Wol12] Mark A. Wolters. A greedy algorithm for unimodal kernel density estimation by data sharpening. Journal of Statistical Software, 47(6):??, April 2012. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v47/i06.

Wolters:2015:CSG

[Wol15] Mark A. Wolters. Code snippet: A genetic algorithm for selection of fixed-size subsets with application to design problems. Journal of Statistical Software, 68(CS-1):??, ???? 2015.

CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v068c01;
http://www.jstatsoft.org/index.php/jss/article/view/
v068c01/v68c01.pdf.

Woodward:2005:BBC

[Woo05] Phil Woodward. BugsXLA: Bayes for the Common Man. Journal of Statistical Software, 14(5):1–18, January 31, 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/counter.php?id=130&url=v14/i05&ct=2;
 http://www.jstatsoft.org/counter.php?id=130&url=v14/
i05/v14i05.pdf&ct=1.

Wood:2016:JAG

[Woo16] Simon N. Wood. Just another Gibbs additive modeler: Interfacing JAGS and mgcv. Journal of Statistical Software, 75(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v075i07;   https://www.jstatsoft.org/
index.php/jss/article/view/v075i07/v75i07.pdf.

Whitcher:2011:QAD

[WS11] Brandon Whitcher and Volker J. Schmid. Quantitative analysis of dynamic contrast-enhanced and diffusion-weighted magnetic resonance imaging for oncology in R. Journal of Statistical Software, 44(5):??, October 2011. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v44/i05.

Whitcher:2011:WDN

[WST11] Brandon Whitcher, Volker J. Schmid, and Andrew Thornton. Working with the DICOM and NIfTI data standards in R. Journal of Statistical Software, 44(6):??, October 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v44/i06.

Wickelmaier:2012:PPC

[WSZ12] Florian Wickelmaier, Carolin Strobl, and Achim Zeileis. Psychoco: Psychometric computing in R. Journal of Statistical Software, 48(1):??, May 2012. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v48/i01.

Wang:2011:DEM

[WW11] Xiao-Feng Wang and Bin Wang. Deconvolution estimation in measurement error models: The R package decon. Journal of Statistical Software, 39(10):??, March 2011. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v39/i10.

Wang:2009:CSC

[WWG09] Bing-Shun Wang, Xiao-Jin Wang, and Li-Kun Gong. Code snippet: The construction of a Williams design and randomization in cross-over clinical trials using SAS. Journal of Statistical Software, 29(CS-1):??, February 2009. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v29/c01.

West:2004:IS

[WWH04] R. Webster West, Yuping Wu, and Duane Heydt. An introduction to StatCrunch 3.0. Journal of Statistical Software, 9(5):??, ???? 2004. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/counter.php?id=
86&url=v09/i05/scjss/&ct=1.

Wang:2016:CSC

[WZ16]  Hong Wang and Xu Zhang.  Code snippet: Confi-

dence band for the differences between two direct adjusted survival curves. Journal of Statistical Software, 70 (??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v070c02;   https://www.jstatsoft.org/
index.php/jss/article/view/v070c02/v70c02.pdf.

Wright:2017:RFI

[WZ17] Marvin N. Wright and Andreas Ziegler. ranger: A fast implementation of random forests for high dimensional data in C++ and R. Journal of Statistical Software, 77 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v077i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v077i01/v77i01.pdf.

Wang:2020:ARP

[WZL20] Lin Wang, Wei Zhang, and Qizhai Li. AssocTests: an R package for genetic association studies. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v094i05;   https://www.jstatsoft.
org/index.php/jss/article/view/v094i05/v94i05.pdf.

Wen:2020:BRP

[WZQW20] Canhong Wen, Aijun Zhang, Shijie Quan, and Xueqin Wang. BeSS: an R package for best subset selection in linear, logistic and Cox proportional hazards models. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v094i04;   https://www.jstatsoft.
org/index.php/jss/article/view/v094i04/v94i04.pdf.

Xiao:2007:CPC

[XGY06] Yuanhui Xiao, Alexander Gordon, and Andrei Yakovlev. A C++ program for the Cramér–Von Mises two-sample test. Journal of Statistical Software, 17(8):1–15, December 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v17/i08.

Xu:2020:SPJ

[XHW20] Cong Xu, Pantelis Z. Hadjipantelis, and Jane-Ling Wang. Semi-parametric joint modeling of survival and longitudinal data: The R package JSM. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v093i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v093i02/v93i02.pdf.

Xie:2013:ARP

[Xie13] Yihui Xie. animation: An R package for creating animations and demonstrating statistical methods. Journal of Statistical Software, 53(1):??, April 2013. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v53/i01.

Xie:2014:PHH

[XM14] Liang Xie and Laurence V. Madden. %HPGLIMMIX: A highperformance SAS macro for GLMM estimation. Journal of Statistical Software, 58(8):??, June 2014. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v58/i08.

Xia:2010:CSP

[XMW10] Xiao-Qin Xia, Michael McClelland, and Yipeng Wang. Code snippet: PypeR, a Python package for using R in Python. Journal of Statistical Software, 35(CS-2):??, July 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.jstatsoft.

org/v35/c02.

Xi:2013:LTB

[XTL13] Bowei Xi, Kean Ming Tan, and Chuanhai Liu. Logarithmic transformation-based gamma random number generators. Journal of Statistical Software, 55(4):1–17, October 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v55/i04.

Xu:2006:BRD

[Xu06] Hongquan Xu. Book review: Discovering Knowledge in Data: An Introduction to Data Mining. Journal of Statistical Software, 16(BR-1):1–2, May 2006. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v16/b01.

Xiao:2015:RPT

[XWHL15] Tao Xiao, G. A. Whitmore, Xin He, and Mei-Ling Ting Lee. The R package threg to implement threshold regression models. Journal of Statistical Software, 66(8):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v066i08;
http://www.jstatsoft.org/index.php/jss/article/view/
v066i08/v66i08.pdf.

Yaffee:2007:SRS

[Yaf07] Robert Alan Yaffee. Software review: Stata 10 (Time Series and Forecasting). Journal of Statistical Software, 23(SR-1):1– 18, December 2007. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v23/s01.

Yalta:2010:BRB

[Yal10] A. Talha Yalta. Book review: Hands-On Intermediate Econometrics Using R. Journal of Statistical Software, 33(BR-1): ??, February 2010. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v33/b01.

Yan:2007:EJC

[Yan07] Jun Yan. Enjoy the joy of copulas: With a package copula. Journal of Statistical Software, 21(4):1–21, October 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v21/i04.

Yee:2010:VPC

[Yee10] Thomas W. Yee. The VGAM package for categorical data analysis. Journal of Statistical Software, 32(10):??, January 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v32/i10.

Yassouridis:2018:GCE

[YEL18] Christina Yassouridis, Dominik Ernst, and Friedrich Leisch. Generalization, combination and extension of functional clustering algorithms: The R package funcy. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v085i09;   https://www.jstatsoft.
org/index.php/jss/article/view/v085i09/v85i09.pdf.

Yaveroglu:2015:EGP

[YFK^{+}15] Ömer Nebil Yaveroglu, Sean M. Fitzhugh, Maciej Kurant, Athina Markopoulou, Carter T. Butts, and Natasa Przulj. ergm.graphlets: A package for ERG modeling based on graphlet statistics. Journal of Statistical Software, 65(12):??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v65/i12.

Ylvisaker:2016:JLS

[Ylv16] Donald Ylvisaker. Jan de Leeuw and statistics at UCLA. Journal of Statistical Software, 73(??):??, ???? 2016. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v073i09;   https://www.
jstatsoft.org/index.php/jss/article/view/v073i09/v73i09.
pdf.

Yu:2020:BRP

[YMB20] Han Yu, Janhavi Moharil, and Rachael Hageman Blair. BayesNetBP: an R package for probabilistic reasoning in Bayesian networks. Journal of Statistical Software, 94 (??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v094i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v094i03/v94i03.pdf.

Young:2010:TRP

[You10] Derek S. Young. tolerance: An R package for estimating tolerance intervals. Journal of Statistical Software, 36(5):??, August 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v36/i05.

Yang:2019:BMT

[YR19] Jingjing Yang and Peng Ren. BFDA: a MATLAB toolbox for Bayesian functional data analysis. Journal of Statistical Software, 89(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v089i02;   https://www.jstatsoft.org/
index.php/jss/article/view/v089i02/v89i02.pdf.

Yalta:2012:CSR

[YS12] A. Talha Yalta and Sven Schreiber. Code snippet: Random number generation in gretl. Journal of Statistical Software, 50(CS-1):??, August 2012. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v50/c01.

Yee:2015:VPC

[YSH15] Thomas W. Yee, Jakub Stoklosa, and Richard M. Huggins. The VGAM package for capture-recapture data using the conditional likelihood. Journal of Statistical Software, 65(5):??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v65/i05.

Yuan:2007:MLM

[Yua07] Lester L. Yuan. Maximum likelihood method for predicting environmental conditions from assemblage composition: The R package bio.infer. Journal of Statistical Software, 22(3): 1–20, September 2007. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/v22/i03.

Yuan:2011:MIU

[Yua11] Yang Yuan. Multiple imputation using SAS software. Journal of Statistical Software, 45(6):??, December 2011. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v45/i06.

Yucel:2011:SMI

[Yuc11] Recai M. Yucel. State of the multiple imputation software. Journal of Statistical Software, 45(1):??, December 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v45/i01.

Yan:2020:BRP

[YZZ^{+}20] Fangrong Yan, Liangcai Zhang, Yanhong Zhou, Haitao Pan, Suyu Liu, and Ying Yuan. BOIN: an R package for designing single-agent and drug-combination dose-finding trials using Bayesian optimal interval designs. Journal of Statistical Software, 94(??):??, ???? 2020. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v094i13;   https://www.jstatsoft.
org/index.php/jss/article/view/v094i13/v94i13.pdf.

Zambom:2017:NRP

[ZA17] Adriano Zanin Zambom and Michael G. Akritas. NonpModelCheck: An R package for nonparametric lack-of-fit testing and variable selection. Journal of Statistical Software, 77 (??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v077i10;   https://www.jstatsoft.org/
index.php/jss/article/view/v077i10/v77i10.pdf.

Zeileis:2010:EMF

[ZC10] Achim Zeileis and Yves Croissant. Extended model formulas in R: Multiple parts and multiple responses. Journal of Statistical Software, 34(1):??, April 2010. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v34/i01.

Zhang:2016:JSM

[ZCI^{+}16] Danjie Zhang, Ming-Hui Chen, Joseph G. Ibrahim, Mark E. Boye, and Wei Shen. JMFit: A SAS macro for joint models of longitudinal and survival data. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v071i03;   https://www.jstatsoft.
org/index.php/jss/article/view/v071i03/v71i03.pdf.

Zumbrunnen:2017:PPR

[ZD17] Niki Zumbrunnen and Lutz Dümbgen. pvclass: An R package for p values for classification. Journal of Statistical Software, 78(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v078i04;   https://www.jstatsoft.org/
index.php/jss/article/view/v078i04/v78i04.pdf.

Zeileis:2004:ECH

[Zei04] Achim Zeileis. Econometric computing with HC and HAC covariance matrix estimators. Journal of Statistical Software, 11(10):1–17, ???? 2004. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/counter.php?id=
106&url=v11/i10/v11i10.pdf&ct=1.

Zeileis:2006:OOC

[Zei06] Achim Zeileis. Object-oriented computation of sandwich estimators. Journal of Statistical Software, 16(9):1–16, August 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v16/i09.

Zeitler:2015:BRG

[Zei15] David W. Zeitler. Book review: Graphical Data Analysis with R, Antony Unwin, CRC Press, Boca Raton, 2015. ISBN 978-1-

4987-1523-2. 310 pp. USD 59.46. http://www.crcpress.com/9781498715232.

Journal of Statistical Software, 68(BR-5):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v068b05;
http://www.jstatsoft.org/index.php/jss/article/view/
v068b05/v68b05.pdf.

Zeitler:2016:BRA

[Zei16] David W. Zeitler. Book review: Applied Multivariate Statistics with R, Daniel Zelterman, Springer-Verlag, Switzerland, 2015. ISBN 978-3-319-14092-6. 393 pp. USD 99.00 (Hardcover), USD

69.99 (eBook). http://www.springer.com/9783319140926.

Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://www.jstatsoft.

org/index.php/jss/article/view/v071b02;   https://
www.jstatsoft.org/index.php/jss/article/view/v071b02/
v71b02.pdf.

Zeugner:2015:BMA

[ZF15] Stefan Zeugner and Martin Feldkircher. Bayesian model averaging employing fixed and flexible priors: The BMS package for R. Journal of Statistical Software, 68(4):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/jss/article/view/v068i04;
http://www.jstatsoft.org/index.php/jss/article/view/
v068i04/v68i04.pdf.

Zeileis:2020:CTM

[ZFH^{+}20] Achim Zeileis, Jason C. Fisher, Kurt Hornik, Ross Ihaka, Claire D. McWhite, Paul Murrell, Reto Stauffer, and Claus O. Wilke. colorspace: a toolbox for manipulating and assessing colors and palettes. Journal of Statistical Software, 96(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v096i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v096i01/v96i01.pdf.

Zang:2010:SAC

[ZFZ10] Yong Zang, Wing Kam Fung, and Gang Zheng. Simple algorithms to calculate asymptotic null distributions of robust tests in case-control genetic association studies in R. Journal of Statistical Software, 33(8):??, February 2010. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/v33/i08.

Zeileis:2005:PZS

[ZG05] Achim Zeileis and Gabor Grothendieck. zoo: S3 infrastructure for regular and irregular time series. Journal of Statistical Software, 14(6):1–27, May 21, 2005. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
counter.php?id=131&url=v14/i06&ct=2;   http://www.
jstatsoft.org/counter.php?id=131&url=v14/i06/v14i06.
pdf&ct=1.

Zarraga:2011:SAP

[ZG11] Amaya Zárraga and Beatriz Goitisolo. Simultaneous analysis in S-PLUS: The SimultAn package. Journal of Statistical Software, 40(11):??, April 2011. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v40/i11.

Zhao:2007:GGA

[Zha07] Jing Hua Zhao. gap: Genetic analysis package. Journal of Statistical Software, 23(8):1–18, December 2007. CODEN JS-

SOBK. ISSN 1548-7660. URL http://www.jstatsoft.org/
v23/i08.

Zhang:2014:PCB

[Zha14] Zhiyong Zhang. WebBUGS: Conducting Bayesian statistical analysis online. Journal of Statistical Software, 61(7):??, November 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v61/i07.

Zhou:2010:BRB

[Zho10] Qing Zhou. Book review: A Guide to QTL Mapping with R/qtl. Journal of Statistical Software, 32(BR-5):??, January 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v32/b05.

Zhou:2015:PVP

[Zho15] Yan Zhou. vSMC: Parallel sequential Monte Carlo in C++. Journal of Statistical Software, 62(9):??, January 2015. CO-DEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v62/i09.

Zhou:2020:SFB

[ZHZ20] Haiming Zhou, Timothy Hanson, and Jiajia Zhang. spBayesSurv: Fitting Bayesian spatial survival models using R. Journal of Statistical Software, 92(??):??, ???? 2020. CODEN JS-

SOBK. ISSN 1548-7660. URL https://www.jstatsoft.org/
index.php/jss/article/view/v092i09;   https://www.
jstatsoft.org/index.php/jss/article/view/v092i09/v92i09.
pdf.

Zivot:2011:SSM

[Ziv11] Eric W. Zivot. State space modeling using SsfPack in S+FinMetrics 3.0. Journal of Statistical Software, 41(5):??, May 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v41/i05.

Zeileis:2008:ERP

[ZK08] Achim Zeileis and Roger Koenker. Econometrics in R: Past, present, and future. Journal of Statistical Software, 27(1):??, July 2008. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v27/i01.

Zink:2012:NSI

[ZK12] Richard C. Zink and Gary G. Koch. NParCov3: a SAS/ IML macro for nonparametric randomization-based analysis of covariance. Journal of Statistical Software, 50(3):??, July 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v50/i03.

Zeileis:2020:VVV

[ZKG20] Achim Zeileis, Susanne Köll, and Nathaniel Graham. Various versatile variances: an object-oriented implementation of clustered covariances in R. Journal of Statistical Software, 95(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v095i01;   https://www.jstatsoft.org/
index.php/jss/article/view/v095i01/v95i01.pdf.

Zeileis:2008:RMC

[ZKJ08] Achim Zeileis, Christian Kleiber, and Simon Jackman. Regression models for count data in R. Journal of Statistical Software, 27(8):??, July 2008. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v27/i08.

Zhao:2018:BLM

[ZLC18] Jing Hua Zhao, Jian’an Luan, and Peter Congdon. Bayesian linear mixed models with polygenic effects. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v085i06;   https://www.jstatsoft.
org/index.php/jss/article/view/v085i06/v85i06.pdf.

Zeileis:2002:PSR

[ZLHK02] Achim Zeileis, Friedrich Leisch, Kurt Hornik, and Christian Kleiber. strucchange: an R package for testing for structural change in linear regression models. Journal of Statistical Software, 7(2):1–38, ???? 2002. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v07/i02;
http://www.jstatsoft.org/v07/i02/strucchange.pdf;
http://www.jstatsoft.org/v07/i02/strucchange_0.9-5.
tar.gz;   http://www.jstatsoft.org/v07/i02/updates.

Zammit-Mangion:2021:FRP

[ZMC21] Andrew Zammit-Mangion and Noel Cressie. FRK: An R package for spatial and spatio-temporal prediction with large datasets. Journal of Statistical Software, 98(??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/jss/article/view/v098i04;
https://www.jstatsoft.org/index.php/jss/article/view/
v098i04/v98i04.pdf.

Zamar:2007:ESI

[ZMG07] David Zamar, Brad McNeney, and Jinko Graham. elrm: Software implementing exact-like inference for logistic regression models. Journal of Statistical Software, 21(3):1–18, October 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v21/i03.

Zhu:2013:ERP

[ZP13] Weicheng Zhu and Changsoon Park. edcc: An R package for the economic design of the control chart. Journal of Statistical Software, 52(9):??, February 2013. CODEN JSSOBK. ISSN

1548-7660. URL http://www.jstatsoft.org/v52/i09.

Zhang:2016:RPR

[ZPC^{+}16] Zhengzheng Zhang, Richard M. A. Parker, Christopher M. J. Charlton, George Leckie, and William J. Browne. R2MLwiN: A package to run MLwiN from within R. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN

1548-7660. URL https://www.jstatsoft.org/index.php/
jss/article/view/v072i10;   https://www.jstatsoft.
org/index.php/jss/article/view/v072i10/v72i10.pdf.

Zhu:2021:BRP

[ZPZW21] Jin Zhu, Wenliang Pan, Wei Zheng, and Xueqin Wang. Ball: an R package for detecting distribution difference and association in metric spaces. Journal of Statistical Software, 97(??):??, ???? 2021. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v097i06;   https://www.jstatsoft.org/
index.php/jss/article/view/v097i06/v97i06.pdf.

Zhang:2016:SRP

[ZQS16] Peng Zhang, Zhenguo Qiu, and Chengchun Shi. simplexreg: An R package for regression analysis of proportional data using the simplex distribution. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v071i11;   https://www.jstatsoft.org/
index.php/jss/article/view/v071i11/v71i11.pdf.

Zeileis:2014:FGL

[ZUL14] Achim Zeileis, Nikolaus Umlauf, and Friedrich Leisch. Flexible generation of e-learning exams in R: Moodle quizzes, OLAT assessments, and beyond. Journal of Statistical Software, 58 (1):??, June 2014. CODEN JSSOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/v58/i01.

Zhao:2017:WRP

[ZWH17] He Zhao, Graham J. Williams, and Joshua Zhexue Huang. wsrf: An R package for classification with scalable weighted subspace random forests. Journal of Statistical Software, 77(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.jstatsoft.org/index.php/jss/
article/view/v077i03;   https://www.jstatsoft.org/
index.php/jss/article/view/v077i03/v77i03.pdf.